### v1

# The Effect of Personalization in FedProx: A Fine-grained Analysis on Statistical Accuracy and Communication Efficiency

Abstract
FedProx is a simple yet effective federated learning method that enables model personalization via regularization. Despite remarkable success in practice, a rigorous analysis of how such a regularization provably improves the statistical accuracy of each client’s local model hasn’t been fully established. Setting the regularization strength heuristically presents a risk, as an inappropriate choice may even degrade accuracy. This work fills in the gap by analyzing the effect of regularization on statistical accuracy, thereby providing a theoretical guideline for setting the regularization strength for achieving personalization. We prove that by adaptively choosing the regularization strength under different statistical heterogeneity, FedProx can consistently outperform pure local training and achieve a nearly minimax-optimal statistical rate. In addition, to shed light on resource allocation, we design an algorithm, provably showing that stronger personalization reduces communication complexity without increasing the computation cost overhead. Finally, our theory is validated on both synthetic and real-world datasets and its generalizability is verified in a non-convex setting.
Keywords: Personalized Federated Learning, Statistical Complexity, Communication Complexity, Minimax Optimality
1 Introduction
Federated Learning (FL) has emerged as an attractive framework for aggregating distributed data, enabling clients to collaboratively train a shared global model while preserving data privacy. In the currently prevalent paradigm (McMahan et al., 2017), FL is formulated as a finite sum minimization problem focusing on a single shared model. Nevertheless, it has been well recognized that one of the key challenges in FL is the statistical heterogeneity of the client datasets. As the participants collect their own local data, it often reflects client-specific characteristics and is not identically distributed. With high statistical heterogeneity, training a single model for all clients by minimizing their average in-sample loss becomes questionable.
To address this challenge, one solution is to relax the common model constraint and solve alternatively the following objective in FedProx (Li et al., 2020a):
$\min_{{{\bm{w}}^{(g)}},\{{{\bm{w}}^{(i)}}\}_{i\in[m]}}\sum_{i\in[m]}p_{i}\left({L}_{i}({{\bm{w}}^{(i)}},S_{i})+\frac{\lambda}{2}\|{{\bm{w}}^{(g)}}-{{\bm{w}}^{(i)}}\|^{2}\right),$ (1)
where $L_{i}({\bm{w}}^{(i)},S_{i})$ is the empirical risk of the $i$-th client, and $S_{i}$ is the local dataset of the $i$-th client. The set $\{p_{i}\}_{i\in[m]}$ is a collection of weights for the clients satisfying $p_{i}\geq 0$ and $\sum_{i}p_{i}=1$. In (1), each client $i$ learns a local personalized model ${\bm{w}}^{(i)}$ by fitting its local data $S_{i}$, while enabling collaborative learning by shrinking all local models towards a common global one ${\bm{w}}^{(g)}$ with intensity controlled by $\lambda$. The smaller $\lambda$ is, the weaker the coupling of the local models the formulation would enforce thus the higher personalization is. Adjusting the strength of regularization controls the extent to which the ${\bm{w}}^{(i)}$’s is adapted to client $i$’s local data, thus the degree of personalization.
The promising empirical performance of FedProx in learning under heterogeneous data (Li et al., 2020a; Donevski et al., 2021; He et al., 2021; An et al., 2023) has led to a sequence of works studying the problem from a pure computation perspective, aiming at developing efficient algorithms under various problem assumptions to reach a (local) solution (Hanzely & Richtárik, 2020; T Dinh et al., 2020; Li et al., 2021; Acar et al., 2021; Yuan & Li, 2022). However, the fundamental question of how to characterize the accuracy of the models produced by (1) is largely unaddressed. As such, for a given set of clients, it is not even clear if introducing personalization will be beneficial, outperforming the baselines given by training a single global model or pure local training. In addition, communication cost in FL is a critical bottleneck in resource-constrained environments. Therefore, understanding when personalization will improve both training efficiency and statistical accuracy is of paramount importance. We address these questions by providing a fine-grained analysis of the effect of personalization in FedProx. Our contribution is summarized below:
$\bullet$ Nearly Optimal Statistical Accuracy. We analyze the statistical accuracy of the objective in (1), taking statistical heterogeneity into account. For each of the models, we establish a state-of-the-art convergence rate $\mathcal{O}(1/(mn))+\mathcal{O}(1/n\wedge R/n^{1/2})$, where $R$ characterizes the statistical heterogeneity and $n$ is the sample size on each client. The first term $\mathcal{O}(1/(mn))$ represents the sample complexity of training a single global model under homogeneity, which is attained by (1) when $R$ is close to zero, highlighting the benefit of FL. For the second term, the first component $\mathcal{O}(1/n)$ is the sample complexity of purely local training and the second component $\mathcal{O}(R/n^{1/2})$ demonstrates the flexibility of FedProx under different statistical heterogeneity, implying the statistical accuracy of FedProx never gets worse than purely local training’s. The fine-grained analysis demonstrates that with an adaptive choice of personalization degree characterized by $\lambda$, FedProx achieves a nearly minimax-optimal statistical accuracy across different levels of statistical heterogeneity.
$\bullet$ Provably Enhanced Communication Efficiency. As the objective (1) could be reformulated as a bi-level problem, we propose an algorithm for (1) with an explicit computation cost assessment. We establish the communication complexity as $\mathcal{O}(\kappa\frac{\lambda+\mu}{\lambda+L}\log\frac{1}{\varepsilon})$ and the local computation complexity in each communication round as $\mathcal{\widetilde{O}}({\frac{\lambda+L}{\lambda+\mu}})$. With higher personalization degree, communication complexity would decrease and the computation complexity per round would increase, demonstrating that there is a trade-off between computation and communication costs. Our analysis concludes that model personalization can reduce communication overhead without incurring additional computation expense.
$\bullet$ Empirical Validation. We conduct numerical studies on logistic regression with synthetic data and real-world data under our problem assumptions. The results corroborate our theoretical findings. We further test the numerical performance of (1) for training CNNs, showing empirically the results, although proved under the convexity assumption, generalize to nonconvex losses.
2 Related Works
In recent years, federated learning has become an attractive solution for training models locally on distributed clients, rather than transferring data to centralized data centers (Mammen, 2021; Wen et al., 2023; Beltrán et al., 2023). As each client generates its local data, statistical heterogeneity is common with data being non-identically distributed between clients (Li et al., 2020a; Ye et al., 2023). Given the variability of data in network, model personalization is a natural strategy used to improve learning accuracy for each client. Formulations enabling model personalization have been studied independently from multiple fields. For example, meta-learning (Chen et al., 2018; Jiang et al., 2019; Khodak et al., 2019; Fallah et al., 2020) assumes all models follow a common distribution. By minimizing the average loss, these methods aim to learn a meta-model that generalizes well to unseen new tasks. Multi-task learning methods (Agarwal et al., 2020; Liang et al., 2020) and transfer learning methods (Li et al., 2022; He et al., 2024a; b) study formulation similar to (1), but the statistical rate is established for either the average of all models or only the target model. Therefore, it is not clear if all clients can benefit from the procedure and learn a more accurate model than purely local training. Representation learning (Zhou et al., 2020; Wang et al., 2024), on the other hand, focuses on a different model structure where the local models can be represented as the composition of two parts, with one common to all clients and the other specific to each client.
In the FL community, personalized FL methods can be broadly divided into two main strands. Different from the works mentioned previously, studies here focus primarily on the properties of the iterates computed by the algorithms. One line of works (Arivazhagan et al., 2019; Liang et al., 2020; Singhal et al., 2021; Collins et al., 2021) is based on the representation learning formulation, where both the algorithmic convergence and statistical accuracy of the iterates have been investigated. Another line of works (Smith et al., 2017; Li et al., 2020a; Hanzely & Richtárik, 2020; Hanzely et al., 2020; 2023), including FedProx, achieves personalization by relaxing the requirement of learning a common global model through regularization techniques. In particular, algorithms and complexity lower bounds specific to the FedProx objective were studied in works (Li et al., 2020a; Hanzely et al., 2020; T Dinh et al., 2020; Hanzely & Richtárik, 2020; Li et al., 2021), see Table 2 for a detailed comparison. These works study the problem from a pure optimization perspective and cannot show personalization improves learning accuracy.
There are only a few recent works we are aware of studying the statistical accuracy of regularization-based PFL in the form of (1). Specifically, Cheng et al. (2023) investigates the asymptotic behavior of the FedProx estimator under an over-parameterized linear regression model; neither a finite-sample rate nor algorithm is provided. Chen et al. (2023b) studies the FedProx estimator under rather strong assumptions like bounded parameter space. By defining a new notion of algorithmic stability, they establish the non-asymptotic sample complexity for FedProx and induce minimax bound. However, even with some overly strong assumptions, their analysis cannot match the statistical lower bound in most cases.
Notation. $[n]$ denotes the set of integers $\{1,2,\cdots,n\}$ for any positive integer $n$. Let $\|\cdot{}\|_{2}$ denote the $\ell_{2}$ vector norm and the Frobenius matrix norm. For two non-negative sequences $a_{n}$, $b_{n}$, we denote $a_{n}=\mathcal{O}(b_{n})$ as $a_{n}/b_{n}\xrightarrow{}{1}$. $\tilde{\mathcal{O}}(\cdot)$ hides logarithmic factors.
3 Preliminaries
Considering there are $m$ clients, each client has a collection of $i.i.d$ data $S_{i}=\{z_{ij}\}_{j\in[n_{i}]}$ over possibly different data distributions $\mathcal{D}_{i}$. For each client, the ultimate goal is to find the solution that minimizes the local risk, i.e. ${{\bm{w}}_{\star}^{(i)}}\in\arg\min_{{\bm{w}}}\>\mathbb{E}_{z\sim\mathcal{D}_{i}}\ell({\bm{w}},z)$. Here, $\ell({\bm{w}},z)$ denotes the loss function of interest and ${{\bm{w}}_{\star}^{(i)}}\in\mathbb{R}^{d}$ is called the ground truth local model for client $i$. In this paper, we assume the model dimension $d$ is finite. Under such a setup, two commonly adopted approaches are pure local training (LocalTrain) and pure global training (GlobalTrain):
[…]
where $L_{i}({\bm{w}},S_{i}):=\sum_{j=1}^{n_{i}}\ell({\bm{w}},z_{ij})/n_{i}$ denotes the empirical risk of client $i$. In LocalTrain, each client individually minimizes its loss, whereas in GlobalTrain, all clients collaboratively minimize their average loss. Compared to LocalTrain, GlobalTrain integrates information from all clients to improve the learning performance of each client. When there are sufficient similarities between the data distribution $\mathcal{D}_{i}$ of different clients, GlobalTrain would be a better choice as more samples are included. However, if $\mathcal{D}_{i}$’s are highly unrelated, LocalTrain will be a more robust choice.
In practice, the data distribution is neither totally the same nor completely unrelated across clients. To handle the wide middle ground of the two extremes, FedProx solves the following problem
$\min_{{{\bm{w}}^{(g)}},\{{{\bm{w}}^{(i)}}\}_{i\in[m]}}\sum_{i=1}^{m}p_{i}\left({L}_{i}({{\bm{w}}^{(i)}},S_{i})+\frac{\lambda}{2}\|{{\bm{w}}^{(g)}}-{{\bm{w}}^{(i)}}\|^{2}\right).$
The key feature of Fedprox is the incorporation of a proximal (regularization) term, which controls the personalization degree. As $\lambda\rightarrow 0$, the FedProx objective increasingly behaves like LocalTrain with higher personalization degree and reduces to $m$ separate LocalTrain problems in the end. On the other hand, as $\lambda\rightarrow\infty$, the local models ${\bm{w}}^{(i)}$ are shrunk towards the global model ${\bm{w}}^{(g)}$ with lower personalization degree and eventually degenerate to optimize a single shared model. Thus, by adjusting $\lambda$, FedProx interpolates between LocalTrain and GlobalTrain, allowing for a flexible choice of personalization degree to achieve higher modeling accuracy. It’s noticeable that this choice of $\lambda$ also affects the optimization process of FedProx. With larger $\lambda$, more frequent communication between clients is expected to ensure faster convergence, as local updates are more tightly coupled to the global model. In contrast, smaller $\lambda$ value allows for more independent updates, reducing the need for communication. In the next few sections, we provide a theoretical guarantee for the intuition provided above.
In Section 4, we analyze when the personalization level, characterized by $\lambda$, would improve the statistical accuracy of both the global model ${\bm{w}}^{(g)}$ and the local models $\{{{\bm{w}}^{(i)}}\}_{i\in[m]}$. This is in sharp contrast to most of the existing works where the error is established for the sum of them. In Section 5, we explore how such a degree of personalization will affect communication complexity.
4 When will Personalization Improve Statistical Accuracy
In this section, we analyze how the choice of $\lambda$, or the degree of personalization, would improve the statistical accuracy of the solution obtained from the FedProx problem (1). Following the existing literature (Li et al., 2023; Chen et al., 2023a; b; Duan & Wang, 2023), we define the statistical heterogeneity as the Euclidean distance between the true local models ${{\bm{w}}_{\star}^{(i)}}$ and the true global model ${{\bm{w}}_{\star}^{(g)}}:=\sum_{i=1}^{m}p_{i}{{\bm{w}}_{\star}^{(i)}}$. Specifically, we adopt the following assumption:
Assumption 1 (Statistical Heterogeneity).
There exists a nonnegative constant $R$ such that we have $\|{{\bm{w}}_{\star}^{(i)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}\leq R^{2},\forall i\in[m]$.
Intuitively, a larger $R$ indicates a greater divergence between the clients’ data distributions, thus stronger statistical heterogeneity, and vice versa. Therefore, in the FedProx objective, the choice of $\lambda$ should be adjusted according to $R$. Next, we introduce two standard assumptions.
Assumption 2 (Strong Convexity).
The loss function $\ell(\cdot,z)$ is $\mu$-strongly convex for each local sample average, i.e. $\sum_{j\in[n_{i}]}\ell(x,z_{ij})/n_{i}\geq\sum_{j\in[n_{i}]}\ell(y,z_{ij})/n_{i}+\langle\sum_{j\in[n_{i}]}\nabla\ell(y,z_{ij})/n_{i},y-x\rangle+\frac{\mu}{2}\|x-y\|^{2},\quad\forall x,y\in\mathbb{R}^{d},i\in[m].$
Assumption 3 (Bound Gradient Variance at Optimum).
There exists a nonnegative constant $\rho$ such that $\forall i\in[m]$, we have $\mathbb{E}_{z\sim\mathcal{D}_{i}}\|\nabla\ell({{\bm{w}}_{\star}^{(i)}},z)\|^{2}\leq\rho^{2}$.
Here, the strongly convex assumption is used to establish the estimation error and is widely adopted in the statistical analysis of regularization-based personalized federated learning (T Dinh et al., 2020; Deng et al., 2020; Hanzely & Richtárik, 2020; Hanzely et al., 2020) and statistical analysis in federated learning (Chen et al., 2023b; Cheng et al., 2023). Assumption 3 is used to quantify the observation noise, and is standard in the literature (Li et al., 2020b; Chen et al., 2023b).
Next, we establish a statistical rate of the optimal solution to the Fedprox problem (1) with a fine-grained analysis. To facilitate the analysis, we first define some notations. Define $\{\widetilde{\bm{w}}^{(i)}\}_{i=1}^{m}$ as the optimal solution for local models. A check on the optimality condition implies that the optimal solution for the global model is $\widetilde{\bm{w}}^{(g)}=\sum^{m}_{i=1}p_{i}\widetilde{\bm{w}}^{(i)}$. In this paper, we use global statistical error $\mathbb{E}\|\widetilde{\bm{w}}^{(g)}-{{\bm{w}}_{\star}^{(g)}}\|^{2}$ and local statistical error $\mathbb{E}\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}$ as the measure of statstical accuracy. Here the expectation is taken with respect to all data. Note that some existing literature adopts alternative metrics, such as average excess risk and individual excess risk (Chen et al., 2023b) to measure the statistical error. They are equivalent to our metrics under the assumptions used in this paper. We first provide a lemma to cast light on the effect of $\lambda$ on the statistical error. See Appendix A.2 for the proof.
Lemma 1.
Suppose Assumptions 1 and 2 hold. We have
$\underbrace{\sum_{i\in[m],j\in[n_{i}]}\frac{p_{i}}{n_{i}}\left(\mu\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}+\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\rangle\right)}_{\text{model estimation error}}\leq\frac{\lambda}{2}\underbrace{(R^{2}-\sum_{i\in[m]}p_{i}\|\widetilde{\bm{w}}^{(i)}-\widetilde{\bm{w}}^{(g)}\|^{2})}_{\text{heterogeneity estimation error}}.$ (3)
Lemma 1 establishes a relationship between the local model estimation error and the statistical heterogeneity. On the right-hand side of (3), the first term $R$, which is larger than $\|{{\bm{w}}_{\star}^{(i)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}$, measures the statistical heterogeneity; the second term, $\sum_{i=1}^{m}p_{i}\|\widetilde{\bm{w}}^{(i)}-\widetilde{\bm{w}}^{(g)}\|^{2}$, can be understood as the estimator of such heterogeneity. Lemma 1 implies that to achieve a tight bound on the estimation error, we should choose a $\lambda$ that “minimizes” the right-hand side of (3). When $R$ is large, one straightforward choice is $\lambda=\mathcal{O}(1/(nR^{2}))$, which will be shown to yield an error bound of order $\mathcal{O}(1/n)$, matching the performance of purely local training and is independent of $R$. When $R$ is small, $\lambda$ should be carefully chosen to obtain the full sample complexity $\mathcal{O}(1/(mn))$. In the following theorem, we provide such a choice and establish a state-of-the-art statistical rate.
Theorem 1.
Suppose Assumptions 1, 2 and 3 hold and $n_{i}=n,\forall i\in[m]$. Consider the FedProx objective in (1) with $p_{i}=\frac{1}{m},\forall i\in[m]$. If $\lambda=\frac{\rho}{\sqrt{n}R}\mathbb{I}\{R\leq n^{-1/2}\}+\frac{\rho^{2}}{nR^{2}}\mathbb{I}\{R>n^{-1/2}\}$, the global model $\widetilde{\bm{w}}^{(g)}$ and local models $\{\widetilde{\bm{w}}^{(i)}\}_{i=1}^{m}$ obtained from solving the objective (1) satisfy
$\mathbb{E}\left\|\widetilde{\bm{w}}^{(g)}-{{\bm{w}}_{\star}^{(g)}}\right\|^{2}\leq\frac{C_{1}}{N}+C_{2}\left(\frac{1}{n}\wedge\frac{R}{\sqrt{n}}\right),$ (4)
$\mathbb{E}\left\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\right\|^{2}\leq\frac{C_{3}}{N}+C_{4}(\frac{1}{n}\wedge\frac{R}{\sqrt{n}})+16(\frac{\rho^{2}}{\mu^{2}n}\wedge R^{2})\quad\text{ for all }i\in[m],$ (5)
where $N=mn$ and
$\begin{split}C_{1}&=\frac{\rho^{2}}{\mu^{2}}=\frac{C_{3}}{8},C_{2}=\big{(}\frac{(1+\sqrt{1+2\mu})\rho}{2\mu}\big{)}^{2}\vee\frac{2\rho}{\mu},C_{4}=\big{(}\frac{\rho(1+(1+4\mu(C_{2}+1))^{\frac{1}{2}})}{2\mu}\big{)}^{2}\wedge(8C_{2}+\frac{2\rho}{\mu}).\end{split}$ (6)
Proof.
See Appendix A.2 for (4) and Appendix A.3 for (5). ∎
Theorem 1 shows that with an appropriate choice of $\lambda$, both the global and local models achieve a fast statistical rate across a wide range of $R$. To interpret the result, first we analyze the error bound provided in (4) and (5). The two error bounds have the same order and can be decomposed into two key components. The first term, of order $\mathcal{O}(N^{-1})$, captures the benefit of collaborative learning, as $N=mn$ is the total sample size across $m$ clients. This term reflects the sample efficiency gained through collaborative learning. The second term, of order $\mathcal{O}(n^{-1}\wedge n^{-1/2}R)$, measures the cost of statistical heterogeneity. As $R$ increases, the convergence rate gradually degrades, transitioning from $\mathcal{O}(N^{-1})$ to $\mathcal{O}(n^{-1})$, showing that the error bound of FedProx interpolates between the two extremes: LocalTrain and GlobalTrain.
Next, to better understand how such an interpolation is achieved through an adaptive personalization degree, we may examine the two extreme cases: as the statistical heterogeneity $R\to\infty$, and $\lambda\to 0$, leading to a high degree of personalization. The statistical error bound converges to $\mathcal{O}(n^{-1})$, matching the rate of LocalTrain; as the statistical heterogeneity $R\to 0$, and $\lambda\to\infty$, leading to a low degree of personalization. The statistical error bound converges to $\mathcal{O}(N^{-1})$, matching the rate of GlobalTrain. As $R\in(0,\infty)$, $\lambda$ is adjusted correspondingly, making FedProx always perform no worse than LocalTrain while benefiting from global training when clients share similarities.
Table 1: Results Comparison. $D:=\sup_{\bm{w},\bm{w}^{\prime}\in\mathcal{D}}\|\bm{w}-\bm{w}^{\prime}\|$ is diameter of the parameter space ($D\geq R$) and $\|\ell\|_{\infty}=\inf\{M\in\mathbb{R}:\ell(\cdot,z)\leq M,\text{for any }z\}$ is an uniform upper bound on the loss function. Statistical rate refers to $\mathbb{E}\|{{\bm{w}}_{\star}^{(i)}}-\widetilde{\bm{w}}^{(i)}\|^{2}$, with constants neglected.
Source Assumption Paradigm Statistical Rate
$R>\frac{1}{\sqrt{n}}$ $\frac{1}{m\sqrt{n}}<R\leq\frac{1}{\sqrt{n}}$ $R\leq\frac{1}{m\sqrt{n}}$
(Chen et al., 2023b) A1,2,3, $D\vee\|\ell\|_{\infty}\leq C$ Two-stage $\frac{D^{2}+\|\ell\|_{\infty}}{n}$ 1 $\frac{\|\ell\|_{\infty}}{\sqrt{n}}R$ $\frac{\|\ell\|_{\infty}}{\sqrt{m}n}$
Ours A1,2,3 One-stage $\frac{1}{n}$ $\frac{1}{\sqrt{n}}R$ $\frac{1}{mn}$
Lower Bound - - $\frac{1}{n}$ $R^{2}\vee\frac{1}{mn}$ $\frac{1}{mn}$
We now compare our theoretical results with the result up to date (Chen et al., 2023b), as summarized in Table 1. Our contribution can be summarized into three aspects:
$\bullet$ Our results are derived under more realistic assumptions. Chen et al. (2023b) imposes additional stringent conditions including a bounded parameter space $D$ and a uniform bound on the loss function $\ell(\cdot,z)$. Such conditions are difficult to satisfy in practice. For example, even a simple linear regression model with sub-Gaussian covariates would violate these assumptions.
$\bullet$ Our result is established for both global and local models with the same choice of $\lambda$. In contrast, Chen et al. (2023b) proposes a two-stage process where the global model is first estimated, followed by an additional local training phase with a different choice of $\lambda$. This two-step approach is likely the artifact of their analysis, as existing work on FedProx also shows that a single-stage optimization with a chosen $\lambda$ suffices to achieve the desired performance (T Dinh et al., 2020).
$\bullet$ We establish a state-of-the-art nearly minimax statistical rate. As shown in Table 1, in a heterogeneous case when $R\geq n^{-1/2}$, Chen et al. (2023b) establishes a rate that is at least of the order $\mathcal{O}\left(n^{-1}R^{2}\right)$ as $D\geq R$. This bound becomes increasingly loose with larger $R$. In contrast, our bound is $\mathcal{O}\left(n^{-1}\right)$, which shows that properly-tuned FedProx is always no worse than LocalTrain, independent of $R$. In a homogeneous case when $R\leq m^{-1}n^{-1/2}$, they can only establish a rate slower than $\mathcal{O}\left(1/(\sqrt{m}n)\right)$, while our result achieves a rate of $\mathcal{O}\left(1/(mn)\right)$, leveraging all $mn$ samples and matching the rate of GlobalTrain on the IID data. Compared to the established lower bound (Chen et al., 2018), our result is nearly minimax optimal. Specifically, when $R>n^{-1/2}$ and $R\leq m^{-1}n^{-1/2}$, our result matches the lower bound. There is a small gap exists in the case when $m^{-1}n^{-1/2}<R\leq n^{-1/2}$, where we achieve a rate of $\mathcal{O}(n^{-1/2}R)$ whereas the lower bound is $\mathcal{O}(R^{2})$. Such a gap is due to the cost of estimating the model difference $\{{{\bm{w}}^{(g)}}-{{\bm{w}}^{(i)}}\}_{i=1}^{m}$ when $\lambda$ is finite (c.f. Lemma 1). The improvement of this gap is non-trivial, which we defer to future work.
5 When will Personalization Improve Communication Efficiency
Notice that if we define the local objective of each client $i$ as $h_{i}({{\bm{w}}^{(i)}},{\bm{w}}^{(g)}):={L}_{i}({{\bm{w}}^{(i)}},S_{i})+\frac{\lambda}{2}\|{{\bm{w}}^{(g)}}-{{\bm{w}}^{(i)}}\|^{2}$, then the FedProx problem given by (1) can be rewritten in the following bilevel (iterated minimization) form:
$\min_{{\bm{w}}^{(g)}}F({\bm{w}}^{(g)}):=\frac{1}{m}\sum_{i=1}^{m}F_{i}({\bm{w}}^{(g)}),\quad\text{where}\quad F_{i}({\bm{w}}^{(g)}):=\min_{{{\bm{w}}^{(i)}}}~{}h_{i}({{\bm{w}}^{(i)}},{\bm{w}}^{(g)}).$ (7)
The reformulation (7) is a finite-sum minimization problem, with each component $F_{i}$ being the Moreau envelope (Moreau, 1965; Yosida, 1964) of $L_{i}$. Let ${\bm{w}}_{\star}^{(i)}({\bm{w}}^{(g)})$ be the minimizer of $h_{i}(\,\cdot\,,{\bm{w}}^{(g)})$, i.e.,
${\bm{w}}_{\star}^{(i)}({\bm{w}}^{(g)}):=\operatorname{prox}_{L_{i}/\lambda}({\bm{w}}^{(g)})=\operatorname*{arg\,min}_{{{\bm{w}}^{(i)}}}~{}h_{i}({{\bm{w}}^{(i)}},{\bm{w}}^{(g)}).$ (8)
As a standard argument about the convergence rate (T Dinh et al., 2020; Li et al., 2021; Chen et al., 2023b), we assume the loss function $\ell$ is $L$-smooth.
Assumption 4 (Smoothness).
The loss function $\ell(\cdot,z)$ is L-smooth, i.e. for any $x,y\in\mathbb{R}^{d}$
$\|\nabla\ell(x,z)-\nabla\ell(y,z)\|\leq L\|x-y\|,\quad\forall z.$ (9)
Based on Assumption 4 and the new notations in (7) and (8), the following lemma describes the optimizing properties for $F_{i}$, $h_{i}$ and ${\bm{w}}_{\star}^{(i)}$. The proof can be found in Appendix B.2.
Lemma 2.
Under Assumption 2 and 4, $F_{i}$ is $\mu_{g}$-strongly convex and $L_{g}$-smooth, with $\mu_{g}=\frac{\lambda\mu}{\lambda+\mu}$ and $L_{g}=\frac{\lambda L}{\lambda+L}$, each $h_{i}$ is $\mu_{\ell}$-strongly convex and $L_{\ell}$-smooth, with $\mu_{\ell}=\mu+\lambda$ and $L_{\ell}=L+\lambda$, and the mapping $\bm{w}_{\star}^{(i)}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is $L_{w}$-Lipschitz with $L_{w}=\frac{\lambda}{\lambda+\mu}$.
Next, we recall the properties of the Moreau envelope to guide the algorithm design.
Lemma 3 (Lemaréchal & Sagastizábal (1997)).
Under Assumption 2, each $F_{i}:\mathbb{R}^{d}\to\mathbb{R}$ is continuously differentiable, and the gradient is given by $\nabla F_{i}({\bm{w}}^{(g)})=\lambda({\bm{w}}^{(g)}-{\bm{w}}_{\star}^{(i)}({\bm{w}}^{(g)})).$
From Lemma 2 and Lemma 3, it suggests applying a simple gradient algorithm to optimize ${\bm{w}}^{(g)}$:
$\begin{split}{\bm{w}}_{t+1}^{(g)}&={\bm{w}}_{t}^{(g)}-\gamma\cdot\nabla F({\bm{w}}_{t}^{(g)})={\bm{w}}_{t}^{(g)}-\gamma\lambda\cdot\frac{1}{m}\sum_{i=1}^{m}\left({\bm{w}}_{t}^{(g)}-{\bm{w}}_{\star}^{(i)}({\bm{w}}_{t}^{(g)})\right),\end{split}$ (10)
where $\bm{w}_{t}^{(g)}$ is the update at the communication round $t$, $\gamma>0$ is a step size to be properly set (cf. Theorem 2). To implement (10), in each communication round $t$ the server broadcasts ${\bm{w}}_{t}^{(g)}$ to all clients, then each client $i$ updates its local model ${\bm{w}}_{\star}^{(i)}({\bm{w}}_{t}^{(g)})$ and upload to the server.
Note that executing (10) requires each client $i$ computing the minimizer ${\bm{w}}_{\star}^{(i)}({\bm{w}}_{t}^{(g)})$. In general, the subproblem (8) does not have a closed-form solution. Therefore, computing the exact gradient $\nabla F$ will incur a high computation cost as well as high latency for the learning process. Leveraging recent advancements in bilevel optimization algorithm design (Ji et al., 2022), we address this issue by approximating ${\bm{w}}_{\star}^{(i)}({\bm{w}}_{t}^{(g)})$ with a finite number of $K$ gradient steps. Specifically, we let each client $i$ maintain a local model ${\bm{w}}_{t,k}^{(i)}$. Per communication round $t$, ${\bm{w}}_{t,k}^{(i)}$ is initialized to be ${\bm{w}}_{t,0}^{(i)}={\bm{w}}_{t-1,K}^{(i)}$ as a warm start, and updated according to
$\begin{split}{\bm{w}}_{t,k+1}^{(i)}={\bm{w}}_{t,k}^{(i)}-\eta\nabla h_{i}({\bm{w}}_{t,k}^{(i)},{\bm{w}}_{t}^{(g)}),\quad\forall\,k=0,\ldots,K-1,\end{split}$ (11)
where $\nabla h_{i}({\bm{w}}_{t,k}^{(i)},{\bm{w}}_{t}^{(g)})$, for notation simplicity, denotes the partial gradient of $h_{i}$ with respect to ${\bm{w}}^{(i)}$. The overall procedure is summarized in Alg. 1.
Input: Initial global model ${\bm{w}}_{1}^{(g)}$, initial local models $\{{\bm{w}}_{0,K}^{(i)}\}_{i\in[m]}$, global rounds $T$, global step sizes $\gamma$, local rounds $K$, local step sizes $\eta$
Output: Local models $\{{\bm{w}}_{T,K}^{(i)}\}_{i\in[m]}$ and global model ${\bm{w}}_{T}^{(g)}$
for $t=1,\dots,T$ do
       The server sends ${\bm{w}}_{t}^{(g)}$ to client $i$, $\forall i\in[m]$
       Set ${\bm{w}}_{t,0}^{(i)}={\bm{w}}_{t-1,K}^{(i)}$;
       for $k=0,\dots,K-1$ do
             ${\bm{w}}_{t,k+1}^{(i)}={\bm{w}}_{t,k}^{(i)}-\frac{\eta}{n_{i}}\sum_{j\in[n_{i}]}\left\{\nabla\ell({\bm{w}}_{t,k}^{(i)},z_{i,j})+\lambda\left({\bm{w}}_{t,k}^{(i)}-\bm{w}_{t}^{(g)}\right)\right\}$
      Push ^$\widehat{\nabla}F_{i}(\bm{w}_{t}^{(g)})=\lambda(\bm{w}_{t}^{(g)}-{\bm{w}}_{t,K}^{(i)})$ to the server;
      $\bm{w}_{t+1}^{(g)}\leftarrow\bm{w}_{t}^{(g)}-\frac{\gamma}{m}\sum_{i\in[m]}\widehat{\nabla}F_{i}(\bm{w}_{t}^{(g)})$;
Algorithm 1 One-stage FedProx
Remark 1.
Alg. 1 is a customization of the general bi-level problem (Ji et al., 2022). While the algorithm design is not new, our contribution lies in analyzing how the personalization degree influences the trade-off between communication and local computation leveraging the FedProx problem structure. For simplicity, we focus on the full gradient update in our analysis; both the algorithm and the analysis can be extended to the stochastic setting with mini-batch gradients, with partial participation and incorporating variance reduction techniques.
Theorem 2.
Suppose Assumptions 2 and 4 hold. Let $\{{\bm{w}}_{t}^{(g)}\}_{t\geq 0}$ and $\{{\bm{w}}_{t,K}^{(i)}\}_{t\geq 0}$ be the sequence generated by Algorithm 1 with $\gamma<1/L_{g}$, $\eta\leq 1/L_{\ell}$, and the inner loop iteration number satisfying
$\Big{(}2+64L_{w}^{2}(1/\mu_{g})^{2}\lambda^{2}\Big{)}(1-\eta\mu_{\ell})^{K}\leq(1-\gamma L_{g})^{4},$ (12)
then ${\bm{w}}_{t}^{(g)}$ converges to $\widetilde{\bm{w}}^{(g)}$ linearly at rate $21-(\gamma\mu_{g})/2-(\gamma\mu_{g})^{2}/$ 2, and ${\bm{w}}_{t,0}^{(i)}$ converges to $\widetilde{\bm{w}}^{(i)}$ linearly at the same rate for any $i\in[m]$.
Proof.
See Appendix B.3 and B.5. ∎
Theorem 2 shows under the step size constraint $\gamma<1/L_{g}$, ${\bm{w}}_{t}^{(g)}$ converges to $\tilde{{\bm{w}}}^{(g)}$ at the same rate as the iteration (10) with exact gradient computation. Furthermore, the larger $\gamma$ is, the larger $K$ should be chosen so as to satisfy (12). This indicates a tradeoff between communication and computation: the fewer communication rounds are performed, then the more local computation is required for the algorithm to converge.
To better interpret the result and obtain insights on the influence of $\lambda$ on the convergence, we provide the communication and computation complexity under specific choices the algorithm’s tuning parameters, as described in the following Corollary 1. See Appendix B.4 for the proof.
Corollary 1.
In the same setting as Theorem 2, if choosing the step size $\eta=1/L_{\ell}=(\lambda+L)^{-1}$, $\gamma=1/(2L_{g})=\frac{\lambda+L}{2\lambda L}$, and the inner loop iteration number $K=\tilde{\mathcal{O}}\left(\frac{\lambda+L}{\lambda+\mu}\right)$, then the communication complexity and computation complexity for Algorithm 1 to find an $\varepsilon$-solution (i.e., $\|{\bm{w}}_{T}^{(g)}-\hat{{\bm{w}}}^{(g)}\|^{2}\leq\varepsilon$ and $\|{\bm{w}}_{T,0}^{(i)}-\hat{{\bm{w}}}^{(i)}\|^{2}\leq\varepsilon$) are as follows ($\kappa:=L/\mu$):
(i) # communication $\begin{split}&\text{(i) \# communication rounds}=\mathcal{O}\big{(}\kappa_{g}\log(1/\varepsilon)\big{)}=\mathcal{O}\left(\kappa\cdot\frac{\lambda+\mu}{\lambda+L}\cdot\log\frac{1}{\varepsilon}\right),\end{split}$ (13)
(ii) # gradient $\begin{split}&\text{(ii) \# gradient evaluations}=\tilde{\mathcal{O}}\left(\kappa\cdot\log\frac{1}{\varepsilon}\right).\end{split}$ (14)
$\bullet$ Communication cost. Corollary 1 shows as $\lambda$ increases from 00 to $\infty$, the communication complexity increases from $\tilde{\mathcal{O}}(1)$ to $\tilde{\mathcal{O}}(\kappa)$. The result reveals a higher personalization degree requires less communication cost to collaboratively learn the models. In the full personalization limit with $\lambda\rightarrow 0$, the order of the communication cost reduces to near constant $\tilde{\mathcal{O}}(1)$, which corresponds to pure local training. It is worth noting that even personalization can reduce communication complexity, as our ultimate goal is to accurately estimate each model so we can not simply pursue higher communication efficiency.
$\bullet$ Computation cost. In contrast to the communication complexity, the gradient evaluations per round $K$ decrease with $\lambda$ increasing. Therefore, a higher degree of personalization requires a more intensive update of the local models once receiving a new global model ${\bm{w}}^{(g)}$, so as to adapt to each client’s local data set. Note that although a small $\lambda$ requires a larger $K$, the overall computation cost given by (LABEL:eq:comp-cost) is independent of $\lambda$. As such, model personalization (small $\lambda$) can provably reduce the communication cost without any extra computation overhead.
[b] Methods Convexity Bounded Domain Computation Cost per Round Communication Cost FedProx (Li et al., 2020a) cvx (1) $\mathcal{O}(\Delta/(\rho\varepsilon))$ pFedMe (T Dinh et al., 2020) s-cvx (2) $\mathcal{O}(1/{(mR\varepsilon)})$ Ditto (Li et al., 2021) s-cvx $\mathcal{O}(1)$ $\mathcal{O}(1/\varepsilon)$ L2GD (Hanzely & Richtárik, 2020) s-cvx $\mathcal{O}(1+\frac{L}{\lambda})$ $\mathcal{O}({\frac{2\lambda\kappa}{\lambda+L}\log(\frac{1}{\varepsilon})})$ Algorithm 3 (Chen et al., 2023b) s-cvx $\mathcal{O}(\lambda/\varepsilon)$ $\mathcal{O}((\lambda\vee 1)/\varepsilon)$ Our Result s-cvx $\mathcal{O}(\frac{\lambda+L}{\lambda+\mu})$ $\mathcal{O}(\kappa\frac{\lambda+\mu}{\lambda+L}\log{\frac{1}{\varepsilon}})$
Table 2: Comparison of the communication and computation complexity in existing works of analyzing the objective in (1) (cvx denotes convexity and s-cvx denotes strong convexity).
  • (1)
    Controlled by the precision of inexact solution. $\rho>0$ measures the subproblem solution accuracy.
  • (2)
    Related to the subiteration number $R$ and the precision $v$.
We end this section with a comparison between our results and the existing algorithms for solving FedProx, as reported in Table 2. Our result shows how the personalization degree trades off the communication cost and local computation cost, whereas most of the existing works either don’t explicitly analyze the local computational cost (Li et al., 2020a; T Dinh et al., 2020), or fail to demonstrate such a trade-off (Chen et al., 2023b). Hanzely & Richtárik (2020) also reveals similar tradeoffs, but their work is a randomized method that requires more coordination between server and clients, and more importantly, when $\lambda$ is small, the computation cost per round is unbounded.
6 Simulation
In this section, we provide empirical validation for our theoretical results, evaluating FedProx on both synthetic and real datasets with convex and non-convex loss functions. We first examine the effect of personalization on statistical accuracy under different levels of heterogeneity (Section 6.2), followed by an analysis of its impact on communication efficiency (Section 6.3). Detailed descriptions of the experimental setup are available in Section 6.1, and the complete anonymized codebase is accessible at https://anonymous.4open.science/r/fedprox-bilevel.
6.1 Experimental Details
Synthetic Dataset. To generate the synthetic data, we follow a similar procedure to prior works (Li et al., 2020a) but with some modifications to align with the setup in this paper. Specifically, for each client we generate samples $(\bm{X}_{k},\bm{y}_{k})$, where the labels $\bm{y}_{k}$ are produced by a logistic regression model $\bm{y}_{k}=\text{argmax}(\sigma({\bm{w}}_{k}^{\top}\bm{X}_{k}))$, with $\sigma$ being the sigmoid function. The feature vectors $\bm{X}_{k}$ are drawn from a multivariate normal distribution $\mathcal{N}(\bm{v}_{k},\bm{\Sigma})$, where $v_{k}\sim\mathcal{N}(0,1)$ and the covariance matrix $\bm{\Sigma}$ follows a diagonal structure $\bm{\Sigma}_{j,j}=j^{-1.2}$. The heterogeneity is introduced by sampling the model weights ${\bm{w}}_{k}$ for each client from a normal distribution $\mathcal{N}(0,R)$, where $R\in[0,3]$ controls the degree of variation (statistical heterogeneity) across clients’ data distributions.
Real Dataset. We use the MNIST dataset for the real dataset analysis. We first formulate the problem as a convex classification task using multinomial logistic regression, then test the generalization of our theory to a more general, non-convex scenario by conducting experiments using a CNN classifier. Following Li et al. (2020a), to impose statistical heterogeneity, we distribute the MNIST data across clients in the way that each client only has access to $s\in\{2,4,6,8,10\}$ classes. The fewer classes each client has access to, the higher the statistical heterogeneity. To be consistent with our theory, for both the synthetic and real datasets, we sample 2,000 samples and distribute them evenly among 10 clients, and all clients fully participate in each communication round. Additional results with 100 and 1,000 clients under partial participation are provided in Appendix C.
Selection of $\lambda$. For the logistic regression on the synthetic dataset, we selected three values for $\lambda$: small $\lambda=0.02$, medium $\lambda=0.1$, and large $\lambda=0.5$. For logistic regression on the MNIST dataset, we used small $\lambda=0.5$, medium $\lambda=1.5$, and large $\lambda=2.5$. For CNN on the MNIST dataset, we used small $\lambda=0.2$, medium $\lambda=0.5$, and large $\lambda=2.5$. Such a choice is to make sure that each line in the figure is clearly separable. Additionally, we implemented an adaptive choice of $\lambda=\frac{\rho}{\sqrt{n}R}\mathbb{I}\{R\leq n^{-1/2}\}+\frac{\rho^{2}}{nR^{2}}\mathbb{I}\{R>n^{-1/2}\}$ following the formula outlined in Theorem 1, with $\rho$ set to 2. The value of $\rho$ was determined through grid search over the range $[1,10]$ to ensure optimal performance.
Selection of Step Size. For the synthetic dataset, aligned with Theorem 2, we set the global step size $\gamma=\frac{\lambda+L}{\lambda L}$ and the local step size $\eta=\frac{1}{L+\lambda}$. The smoothness constant $L$ is computed as the upper bound of the $L$-smoothness constant for logistic regression, specifically $4^{-1}\Lambda_{\text{max}}$, where $\Lambda_{\text{max}}$ denotes the largest eigenvalue of the Gram matrix $\bm{X}^{\top}\bm{X}$. The local step size for the synthetic dataset was further tuned through grid search in $\{0.001,0.01,0.1\}$. For the MNIST dataset, the global learning rate was set to $\frac{1}{\lambda}$, while the local learning rate was set to 0.01, chosen via grid search. For both GlobalTrain and LocalTrain training baselines, the learning rate was also chosen via grid search within the range $\{0.001,0.01,0.1\}$. LocalTrain is implemented using a local (S)GD, and GlobalTrain is implemented using FedAvg (McMahan et al., 2017).
Selection of Hyperparameters. For the MNIST dataset, we implemented logistic regression using the SGD solver with 5 epochs, a batch size of 32, and 20 total runs. For the CNN model on the MNIST dataset, the same hyperparameter setup was used, except the batch size was reduced to 16 to ensure stable convergence.
Optimizer. For the synthetic dataset, following the analysis in the paper, the loss function is optimized using gradient descent (GD). For the MNIST dataset, all models are implemented in PyTorch, and optimization is performed using the stochastic gradient descent (SGD) solver.
Evaluation Criterion. In terms of evaluation, for the synthetic dataset, we track the ground-truth global model ${{\bm{w}}_{\star}^{(g)}}$ and local model ${{\bm{w}}_{\star}^{(i)}}$. We compute the global statistical error as $\|\widetilde{\bm{w}}^{(g)}-{{\bm{w}}_{\star}^{(g)}}\|^{2}$ and the local statistical error as $\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}$. To capture the total error at the $t$-th iteration, we report both the global total error $\|\bm{w}_{t}^{(g)}-{{\bm{w}}_{\star}^{(g)}}\|^{2}$ and the local total error $\|{\bm{w}}_{t,0}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}$, which combine both optimization and statistical errors. For the MNIST dataset, we report test accuracy on the MNIST testing set as a measure of the statistical error, and training loss for each client’s training set as total error at the $t$-th iteration.
Figure 1: Statistical and total error of different methods. In the top row, subfigures (a), (b), and (c) represent statistical error, showing the difference between the minimizer and the underlying true model for Synthetic-Logit (logistic regression on synthetic data), MNIST-Logit (logistic regression on MNIST), and MNIST-CNN (CNN on MNIST), respectively. In the bottom row, subfigures (d), (e), and (f) represent total error, showing the error between the model at each round and the true model for the same datasets. (1) FedProx adapts between GlobalTrain and LocalTrain, achieving low error across varying levels of heterogeneity with a properly chosen $\lambda$; (2) Larger $\lambda$ (less personalization) requires more communication to converge, and vice versa.
6.2 Statistical Aspect: Balancing Local and Global Learning
We now explore how FedProx balances between GlobalTrain and LocalTrain by achieving different levels of personalization. For comparison, in addition to implementing FedProx with different choices of $\lambda$, we also implemented LocalTrain and GlobalTrain, with details specified in Appendx LABEL:sec:additional_exp. We run all methods to a stable convergence and measure the statistical error. The top row of Figure 1 shows the performance of these methods under different impact of statistical heterogeneity.
From Figure 1(a), in a controlled convex setting, we observe that FedProx consistently interpolates between GlobalTrain and LocalTrain. As $\lambda$ increases, FedProx transitions from behaving similarly to LocalTrain towards GlobalTrain. When statistical heterogeneity is low, both GlobalTrain and FedProx with larger $\lambda$ values achieve better statistical performance. However, as heterogeneity increases, LocalTrain and FedProx with smaller $\lambda$ values show better results. The figure also demonstrates that with an adaptive choice of $\lambda$ outlined in Theorem 1, FedProx consistently achieves a low error across varying levels of heterogeneity. See more results and discussion in Appendix C. Note that as statistical heterogeneity increases, LocalTrain experiences a slight rise in error due to the increased difficulty of the statistical problem. However, it remains the least affected by heterogeneity since it relies solely on local data. In Figure 1(b) and 1(c), for the MNIST dataset, which includes both convex logistic regression and non-convex CNN settings, we observe similar patterns. FedProx again strikes a balance between GlobalTrain and LocalTrain, adapting its performance as $\lambda$ changes. While GlobalTrain performs best in this scenario due to the relatively weak statistical heterogeneity, FedProx, with increasing $\lambda$, moves closer to GlobalTrain performance. In Figure 4 in Appendix C, we also evaluate the local model’s statistical error and observe that as statistical heterogeneity increases, FedProx with smaller $\lambda$ values can surpass GlobalTrain.
Figure 2: Effect of local updates on FedProx performance. Smaller $\lambda$ reduces communication with more local updates, while larger $\lambda$ benefits marginally from increasing local updates and requires more frequent communication.
6.3 Optimization Aspect: Communication-Computation Trade-off
In this section, we explore the trade-off between communication and computation in FedProx, controlled by varying the personalization degree.
Communication Cost. In the bottom row of Figure 1, we show the change in the total error of the algorithm (i.e., the difference of the model in $t$th round between the true underlying model) over communication rounds. The first observation is that as $\lambda$ increases, the convergence rate w.r.t. communication rounds slows across all settings. FedProx with a large $\lambda$ (closer to GlobalTrain) needs the most communication rounds to converge, while smaller $\lambda$ values (closer to LocalTrain) result in faster convergence. This result aligns with our theory (c.f. Corollary 1), showing that Fedprox with less personalization relies more heavily on communication. On the statistical side, Figure 1(d) shows that when statistical heterogeneity is low, larger $\lambda$ leads to the smallest total error. As $R$ increases, smaller $\lambda$ (more personalization) results in better performance.
Computation Cost. In Figure 2, we investigate how the number of local updates per communication round impacts the performance of FedProx at different personalization degree. Additional results in the non-convex setting can be found in Appendix C. Notice that for models with smaller $\lambda$ (higher personalization), increasing the number of local updates significantly improves convergence, making the algorithm less dependent on frequent communication. This aligns with our analysis: more personalized models rely less on communication and benefit more from local computation.
7 Conclusion
In this paper, we analyzed the effect of personalization in FedProx. With an adaptive choice of personalization degree, FedProx achieves a nearly minimax-optimal statistical accuracy under different statistical heterogeneity degrees and consistently outperforms pure local training. As a minor contribution, we provably show that personalization reduces communication costs without increasing the computation overhead. The results are validated through extensive experiments.

### v2

@@
 Assumption 2 (Strong Convexity).
-The loss function $\ell(\cdot,z)$ is $\mu$-strongly convex for each local sample average, i.e.
-$\sum_{j\in[n_{i}]}\ell(x,z_{ij})/n_{i}\geq\sum_{j\in[n_{i}]}\ell(y,z_{ij})/n_{i}+\langle\sum_{j\in[n_{i}]}\nabla\ell(y,z_{ij})/n_{i},y-x\rangle+\frac{\mu}{2}\|x-y\|^{2},\quad\forall x,y\in\mathbb{R}^{d},i\in[m].$
-Assumption 3 (Bound Gradient Variance at Optimum).
+Each loss function $L_{i}$ is $\mu$-strongly convex, i.e., $L_{i}({\bm{w}},S_{i})\geq L_{i}({\bm{w}}^{\prime},S_{i})+\langle\nabla L_{i}({\bm{w}}^{\prime},S_{i}),{\bm{w}}-{\bm{w}}^{\prime}\rangle+\frac{\mu}{2}\|{\bm{w}}-{\bm{w}}^{\prime}\|^{2},\quad\forall{\bm{w}},{\bm{w}}^{\prime}\in\mathbb{R}^{d},i\in[m].$
+Assumption 3 (Bound Gradient Variance at the Optimum).
 There exists a nonnegative constant $\rho$ such that $\forall i\in[m]$, we have $\mathbb{E}_{z\sim\mathcal{D}_{i}}\|\nabla\ell({{\bm{w}}_{\star}^{(i)}},z)\|^{2}\leq\rho^{2}$.
@@
 We first provide a lemma to cast light on the effect of $\lambda$ on the statistical error.
-See Appendix A.2 for the proof.
+See A.2 for the proof.
 Lemma 1.
@@
 $\|\nabla\ell(x,z)-\nabla\ell(y,z)\|\leq L\|x-y\|,\quad\forall z.$ (9)
-Based on Assumption 4 and the new notations in (7) and (8), the following lemma describes the optimizing properties for $F_{i}$, $h_{i}$ and ${\bm{w}}_{\star}^{(i)}$.
-The proof can be found in Appendix B.2.
+Based on Assumption 4 and the new notations in (7) and (8), the following lemma describes the optimizing properties for $F_{i}$, $h_{i}$ and ${\bm{w}}_{\star}^{(i)}$. The proof can be found in B.1.
 Lemma 2.
@@
 Proof.
-See Appendix B.3 and B.5. ∎
+See Appendix B.2 and B.4. ∎
 Theorem 2 shows under the step size constraint $\gamma<1/L_{g}$, ${\bm{w}}_{t}^{(g)}$ converges to $\tilde{{\bm{w}}}^{(g)}$ at the same rate as the iteration (10) with exact gradient computation.
@@
 We first examine the effect of personalization on statistical accuracy under different levels of heterogeneity (Section 6.2), followed by an analysis of its impact on communication efficiency (Section 6.3).
-Detailed descriptions of the experimental setup are available in Section 6.1, and the complete anonymized codebase is accessible at https://anonymous.4open.science/r/fedprox-bilevel.
+Detailed descriptions of the experimental setup are available in Section 6.1.
 6.1 Experimental Details
@@
 We now explore how FedProx balances between GlobalTrain and LocalTrain by achieving different levels of personalization.
-For comparison, in addition to implementing FedProx with different choices of $\lambda$, we also implemented LocalTrain and GlobalTrain, with details specified in Appendx LABEL:sec:additional_exp.
+For comparison, in addition to implementing FedProx with different choices of $\lambda$, we also implemented LocalTrain and GlobalTrain, with details specified in Appendx C.
 We run all methods to a stable convergence and measure the statistical error.
@@
 The results are validated through extensive experiments.
+Appendix A Proof of statistical convergence
+To facilitate our analysis, we denote
+$\tilde{\Delta}^{(g)}_{\text{stat}}={{\bm{w}}_{\star}^{(g)}}-\widetilde{\bm{w}}^{(g)},$ (15)
+$\tilde{\Delta}^{(i)}_{\text{stat}}={{\bm{w}}_{\star}^{(i)}}-\widetilde{\bm{w}}^{(i)}\quad\forall i\in[m],$ (16)
+where $\tilde{\Delta}^{(g)}_{\text{stat}}$ is the difference between the true global model and the optimal solution of the global model in (1) and $\tilde{\Delta}^{(i)}_{\text{stat}}$ is the difference between the true local model and the optimal solution of the local model in (1) for any $i\in[m]$.
+The statistical error bound is established as $\mathbb{E}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}$ and $\mathbb{E}\|\tilde{\Delta}^{(i)}_{\text{stat}}\|^{2}$, where the expectation is taken over all the data.
+A.1 Discussion on Assumption 1
+When assuming the statistical heterogeneity of each client is different ($R$ is related to the client index $i$), it would provide a more delicate description of statistical heterogeneity, and it also requires different personalization degree per client.
+Therefore, it is better to consider the objective (17), where the $i$-th client will be shrunk to the global with strength $\lambda_{i}$.
+Therefore, solving (17) will consider different personalization degree for different clients.
+Solving the objective (17) would be open for future work.
+$\min_{{{\bm{w}}^{(g)}},\{{{\bm{w}}^{(i)}}\}_{i\in[m]}}\sum_{i=1}^{m}p_{i}\left({L}_{i}({{\bm{w}}^{(i)}},S_{i})+\frac{\lambda_{i}}{2}\|{{\bm{w}}^{(g)}}-{{\bm{w}}^{(i)}}\|^{2}\right),$ (17)
+Under Assumptions 2 and 4 as we assume, Assumption 1 about statistical heterogeneity would be equivalent to many other similar assumptions for statistical heterogeneity, like $B$-dissimilarity (Li et al., 2020a), parameter difference (Chen et al., 2023b) and gradient diversity (T Dinh et al., 2020; Deng et al., 2020).
+A.2 Proof of global statistical Error Bound
+In this part, we will establish the statistical error bound for the global model with fewer assumptions: firstly, in the objective (1), we assume the vector $\mathcal{P}=(p_{1},\cdots,p_{m})$ just under the constraint: $p_{i}\geq 0,\forall i\in[m]$ and $\sum_{i\in[m]}p_{i}=1$; secondly, we assume the data size of the client $i$ is $n_{i}$ for any $i\in[m]$.
+For convenience, we neglect the notation of taking an expectation over all data during the inducing process.
+First, we will prove Lemma 1, which is a crucial part to prompt the analysis of the global statistical error bound.
+Proof.
+We introduce some basic results from the optimality condition of the objective (1), which will be utilized later.
+Recall that ${\tilde{\bm{w}}^{(g)}}$ and $\{{\tilde{\bm{w}}^{(i)}}\}_{i\in[m]}$ are the minimizers of the objective (1). According to the first-order condition, we have
+$\nabla_{{{\bm{w}}^{(i)}}}L_{i}\left({{\bm{w}}^{(i)}},S_{i}\right)\big{|}_{{{\bm{w}}^{(i)}}={\tilde{\bm{w}}^{(i)}}}=\lambda\left({\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\right),$ (18)
+${\tilde{\bm{w}}^{(g)}}=\sum_{i}p_{i}{\tilde{\bm{w}}^{(i)}}.$ (19)
+To facilitate our analysis, we denote
+$\bm{\delta}_{\star}^{(i)}:={{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}},\quad\tilde{\bm{\delta}}^{(i)}:={\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}.$ (20)
+that is, $\bm{\delta}_{\star}^{(i)}$ measures the difference between the true global model ${{\bm{w}}_{\star}^{(g)}}$ and the true local model ${{\bm{w}}_{\star}^{(i)}}$, and $\tilde{\bm{\delta}}^{(i)}$ is the estimator of such a difference.
+To prove the global statistical error bound, we start with the optimality condition of ${\tilde{\bm{w}}^{(i)}}$ and ${\tilde{\bm{w}}^{(g)}}$, which yields
+00 0 $\geq\sum_{i\in[m]}p_{i}\left[L_{i}({\tilde{\bm{w}}^{(i)}},S_{i})+\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\|^{2}\right]-\sum_{i\in[m]}p_{i}\left[L_{i}({{\bm{w}}_{\star}^{(i)}},S_{i})+\frac{\lambda}{2}\|{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}\right].$ (21)
+Reorganizing the terms, we obtain
+00 0 $\geq\sum_{i\in[m]}\sum_{j\in[n_{i}]}\frac{p_{i}}{n_{i}}\left(\ell({\tilde{\bm{w}}^{(i)}},z_{ij})-\ell({{\bm{w}}_{\star}^{(i)}},z_{ij})\right)+\sum_{i\in[m]}p_{i}\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\|^{2}-\sum_{i\in[m]}p_{i}\frac{\lambda}{2}\|{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}.$ (22)
+For the first term on the RHS in (22), apply the $\mu$-strongly convexity of the loss function $\ell$ (Assumption 2)
+$\begin{split}\sum_{i\in[m]}\sum_{j\in[n_{i}]}\frac{p_{i}}{n_{i}}\left(\ell({\tilde{\bm{w}}^{(i)}},z_{ij})-\ell({{\bm{w}}_{\star}^{(i)}},z_{ij})\right)&\geq-\sum_{i\in[m]}\sum_{j\in[n_{i}]}\frac{p_{i}}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle\\ &+\mu\sum_{i\in[m]}\sum_{j\in[n_{i}]}\frac{p_{i}}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2},\\ \end{split}$ (23)
+where we use the definition of $\tilde{\Delta}^{(i)}_{\text{stat}}$ in (16).
+Plugging (23) back into (22), it yields
+$\begin{split}0&\geq-\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle+\mu\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\\ &+\sum_{i\in[m]}p_{i}\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\|^{2}-\sum_{i\in[m]}p_{i}\frac{\lambda}{2}\|{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}.\end{split}$ (24)
+Recall the definition in (20) and the Assumption 1 about statistical heterogeneity
+00 0 $\geq-\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle+\mu\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\sum_{i}p_{i}\left\|\tilde{\bm{\delta}}^{(i)}\right\|^{2}-\frac{\lambda}{2}R^{2}.$ (25)
+Utilizing the result (25) in Lemma 1, we are ready to proof Theorem 1
+Proof.
+Based on the result in (25), we will discuss the upper bound of $\tilde{\Delta}^{(g)}_{\text{stat}}$ with two cases of statistical heterogeneoity:
+Case 1: for the first case, if $R$ is large, i.e, $R>\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$.
+First, applying Cauchy inequality on the first term of (25), it yields
+$\begin{split}\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle&=\sum_{i}p_{i}\left\langle\sum_{j}\frac{1}{n_{i}}\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle\\ &\leq\sum_{i}p_{i}\left\|\sum_{j}\frac{1}{n_{i}}\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij})\right\|_{2}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}.\end{split}$ (26)
+Consider the Assumption 3 and $z_{ij}$ are i.i.d data
+$\left(\mathbb{E}\left\|\sum_{j}\frac{1}{n_{i}}\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij})\right\|_{2}\right)^{2}\leq\mathbb{E}\left\|\sum_{j}\frac{1}{n_{i}}\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij})\right\|^{2}\leq\frac{\rho^{2}}{n_{i}}.$ (27)
+Combining (26) and (27), it yields
+$\begin{split}\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\mathbb{E}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle\leq\sum_{i}\rho\frac{p_{i}}{\sqrt{n_{i}}}\|\tilde{\Delta}^{(i)}_{\text{stat}}\|_{2}.\\ \end{split}$ (28)
+Plugging (28) back to (25) yields
+$0\geq-\rho\sum_{i}\frac{p_{i}}{\sqrt{n_{i}}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\mu\sum_{i}p_{i}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\sum_{i}p_{i}\left\|\tilde{\bm{\delta}}^{(i)}\right\|^{2}-\frac{\lambda}{2}R^{2}.$ (29)
+After dropping the third term and applying Cauchy inequality, we obtain
+$0\geq-\rho\left(\Big{(}\sum_{i\in[m]}\frac{p_{i}}{n_{i}}\Big{)}\Big{(}\sum_{i}p_{i}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\Big{)}\right)^{\frac{1}{2}}+\mu\sum_{i}p_{i}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}-\frac{\lambda}{2}R^{2}.$ (30)
+The equation (30) would be regarded as a quadratic function of $\sum_{i}p_{i}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}$
+Set the $\lambda$ as
+$\lambda=\frac{\rho^{2}}{R^{2}}\sum_{i\in[m]}\frac{p_{i}}{n_{i}}.$ (31)
+Solving the inequality in (30) w.r.t $\sum_{i}p_{i}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}$, it yields
+$\begin{split}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2}&\leq\left(\sum_{i}p_{i}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\right)^{\frac{1}{2}}\leq\frac{1}{2\mu}\left(\rho\left(\sum_{i\in[m]}\frac{p_{i}}{n_{i}}\right)^{\frac{1}{2}}+\left(\rho^{2}\sum_{i\in[m]}\frac{p_{i}}{n_{i}}(1+2\mu)\right)^{\frac{1}{2}}\right)\\ &=\frac{(1+\sqrt{1+2\mu})\rho}{2\mu}\left(\sum_{i\in[m]}\frac{p_{i}}{n_{i}}\right)^{\frac{1}{2}}.\\ \end{split}$ (32)
+Case 2: for the second case, if $R$ is relatively small, which means $R\leq\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$, based on the result (25), we know
+$\begin{split}0&\geq\sum_{i\in[m]}\sum_{j\in[n_{i}]}\frac{p_{i}}{n_{i}}\left[\ell\left({\tilde{\bm{w}}^{(i)}},x_{ij}\right)-\ell({{\bm{w}}_{\star}^{(i)}},x_{ij})\right]+\frac{\lambda}{2}\sum_{i\in[m]}p_{i}\left(\left\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\right\|^{2}-\left\|{{\bm{w}}_{\star}^{(g)}}-\bm{w}_{\star}^{(i)}\right\|^{2}\right)\\ &\geq-\underbrace{\sum_{i\in[m]}\sum_{j\in[n_{i}]}\frac{p_{i}}{n_{i}}\left\langle\nabla\ell({\tilde{\bm{w}}^{(i)}},x_{ij}),\tilde{\Delta}^{(g)}_{\text{stat}}\right\rangle}_{T_{1}}+\underbrace{\sum_{i\in[m]}\sum_{j\in[n_{i}]}\frac{p_{i}}{n_{i}}\left\langle\nabla\ell({\tilde{\bm{w}}^{(i)}},x_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}-\tilde{\Delta}^{(g)}_{\text{stat}}\right\rangle}_{T_{2}}\\ &\quad+\sum_{i\in[m]}\sum_{j\in[n_{i}]}\frac{p_{i}}{n_{i}}\mu\|\tilde{\Delta}^{(i)}_{\text{stat}}\|^{2}+\frac{\lambda}{2}\sum_{i\in[m]}p_{i}\|\tilde{\bm{\delta}}^{(i)}\|^{2}-\frac{\lambda}{2}\sum_{i\in[m]}p_{i}\|\bm{\delta}_{\star}^{(i)}\|^{2}.\end{split}$ (33)
+It is necessary to remark that in the case of $R\leq\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$, we should take the GlobalTrain (2) into account so that our analysis different from the case 1.
+For $T_{1}$
+$T_{1}\leq\big{\|}\sum_{ij}\frac{p_{i}}{n_{i}}\nabla\ell({\tilde{\bm{w}}^{(i)}},z_{ij})\big{\|}_{2}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2}\leq\rho\sqrt{\sum_{i}\frac{p_{i}^{2}}{n_{i}}}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2},$ (34)
+where the first inequality comes from Cauchy inequality and to obtain the second inequality, we need to note the fact that from Assumption 3, we have
+$\left(\mathbb{E}\left\|\sum_{ij}\frac{p_{i}}{n_{i}}\nabla\ell(\widetilde{\bm{w}}^{(i)},z_{ij})\right\|_{2}\right)^{2}\leq\mathbb{E}\left\|\sum_{ij}\frac{p_{i}}{n_{i}}\nabla\ell(\widetilde{\bm{w}}^{(i)},z_{ij})\right\|_{2}^{2}\leq\sum_{ij}\frac{p_{i}^{2}}{n_{i}^{2}}\rho^{2}.$ (35)
+Analogy to the induction of $T_{1}$, we can induce $T_{2}$ in a similar way
+$\begin{split}T_{2}&\leq\sum_{i}p_{i}\left\langle\frac{1}{n_{i}}\sum_{j}\nabla\ell({\tilde{\bm{w}}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}-\tilde{\Delta}^{(g)}_{\text{stat}}\right\rangle\\ &\leq\sum_{i}p_{i}\left\|\frac{1}{n_{i}}\sum_{j}\nabla\ell({\tilde{\bm{w}}^{(i)}},z_{ij})\right\|_{2}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}-\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2}\\ &\leq\sum_{i}p_{i}\frac{\rho}{\sqrt{n_{i}}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}-\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2}.\\ \end{split}$ (36)
+Note the fact
+$\left\|\tilde{\Delta}^{(i)}_{\text{stat}}-\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2}=\left\|{{\bm{w}}_{\star}^{(i)}}-{\tilde{\bm{w}}^{(i)}}-{{\bm{w}}_{\star}^{(g)}}+{\tilde{\bm{w}}^{(g)}}\right\|_{2}=\left\|\bm{\delta}_{\star}^{(i)}-\tilde{\bm{\delta}}^{(i)}\right\|_{2}.$ (37)
+Combining (37) back to (36), it yields
+$T_{2}\leq\sum_{i}p_{i}\frac{\rho}{\sqrt{n_{i}}}\left\|\bm{\delta}_{\star}^{(i)}-\tilde{\bm{\delta}}^{(i)}\right\|_{2}\leq\sum_{i}p_{i}\frac{\rho}{\sqrt{n_{i}}}\left(\left\|\bm{\delta}_{\star}^{(i)}\right\|_{2}+\left\|\tilde{\bm{\delta}}^{(i)}\right\|_{2}\right).$ (38)
+Plugging (34) and (38) into (33), it yields
+$\begin{split}0&\geq-\rho\sqrt{\sum_{i}\frac{p_{i}^{2}}{n_{i}}}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|_{2}-\sum_{i}p_{i}\frac{\rho}{\sqrt{n_{i}}}(\|\bm{\delta}_{\star}^{(i)}\|_{2}+\|\tilde{\bm{\delta}}^{(i)}\|_{2})+\mu\sum_{i}p_{i}\|\tilde{\Delta}^{(i)}_{\text{stat}}\|^{2}\\ &\quad\quad+\frac{\lambda}{2}\sum_{i}p_{i}\|\tilde{\bm{\delta}}^{(i)}\|^{2}-\frac{\lambda}{2}\sum_{i\in[m]}p_{i}\|\bm{\delta}_{\star}^{(i)}\|^{2}\\ &\geq\mu\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}-\rho\sqrt{\sum_{i}\frac{p_{i}^{2}}{n_{i}}}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|_{2}-\rho\sum_{i}\frac{p_{i}}{\sqrt{n_{i}}}\|\tilde{\bm{\delta}}^{(i)}\|_{2}+\frac{\lambda}{2}\sum_{i}p_{i}\|\tilde{\bm{\delta}}^{(i)}\|^{2}-R\rho\sum_{i}\frac{p_{i}}{\sqrt{n_{i}}}-\frac{\lambda}{2}R^{2},\end{split}$ (39)
+where the last line uses Assumption 1. Then focusing on the right-hand side of the above inequality and regarding it as a function of $\tilde{\bm{\delta}}^{(i)}$, it can be written as
+$\begin{split}&\quad-\rho\sum_{i}\frac{p_{i}}{\sqrt{n_{i}}}\|\tilde{\bm{\delta}}^{(i)}\|_{2}+\frac{\lambda}{2}\sum_{i}p_{i}\|\tilde{\bm{\delta}}^{(i)}\|^{2}\\ &=\sum_{i}\left(\frac{\lambda}{2}p_{i}\|\tilde{\bm{\delta}}^{(i)}\|^{2}-\rho\frac{p_{i}}{\sqrt{n_{i}}}\|\tilde{\bm{\delta}}^{(i)}\|_{2}\right)\\ &\geq-\frac{\rho^{2}}{2\lambda}\sum_{i}{\frac{p_{i}}{n_{i}}},\end{split}$ (40)
+where we choose the optimal $\|\tilde{\bm{\delta}}^{(i)}\|_{2}^{\star}$ as $\frac{\rho}{\lambda\sqrt{n_{i}}}$.
+Plugging (LABEL:global:_case2_opt5) into (39), it yields
+$0\geq\mu\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}-\rho\sqrt{\sum_{i}\frac{p_{i}^{2}}{n_{i}}}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|_{2}-\frac{\rho^{2}}{2\lambda}\sum_{i}\frac{p_{i}}{n_{i}}-R\rho\sum_{i}\frac{p_{i}}{\sqrt{n_{i}}}-\frac{\lambda}{2}R^{2}.$ (41)
+Then we maximize the RHS of (41) w.r.t $\lambda$, the optimal solution of $\lambda$ is $\frac{\rho}{R}\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$. We add the result into (41) and it yields
+00 0 $\geq\mu\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}-\rho\sqrt{\sum_{i}\frac{p_{i}^{2}}{n_{i}}}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|_{2}-\rho R\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}-\rho R\sum_{i}\frac{p_{i}}{\sqrt{n_{i}}}.$ (42)
+We can regard (42) as a function of $\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}$ and induce
+$\begin{split}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|_{2}&\leq\frac{1}{2\mu}\left\{\rho\sqrt{\sum_{i}\frac{p_{i}^{2}}{n_{i}}}+\left(\rho^{2}\sum_{i}\frac{p_{i}^{2}}{n_{i}}+4\mu\rho R\left(\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}+\sum_{i}\frac{p_{i}}{\sqrt{n_{i}}}\right)\right)^{\frac{1}{2}}\right\}.\end{split}$ (43)
+Thus
+$\begin{split}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}\leq\frac{\rho^{2}}{\mu^{2}}\sum_{i}\frac{p^{2}_{i}}{n_{i}}+\frac{2\rho R}{\mu}\left(\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}+\sum_{i}\frac{p_{i}}{\sqrt{n_{i}}}\right).\end{split}$ (44)
+Combine two-case arguments, we have:
+$\Rightarrow\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|^{2}\leq\begin{cases}\left(\frac{(1+\sqrt{1+2\mu})\rho}{2\mu}\right)^{2}\sum_{i\in[m]}\frac{p_{i}}{n_{i}}&R>\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}\\ \frac{\rho^{2}}{\mu^{2}}\sum_{i}\frac{p^{2}_{i}}{n_{i}}+\frac{2\rho R}{\mu}\left(\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}+\sum_{i}\frac{p_{i}}{\sqrt{n_{i}}}\right)&R\leq\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}.\end{cases}$ (45)
+Thus it yields a one-line rate
+$\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|^{2}\leq C_{1}\sum_{i}\frac{p_{i}^{2}}{n_{i}}+C_{2}\left(\sum_{i\in[m]}\frac{p_{i}}{n_{i}}\wedge R\left(\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}+\sum_{i}\frac{p_{i}}{\sqrt{n_{i}}}\right)\right),$ (46)
+where
+$\begin{split}C_{1}&=\frac{\rho^{2}}{\mu^{2}},\\ C_{2}&=\left(\frac{(1+\sqrt{1+2\mu})\rho}{2\mu}\right)^{2}\vee\frac{2\rho}{\mu}.\end{split}$ (47)
+Moreover, if we assume data balanced, i.e $n_{i}=n,\forall i\in[m]$ and $p_{i}=\frac{1}{m}$, we obtain
+$\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|^{2}\leq C_{1}\frac{1}{N}+C_{2}\left(\frac{1}{n}\wedge\frac{R}{\sqrt{n}}\right).$ (48)
+A.3 Proof of local statistical Error Bound
+Proof.
+Firstly, we will provide some basic results for the local optimality.
+To prove the local statistical error bound, we start with the optimality condition of ${\tilde{\bm{w}}^{(i)}}$ and ${\tilde{\bm{w}}^{(g)}}$ for a single client, which yields
+00 0 $\geq L_{i}\left({\tilde{\bm{w}}^{(i)}},S_{i}\right)+\frac{\lambda}{2}\left\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\right\|^{2}-L_{i}\left({{\bm{w}}_{\star}^{(i)}},S_{i}\right)-\frac{\lambda}{2}\left\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\right\|^{2}.$ (49)
+Reorganized the terms, we obtain
+00 0 $\geq\sum_{j\in[n_{i}]}\frac{1}{n_{i}}\left(\ell\left({\tilde{\bm{w}}^{(i)}},z_{ij}\right)-\ell\left({{\bm{w}}_{\star}^{(i)}},z_{ij}\right)\right)+\frac{\lambda}{2}\left\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\right\|^{2}-\frac{\lambda}{2}\left\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\right\|^{2}.$ (50)
+For the first term on the RHS, apply the $\mu$-strongly convex of the loss function $\ell$ (Assumption 2)
+$\begin{split}\sum_{j\in[n_{i}]}\frac{1}{n_{i}}\left(\ell({\tilde{\bm{w}}^{(i)}},z_{ij})-\ell({{\bm{w}}_{\star}^{(i)}},z_{ij})\right)&\geq-\sum_{j\in[n_{i}]}\frac{1}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle+\mu\sum_{j\in[n_{i}]}\frac{1}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2},\end{split}$ (51)
+where we use the definition of $\tilde{\Delta}^{(i)}_{\text{stat}}$ in (16).
+Plugging (51) back into (50), it yields
+$\begin{split}0&\geq-\sum_{j}\frac{1}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle+\mu\sum_{j}\frac{1}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\|^{2}-\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}.\end{split}$ (52)
+Apply Cauchy inequality on the first term of (52)
+$\begin{split}\sum_{j}\frac{1}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle&=\left\langle\sum_{j}\frac{1}{n_{i}}\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle\\ &\leq\left\|\sum_{j}\frac{1}{n_{i}}\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij})\right\|_{2}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}.\end{split}$ (53)
+Consider the Assumption 3 and $z_{ij}$ are i.i.d data
+$\left(\mathbb{E}\left\|\sum_{j}\frac{1}{n_{i}}\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij})\right\|_{2}\right)^{2}\leq\mathbb{E}\left\|\sum_{j}\frac{1}{n_{i}}\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij})\right\|^{2}\leq\frac{\rho^{2}}{n_{i}}.$ (54)
+Combining (53) and (54), it yields
+$\begin{split}\sum_{j}\frac{1}{n_{i}}\mathbb{E}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle\leq\rho\frac{1}{\sqrt{n_{i}}}\|\tilde{\Delta}^{(i)}_{\text{stat}}\|_{2}.\\ \end{split}$ (55)
+Plugging (55) back to (52) yields
+$0\geq-\rho\frac{1}{\sqrt{n_{i}}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\mu\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\|^{2}-\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}.$ (56)
+To utilize the statistical global convergence rate shown in Appendix A.2, we need to argue $R$ in the same way.
+Recalling the upper bound of $\tilde{\Delta}^{(g)}_{\text{stat}}$ if data is balanced (48)
+$\Rightarrow\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}\leq\begin{cases}\frac{C_{2}}{n}&R>\frac{1}{\sqrt{n}}\\ \frac{C_{1}}{N}+\frac{C_{2}R}{{n}^{\frac{1}{2}}}&R\leq\frac{1}{\sqrt{n}}.\end{cases}$ (57)
+For simplicity, we denote the upper bound of $\|\tilde{\Delta}^{(g)}_{\text{stat}}\|_{2}$ as $U_{0}$
+Secondly, we need to discuss the upper bound of $\tilde{\Delta}^{(i)}_{\text{stat}}$ with two cases below.
+Case 1: For the first case, we consider the statistical heterogeneity is relative, i.e., $R>\frac{1}{\sqrt{n_{i}}}$.
+Note that
+$\begin{split}-\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}&=-\left\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}+{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\right\|^{2}\\ &=-\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}-\|{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}-2\|{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|_{2}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2}\\ &\geq-\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}-R^{2}-2R\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2}.\end{split}$ (58)
+Plug (58) into (56)
+$\begin{split}0\geq&-\frac{\rho}{\sqrt{n_{i}}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\mu\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\left\|\tilde{\bm{\delta}}^{(i)}\right\|^{2}-\frac{\lambda}{2}\left\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}+{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\right\|^{2}\\ =&-\frac{\rho}{\sqrt{n_{i}}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\mu\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\left\|\tilde{\bm{\delta}}^{(i)}\right\|^{2}-\frac{\lambda}{2}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|^{2}-\frac{\lambda}{2}R^{2}-\lambda R\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2}.\\ \end{split}$ (59)
+Put (57) into (59) and reorganize
+$\begin{split}0\geq&-\frac{\rho}{\sqrt{n_{i}}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\mu\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\left\|\tilde{\bm{\delta}}^{(i)}\right\|^{2}-\frac{\lambda}{2}U_{0}^{2}-\frac{\lambda}{2}R^{2}-\lambda R\;U_{0}\\ \geq&\mu\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}-\frac{\rho}{\sqrt{n_{i}}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}-\frac{\lambda}{2}\left(U_{0}+R\right)^{2}.\end{split}$ (60)
+In addition, recall that in this case we set $\lambda^{\star}=\frac{\rho^{2}}{R^{2}}\sum_{i\in[m]}\frac{p_{i}}{n_{i}}$ in the statistical convergence analysis of the global model.
+If we assume data balanced, i.e.
+$p_{i}=\frac{1}{m}$ and $n_{i}=n,\forall i\in[m]$, it yields $\lambda^{\star}=\frac{1}{nR^{2}}$
+$\begin{split}\|\tilde{\Delta}^{(i)}_{\text{stat}}\|_{2}&\leq\frac{1}{2\mu}\left[\frac{\rho}{\sqrt{n}}+\left(\frac{\rho^{2}}{n}+\frac{2\mu\rho^{2}}{nR^{2}}\left(U_{0}+R\right)^{2}\right)^{\frac{1}{2}}\right].\\ &=\frac{1}{\sqrt{n}}\frac{1}{2\mu}\left(\rho+\left(\rho^{2}+\frac{2\mu\rho^{2}}{R^{2}}(U_{0}+R)^{2}\right)^{\frac{1}{2}}\right)\\ &\leq\frac{1}{\sqrt{n}}\frac{\rho+(\rho^{2}+4\mu\rho^{2}(C_{2}+1))^{\frac{1}{2}}}{2\mu}\\ &\leq\frac{1}{\sqrt{n}}\frac{\rho(1+(1+4\mu(C_{2}+1))^{\frac{1}{2}})}{2\mu}.\end{split}$ (61)
+Case 2: On the other hand, we will focus on the case that $R$ is relatively small, which means $R\leq\sqrt{\frac{m}{N}}$. From (56), note the definition of $\bm{\delta}_{\star}^{(i)}$ in (20)
+$\begin{split}0&\geq-\frac{\rho}{\sqrt{n_{i}}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\mu\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\left\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}+\delta_{\star}^{(i)}+{{\bm{w}}_{\star}^{(i)}}-{\tilde{\bm{w}}^{(i)}}\right\|^{2}\\ &-\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}.\end{split}$ (62)
+Note
+$2\left\|-\tilde{\Delta}^{(g)}_{\text{stat}}+\tilde{\Delta}^{(i)}_{\text{stat}}+\delta_{\star}^{(i)}\right\|^{2}+2\left\|\tilde{\Delta}^{(g)}_{\text{stat}}-\delta_{\star}^{(i)}\right\|^{2}\geq\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}.$ (63)
+Thus
+$\left\|-\tilde{\Delta}^{(g)}_{\text{stat}}+\tilde{\Delta}^{(i)}_{\text{stat}}+\delta_{\star}^{(i)}\right\|^{2}\geq\frac{1}{2}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}-\left\|\tilde{\Delta}^{(g)}_{\text{stat}}-\delta_{\star}^{(i)}\right\|^{2}.$ (64)
+Furthermore, note
+$\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}+\|{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}=\|\tilde{\Delta}^{(i)}_{\text{stat}}\|^{2}+\|\bm{\delta}_{\star}^{(i)}\|^{2}\geq\frac{1}{2}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}.$ (65)
+Plugging (64) and (66) into (62), it yields
+$0\geq-\frac{\rho}{\sqrt{n_{i}}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\mu\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{4}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}-\frac{\lambda}{2}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}-\delta_{\star}^{(i)}\right\|^{2}-\lambda\|\tilde{\Delta}^{(i)}_{\text{stat}}\|^{2}-\lambda\|\bm{\delta}_{\star}^{(i)}\|^{2},$ (66)
+which implies
+$0\geq-\frac{\rho}{\sqrt{n_{i}}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\mu\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{4}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}-2\lambda\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|^{2}-2\lambda\|\bm{\delta}_{\star}^{(i)}\|^{2}.$ (67)
+Apply Assumption 1 and (57) to (67)
+$0\geq\left(\mu+\frac{\lambda}{4}\right)\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}-\frac{\rho}{\sqrt{n_{i}}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}-2\lambda U_{0}^{2}-2\lambda R^{2}.$ (68)
+Solving (68) as a function of $\|\tilde{\Delta}^{(i)}_{\text{stat}}\|_{2}$, it yields
+$\begin{split}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}&\leq\frac{1}{2\mu+\lambda/2}\left\{\frac{\rho}{\sqrt{n_{i}}}+\left[\frac{\rho^{2}}{n_{i}}+4\left(\mu+\frac{\lambda}{4}\right)\lambda\left(U_{0}^{2}+R^{2}\right)\right]^{\frac{1}{2}}\right\}.\end{split}$ (69)
+If we assume data size is same among different clients, i.e.
+$n=n_{i},\forall i\in[m]$ and choose $\lambda=\frac{\rho}{R}\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}=\frac{\rho}{R}\frac{1}{\sqrt{n}}$, (69) can be reformulated as
+$\begin{split}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}&\leq\frac{2}{4\mu+\frac{\rho}{\sqrt{n}R}}\left\{\frac{\rho}{\sqrt{n}}+\left[\frac{\rho^{2}}{n}+\left(4\mu+\frac{\rho}{\sqrt{n}R}\right)\frac{\rho}{\sqrt{n}R}\left(U_{0}^{2}+R^{2}\right)\right]^{\frac{1}{2}}\right\}.\end{split}$ (70)
+Therefore, we have
+$\|\tilde{\Delta}^{(i)}_{\text{stat}}\|^{2}\leq\frac{16\rho^{2}}{\left((4\mu+\frac{\rho}{R\sqrt{n}})\sqrt{n}\right)^{2}}+\frac{8}{4\mu+\frac{\rho}{R\sqrt{n}}}\frac{\rho}{R\sqrt{n}}\left(\frac{C_{1}}{N}+\frac{C_{2}R}{\sqrt{n}}+R^{2}\right).$ (71)
+If $\mu<\frac{\rho}{R\sqrt{n}}$
+$\frac{\rho}{(4\mu+\frac{\rho}{R\sqrt{n}})\sqrt{n}}\leq\frac{\rho}{\frac{\rho}{R\sqrt{n}}\sqrt{n}}\leq R.$ (72)
+If $\mu\geq\frac{\rho}{R\sqrt{n}}$
+$\frac{\rho}{(4\mu+\frac{\rho}{R\sqrt{n}})\sqrt{n}}\leq\frac{\rho}{\mu\sqrt{n}}.$ (73)
+Combining the discussion in (72) and (73), it yields
+$\frac{16\rho^{2}}{\left((4\mu+\frac{\rho}{R\sqrt{n}})\sqrt{n}\right)^{2}}\leq 16\min{\{R^{2},\frac{\rho^{2}}{\mu^{2}n}\}}.$ (74)
+Then for the second term on the RHS of (71)
+$\begin{split}\frac{8}{4\mu+\frac{\rho}{R\sqrt{n}}}\frac{\rho}{R\sqrt{n}}\left(\frac{C_{1}}{N}+\frac{C_{2}R}{\sqrt{n}}+R^{2}\right)\leq\frac{8C_{1}}{N}+\frac{8C_{2}R}{\sqrt{n}}+\frac{8\rho R}{4\mu\sqrt{n}}.\end{split}$ (75)
+Plugging (74) and (75) into (71), we obtain
+$\begin{split}\|\tilde{\Delta}^{(i)}_{\text{stat}}\|^{2}&\leq 16\min{\{R^{2},\frac{\rho^{2}}{\mu^{2}n}\}}+\frac{8C_{1}}{N}+\frac{8C_{2}R}{\sqrt{n}}+\frac{8\rho R}{4\mu\sqrt{n}}\\ &\leq 16\min{\{R^{2},\frac{\rho^{2}}{\mu^{2}n}\}}+\frac{8C_{1}}{N}+(8C_{2}+\frac{2\rho}{\mu})\frac{R}{\sqrt{n}}\end{split}$ (76)
+Combining two-case arguments in (61) and (76), we have:
+$\Rightarrow\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\leq\begin{cases}\frac{1}{n}\left(\frac{\rho(1+(1+4\mu(C_{2}+1))^{\frac{1}{2}})}{2\mu}\right)^{2}&R>\sqrt{\frac{m}{N}}\\ 16(R^{2}\wedge\frac{\rho^{2}}{\mu^{2}n})+\frac{8C_{1}}{N}+(8C_{2}+\frac{2\rho}{\mu})\frac{R}{\sqrt{n}}&R\leq\sqrt{\frac{m}{N}}\end{cases}.$ (77)
+And it yields a one-line rate
+$\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\leq\frac{C_{3}}{N}+16(R^{2}\wedge\frac{\rho^{2}}{\mu^{2}n})+C_{4}(\frac{R}{\sqrt{n}}\wedge\frac{1}{n}),$ (78)
+where
+$C_{3}$ $=8C_{1},$ (79)
+$C_{4}$ $=\left(\frac{\rho(1+(1+4\mu(C_{2}+1))^{\frac{1}{2}})}{2\mu}\right)^{2}\wedge(8C_{2}+\frac{2\rho}{\mu}),$ (80)
+and $C_{1},C_{2}$ are specified in (47). ∎
+Appendix B Optimization Convergence Rate
+B.1 Proof of lemmas
+B.1.1 Proof of Lemma 2
+To facilitate our analysis, we state a claim first.
+Claim: If $f$ is $\mu$-strongly-convex, then the proximal operator
+$\text{prox}_{f/\lambda}(x)=\arg\min_{v}\left(f(v)+\frac{\lambda}{2}\|x-v\|^{2}\right),$ (82)
+is $L_{w}$- Lipschitz continuous with $L_{w}=\frac{\lambda}{\lambda+\mu}$.
+Proof.
+Based on the definition of $\text{prox}_{f/\lambda}(x)$, it will satisfy the first order condition
+$\nabla f(\text{prox}_{\frac{1}{\lambda}f}(x))-\lambda(x-\text{prox}_{\frac{1}{\lambda}f}(x))=0.$ (83)
+Then we have
+$\nabla f(x_{1})-\lambda(v_{1}-x_{1})=0\quad\text{where }x_{1}=\text{prox}_{\frac{1}{\lambda}f}(v),$ (84)
+$\nabla f(x_{2})-\lambda(v_{2}-x_{2})=0\quad\text{where }x_{2}=\text{prox}_{\frac{1}{\lambda}f}(v_{2}).$ (85)
+Based on the $\mu$-strongly convexity of $f$, we have
+$\langle\nabla f(x_{1})-\nabla f(x_{2}),x_{1}-x_{2}\rangle$ $\geq\mu\|x_{1}-x_{2}\|^{2}$ (86)
+$\Rightarrow\|\text{prox}_{\frac{1}{\lambda}f}(x_{1})-\text{prox}_{\frac{1}{\lambda}f}(x_{2})\|$ $\leq\frac{1}{1+\frac{\mu}{\lambda}}\|x_{1}-x_{2}\|.$ (87)
+Now, we are ready to prove Lemma 2.
+Proof.
+First, from mishchenko2023convergence, we can directly obtain the analytic properties of $F_{i}$.
+Next, the analytic properties of $h_{i}$ are direct results from Assumption 2 and 4.
+Finally, to analyze the property of the mapping $w_{\star}^{(i)}$, we just need to use the result from the above claim as we know the definition of ${\bm{w}}_{\star}^{(i)}(\cdot)$ in (8).
+B.1.2 Inner Loop Convergence
+Lemma 4 ((Ji et al., 2022), Inner Loop Error Contraction).
+Suppose Assumption 2 and 4 hold, then if $\eta\leq 1/L_{\ell}$, for all $t\geq 0$ and $\epsilon>0$, the inner loop error bound of all clients can be formulated as:
+$e_{t+1}^{(i)}:=\|{\bm{w}}_{t+1,K}^{(i)}-{\bm{w}}_{t+1,\star}^{(i)}\|^{2}\leq(1+\epsilon)(1-\eta\mu_{\ell})^{K}e_{t}^{(i)}+\left(1+\frac{1}{\epsilon}\right)(1-\eta\mu_{\ell})^{K}\|\bm{w}^{(i)}_{t,\star}-\bm{w}^{(i)}_{t-1,\star}\|^{2}.$ (88)
+Proof.
+The proof can be found in Lemma 2 (Ji et al., 2022).
+B.1.3 Outer Loop convergence
+Lemma 5 (Outer Loop Error Contraction).
+Suppose Assumption 2 and 4 hold, then if $\gamma\leq L_{g}$, we have for all $t>0$ and $\epsilon_{g}>0$:
+$\begin{split}\|{\bm{w}}_{t+1}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}&\leq(1-\gamma\mu_{g}+\gamma\epsilon_{g})\|{\bm{w}}_{t}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}-\gamma^{2}\|\nabla F({\bm{w}}_{t}^{(g)})\|^{2}\\ &+(\gamma^{3}L_{g})\|\widehat{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}+\left(\frac{\gamma}{\epsilon_{g}}+\gamma^{2}\right)\|\nabla F({\bm{w}}_{t}^{(g)})-\widehat{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}.\end{split}$ (89)
+Proof.
+Note
+^$\widehat{\nabla}F_{i}(\bm{w}_{t}^{(g)})=\lambda(\bm{w}_{t}^{(g)}-{\bm{w}}_{t,K}^{(i)}).$ (90)
+We define
+^$\widehat{\nabla}F(\bm{w}_{t}^{(g)}):=\frac{1}{m}\sum_{i\in[m]}\widehat{\nabla}F_{i}(\bm{w}_{t}^{(g)})=\frac{1}{m}\sum_{i\in[m]}\lambda(\bm{w}_{t}^{(g)}-{\bm{w}}_{t,K}^{(i)}).$ (91)
+Thus the global update rule can be written as ${\bm{w}}_{t+1}^{(g)}={\bm{w}}_{t}^{(g)}-\gamma\widehat{\nabla}F({\bm{w}}_{t}^{(g)})$. For the outer loop,
+$\begin{split}&\|{\bm{w}}_{t+1}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}\\ =&\|{\bm{w}}_{t}^{(g)}-\widetilde{\bm{w}}^{(g)}-\gamma\widehat{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}\\ =&\|{\bm{w}}_{t}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}-2\gamma\langle{\bm{w}}_{t}^{(g)}-\widetilde{\bm{w}}^{(g)},\widehat{\nabla}F({\bm{w}}_{t}^{(g)})\rangle+\gamma^{2}\|\widehat{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}\\ \leq&(1+\epsilon_{g}\gamma)\|{\bm{w}}_{t}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}-2\gamma\langle{\bm{w}}_{t}^{(g)}-\widetilde{\bm{w}}^{(g)},{\nabla}F({\bm{w}}_{t}^{(g)})\rangle+\gamma^{2}\|\widehat{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}\\ &+\left(\frac{\gamma}{\epsilon_{g}}\right)\|\widehat{\nabla}F({\bm{w}}_{t}^{(g)})-\nabla F({\bm{w}}_{t}^{(g)})\|^{2},\end{split}$ (92)
+where we apply the Young’s inequality with $\epsilon_{g}>0$ in the last line.
+Next, we bound the inner product term in the last line using the $\mu_{g}$-strong convexity of $F$:
+$\begin{split}-\gamma\langle{\bm{w}}_{t}^{(g)}-\widetilde{\bm{w}}^{(g)},{\nabla}F({\bm{w}}_{t}^{(g)})\rangle&\leq-\gamma\left(F({\bm{w}}_{t}^{(g)})-\widehat{F}+\frac{\mu_{g}}{2}\|{\bm{w}}_{t}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}\right)\\ &=-\gamma(F({\bm{w}}_{t+1}^{(g)})-\widehat{F})-\frac{\gamma\mu_{g}}{2}\|{\bm{w}}_{t}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}+\gamma\big{(}F({\bm{w}}_{t+1}^{(g)})-F({\bm{w}}_{t}^{(g)})\big{)}.\end{split}$ (93)
+where $\hat{F}:=F(\widetilde{\bm{w}}^{(g)})$ and the last term $F({\bm{w}}_{t+1}^{(g)})-F({\bm{w}}_{t}^{(g)})$ can be upper bounded using the $L_{g}$-smoothness of $F$:
+$\begin{split}F({\bm{w}}_{t+1}^{(g)})\leq&F({\bm{w}}_{t}^{(g)})-\gamma\langle\nabla F({\bm{w}}_{t}^{(g)}),\widehat{\nabla}F({\bm{w}}_{t}^{(g)})\rangle+\frac{L_{g}\gamma^{2}}{2}\|\widehat{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}\\ \overset{(a)}{=}&F({\bm{w}}_{t}^{(g)})+\frac{\gamma}{2}\left(\|\nabla F({\bm{w}}_{t}^{(g)})-\widehat{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}-\|\nabla F({\bm{w}}_{t}^{(g)})\|^{2}-\|\widehat{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}\right)\\ &\quad+\frac{L_{g}\gamma^{2}}{2}\|\widehat{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}\\ =&F({\bm{w}}_{t}^{(g)})-\frac{\gamma}{2}\|\nabla F({\bm{w}}_{t}^{(g)})\|^{2}-\left(\frac{\gamma}{2}-\frac{L_{g}\gamma^{2}}{2}\right)\|\widehat{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}+\frac{\gamma}{2}\|\nabla F({\bm{w}}_{t}^{(g)})-\widehat{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2},\end{split}$ (94)
+where the step (a) uses the fact $-\langle a,b\rangle=\frac{1}{2}\|a-b\|^{2}-\frac{1}{2}\|a\|^{2}-\frac{1}{2}\|b\|^{2}$.
+Substituting (93) and (94) into (92) leads to
+$\begin{split}&\|{\bm{w}}_{t+1}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}\\ \leq&(1+\epsilon_{g}\gamma)\|{\bm{w}}_{t}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}+\gamma^{2}\|\widetilde{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}+\left(\frac{\gamma}{\epsilon_{g}}\right)\|\widetilde{\nabla}F({\bm{w}}_{t}^{(g)})-\nabla F({\bm{w}}_{t}^{(g)})\|^{2}\\ &-2\gamma(F({\bm{w}}_{t+1}^{(g)})-\widehat{F})-\gamma\mu_{g}\|{\bm{w}}_{t}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}\\ &-\gamma^{2}\|\nabla F({\bm{w}}_{t}^{(g)})\|^{2}-\gamma\left(\gamma-L_{g}\gamma^{2}\right)\|\widetilde{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}+\gamma^{2}\|\nabla F({\bm{w}}_{t}^{(g)})-\widetilde{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}\\ \leq&(1-\gamma\mu_{g}+\gamma\epsilon_{g})\|{\bm{w}}_{t}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}-\gamma^{2}\|\nabla F({\bm{w}}_{t}^{(g)})\|^{2}+(\gamma^{3}L_{g})\|\widetilde{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}\\ &+\left(\frac{\gamma}{\epsilon_{g}}+\gamma^{2}\right)\|\nabla F({\bm{w}}_{t}^{(g)})-\widetilde{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}.\end{split}$ (95)
+B.2 Proof of Theorem 2
+Recall the definition of $e_{t}^{(i)}$ in Lemma 4, we define
+$e_{t}=\frac{1}{m}\sum_{i\in[m]}e_{t}^{(i)}.$ (96)
+Recall the definition $\nabla F({\bm{w}}_{t}^{(g)}),\widehat{\nabla}F({\bm{w}}_{t}^{(g)})$ in Lemma and eq (91), the gradient approximation error is
+$\begin{split}\|\nabla F({\bm{w}}_{t}^{(g)})-\widehat{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}=\lambda^{2}\Big{\|}\frac{1}{m}\sum_{i=1}^{m}\big{(}{\bm{w}}_{t,K}^{(i)}-{\bm{w}}_{t,\star}^{(i)}\big{)}\Big{\|}^{2}\leq\lambda^{2}e_{t}.\\ \end{split}$ (97)
+Therefore, the recursion of optimization error given by Lemma 5 can be further bounded as:
+$\begin{split}&\|{\bm{w}}_{t+1}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}\\ \overset{(a)}{\leq}&(1-\gamma\mu_{g}+\gamma\epsilon_{g})\|{\bm{w}}_{t}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}-\left(\gamma^{2}-\gamma^{3}L_{g}(1+\zeta)\right)\|\nabla F({\bm{w}}_{t}^{(g)})\|^{2}\\ &+\left(\frac{\gamma}{\epsilon_{g}}+\gamma^{2}+L_{g}\gamma^{3}(1+\zeta^{-1})\right)\|\nabla F({\bm{w}}_{t}^{(g)})-\widehat{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}\\ \overset{(b)}{\leq}&(1-\gamma\mu_{g}+\gamma\epsilon_{g})\|{\bm{w}}_{t}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}-\gamma^{2}\Big{(}1-\gamma L_{g}(1+\zeta)\Big{)}\|\nabla F({\bm{w}}_{t}^{(g)})\|^{2}\\ &+\left(\frac{\gamma}{\epsilon_{g}}+\gamma^{2}(2+\zeta^{-1})\right)\lambda^{2}e_{t},\end{split}$ (98)
+where step (a) comes from Young’s inequality with $\zeta>0$ to be chosen and step (b) follows from the step size condition $\gamma\leq 1/L_{g}$.
+Invoking Lemma 2 and Lemma 4, the inner loop optimization error is updated as
+$\begin{split}e_{t+1}&\leq(1+\epsilon)(1-\eta\mu_{\ell})^{K}e_{t}+\left(1+\frac{1}{\epsilon}\right)(1-\eta\mu_{\ell})^{K}\frac{1}{m}\sum_{i=1}^{m}\|{\bm{w}}_{t+1,\star}^{(i)}-{\bm{w}}_{t,\star}^{(i)}\|^{2}\\ &\stackrel{{\scriptstyle(a)}}{{\leq}}(1+\epsilon)(1-\eta\mu_{\ell})^{K}e_{t}+\left(1+\frac{1}{\epsilon}\right)(1-\eta\mu_{\ell})^{K}L_{w}^{2}\|{\bm{w}}_{t}^{(g)}-{\bm{w}}_{t+1}^{g}\|^{2}\\ &\stackrel{{\scriptstyle(b)}}{{\leq}}(1+\epsilon)(1-\eta\mu_{\ell})^{K}e_{t}+\left(1+\frac{1}{\epsilon}\right)(1-\eta\mu_{\ell})^{K}L_{w}^{2}\gamma^{2}\|\widehat{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}\\ &\stackrel{{\scriptstyle(c)}}{{\leq}}(1+\epsilon)(1-\eta\mu_{\ell})^{K}e_{t}+\left(1+\frac{1}{\epsilon}\right)(1-\eta\mu_{\ell})^{K}L_{w}^{2}\gamma^{2}\big{(}2\|{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}+2\lambda^{2}e_{t}\big{)}\\ &=\underbrace{\Big{(}1+\epsilon+2(1+\epsilon^{-1})L_{w}^{2}\gamma^{2}\lambda^{2}\Big{)}(1-\eta\mu_{\ell})^{K}}_{:=q}e_{t}+2\left(1+\frac{1}{\epsilon}\right)(1-\eta\mu_{\ell})^{K}L_{w}^{2}\gamma^{2}\|{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2},\end{split}$ (99)
+where step (a) is due to Lemma 2, step (b) uses the global update: ${\bm{w}}_{t+1}^{(g)}-{\bm{w}}_{t}^{(g)}=-\gamma\widehat{\nabla}F({\bm{w}}_{t}^{(g)})$ and in step (c), we just apply Cauchy inequality instead of Young’s inequality.
+Under $\eta\leq 1/L_{\ell}$, we have $11-\eta\mu_{\ell}<$ 1. Therefore, by choosing a large $K$ we can always drive $q$ arbitrarily small. Let $K$ be such that
+$q\leq 1-\gamma\mu_{g}\leq 1-\gamma\mu_{g}+\gamma\epsilon_{g}.$ (100)
+Then, we can rewrite (99) as:
+$\begin{split}e_{t+1}\leq&\frac{1-\gamma\mu_{g}+\gamma\epsilon_{g}+q}{2}e_{t}-\frac{1-\gamma\mu_{g}+\gamma\epsilon_{g}-q}{2}e_{t}\\ &+2\left(1+\frac{1}{\epsilon}\right)(1-\eta\mu_{\ell})^{K}L_{w}^{2}\gamma^{2}\|{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}.\end{split}$ (101)
+Letting $c_{1}=\frac{2}{1-\gamma\mu_{g}+\gamma\epsilon_{g}-q}\cdot\left(\frac{\gamma}{\epsilon_{g}}+\gamma^{2}(2+\zeta^{-1})\right)\lambda^{2}$, we can combine (101) with (98) and obtain
+$\begin{split}&\|{\bm{w}}_{t+1}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}+c_{1}\cdot e_{t+1}\\ \leq&(1-\gamma\mu_{g}+\gamma\epsilon_{g})\left(\|{\bm{w}}_{t}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}+c_{1}\cdot e_{t}\right)-\gamma^{2}\Big{(}1-\gamma L_{g}(1+\zeta)\Big{)}\|\nabla F({\bm{w}}_{t}^{(g)})\|^{2}\\ &+\left(\frac{2}{1-\gamma\mu_{g}+\gamma\epsilon_{g}-q}\cdot\left(\frac{\gamma}{\epsilon_{g}}+\gamma^{2}(2+\zeta^{-1})\right)\lambda^{2}\right)\cdot 2\left(1+\frac{1}{\epsilon}\right)(1-\eta\mu_{\ell})^{K}L_{w}^{2}\gamma^{2}\|{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}.\end{split}$ (102)
+Therefore, if the condition (100) and
+$\begin{split}\frac{4\lambda^{2}\gamma}{1-\gamma\mu_{g}+\gamma\epsilon_{g}-q}\left(\frac{1}{\epsilon_{g}}+\gamma(2+\zeta^{-1})\right)\left(1+\frac{1}{\epsilon}\right)(1-\eta\mu_{\ell})^{K}L_{w}^{2}\leq 1-\gamma L_{g}(1+\zeta)\end{split}$ (103)
+hold then (102) implies both $\|{\bm{w}}_{t}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}$ and $e_{t}$ converge to zero at rate $1-\gamma\mu_{g}+\gamma\epsilon_{g}$.
+It remains to specify the free parameters $(\epsilon_{g},\epsilon,\zeta$) in (100), (103) and others listed in Lemma 5 and Lemma 4.
+$\bullet$ Convergence Rate For the connection term in (102), if we set
+$\epsilon_{g}=\frac{\mu_{g}}{2}\cdot(1-\gamma\mu_{g})>0,$ (104)
+which, substituting into $1-\gamma\mu_{g}+\gamma\epsilon_{g}$, gives the convergence rate
+$r:=1-\frac{\gamma\mu_{g}}{2}-\frac{(\gamma\mu_{g})^{2}}{2}.$ (105)
+Under condition $\gamma\leq 1/L_{g}$, one can verify that $r\in(0,1)$.
+$\bullet$ Step size conditions. Under the requirement (100), we have
+$1-\gamma\mu_{g}+\gamma\epsilon_{g}-q\geq\gamma\epsilon_{g},$ (106)
+which gives the following sufficient condition for (103):
+$\begin{split}\frac{2}{\epsilon_{g}}\cdot\left(\frac{1}{\epsilon_{g}}+\gamma(2+\zeta^{-1})\right)\lambda^{2}\cdot 2\left(1+\frac{1}{\epsilon}\right)(1-\eta\mu_{\ell})^{K}L_{w}^{2}\leq 1-\gamma L_{g}(1+\zeta).\end{split}$ (107)
+Further restricting $\gamma$ such that
+$\gamma(2+\zeta^{-1})\frac{\mu_{g}}{2}\leq 1\quad\implies\gamma(2+\zeta^{-1})\leq\frac{1}{\epsilon_{g}},$ (108)
+it suffices to require the following for (107) to hold:
+$\begin{split}\left(\frac{2}{\epsilon_{g}}\right)^{2}\lambda^{2}\cdot 2\left(1+\frac{1}{\epsilon}\right)(1-\eta\mu_{\ell})^{K}L_{w}^{2}\leq 1-\gamma L_{g}(1+\zeta).\end{split}$ (109)
+Letting $\epsilon=1$ and $\zeta=1-\gamma L_{g}>0$, we collect all the conditions (104), (109) and (100) on $\gamma$ respectively as follows:
+$1-\gamma L_{g}>0,\quad 1-\gamma\mu_{g}>0,$ (110)
+$\left(1-\gamma\mu_{g}\right)^{2}(1-\gamma L_{g})^{2}\geq 4\lambda^{2}\cdot 4(1-\eta\mu_{\ell})^{K}L_{w}^{2}\cdot\left(\frac{2}{\mu_{g}}\right)^{2},$ (111)
+$1-\gamma\mu_{g}\geq\Big{(}2+4L_{w}^{2}\gamma^{2}\lambda^{2}\Big{)}(1-\eta\mu_{\ell})^{K}.$ (112)
+Using the fact that $\mu_{g}\leq L_{g}$, the above conditions simply to
+$\gamma<1/L_{g},\quad\Big{(}2+64L_{w}^{2}(1/\mu_{g})^{2}\lambda^{2}\Big{)}(1-\eta\mu_{\ell})^{K}\leq(1-\gamma L_{g})^{4}.$ (113)
+B.3 Proof of Corollary 1
+With $\eta=1/L_{\ell}=(\lambda+L)^{-1}$, $\gamma=1/(2L_{g})=\frac{\lambda+L}{2\lambda L}$, and using the fact that $L_{w}=\frac{\lambda}{\lambda+\mu}$ and $\mu_{g}=\frac{\lambda\mu}{\lambda+\mu}$ (cf.
+Lemma 2), condition (113) for $K$ becomes
+$\begin{split}&\Big{(}2+64L_{w}^{2}(1/\mu_{g})^{2}\lambda^{2}\Big{)}(1-\eta\mu_{\ell})^{K}\\ =&\left(2+64\frac{\lambda^{2}}{(\lambda+\mu)^{2}}\frac{(\lambda+\mu)^{2}}{(\lambda\mu)^{2}}\lambda^{2}\right)\left(1-\frac{\lambda+\mu}{\lambda+L}\right)^{K}\\ =&\left(2+64\frac{\lambda^{2}}{\mu^{2}}\right)\left(\frac{L-\mu}{\lambda+L}\right)^{K}\leq\frac{1}{16}.\end{split}$ (114)
+If $L=\mu$, then the above condition trivially holds. Otherwise, a sufficient condition for it is
+$66\kappa^{2}\left(\frac{L-\mu}{\lambda+L}\right)^{K-2}\leq\frac{1}{16},$ (115)
+where we have used the fact that
+$\frac{\lambda^{2}}{\mu^{2}}\left(\frac{L-\mu}{\lambda+L}\right)^{K}=\left(\frac{\lambda}{\lambda+L}\right)^{2}\cdot\left(\frac{L-\mu}{\mu}\right)^{2}\left(\frac{L-\mu}{\lambda+L}\right)^{K-2}\leq\kappa^{2}\left(\frac{L-\mu}{\lambda+L}\right)^{K-2}$ (116)
+and $\kappa:=L/\mu\geq 1$. Finally, using the inequality $\log(1/x)\geq 1-x$ for $0<x\leq 1$ we obtain
+$K\geq 2+\frac{\lambda+L}{\lambda+\mu}\cdot\log(1056\kappa^{2}).$ (117)
+B.4 Proof of The Local Model Convergence Rate in Theorem 2
+Proof.
+Note that
+$\|{\bm{w}}_{t,K}^{(i)}-\widetilde{\bm{w}}^{(i)}\|^{2}\leq 2\|{\bm{w}}_{t,K}^{(i)}-\bm{w}^{(i)}_{t,\star}\|^{2}+2\|\bm{w}^{(i)}_{t,\star}-\widetilde{\bm{w}}^{(i)}\|^{2}.$ (118)
+For the first part on the right-hand side, recalling the definition of $e_{t}^{(i)}$ in Lemma 4
+$2\|{\bm{w}}_{t,K}^{(i)}-\bm{w}^{(i)}_{t,\star}\|^{2}=2\|e_{t}^{(i)}\|^{2}.$ (119)
+For the second term on the right-hand side, first, note the property in (8)
+$\bm{w}^{(i)}_{t,\star}=\operatorname{prox}_{L_{i}/\lambda}({\bm{w}}_{t}^{(g)}).$ (120)
+Based on the first order condition of (1)
+$\widetilde{\bm{w}}^{(i)}=\operatorname{prox}_{L_{i}/\lambda}(\widetilde{\bm{w}}^{(g)}).$ (121)
+Next, plugging (120) and (121) into (118), it yields
+$\begin{split}\|{\bm{w}}_{t,K}^{(i)}-\widetilde{\bm{w}}^{(i)}\|^{2}&\leq 2\|e_{t}^{(i)}\|^{2}+2\|\operatorname{prox}_{L_{i}/\lambda}({\bm{w}}^{(g)})-\operatorname{prox}_{L_{i}/\lambda}(\widetilde{{\bm{w}}}^{(g)})\|^{2}\\ &\leq 2\|e_{t}^{(i)}\|^{2}+2L_{w}^{2}\|\bm{w}^{(i)}_{t,\star}-\widetilde{\bm{w}}^{(i)}\|^{2},\end{split}$ (122)
+where the last we use Lemma 2.
+In the proof of Theorem 2, we have established that both $\|\bm{w}_{t}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}$ and $e_{t}$ converges to zero linearly at rate $21-\gamma\mu_{g}/2-(\gamma\mu_{g})^{2}/$ 2, combining with (122), the proof is finished.
+∎
+Appendix C Additional Experiment Details and Results
+C.1 Additional Experiment Details
+In this section, we provide more details on the implementation of the methods used in the empirical analysis.
+Selection of $\lambda$.
+For the logistic regression on the synthetic dataset, we selected three values for $\lambda$: small $\lambda=0.02$, medium $\lambda=0.1$, and large $\lambda=0.5$.
+For logistic regression on the MNIST dataset, we used small $\lambda=0.5$, medium $\lambda=1.5$, and large $\lambda=2.5$.
+For CNN on the MNIST dataset, we used small $\lambda=0.2$, medium $\lambda=0.5$, and large $\lambda=2.5$.
+Such a choice is to make sure that each line in the figure is clearly separable.
+Additionally, we implemented an adaptive choice of $\lambda=\frac{\rho}{\sqrt{n}R}\mathbb{I}\{R\leq n^{-1/2}\}+\frac{\rho^{2}}{nR^{2}}\mathbb{I}\{R>n^{-1/2}\}$ following the formula outlined in Theorem 1, with $\rho$ set to 2.
+The value of $\rho$ was determined through grid search over the range $[1,10]$ to ensure optimal performance.
+Selection of Step Size.
+For the synthetic dataset, aligned with Theorem 2, we set the global step size $\gamma=\frac{\lambda+L}{\lambda L}$ and the local step size $\eta=\frac{1}{L+\lambda}$.
+The smoothness constant $L$ is computed as the upper bound of the $L$-smoothness constant for logistic regression, specifically $4^{-1}\Lambda_{\text{max}}$, where $\Lambda_{\text{max}}$ denotes the largest eigenvalue of the Gram matrix $\bm{X}^{\top}\bm{X}$.
+The local step size for the synthetic dataset was further tuned through grid search in $\{0.001,0.01,0.1\}$.
+For the MNIST dataset, the global learning rate was set to $\frac{1}{\lambda}$, while the local learning rate was set to 0.01, chosen via grid search.
+For both GlobalTrain and LocalTrain training baselines, the learning rate was also chosen via grid search within the range $\{0.001,0.01,0.1\}$.
+LocalTrain is implemented using a local (S)GD, and GlobalTrain is implemented using FedAvg (McMahan et al., 2017).
+Selection of Hyperparameters.
+For the MNIST dataset, we implemented logistic regression using the SGD solver with 5 epochs, a batch size of 32, and 20 total runs.
+For the CNN model on the MNIST dataset, the same hyperparameter setup was used, except the batch size was reduced to 16 to ensure stable convergence.
+Optimizer.
+For the synthetic dataset, following the analysis in the paper, the loss function is optimized using gradient descent (GD).
+For the MNIST dataset, all models are implemented in PyTorch, and optimization is performed using the stochastic gradient descent (SGD) solver.
+Evaluation Criterion.
+In terms of evaluation, for the synthetic dataset, we track the ground-truth global model ${{\bm{w}}_{\star}^{(g)}}$ and local model ${{\bm{w}}_{\star}^{(i)}}$.
+We compute the global statistical error as $\|\widetilde{\bm{w}}^{(g)}-{{\bm{w}}_{\star}^{(g)}}\|^{2}$ and the local statistical error as $\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}$.
+To capture the total error at the $t$-th iteration, we report both the global total error $\|\bm{w}_{t}^{(g)}-{{\bm{w}}_{\star}^{(g)}}\|^{2}$ and the local total error $\|{\bm{w}}_{t,0}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}$, which combine both optimization and statistical errors.
+For the MNIST dataset, we report test accuracy on the MNIST testing set as a measure of the statistical error, and training loss for each client’s training set as total error at the $t$-th iteration.
+C.2 Additional Experiment Results
+Figure 3: Performance of the FedProx algorithm with an adaptive choice of $\lambda$ (from Theorem 1) under varying levels of statistical heterogeneity.
+In Figure 3, Statistical error of the FedProx algorithm with an adaptive choice of $\lambda$ (from Theorem 1) under varying levels of statistical heterogeneity.
+With the adaptive choice of $\lambda$ as defined in Theorem 1, FedProx consistently achieves the lowest statistical error across most scenarios, outperforming both GlobalTrain and LocalTrain baselines.
+This is particularly evident under moderate and high levels of statistical heterogeneity.
+However, as $R$ approaches very small values, the global baseline performs slightly better, which aligns with the slower rate of FedProx with small $R$ described in our theoretical results.
+Overall, these results validate our choice of adaptive $\lambda$ and confirm the statistical rate provided by our analysis.
+Figure 4: (a) Testing accuracy of different methods (FedProx, LocalTrain, GlobalTrain) versus statistical heterogeneity levels on the MNIST dataset (logistic regression).
+(b) Training loss versus communication rounds for the same methods.
+In Figure 4(a), we compare FedProx, LocalTrain, and GlobalTrain on the MNIST dataset by plotting testing accuracy against varying levels of statistical heterogeneity.
+As heterogeneity increases, each client has access to fewer classes, resulting in fewer classification tasks.
+This change simplifies the problem, which explains why all methods show improved evaluation accuracy as the number of classes decreases.
+However, this does not diminish the insights from the relative comparison between methods.
+GlobalTrain and FedProx with larger $\lambda$ perform better at lower heterogeneity levels, while LocalTrain and FedProx with smaller $\lambda$ perform better as heterogeneity increases.
+In Figure 4(b), we track testing loss as a function of communication rounds.
+Across all sub-figures, FedProx with smaller $\lambda$ (higher personalization) demonstrates faster convergence, especially in the early stages of training.
+The final training loss achieved by different methods also aligns with the results shown in Figure 4(a) and Figure 1.
+Figure 5: Performance of FedProx with 100 clients (top row) and 1,000 clients (bottom row) with 10% client sampling under different levels of statistical heterogeneity ($R$).
+In Figure 5, the top row shows results for 100 clients, and the bottom row shows results for 1,000 clients, and in both results we adopt a client sampling of 10% in each communication round.
+This tests the scalability and generalizability of the our results under typical federated learning setups.
+Consistent with the findings from Figure 1, FedProx with smaller $\lambda$ (greater personalization) consistently achieves faster convergence at the early stages of training, while FedProx with larger $\lambda$ (less personalization) takes more communication rounds to converge.
+As statistical heterogeneity increases (measured by $R$), FedProx with smaller $\lambda$ gradually outperforms models with larger $\lambda$, showcasing the adaptability of FedProx to different levels of heterogeneity.
+These results further validate the theoretical insights from Theorem 1 and Theorem 2, highlighting the influence of personalization on statistical and optimization performance in federated setups with larger numbers of clients.
+Figure 6: Effect of local updates on FedProx performance for local models in MNIST logistic regression and global models ni MNIST CNN, complementing the results shown in Figure 2.
+Figure 6 demonstrates the impact of varying the number of local updates on FedProx performance, complementing the results shown in Figure 2.
+For FedProx with smaller $\lambda$ (higher personalization), increasing the number of local updates notably reduces the need for communication, leading to faster convergence.
+In contrast, for FedProx with larger $\lambda$ (lower personalization), increasing local updates yields minimal benefit and only adds to computation costs without reducing communication demands.
+The results indicate that for larger $\lambda$, the optimal strategy is to prioritize frequent communication rather than increasing local computation.
+This highlights the trade-off between local computation and communication efficiency under different levels of personalization, validating our theoretical findings in Corollary 1.

### v3

@@
 This work fills in the gap by analyzing the effect of regularization on statistical accuracy, thereby providing a theoretical guideline for setting the regularization strength for achieving personalization.
-We prove that by adaptively choosing the regularization strength under different statistical heterogeneity, FedProx can consistently outperform pure local training and achieve a nearly minimax-optimal statistical rate.
+We prove that by adaptively choosing the regularization strength under different statistical heterogeneity, FedProx can consistently outperform pure local training and achieve a minimax-optimal statistical rate.
 In addition, to shed light on resource allocation, we design an algorithm, provably showing that stronger personalization reduces communication complexity without increasing the computation cost overhead.
 Finally, our theory is validated on both synthetic and real-world datasets and its generalizability is verified in a non-convex setting.
-Keywords: Personalized Federated Learning, Statistical Complexity, Communication Complexity, Minimax Optimality
 1 Introduction
@@
 Our contribution is summarized below:
-$\bullet$ Nearly Optimal Statistical Accuracy.
+$\bullet$ Minimax-optimal Statistical Accuracy.
 We analyze the statistical accuracy of the objective in (1), taking statistical heterogeneity into account.
-For each of the models, we establish a state-of-the-art convergence rate $\mathcal{O}(1/(mn))+\mathcal{O}(1/n\wedge R/n^{1/2})$, where $R$ characterizes the statistical heterogeneity and $n$ is the sample size on each client.
+For each of the models, we establish a state-of-the-art convergence rate $\mathcal{O}(1/(mn))+\mathcal{O}(1/n\wedge R^{2})$, where $R$ characterizes the statistical heterogeneity and $n$ is the sample size on each client.
 The first term $\mathcal{O}(1/(mn))$ represents the sample complexity of training a single global model under homogeneity, which is attained by (1) when $R$ is close to zero, highlighting the benefit of FL.
-For the second term, the first component $\mathcal{O}(1/n)$ is the sample complexity of purely local training and the second component $\mathcal{O}(R/n^{1/2})$ demonstrates the flexibility of FedProx under different statistical heterogeneity, implying the statistical accuracy of FedProx never gets worse than purely local training’s.
-The fine-grained analysis demonstrates that with an adaptive choice of personalization degree characterized by $\lambda$, FedProx achieves a nearly minimax-optimal statistical accuracy across different levels of statistical heterogeneity.
+For the second term, the first component $\mathcal{O}(1/n)$ is the sample complexity of purely local training, and the second component $\mathcal{O}(R^{2})$ demonstrates the flexibility of FedProx under different statistical heterogeneity, implying the statistical accuracy of FedProx never gets worse than purely local training’s.
+Such a rate matches the existing lower bound for the problem.
+Therefore, to the best of our knowledge, this is the first work that demonstrates that with a theoretically-informed adaptive choice of personalization degree characterized by $\lambda$, the objective in (1) achieves a minimax-optimal statistical accuracy across different levels of statistical heterogeneity.
+The theoretical analysis we use for objective (1) would extend to different objectives in regularization-based personalized federated learning (Smith et al., 2017; Li et al., 2020a; Hanzely & Richtárik, 2020; Hanzely et al., 2020; 2023).
 $\bullet$ Provably Enhanced Communication Efficiency.
@@
 With higher personalization degree, communication complexity would decrease and the computation complexity per round would increase, demonstrating that there is a trade-off between computation and communication costs.
-Our analysis concludes that model personalization can reduce communication overhead without incurring additional computation expense.
+Our analysis concludes that model personalization can reduce communication overhead without incurring additional computation expenses.
 $\bullet$ Empirical Validation.
@@
 Therefore, in the FedProx objective, the choice of $\lambda$ should be adjusted according to $R$.
-Next, we introduce two standard assumptions.
+More discussion about Assumption 1 can be found in Appendix A.1.
+Next, we introduce three standard assumptions.
+Assumption 2 (Smoothness).
+The loss function $\ell(\cdot,z)$ is L-smooth, i.e. for any $x,y\in\mathbb{R}^{d}$
+$\|\nabla\ell(x,z)-\nabla\ell(y,z)\|\leq L\|x-y\|,\quad\forall z.$ (3)
 Assumption 3 (Strong Convexity).
-Each loss function $L_{i}$ is $\mu$-strongly convex, i.e., $L_{i}({\bm{w}},S_{i})\geq L_{i}({\bm{w}}^{\prime},S_{i})+\langle\nabla L_{i}({\bm{w}}^{\prime},S_{i}),{\bm{w}}-{\bm{w}}^{\prime}\rangle+\frac{\mu}{2}\|{\bm{w}}-{\bm{w}}^{\prime}\|^{2},\quad\forall{\bm{w}},{\bm{w}}^{\prime}\in\mathbb{R}^{d},i\in[m].$
-Assumption 3 (Bound Gradient Variance at the Optimum).
+The loss function $\ell(\cdot,z)$ is almost surely $\mu$-strongly convex: $\ell(x,z)\geq\ell(y,z)+\langle\nabla\ell(y,z),y-x\rangle+\frac{\mu}{2}\|x-y\|^{2},\quad\forall x,y\in\mathbb{R}^{d}.$
+Assumption 4 (Bound Gradient Variance at Optimum).
 There exists a nonnegative constant $\rho$ such that $\forall i\in[m]$, we have $\mathbb{E}_{z\sim\mathcal{D}_{i}}\|\nabla\ell({{\bm{w}}_{\star}^{(i)}},z)\|^{2}\leq\rho^{2}$.
-Here, the strongly convex assumption is used to establish the estimation error and is widely adopted in the statistical analysis of regularization-based personalized federated learning (T Dinh et al., 2020; Deng et al., 2020; Hanzely & Richtárik, 2020; Hanzely et al., 2020) and statistical analysis in federated learning (Chen et al., 2023b; Cheng et al., 2023).
-Assumption 3 is used to quantify the observation noise, and is standard in the literature (Li et al., 2020b; Chen et al., 2023b).
+Here, the strongly convex and smooth assumption is used to establish the estimation error and is widely adopted in the theoretical analysis of regularization-based personalized federated learning (T Dinh et al., 2020; Deng et al., 2020; Hanzely & Richtárik, 2020; Hanzely et al., 2020) and theoretical analysis in federated learning (Chen et al., 2023b; Cheng et al., 2023).
+Assumption 4 is used to quantify the observation noise, and is also standard in the literature (Duan & Wang, 2023; Chen et al., 2023b).
 Next, we establish a statistical rate of the optimal solution to the Fedprox problem (1) with a fine-grained analysis.
 To facilitate the analysis, we first define some notations.
-Define $\{\widetilde{\bm{w}}^{(i)}\}_{i=1}^{m}$ as the optimal solution for local models.
+Define $\{\widetilde{\bm{w}}^{(i)}\}_{i=1}^{m}$ as the optimal solution for local models $\bm{w}^{(i)}$ in objective (1).
 A check on the optimality condition implies that the optimal solution for the global model is $\widetilde{\bm{w}}^{(g)}=\sum^{m}_{i=1}p_{i}\widetilde{\bm{w}}^{(i)}$.
@@
 Note that some existing literature adopts alternative metrics, such as average excess risk and individual excess risk (Chen et al., 2023b) to measure the statistical error.
-They are equivalent to our metrics under the assumptions used in this paper.
+The convergence rate of these two definitions is closely related up to the smoothness constant and strong convexity constant.
+Define $\tilde{\bm{w}}_{\text{GT}}$ as the optimal solution for the GlobalTrain problem in (2).
 We first provide a lemma to cast light on the effect of $\lambda$ on the statistical error.
-See A.2 for the proof.
+See Appendix A.2 for the proof.
 Lemma 1.
-Suppose Assumptions 1 and 2 hold. We have
-$\underbrace{\sum_{i\in[m],j\in[n_{i}]}\frac{p_{i}}{n_{i}}\left(\mu\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}+\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\rangle\right)}_{\text{model estimation error}}\leq\frac{\lambda}{2}\underbrace{(R^{2}-\sum_{i\in[m]}p_{i}\|\widetilde{\bm{w}}^{(i)}-\widetilde{\bm{w}}^{(g)}\|^{2})}_{\text{heterogeneity estimation error}}.$ (3)
-Lemma 1 establishes a relationship between the local model estimation error and the statistical heterogeneity.
-On the right-hand side of (3), the first term $R$, which is larger than $\|{{\bm{w}}_{\star}^{(i)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}$, measures the statistical heterogeneity; the second term, $\sum_{i=1}^{m}p_{i}\|\widetilde{\bm{w}}^{(i)}-\widetilde{\bm{w}}^{(g)}\|^{2}$, can be understood as the estimator of such heterogeneity.
-Lemma 1 implies that to achieve a tight bound on the estimation error, we should choose a $\lambda$ that “minimizes” the right-hand side of (3).
-When $R$ is large, one straightforward choice is $\lambda=\mathcal{O}(1/(nR^{2}))$, which will be shown to yield an error bound of order $\mathcal{O}(1/n)$, matching the performance of purely local training and is independent of $R$.
-When $R$ is small, $\lambda$ should be carefully chosen to obtain the full sample complexity $\mathcal{O}(1/(mn))$.
+Suppose Assumptions 1, 2 and 3 hold. We have
+$\|\widetilde{\bm{w}}^{(g)}-\tilde{\bm{w}}_{\text{GT}}\|^{2}\leq\frac{4\left(1+L/\mu\right)}{\mu(\mu+\lambda)}\sum_{i\in[m]}p_{i}\|\nabla L_{i}(\widetilde{\bm{w}}^{(g)},S_{i})\|^{2},$ (4)
+$\sum_{i\in[m],j\in[n_{i}]}\frac{p_{i}}{n_{i}}$ $\left(\mu\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}+\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\rangle\right)\leq\frac{\lambda}{2}R^{2}.$ (5)
+Lemma 1 establishes a relationship between the model estimation error and the tuning parameter $\lambda$.
+From the bound (4), we can see that a larger $\lambda$ will push the global model solution $\widetilde{\bm{w}}^{(g)}$ towards the GlobalTrain solution.
+On the other hand, from the bound (5), we can set a smaller $\lambda\propto 1/nR^{2}$ to obtain a rate $O(1/n)$, independent of $R$.
+This suggests that $\lambda$ can be chosen adaptively w.r.t.
+$R$ to achieve the optimal statistical rate.
+A detailed discussion can be found in Appendix A.3.
 In the following theorem, we provide such a choice and establish a state-of-the-art statistical rate.
 Theorem 1.
-Suppose Assumptions 1, 2 and 3 hold and $n_{i}=n,\forall i\in[m]$.
+Suppose Assumptions 1, 2, 3 and 4 hold and $n_{i}=n,\forall i\in[m]$.
 Consider the FedProx objective in (1) with $p_{i}=\frac{1}{m},\forall i\in[m]$.
-If $\lambda=\frac{\rho}{\sqrt{n}R}\mathbb{I}\{R\leq n^{-1/2}\}+\frac{\rho^{2}}{nR^{2}}\mathbb{I}\{R>n^{-1/2}\}$, the global model $\widetilde{\bm{w}}^{(g)}$ and local models $\{\widetilde{\bm{w}}^{(i)}\}_{i=1}^{m}$ obtained from solving the objective (1) satisfy
-$\mathbb{E}\left\|\widetilde{\bm{w}}^{(g)}-{{\bm{w}}_{\star}^{(g)}}\right\|^{2}\leq\frac{C_{1}}{N}+C_{2}\left(\frac{1}{n}\wedge\frac{R}{\sqrt{n}}\right),$ (4)
-$\mathbb{E}\left\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\right\|^{2}\leq\frac{C_{3}}{N}+C_{4}(\frac{1}{n}\wedge\frac{R}{\sqrt{n}})+16(\frac{\rho^{2}}{\mu^{2}n}\wedge R^{2})\quad\text{ for all }i\in[m],$ (5)
+If it takes $\lambda\geq\max\left\{64\kappa^{2}L,\left(2\kappa\vee 5\right)\mu\frac{2L^{2}R^{2}+\rho^{2}/n}{L^{2}R^{2}+\rho^{2}/N}\right\}-\mu$ when $R\leq\frac{1}{\sqrt{n}}$, and $\lambda\leq\frac{\rho^{2}}{n\mu R^{2}}$ when $R>\frac{1}{\sqrt{n}}$, the global model $\widetilde{\bm{w}}^{(g)}$ and local models $\{\widetilde{\bm{w}}^{(i)}\}_{i=1}^{m}$ obtained from solving the objective (1) satisfy
+$\mathbb{E}\left\|\widetilde{\bm{w}}^{(g)}-{{\bm{w}}_{\star}^{(g)}}\right\|^{2}\leq\frac{C_{1}}{N}+C_{2}\left(\frac{1}{n}\wedge R^{2}\right),$ (6)
+$\mathbb{E}\left\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\right\|^{2}\leq\frac{C_{3}}{N}+C_{4}(\frac{1}{n}\wedge R^{2})\quad\text{ forall }i\in[m],$ (7)
 where $N=mn$ and
-$\begin{split}C_{1}&=\frac{\rho^{2}}{\mu^{2}}=\frac{C_{3}}{8},C_{2}=\big{(}\frac{(1+\sqrt{1+2\mu})\rho}{2\mu}\big{)}^{2}\vee\frac{2\rho}{\mu},C_{4}=\big{(}\frac{\rho(1+(1+4\mu(C_{2}+1))^{\frac{1}{2}})}{2\mu}\big{)}^{2}\wedge(8C_{2}+\frac{2\rho}{\mu}).\end{split}$ (6)
+$C_{1}=32\frac{\rho^{2}}{\mu^{2}},C_{2}=\big{(}\frac{(1+\sqrt{3})\rho}{\mu}\big{)}^{2}\vee\frac{32L^{2}}{\mu^{2}},C_{3}=\frac{192\rho^{2}}{\mu^{2}},C_{4}=\left(\frac{\rho(1+(4C_{2}+5)^{\frac{1}{2}})}{\mu}\right)^{2}\vee\frac{198L^{2}}{\mu^{2}}.$
 Proof.
-See Appendix A.2 for (4) and Appendix A.3 for (5). ∎
+See Appendix A.3. ∎
 Theorem 1 shows that with an appropriate choice of $\lambda$, both the global and local models achieve a fast statistical rate across a wide range of $R$.
-To interpret the result, first we analyze the error bound provided in (4) and (5).
+To interpret the result, we first analyze the error bound provided in (6) and (7).
 The two error bounds have the same order and can be decomposed into two key components.
@@
 This term reflects the sample efficiency gained through collaborative learning.
-The second term, of order $\mathcal{O}(n^{-1}\wedge n^{-1/2}R)$, measures the cost of statistical heterogeneity.
-As $R$ increases, the convergence rate gradually degrades, transitioning from $\mathcal{O}(N^{-1})$ to $\mathcal{O}(n^{-1})$, showing that the error bound of FedProx interpolates between the two extremes: LocalTrain and GlobalTrain.
-Next, to better understand how such an interpolation is achieved through an adaptive personalization degree, we may examine the two extreme cases: as the statistical heterogeneity $R\to\infty$, and $\lambda\to 0$, leading to a high degree of personalization.
-The statistical error bound converges to $\mathcal{O}(n^{-1})$, matching the rate of LocalTrain; as the statistical heterogeneity $R\to 0$, and $\lambda\to\infty$, leading to a low degree of personalization.
-The statistical error bound converges to $\mathcal{O}(N^{-1})$, matching the rate of GlobalTrain.
-As $R\in(0,\infty)$, $\lambda$ is adjusted correspondingly, making FedProx always perform no worse than LocalTrain while benefiting from global training when clients share similarities.
+The second term, of order $\mathcal{O}(n^{-1}\wedge R^{2})$, measures the cost of statistical heterogeneity.
+As $R$ increases, the convergence rate gradually degrades, transitioning from $\mathcal{O}(N^{-1})$ to $\mathcal{O}(n^{-1})$.
+Notably, the guarantee is minimax-optimal, as the derived bound matches the problem’s lower bound under the given assumptions (Chen et al., 2023b).
+Remark 2 in Appendix A.3 discusses about how such an interpolation is achieved through different personalization degrees.
 Table 1: Results Comparison.
@@
 $R>\frac{1}{\sqrt{n}}$ $\frac{1}{m\sqrt{n}}<R\leq\frac{1}{\sqrt{n}}$ $R\leq\frac{1}{m\sqrt{n}}$
-(Chen et al., 2023b) A1,2,3, $D\vee\|\ell\|_{\infty}\leq C$ Two-stage $\frac{D^{2}+\|\ell\|_{\infty}}{n}$ 1 $\frac{\|\ell\|_{\infty}}{\sqrt{n}}R$ $\frac{\|\ell\|_{\infty}}{\sqrt{m}n}$
-Ours A1,2,3 One-stage $\frac{1}{n}$ $\frac{1}{\sqrt{n}}R$ $\frac{1}{mn}$
-Lower Bound - - $\frac{1}{n}$ $R^{2}\vee\frac{1}{mn}$ $\frac{1}{mn}$
-We now compare our theoretical results with the result up to date (Chen et al., 2023b), as summarized in Table 1. Our contribution can be summarized into three aspects:
+Chen et al. (2023b) A1,2,3,4, $D\vee\|\ell\|_{\infty}\leq C$ Two-stage $\frac{D^{2}+\|\ell\|_{\infty}}{n}$ 1 $\frac{\|\ell\|_{\infty}}{\sqrt{n}}R$ $\frac{\|\ell\|_{\infty}}{\sqrt{m}n}$
+Ours A1,2,3,4 One-stage $\frac{1}{n}$ $R^{2}\wedge\frac{1}{mn}$ $\frac{1}{mn}$
+Lower Bound - - $\frac{1}{n}$ $R^{2}\wedge\frac{1}{mn}$ $\frac{1}{mn}$
+We now compare our theoretical results with the result up to date (Chen et al., 2023b), as summarized in Table 1.
+For a fair comparison, it is noticeable that we have transformed their result into our metrics $\mathbb{E}\|{{\bm{w}}_{\star}^{(i)}}-\widetilde{\bm{w}}^{(i)}\|^{2}$, by removing their optimization error and problem-dependent constant.
+Our contribution can be summarized into three aspects:
 $\bullet$ Our results are derived under more realistic assumptions.
@@
 This two-step approach is likely the artifact of their analysis, as existing work on FedProx also shows that a single-stage optimization with a chosen $\lambda$ suffices to achieve the desired performance (T Dinh et al., 2020).
-$\bullet$ We establish a state-of-the-art nearly minimax statistical rate.
-As shown in Table 1, in a heterogeneous case when $R\geq n^{-1/2}$, Chen et al.
+$\bullet$ We establish a state-of-the-art minimax statistical rate.
+As shown in Table 1, the established upper bound matches the lower bound established in Chen et al.
+(2023b).
+In a heterogeneous case when $R\geq n^{-1/2}$, Chen et al.
 (2023b) establishes a rate that is at least of the order $\mathcal{O}\left(n^{-1}R^{2}\right)$ as $D\geq R$.
@@
 In a homogeneous case when $R\leq m^{-1}n^{-1/2}$, they can only establish a rate slower than $\mathcal{O}\left(1/(\sqrt{m}n)\right)$, while our result achieves a rate of $\mathcal{O}\left(1/(mn)\right)$, leveraging all $mn$ samples and matching the rate of GlobalTrain on the IID data.
-Compared to the established lower bound (Chen et al., 2018), our result is nearly minimax optimal.
-Specifically, when $R>n^{-1/2}$ and $R\leq m^{-1}n^{-1/2}$, our result matches the lower bound.
-There is a small gap exists in the case when $m^{-1}n^{-1/2}<R\leq n^{-1/2}$, where we achieve a rate of $\mathcal{O}(n^{-1/2}R)$ whereas the lower bound is $\mathcal{O}(R^{2})$.
-Such a gap is due to the cost of estimating the model difference $\{{{\bm{w}}^{(g)}}-{{\bm{w}}^{(i)}}\}_{i=1}^{m}$ when $\lambda$ is finite (c.f.
-Lemma 1).
-The improvement of this gap is non-trivial, which we defer to future work.
+As the client number could be extremely large ($10^{5}$ devices in chen2023fs), order of $m$ is non-trivial.
+During the transition period when $m^{-1}n^{-1/2}<R<n^{-1/2}$, we achieve a rate of $\mathcal{O}(R^{2})$, again matching the lower bound, and is strictly faster than the rate $\mathcal{O}(n^{-1/2}R)$ established in Chen et al.
+(2023b).
 5 When will Personalization Improve Communication Efficiency
@@
 ${\bm{w}}_{\star}^{(i)}({\bm{w}}^{(g)}):=\operatorname{prox}_{L_{i}/\lambda}({\bm{w}}^{(g)})=\operatorname*{arg\,min}_{{{\bm{w}}^{(i)}}}~{}h_{i}({{\bm{w}}^{(i)}},{\bm{w}}^{(g)}).$ (9)
-As a standard argument about the convergence rate (T Dinh et al., 2020; Li et al., 2021; Chen et al., 2023b), we assume the loss function $\ell$ is $L$-smooth.
-Assumption 4 (Smoothness).
-The loss function $\ell(\cdot,z)$ is L-smooth, i.e. for any $x,y\in\mathbb{R}^{d}$
-$\|\nabla\ell(x,z)-\nabla\ell(y,z)\|\leq L\|x-y\|,\quad\forall z.$ (9)
-Based on Assumption 4 and the new notations in (7) and (8), the following lemma describes the optimizing properties for $F_{i}$, $h_{i}$ and ${\bm{w}}_{\star}^{(i)}$. The proof can be found in B.1.
+Based on Assumption 2 and the new notations in (8) and (9), the following lemma describes the optimizing properties for $F_{i}$, $h_{i}$ and ${\bm{w}}_{\star}^{(i)}$.
+The proof can be found in Appendix B.2.
 Lemma 2.
-Under Assumption 2 and 4, $F_{i}$ is $\mu_{g}$-strongly convex and $L_{g}$-smooth, with $\mu_{g}=\frac{\lambda\mu}{\lambda+\mu}$ and $L_{g}=\frac{\lambda L}{\lambda+L}$, each $h_{i}$ is $\mu_{\ell}$-strongly convex and $L_{\ell}$-smooth, with $\mu_{\ell}=\mu+\lambda$ and $L_{\ell}=L+\lambda$, and the mapping $\bm{w}_{\star}^{(i)}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is $L_{w}$-Lipschitz with $L_{w}=\frac{\lambda}{\lambda+\mu}$.
+Under Assumption 3 and 2, $F_{i}$ is $\mu_{g}$-strongly convex and $L_{g}$-smooth, with $\mu_{g}=\frac{\lambda\mu}{\lambda+\mu}$ and $L_{g}=\frac{\lambda L}{\lambda+L}$, each $h_{i}$ is $\mu_{\ell}$-strongly convex and $L_{\ell}$-smooth, with $\mu_{\ell}=\mu+\lambda$ and $L_{\ell}=L+\lambda$, and the mapping $\bm{w}_{\star}^{(i)}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is $L_{w}$-Lipschitz with $L_{w}=\frac{\lambda}{\lambda+\mu}$.
 Next, we recall the properties of the Moreau envelope to guide the algorithm design.
@@
 The overall procedure is summarized in Alg.
-1.
-Input: Initial global model ${\bm{w}}_{1}^{(g)}$, initial local models $\{{\bm{w}}_{0,K}^{(i)}\}_{i\in[m]}$, global rounds $T$, global step sizes $\gamma$, local rounds $K$, local step sizes $\eta$
-Output: Local models $\{{\bm{w}}_{T,K}^{(i)}\}_{i\in[m]}$ and global model ${\bm{w}}_{T}^{(g)}$
-for $t=1,\dots,T$ do
-       The server sends ${\bm{w}}_{t}^{(g)}$ to client $i$, $\forall i\in[m]$
-       Set ${\bm{w}}_{t,0}^{(i)}={\bm{w}}_{t-1,K}^{(i)}$;
-       for $k=0,\dots,K-1$ do
-             ${\bm{w}}_{t,k+1}^{(i)}={\bm{w}}_{t,k}^{(i)}-\frac{\eta}{n_{i}}\sum_{j\in[n_{i}]}\left\{\nabla\ell({\bm{w}}_{t,k}^{(i)},z_{i,j})+\lambda\left({\bm{w}}_{t,k}^{(i)}-\bm{w}_{t}^{(g)}\right)\right\}$
-      Push ^$\widehat{\nabla}F_{i}(\bm{w}_{t}^{(g)})=\lambda(\bm{w}_{t}^{(g)}-{\bm{w}}_{t,K}^{(i)})$ to the server;
-      $\bm{w}_{t+1}^{(g)}\leftarrow\bm{w}_{t}^{(g)}-\frac{\gamma}{m}\sum_{i\in[m]}\widehat{\nabla}F_{i}(\bm{w}_{t}^{(g)})$;
-Algorithm 1 One-stage FedProx
+1, see Appendix B.1.
 Remark 1.
@@
 While the algorithm design is not new, our contribution lies in analyzing how the personalization degree influences the trade-off between communication and local computation leveraging the FedProx problem structure.
-For simplicity, we focus on the full gradient update in our analysis; both the algorithm and the analysis can be extended to the stochastic setting with mini-batch gradients, with partial participation and incorporating variance reduction techniques.
+Extending to noisy settings to consider stochastic gradient or variance reduction would be interesting future work, which requires additional technical effort.
 Theorem 2.
-Suppose Assumptions 2 and 4 hold.
+Suppose Assumptions 3 and 2 hold.
 Let $\{{\bm{w}}_{t}^{(g)}\}_{t\geq 0}$ and $\{{\bm{w}}_{t,K}^{(i)}\}_{t\geq 0}$ be the sequence generated by Algorithm 1 with $\gamma<1/L_{g}$, $\eta\leq 1/L_{\ell}$, and the inner loop iteration number satisfying
@@
 Proof.
-See Appendix B.2 and B.4. ∎
+See Appendix B.3 and B.5. ∎
 Theorem 2 shows under the step size constraint $\gamma<1/L_{g}$, ${\bm{w}}_{t}^{(g)}$ converges to $\tilde{{\bm{w}}}^{(g)}$ at the same rate as the iteration (10) with exact gradient computation.
@@
 Corollary 1.
-In the same setting as Theorem 2, if choosing the step size $\eta=1/L_{\ell}=(\lambda+L)^{-1}$, $\gamma=1/(2L_{g})=\frac{\lambda+L}{2\lambda L}$, and the inner loop iteration number $K=\tilde{\mathcal{O}}\left(\frac{\lambda+L}{\lambda+\mu}\right)$, then the communication complexity and computation complexity for Algorithm 1 to find an $\varepsilon$-solution (i.e., $\|{\bm{w}}_{T}^{(g)}-\hat{{\bm{w}}}^{(g)}\|^{2}\leq\varepsilon$ and $\|{\bm{w}}_{T,0}^{(i)}-\hat{{\bm{w}}}^{(i)}\|^{2}\leq\varepsilon$) are as follows ($\kappa:=L/\mu$):
+In the same setting as Theorem 2, if choosing the step size $\eta=1/L_{\ell}=(\lambda+L)^{-1}$, $\gamma=1/(2L_{g})=\frac{\lambda+L}{2\lambda L}$, and the inner loop iteration number $K=\tilde{\mathcal{O}}\left(\frac{\lambda+L}{\lambda+\mu}\right)$, then the communication complexity and computation complexity for Algorithm 1 to find an $\varepsilon$-solution (i.e., $\|{\bm{w}}_{T}^{(g)}-\tilde{{\bm{w}}}^{(g)}\|^{2}\leq\varepsilon$ and $\|{\bm{w}}_{T,0}^{(i)}-\tilde{{\bm{w}}}^{(i)}\|^{2}\leq\varepsilon$) are as follows ($\kappa:=L/\mu$):
 (i) # communication $\begin{split}&\text{(i) \# communication rounds}=\mathcal{O}\big{(}\kappa_{g}\log(1/\varepsilon)\big{)}=\mathcal{O}\left(\kappa\cdot\frac{\lambda+\mu}{\lambda+L}\cdot\log\frac{1}{\varepsilon}\right),\end{split}$ (13)
@@
 We first examine the effect of personalization on statistical accuracy under different levels of heterogeneity (Section 6.2), followed by an analysis of its impact on communication efficiency (Section 6.3).
-Detailed descriptions of the experimental setup are available in Section 6.1.
+Detailed descriptions of the experimental setup are available in Section 6.1 and Appendix C, and the complete anonymized codebase is accessible at https://anonymous.4open.science/r/fedprox-bilevel.
 6.1 Experimental Details
@@
 Real Dataset.
-We use the MNIST dataset for the real dataset analysis.
-We first formulate the problem as a convex classification task using multinomial logistic regression, then test the generalization of our theory to a more general, non-convex scenario by conducting experiments using a CNN classifier.
+We use the MNIST, EMNIST, and CIFAR10 datasets for the real dataset analysis.
+We first formulate the problem as a convex classification task using multinomial logistic regression, then test the generalization of our theory to a more general, non-convex scenario by conducting experiments using different CNN classifiers.
 Following Li et al.
-(2020a), to impose statistical heterogeneity, we distribute the MNIST data across clients in the way that each client only has access to $s\in\{2,4,6,8,10\}$ classes.
+(2020a), to impose statistical heterogeneity, we distribute the data across clients in the way that each client only has access to a fixed number of classes.
 The fewer classes each client has access to, the higher the statistical heterogeneity.
-To be consistent with our theory, for both the synthetic and real datasets, we sample 2,000 samples and distribute them evenly among 10 clients, and all clients fully participate in each communication round.
-Additional results with 100 and 1,000 clients under partial participation are provided in Appendix C.
-Selection of $\lambda$.
-For the logistic regression on the synthetic dataset, we selected three values for $\lambda$: small $\lambda=0.02$, medium $\lambda=0.1$, and large $\lambda=0.5$.
-For logistic regression on the MNIST dataset, we used small $\lambda=0.5$, medium $\lambda=1.5$, and large $\lambda=2.5$.
-For CNN on the MNIST dataset, we used small $\lambda=0.2$, medium $\lambda=0.5$, and large $\lambda=2.5$.
-Such a choice is to make sure that each line in the figure is clearly separable.
-Additionally, we implemented an adaptive choice of $\lambda=\frac{\rho}{\sqrt{n}R}\mathbb{I}\{R\leq n^{-1/2}\}+\frac{\rho^{2}}{nR^{2}}\mathbb{I}\{R>n^{-1/2}\}$ following the formula outlined in Theorem 1, with $\rho$ set to 2.
-The value of $\rho$ was determined through grid search over the range $[1,10]$ to ensure optimal performance.
-Selection of Step Size.
-For the synthetic dataset, aligned with Theorem 2, we set the global step size $\gamma=\frac{\lambda+L}{\lambda L}$ and the local step size $\eta=\frac{1}{L+\lambda}$.
-The smoothness constant $L$ is computed as the upper bound of the $L$-smoothness constant for logistic regression, specifically $4^{-1}\Lambda_{\text{max}}$, where $\Lambda_{\text{max}}$ denotes the largest eigenvalue of the Gram matrix $\bm{X}^{\top}\bm{X}$.
-The local step size for the synthetic dataset was further tuned through grid search in $\{0.001,0.01,0.1\}$.
-For the MNIST dataset, the global learning rate was set to $\frac{1}{\lambda}$, while the local learning rate was set to 0.01, chosen via grid search.
-For both GlobalTrain and LocalTrain training baselines, the learning rate was also chosen via grid search within the range $\{0.001,0.01,0.1\}$.
-LocalTrain is implemented using a local (S)GD, and GlobalTrain is implemented using FedAvg (McMahan et al., 2017).
-Selection of Hyperparameters.
-For the MNIST dataset, we implemented logistic regression using the SGD solver with 5 epochs, a batch size of 32, and 20 total runs.
-For the CNN model on the MNIST dataset, the same hyperparameter setup was used, except the batch size was reduced to 16 to ensure stable convergence.
-Optimizer.
-For the synthetic dataset, following the analysis in the paper, the loss function is optimized using gradient descent (GD).
-For the MNIST dataset, all models are implemented in PyTorch, and optimization is performed using the stochastic gradient descent (SGD) solver.
-Evaluation Criterion.
-In terms of evaluation, for the synthetic dataset, we track the ground-truth global model ${{\bm{w}}_{\star}^{(g)}}$ and local model ${{\bm{w}}_{\star}^{(i)}}$.
-We compute the global statistical error as $\|\widetilde{\bm{w}}^{(g)}-{{\bm{w}}_{\star}^{(g)}}\|^{2}$ and the local statistical error as $\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}$.
-To capture the total error at the $t$-th iteration, we report both the global total error $\|\bm{w}_{t}^{(g)}-{{\bm{w}}_{\star}^{(g)}}\|^{2}$ and the local total error $\|{\bm{w}}_{t,0}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}$, which combine both optimization and statistical errors.
-For the MNIST dataset, we report test accuracy on the MNIST testing set as a measure of the statistical error, and training loss for each client’s training set as total error at the $t$-th iteration.
+The detail of data processing to set data heterogeneity for each dataset can be found in Appendix C.
+Implementation and Evaluation.
+For each setup, we evaluate the FedProx algorithm with three different choices of $\lambda$: small, medium, and large, with specific values varying based on the dataset (details provided in Appendix C).
+For the synthetic dataset, in alignment with theorem 2, we implement Algorithm 1 using a global step size $\gamma=(\lambda+L)/(\lambda L)$ and a local step size $\eta=(L+\lambda)^{-1}$.
+For the real dataset, since $L$ is unknown, we implement the same algorithm with a global step size $\gamma$ set to $1/(\lambda)$, while the local step size is determined via grid search.
+In terms of evaluation, for the synthetic dataset, we evaluate by tracking the ground-truth models and measuring error as the distance from these true models.
+For the real dataset, we report training loss and testing accuracy.
+Additional details, including hyperparameter setting and evaluation metric definition, can be found in Appendix C.
 Figure 1: Statistical and total error of different methods.
@@
 However, as heterogeneity increases, LocalTrain and FedProx with smaller $\lambda$ values show better results.
-The figure also demonstrates that with an adaptive choice of $\lambda$ outlined in Theorem 1, FedProx consistently achieves a low error across varying levels of heterogeneity.
+The figure also demonstrates that with an adaptive choice of $\lambda$ in theoretical analysis outlined in Theorem 1, FedProx consistently achieves a low error across varying levels of heterogeneity.
 See more results and discussion in Appendix C.
@@
 7 Conclusion
-In this paper, we analyzed the effect of personalization in FedProx.
-With an adaptive choice of personalization degree, FedProx achieves a nearly minimax-optimal statistical accuracy under different statistical heterogeneity degrees and consistently outperforms pure local training.
+In this paper, we analyzed of the effect of personalization in FedProx.
+With a theoretically-informed adaptive choice of personalization degree, FedProx achieves a minimax-optimal statistical accuracy under different statistical heterogeneity degrees and consistently outperforms pure local training.
 As a minor contribution, we provably show that personalization reduces communication costs without increasing the computation overhead.
 The results are validated through extensive experiments.
-Appendix A Proof of statistical convergence
+Appendix A Proof of statistical converngence
 To facilitate our analysis, we denote
@@
 The statistical error bound is established as $\mathbb{E}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}$ and $\mathbb{E}\|\tilde{\Delta}^{(i)}_{\text{stat}}\|^{2}$, where the expectation is taken over all the data.
+And we also denote
+$\bm{\delta}_{\star}^{(i)}:={{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}},\quad\tilde{\bm{\delta}}^{(i)}:={\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}},$ (17)
+where $\bm{\delta}_{\star}^{(i)}$ measures the difference between the true global model ${{\bm{w}}_{\star}^{(g)}}$ and the true local model ${{\bm{w}}_{\star}^{(i)}}$, and $\tilde{\bm{\delta}}^{(i)}$ is the estimator of such a difference.
 A.1 Discussion on Assumption 1
@@
 $\min_{{{\bm{w}}^{(g)}},\{{{\bm{w}}^{(i)}}\}_{i\in[m]}}\sum_{i=1}^{m}p_{i}\left({L}_{i}({{\bm{w}}^{(i)}},S_{i})+\frac{\lambda_{i}}{2}\|{{\bm{w}}^{(g)}}-{{\bm{w}}^{(i)}}\|^{2}\right),$ (18)
-Under Assumptions 2 and 4 as we assume, Assumption 1 about statistical heterogeneity would be equivalent to many other similar assumptions for statistical heterogeneity, like $B$-dissimilarity (Li et al., 2020a), parameter difference (Chen et al., 2023b) and gradient diversity (T Dinh et al., 2020; Deng et al., 2020).
-A.2 Proof of global statistical Error Bound
-In this part, we will establish the statistical error bound for the global model with fewer assumptions: firstly, in the objective (1), we assume the vector $\mathcal{P}=(p_{1},\cdots,p_{m})$ just under the constraint: $p_{i}\geq 0,\forall i\in[m]$ and $\sum_{i\in[m]}p_{i}=1$; secondly, we assume the data size of the client $i$ is $n_{i}$ for any $i\in[m]$.
-For convenience, we neglect the notation of taking an expectation over all data during the inducing process.
-First, we will prove Lemma 1, which is a crucial part to prompt the analysis of the global statistical error bound.
-Proof.
-We introduce some basic results from the optimality condition of the objective (1), which will be utilized later.
+Under Assumptions 2 and 3 as we assume, Assumption 1 about statistical heterogeneity would be equivalent to many other similar assumptions for statistical heterogeneity, like $B$-dissimilarity (Li et al., 2020a), parameter difference (Chen et al., 2023b) and gradient diversity (T Dinh et al., 2020; Deng et al., 2020).
+A.2 Lemmas and the Proof
+To facilitate the proof of the theorem, we first present two additional lemmas and provide their proof.
+Lemma 4.
+Under Assumption 3, for the optimal solution ${\tilde{\bm{w}}^{(g)}}$ and ${\tilde{\bm{w}}^{(i)}}$’s in Problem (1), we have
+$\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\|_{2}\leq\frac{2\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})\|_{2}}{\mu+\lambda}.$
+By the $\mu$-strongly convexity of $\nabla L_{i}$, we have that for any ${\bm{w}}\in\mathbb{R}^{d}$,
+$L_{i}({\bm{w}},S_{i})+\frac{\lambda}{2}\|{\bm{w}}-{\tilde{\bm{w}}^{(g)}}\|^{2}$ $\geq L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})+\left\langle\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i}),{\bm{w}}-{\tilde{\bm{w}}^{(g)}}\right\rangle+\frac{\mu+\lambda}{2}\|{\bm{w}}-{\tilde{\bm{w}}^{(g)}}\|^{2}$
+$\geq L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})+\|{\bm{w}}-{\tilde{\bm{w}}^{(g)}}\|\left(\frac{\mu+\lambda}{2}\|{\bm{w}}-{\tilde{\bm{w}}^{(g)}}\|-\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})\|\right),$ (19)
+where in the last step we used the Cauchy inequality.
+In addition, notice that ${\tilde{\bm{w}}^{(i)}}$ is the optimal solution for the local model ${{\bm{w}}^{(i)}}$ in the FedProx problem (1), thus we have
+${\tilde{\bm{w}}^{(i)}}=\underset{{\bm{w}}}{\operatorname{argmin}}\left\{L_{i}({\bm{w}},S_{i})+\frac{\lambda}{2}\|{\bm{w}}-{\tilde{\bm{w}}^{(g)}}\|^{2}\right\}.$
+This result together with (A.2) implies that
+$L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})$ $=L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})+\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(g)}}\|^{2}\geq L_{i}({\tilde{\bm{w}}^{(i)}},S_{i})+\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(i)}}-{\tilde{\bm{w}}^{(g)}}\|^{2}$
+$\geq L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})+\|{\tilde{\bm{w}}^{(i)}}-{\tilde{\bm{w}}^{(g)}}\|\left(\frac{\mu+\lambda}{2}\|{\tilde{\bm{w}}^{(i)}}-{\tilde{\bm{w}}^{(g)}}\|-\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})\|\right).$
+Therefore, since $\mu+\lambda>0$ and $\|{\tilde{\bm{w}}^{(i)}}-{\tilde{\bm{w}}^{(g)}}\|\geq 0$, rearranging the terms leads to
+$\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\|\leq\frac{2\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})\|}{\mu+\lambda},$ (20)
+as claimed.
+Lemma 5.
+Under Assumption 1, 2, 3, 4, for $\tilde{\bm{w}}_{\text{GT}}$, the solution of the GlobalTrain problem in (2), we have
+$\mathbb{E}\|\tilde{\bm{w}}_{\text{GT}}-{{\bm{w}}_{\star}^{(g)}}\|\leq\frac{\rho\sqrt{\sum_{i\in[m]}\frac{p_{i}^{2}}{n_{i}}}+LR}{\mu/2}.$
+By the optimality condition of the GlobalTrain problem in (2) and the strong convexity of $L_{i}$’s, we have
+00 0 $\geq\sum_{i\in[m]}p_{i}\left(L_{i}(\tilde{\bm{w}}_{\text{GT}},S_{i})-L_{i}({{\bm{w}}_{\star}^{(g)}},S_{i})\right)$
+$\geq\sum_{i\in[m]}p_{i}\left\langle\nabla L_{i}({{\bm{w}}_{\star}^{(g)}},S_{i}),\tilde{\bm{w}}_{\text{GT}}-{{\bm{w}}_{\star}^{(g)}}\right\rangle+\frac{\mu}{2}\|\tilde{\bm{w}}_{\text{GT}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}$
+$\geq-\Big{\|}\sum_{i\in[m]}p_{i}\nabla L_{i}({{\bm{w}}_{\star}^{(g)}},S_{i})\Big{\|}\Big{\|}\tilde{\bm{w}}_{\text{GT}}-{{\bm{w}}_{\star}^{(g)}}\Big{\|}+\frac{\mu}{2}\|\tilde{\bm{w}}_{\text{GT}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}.$
+Therefore, we have
+$\|\tilde{\bm{w}}_{\text{GT}}-{{\bm{w}}_{\star}^{(g)}}\|$ $\leq\frac{\Big{\|}\sum_{i\in[m]}p_{i}\nabla L_{i}({{\bm{w}}_{\star}^{(g)}},S_{i})\Big{\|}}{\mu/2}$
+$\leq\frac{\Big{\|}\sum_{i\in[m]}p_{i}\nabla L_{i}({{\bm{w}}_{\star}^{(i)}},S_{i})\Big{\|}+\sum_{i\in[m]}p_{i}\Big{\|}\nabla L_{i}({{\bm{w}}_{\star}^{(g)}},S_{i})-\nabla L_{i}({{\bm{w}}_{\star}^{(i)}},S_{i})\Big{\|}}{\mu/2}$ (21)
+By the $L$-smoothness property of $L_{i}$’s in Assumption 2 and the bounded gradient property listed in Assumption 4, we have
+$\sum_{i\in[m]}p_{i}\Big{\|}\nabla L_{i}({{\bm{w}}_{\star}^{(g)}},S_{i})-\nabla L_{i}({{\bm{w}}_{\star}^{(i)}},S_{i})\Big{\|}\leq\sum_{i\in[m]}p_{i}L\|{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|\leq LR,$
+$\mathbb{E}\Big{\|}\sum_{i\in[m]}p_{i}\nabla L_{i}({{\bm{w}}_{\star}^{(i)}},S_{i})\Big{\|}=\mathbb{E}\Big{\|}\sum_{i\in[m]}\sum_{j\in[n_{i}]}\frac{p_{i}}{n_{i}}\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij})\Big{\|}\leq\rho\sqrt{\sum_{i\in[m]}\frac{p_{i}^{2}}{n_{i}}}.$
+These results together with (A.2) implies
+$\mathbb{E}\|\tilde{\bm{w}}_{\text{GT}}-{{\bm{w}}_{\star}^{(g)}}\|\leq\frac{\rho\sqrt{\sum_{i\in[m]}\frac{p_{i}^{2}}{n_{i}}}+LR}{\mu/2}$
+as claimed.
+A.3 Proof of Lemma 1 and Theorem 1
+A.3.1 Proof of Lemma 1
+Lemma 1 is a direct consequence of the proof of Theorem 1. To see this, from (A.3.2) and the fact that $(\mu+\lambda)^{2}\geq\mu(\mu+\lambda)$, we have
+$\|{\tilde{\bm{w}}^{(g)}}-\tilde{\bm{w}}_{\text{GT}}\|^{2}$ $\leq\frac{4}{\mu}\left(\frac{1}{\mu+\lambda}+\frac{L}{(\mu+\lambda)^{2}}\right)\sum_{i}p_{i}\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})\|^{2}$
+$\leq\frac{4(1+L/\mu)}{\mu(\mu+\lambda)}\sum_{i}p_{i}\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})\|^{2},$
+thus (4) holds. In addition, with the fact that $\|\tilde{\bm{\delta}}^{(i)}\|_{2}\geq 0$, from (28) we find
+00 0 $\geq-\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle+\frac{\mu}{2}\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\sum_{i}p_{i}\left\|\tilde{\bm{\delta}}^{(i)}\right\|^{2}-\frac{\lambda}{2}R^{2}$
+$\geq-\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle+\frac{\mu}{2}\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}-\frac{\lambda}{2}R^{2}.$
+which implies the result in (5).
+A.3.2 Proof of Global Statistical Error Bound in Theorem 1
+Case 1: We first consider the case when $R>\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$.
 Recall that ${\tilde{\bm{w}}^{(g)}}$ and $\{{\tilde{\bm{w}}^{(i)}}\}_{i\in[m]}$ are the minimizers of the objective (1). According to the first-order condition, we have
@@
 ${\tilde{\bm{w}}^{(g)}}=\sum_{i}p_{i}{\tilde{\bm{w}}^{(i)}}.$ (23)
-To facilitate our analysis, we denote
-$\bm{\delta}_{\star}^{(i)}:={{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}},\quad\tilde{\bm{\delta}}^{(i)}:={\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}.$ (20)
-that is, $\bm{\delta}_{\star}^{(i)}$ measures the difference between the true global model ${{\bm{w}}_{\star}^{(g)}}$ and the true local model ${{\bm{w}}_{\star}^{(i)}}$, and $\tilde{\bm{\delta}}^{(i)}$ is the estimator of such a difference.
-To prove the global statistical error bound, we start with the optimality condition of ${\tilde{\bm{w}}^{(i)}}$ and ${\tilde{\bm{w}}^{(g)}}$, which yields
+We start with the optimality condition of ${\tilde{\bm{w}}^{(i)}}$ and ${\tilde{\bm{w}}^{(g)}}$, which yields
 00 0 $\geq\sum_{i\in[m]}p_{i}\left[L_{i}({\tilde{\bm{w}}^{(i)}},S_{i})+\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\|^{2}\right]-\sum_{i\in[m]}p_{i}\left[L_{i}({{\bm{w}}_{\star}^{(i)}},S_{i})+\frac{\lambda}{2}\|{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}\right].$ (24)
@@
 00 0 $\geq\sum_{i\in[m]}\sum_{j\in[n_{i}]}\frac{p_{i}}{n_{i}}\left(\ell({\tilde{\bm{w}}^{(i)}},z_{ij})-\ell({{\bm{w}}_{\star}^{(i)}},z_{ij})\right)+\sum_{i\in[m]}p_{i}\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\|^{2}-\sum_{i\in[m]}p_{i}\frac{\lambda}{2}\|{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}.$ (25)
-For the first term on the RHS in (22), apply the $\mu$-strongly convexity of the loss function $\ell$ (Assumption 2)
-$\begin{split}\sum_{i\in[m]}\sum_{j\in[n_{i}]}\frac{p_{i}}{n_{i}}\left(\ell({\tilde{\bm{w}}^{(i)}},z_{ij})-\ell({{\bm{w}}_{\star}^{(i)}},z_{ij})\right)&\geq-\sum_{i\in[m]}\sum_{j\in[n_{i}]}\frac{p_{i}}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle\\ &+\mu\sum_{i\in[m]}\sum_{j\in[n_{i}]}\frac{p_{i}}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2},\\ \end{split}$ (23)
+For the first term on the RHS in (25), apply the $\mu$-strongly convexity of the loss function $\ell$ (cf. Assumption 3), we then have
+$\begin{split}\sum_{i\in[m]}\sum_{j\in[n_{i}]}\frac{p_{i}}{n_{i}}\left(\ell({\tilde{\bm{w}}^{(i)}},z_{ij})-\ell({{\bm{w}}_{\star}^{(i)}},z_{ij})\right)&\geq-\sum_{i\in[m]}\sum_{j\in[n_{i}]}\frac{p_{i}}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle\\ &+\frac{\mu}{2}\sum_{i\in[m]}\sum_{j\in[n_{i}]}\frac{p_{i}}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2},\\ \end{split}$ (26)
 where we use the definition of $\tilde{\Delta}^{(i)}_{\text{stat}}$ in (16).
 Plugging (26) back into (25), it yields
-$\begin{split}0&\geq-\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle+\mu\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\\ &+\sum_{i\in[m]}p_{i}\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\|^{2}-\sum_{i\in[m]}p_{i}\frac{\lambda}{2}\|{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}.\end{split}$ (24)
+$\begin{split}0&\geq-\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle+\frac{\mu}{2}\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\\ &+\sum_{i\in[m]}p_{i}\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\|^{2}-\sum_{i\in[m]}p_{i}\frac{\lambda}{2}\|{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}.\end{split}$ (27)
 Recall the definition in (17) and the Assumption 1 about statistical heterogeneity
-00 0 $\geq-\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle+\mu\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\sum_{i}p_{i}\left\|\tilde{\bm{\delta}}^{(i)}\right\|^{2}-\frac{\lambda}{2}R^{2}.$ (25)
-Utilizing the result (25) in Lemma 1, we are ready to proof Theorem 1
-Proof.
-Based on the result in (25), we will discuss the upper bound of $\tilde{\Delta}^{(g)}_{\text{stat}}$ with two cases of statistical heterogeneoity:
-Case 1: for the first case, if $R$ is large, i.e, $R>\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$.
-First, applying Cauchy inequality on the first term of (25), it yields
+00 0 $\geq-\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle+\frac{\mu}{2}\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\sum_{i}p_{i}\left\|\tilde{\bm{\delta}}^{(i)}\right\|^{2}-\frac{\lambda}{2}R^{2}.$ (28)
+Applying Cauchy inequality on the first term of (28), it yields
 $\begin{split}\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle&=\sum_{i}p_{i}\left\langle\sum_{j}\frac{1}{n_{i}}\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle\\ &\leq\sum_{i}p_{i}\left\|\sum_{j}\frac{1}{n_{i}}\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij})\right\|_{2}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}.\end{split}$ (29)
@@
 Combining (29) and (30), it yields
-$\begin{split}\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\mathbb{E}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle\leq\sum_{i}\rho\frac{p_{i}}{\sqrt{n_{i}}}\|\tilde{\Delta}^{(i)}_{\text{stat}}\|_{2}.\\ \end{split}$ (28)
+$\begin{split}\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\mathbb{E}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle\leq\sum_{i}\rho\frac{p_{i}}{\sqrt{n_{i}}}\mathbb{E}\|\tilde{\Delta}^{(i)}_{\text{stat}}\|_{2}.\\ \end{split}$ (31)
 Plugging (31) back to (28) yields
-$0\geq-\rho\sum_{i}\frac{p_{i}}{\sqrt{n_{i}}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\mu\sum_{i}p_{i}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\sum_{i}p_{i}\left\|\tilde{\bm{\delta}}^{(i)}\right\|^{2}-\frac{\lambda}{2}R^{2}.$ (29)
+$0\geq-\rho\sum_{i}\frac{p_{i}}{\sqrt{n_{i}}}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\frac{\mu}{2}\sum_{i}p_{i}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\sum_{i}p_{i}\mathbb{E}\left\|\tilde{\bm{\delta}}^{(i)}\right\|^{2}-\frac{\lambda}{2}R^{2}.$ (32)
 After dropping the third term and applying Cauchy inequality, we obtain
-$0\geq-\rho\left(\Big{(}\sum_{i\in[m]}\frac{p_{i}}{n_{i}}\Big{)}\Big{(}\sum_{i}p_{i}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\Big{)}\right)^{\frac{1}{2}}+\mu\sum_{i}p_{i}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}-\frac{\lambda}{2}R^{2}.$ (30)
-The equation (30) would be regarded as a quadratic function of $\sum_{i}p_{i}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}$
-Set the $\lambda$ as
-$\lambda=\frac{\rho^{2}}{R^{2}}\sum_{i\in[m]}\frac{p_{i}}{n_{i}}.$ (31)
-Solving the inequality in (30) w.r.t $\sum_{i}p_{i}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}$, it yields
-$\begin{split}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2}&\leq\left(\sum_{i}p_{i}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\right)^{\frac{1}{2}}\leq\frac{1}{2\mu}\left(\rho\left(\sum_{i\in[m]}\frac{p_{i}}{n_{i}}\right)^{\frac{1}{2}}+\left(\rho^{2}\sum_{i\in[m]}\frac{p_{i}}{n_{i}}(1+2\mu)\right)^{\frac{1}{2}}\right)\\ &=\frac{(1+\sqrt{1+2\mu})\rho}{2\mu}\left(\sum_{i\in[m]}\frac{p_{i}}{n_{i}}\right)^{\frac{1}{2}}.\\ \end{split}$ (32)
-Case 2: for the second case, if $R$ is relatively small, which means $R\leq\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$, based on the result (25), we know
-$\begin{split}0&\geq\sum_{i\in[m]}\sum_{j\in[n_{i}]}\frac{p_{i}}{n_{i}}\left[\ell\left({\tilde{\bm{w}}^{(i)}},x_{ij}\right)-\ell({{\bm{w}}_{\star}^{(i)}},x_{ij})\right]+\frac{\lambda}{2}\sum_{i\in[m]}p_{i}\left(\left\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\right\|^{2}-\left\|{{\bm{w}}_{\star}^{(g)}}-\bm{w}_{\star}^{(i)}\right\|^{2}\right)\\ &\geq-\underbrace{\sum_{i\in[m]}\sum_{j\in[n_{i}]}\frac{p_{i}}{n_{i}}\left\langle\nabla\ell({\tilde{\bm{w}}^{(i)}},x_{ij}),\tilde{\Delta}^{(g)}_{\text{stat}}\right\rangle}_{T_{1}}+\underbrace{\sum_{i\in[m]}\sum_{j\in[n_{i}]}\frac{p_{i}}{n_{i}}\left\langle\nabla\ell({\tilde{\bm{w}}^{(i)}},x_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}-\tilde{\Delta}^{(g)}_{\text{stat}}\right\rangle}_{T_{2}}\\ &\quad+\sum_{i\in[m]}\sum_{j\in[n_{i}]}\frac{p_{i}}{n_{i}}\mu\|\tilde{\Delta}^{(i)}_{\text{stat}}\|^{2}+\frac{\lambda}{2}\sum_{i\in[m]}p_{i}\|\tilde{\bm{\delta}}^{(i)}\|^{2}-\frac{\lambda}{2}\sum_{i\in[m]}p_{i}\|\bm{\delta}_{\star}^{(i)}\|^{2}.\end{split}$ (33)
-It is necessary to remark that in the case of $R\leq\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$, we should take the GlobalTrain (2) into account so that our analysis different from the case 1.
-For $T_{1}$
-$T_{1}\leq\big{\|}\sum_{ij}\frac{p_{i}}{n_{i}}\nabla\ell({\tilde{\bm{w}}^{(i)}},z_{ij})\big{\|}_{2}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2}\leq\rho\sqrt{\sum_{i}\frac{p_{i}^{2}}{n_{i}}}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2},$ (34)
-where the first inequality comes from Cauchy inequality and to obtain the second inequality, we need to note the fact that from Assumption 3, we have
-$\left(\mathbb{E}\left\|\sum_{ij}\frac{p_{i}}{n_{i}}\nabla\ell(\widetilde{\bm{w}}^{(i)},z_{ij})\right\|_{2}\right)^{2}\leq\mathbb{E}\left\|\sum_{ij}\frac{p_{i}}{n_{i}}\nabla\ell(\widetilde{\bm{w}}^{(i)},z_{ij})\right\|_{2}^{2}\leq\sum_{ij}\frac{p_{i}^{2}}{n_{i}^{2}}\rho^{2}.$ (35)
-Analogy to the induction of $T_{1}$, we can induce $T_{2}$ in a similar way
-$\begin{split}T_{2}&\leq\sum_{i}p_{i}\left\langle\frac{1}{n_{i}}\sum_{j}\nabla\ell({\tilde{\bm{w}}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}-\tilde{\Delta}^{(g)}_{\text{stat}}\right\rangle\\ &\leq\sum_{i}p_{i}\left\|\frac{1}{n_{i}}\sum_{j}\nabla\ell({\tilde{\bm{w}}^{(i)}},z_{ij})\right\|_{2}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}-\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2}\\ &\leq\sum_{i}p_{i}\frac{\rho}{\sqrt{n_{i}}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}-\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2}.\\ \end{split}$ (36)
-Note the fact
-$\left\|\tilde{\Delta}^{(i)}_{\text{stat}}-\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2}=\left\|{{\bm{w}}_{\star}^{(i)}}-{\tilde{\bm{w}}^{(i)}}-{{\bm{w}}_{\star}^{(g)}}+{\tilde{\bm{w}}^{(g)}}\right\|_{2}=\left\|\bm{\delta}_{\star}^{(i)}-\tilde{\bm{\delta}}^{(i)}\right\|_{2}.$ (37)
-Combining (37) back to (36), it yields
-$T_{2}\leq\sum_{i}p_{i}\frac{\rho}{\sqrt{n_{i}}}\left\|\bm{\delta}_{\star}^{(i)}-\tilde{\bm{\delta}}^{(i)}\right\|_{2}\leq\sum_{i}p_{i}\frac{\rho}{\sqrt{n_{i}}}\left(\left\|\bm{\delta}_{\star}^{(i)}\right\|_{2}+\left\|\tilde{\bm{\delta}}^{(i)}\right\|_{2}\right).$ (38)
-Plugging (34) and (38) into (33), it yields
-$\begin{split}0&\geq-\rho\sqrt{\sum_{i}\frac{p_{i}^{2}}{n_{i}}}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|_{2}-\sum_{i}p_{i}\frac{\rho}{\sqrt{n_{i}}}(\|\bm{\delta}_{\star}^{(i)}\|_{2}+\|\tilde{\bm{\delta}}^{(i)}\|_{2})+\mu\sum_{i}p_{i}\|\tilde{\Delta}^{(i)}_{\text{stat}}\|^{2}\\ &\quad\quad+\frac{\lambda}{2}\sum_{i}p_{i}\|\tilde{\bm{\delta}}^{(i)}\|^{2}-\frac{\lambda}{2}\sum_{i\in[m]}p_{i}\|\bm{\delta}_{\star}^{(i)}\|^{2}\\ &\geq\mu\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}-\rho\sqrt{\sum_{i}\frac{p_{i}^{2}}{n_{i}}}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|_{2}-\rho\sum_{i}\frac{p_{i}}{\sqrt{n_{i}}}\|\tilde{\bm{\delta}}^{(i)}\|_{2}+\frac{\lambda}{2}\sum_{i}p_{i}\|\tilde{\bm{\delta}}^{(i)}\|^{2}-R\rho\sum_{i}\frac{p_{i}}{\sqrt{n_{i}}}-\frac{\lambda}{2}R^{2},\end{split}$ (39)
-where the last line uses Assumption 1. Then focusing on the right-hand side of the above inequality and regarding it as a function of $\tilde{\bm{\delta}}^{(i)}$, it can be written as
-$\begin{split}&\quad-\rho\sum_{i}\frac{p_{i}}{\sqrt{n_{i}}}\|\tilde{\bm{\delta}}^{(i)}\|_{2}+\frac{\lambda}{2}\sum_{i}p_{i}\|\tilde{\bm{\delta}}^{(i)}\|^{2}\\ &=\sum_{i}\left(\frac{\lambda}{2}p_{i}\|\tilde{\bm{\delta}}^{(i)}\|^{2}-\rho\frac{p_{i}}{\sqrt{n_{i}}}\|\tilde{\bm{\delta}}^{(i)}\|_{2}\right)\\ &\geq-\frac{\rho^{2}}{2\lambda}\sum_{i}{\frac{p_{i}}{n_{i}}},\end{split}$ (40)
-where we choose the optimal $\|\tilde{\bm{\delta}}^{(i)}\|_{2}^{\star}$ as $\frac{\rho}{\lambda\sqrt{n_{i}}}$.
-Plugging (LABEL:global:_case2_opt5) into (39), it yields
-$0\geq\mu\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}-\rho\sqrt{\sum_{i}\frac{p_{i}^{2}}{n_{i}}}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|_{2}-\frac{\rho^{2}}{2\lambda}\sum_{i}\frac{p_{i}}{n_{i}}-R\rho\sum_{i}\frac{p_{i}}{\sqrt{n_{i}}}-\frac{\lambda}{2}R^{2}.$ (41)
-Then we maximize the RHS of (41) w.r.t $\lambda$, the optimal solution of $\lambda$ is $\frac{\rho}{R}\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$. We add the result into (41) and it yields
-00 0 $\geq\mu\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}-\rho\sqrt{\sum_{i}\frac{p_{i}^{2}}{n_{i}}}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|_{2}-\rho R\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}-\rho R\sum_{i}\frac{p_{i}}{\sqrt{n_{i}}}.$ (42)
-We can regard (42) as a function of $\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}$ and induce
-$\begin{split}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|_{2}&\leq\frac{1}{2\mu}\left\{\rho\sqrt{\sum_{i}\frac{p_{i}^{2}}{n_{i}}}+\left(\rho^{2}\sum_{i}\frac{p_{i}^{2}}{n_{i}}+4\mu\rho R\left(\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}+\sum_{i}\frac{p_{i}}{\sqrt{n_{i}}}\right)\right)^{\frac{1}{2}}\right\}.\end{split}$ (43)
-Thus
-$\begin{split}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}\leq\frac{\rho^{2}}{\mu^{2}}\sum_{i}\frac{p^{2}_{i}}{n_{i}}+\frac{2\rho R}{\mu}\left(\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}+\sum_{i}\frac{p_{i}}{\sqrt{n_{i}}}\right).\end{split}$ (44)
-Combine two-case arguments, we have:
-$\Rightarrow\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|^{2}\leq\begin{cases}\left(\frac{(1+\sqrt{1+2\mu})\rho}{2\mu}\right)^{2}\sum_{i\in[m]}\frac{p_{i}}{n_{i}}&R>\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}\\ \frac{\rho^{2}}{\mu^{2}}\sum_{i}\frac{p^{2}_{i}}{n_{i}}+\frac{2\rho R}{\mu}\left(\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}+\sum_{i}\frac{p_{i}}{\sqrt{n_{i}}}\right)&R\leq\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}.\end{cases}$ (45)
+$0\geq-\rho\left(\Big{(}\sum_{i\in[m]}\frac{p_{i}}{n_{i}}\Big{)}\Big{(}\sum_{i}p_{i}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\Big{)}\right)^{\frac{1}{2}}+\frac{\mu}{2}\sum_{i}p_{i}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}-\frac{\lambda}{2}R^{2}.$ (33)
+The equation (33) would be regarded as a quadratic function of $\sum_{i}p_{i}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}$. If we set the $\lambda$ as
+$\lambda\leq\frac{\rho^{2}}{\mu R^{2}}\sum_{i\in[m]}\frac{p_{i}}{n_{i}},$ (34)
+then solving the inequality in (33) w.r.t $\sum_{i}p_{i}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}$ yields
+$\begin{split}\mathbb{E}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2}&\leq\left(\sum_{i}p_{i}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\right)^{\frac{1}{2}}\leq\frac{1}{\mu}\left(\rho\left(\sum_{i\in[m]}\frac{p_{i}}{n_{i}}\right)^{\frac{1}{2}}+\left(\rho^{2}\sum_{i\in[m]}\frac{3p_{i}}{n_{i}}\right)^{\frac{1}{2}}\right)\\ &=\frac{(1+\sqrt{3})\rho}{\mu}\left(\sum_{i\in[m]}\frac{p_{i}}{n_{i}}\right)^{\frac{1}{2}}.\\ \end{split}$ (35)
+Case 2: Then we consider $R\leq\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$.
+Recall the definition of the global minimizer of GlobalTraining $\tilde{\bm{w}}_{\text{GT}}=\arg\min_{w}\sum_{i}p_{i}L_{i}(w;S_{i})$, based on the optimality condition of $\widetilde{\bm{w}}_{i},\widetilde{\bm{w}}_{g}$, $i\in[m]$, we obtain
+$\begin{split}\sum_{i}p_{i}\left(L_{i}({\tilde{\bm{w}}^{(i)}},S_{i})+\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(i)}}-{\tilde{\bm{w}}^{(g)}}\|^{2}\right)&\leq\sum_{i}p_{i}\left(L_{i}(\tilde{\bm{w}}_{\text{GT}},S_{i})+\frac{\lambda}{2}\|\ \tilde{\bm{w}}_{\text{GT}}-\tilde{\bm{w}}_{\text{GT}}\|^{2}\right)\\ &=\sum_{i}p_{i}L_{i}(\tilde{\bm{w}}_{\text{GT}},S_{i}).\end{split}$ (36)
+Therefore, we obtain
+$\sum_{i}p_{i}L_{i}(\tilde{\bm{w}}_{\text{GT}},S_{i})$ $\geq\sum_{i}p_{i}L_{i}({\tilde{\bm{w}}^{(i)}},S_{i}).$ (37)
+In addition, applying the smoothness assumption in Assumption 2 yields
+$\left|L_{i}({\tilde{\bm{w}}^{(i)}},S_{i})-L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})-\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})^{T}({\tilde{\bm{w}}^{(i)}}-{\tilde{\bm{w}}^{(g)}})\right|$ $\leq\frac{L}{2}\|{\tilde{\bm{w}}^{(i)}}-{\tilde{\bm{w}}^{(g)}}\|^{2}.$ (38)
+Combining (37) and (38), we have
+$\sum_{i}p_{i}L_{i}(\tilde{\bm{w}}_{\text{GT}},S_{i})$ $\geq\sum_{i}p_{i}L_{i}({\tilde{\bm{w}}^{(i)}},S_{i})$ (39)
+$\geq\sum_{i}p_{i}L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})+\sum_{i}p_{i}\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})^{\top}({\tilde{\bm{w}}^{(i)}}-{\tilde{\bm{w}}^{(g)}})-\sum_{i}p_{i}\frac{L}{2}\|{\tilde{\bm{w}}^{(i)}}-{\tilde{\bm{w}}^{(g)}}|^{2}.$ (40)
+Using Cauchy inequality on the second term of the R.H.S.,
+$\sum_{i}p_{i}L_{i}(\tilde{\bm{w}}_{\text{GT}},S_{i})$ $\geq\sum_{i}p_{i}L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})-\sum_{i}p_{i}\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})\|_{2}\|{\tilde{\bm{w}}^{(i)}}-{\tilde{\bm{w}}^{(g)}}\|_{2}$
+$\quad-\sum_{i}p_{i}\frac{L}{2}\|{\tilde{\bm{w}}^{(i)}}-{\tilde{\bm{w}}^{(g)}}\|^{2}.$ (41)
+On the other hand, note that the optimality condition of $\tilde{\bm{w}}_{\text{GT}}$ is
+$\sum_{i}p_{i}\nabla L_{i}(\tilde{\bm{w}}_{\text{GT}},S_{i})=0,$ (42)
+Using the strong-convexity assumption in Assumption 3, we can show that
+$\sum_{i}p_{i}\left\{L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})-L_{i}(\tilde{\bm{w}}_{\text{GT}},S_{i})-\nabla L_{i}(\tilde{\bm{w}}_{\text{GT}},S_{i})^{T}({\tilde{\bm{w}}^{(g)}}-\tilde{\bm{w}}_{\text{GT}})\right\}$
+$=\sum_{i}p_{i}\left\{L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})-L_{i}(\tilde{\bm{w}}_{\text{GT}},S_{i})\right\}$
+$\geq\sum_{i}p_{i}\frac{\mu}{2}\|{\tilde{\bm{w}}^{(g)}}-\tilde{\bm{w}}_{\text{GT}}\|^{2}$
+$=\frac{\mu}{2}\|{\tilde{\bm{w}}^{(g)}}-\tilde{\bm{w}}_{\text{GT}}\|^{2}.$ (43)
+Combining the results (A.3.2) and (A.3.2), it yields
+$\sum_{i}p_{i}L_{i}(\tilde{\bm{w}}_{\text{GT}},S_{i})$ $\geq\sum_{i}p_{i}L_{i}(\tilde{\bm{w}}_{\text{GT}},S_{i})+\frac{\mu}{2}\|{\tilde{\bm{w}}^{(g)}}-\tilde{\bm{w}}_{\text{GT}}\|^{2}$
+$-\sum_{i}p_{i}\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})\|_{2}\|{\tilde{\bm{w}}^{(i)}}-{\tilde{\bm{w}}^{(g)}}\|_{2}-\sum_{i}p_{i}\frac{L}{2}\|{\tilde{\bm{w}}^{(i)}}-{\tilde{\bm{w}}^{(g)}}\|^{2}.$ (44)
+Reorganizing these terms yields
+$\frac{\mu}{2}\|{\tilde{\bm{w}}^{(g)}}-\tilde{\bm{w}}_{\text{GT}}\|^{2}\leq\sum_{i}p_{i}\left\{\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})\|_{2}\|{\tilde{\bm{w}}^{(i)}}-{\tilde{\bm{w}}^{(g)}}\|_{2}+\frac{L}{2}\|{\tilde{\bm{w}}^{(i)}}-{\tilde{\bm{w}}^{(g)}}\|^{2}\right\}.$ (45)
+Using Lemma 4 to bound the term $\|{\tilde{\bm{w}}^{(i)}}-{\tilde{\bm{w}}^{(g)}}\|^{2}$, we have
+$\|{\tilde{\bm{w}}^{(g)}}-\tilde{\bm{w}}_{\text{GT}}\|^{2}$ $\leq\frac{2}{\mu}\sum_{i}p_{i}\left[\frac{2}{\mu+\lambda}+\frac{2L}{(\mu+\lambda)^{2}}\right]\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})\|^{2}$
+$=\frac{4}{\mu}\left(\frac{1}{\mu+\lambda}+\frac{L}{(\mu+\lambda)^{2}}\right)\sum_{i}p_{i}\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})\|^{2}.$ (46)
+Note that by the smoothness Assumption and triangle inequality,
+$\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})\|_{2}^{2}$ $\leq 2\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})-\nabla L_{i}({{\bm{w}}_{\star}^{(i)}},S_{i})\|^{2}+2\|\nabla L_{i}({{\bm{w}}_{\star}^{(i)}},S_{i})\|^{2}$
+$\leq 2L^{2}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}+2\|\nabla L_{i}({{\bm{w}}_{\star}^{(i)}},S_{i})\|^{2}.$ (47)
+Take expectation w.r.t $S_{i}$ and use Assumption 1 and 4, then we have
+$\mathbb{E}[\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})\|^{2}]$ $\leq 2L^{2}\mathbb{E}[\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}]+2\mathbb{E}\left[\left\|\frac{1}{n_{i}}\sum_{j=1}^{n_{i}}\nabla L({{\bm{w}}_{\star}^{(i)}},z_{ij})\right\|^{2}\right]$
+$\leq 4L^{2}\mathbb{E}[\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}]+4L^{2}R^{2}+2\frac{\rho^{2}}{n},$ (48)
+where the last term comes from (30).
+Plugging (A.3.2) back into (A.3.2), it yields
+$\mathbb{E}[\|{\tilde{\bm{w}}^{(g)}}-\tilde{\bm{w}}_{\text{GT}}\|^{2}]$ $\leq\frac{4}{\mu}\left(\frac{1}{\mu+\lambda}+\frac{L}{(\mu+\lambda)^{2}}\right)\left\{4L^{2}\mathbb{E}[\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}]+4L^{2}R^{2}+2\frac{\rho^{2}}{n}\right\}.$ (49)
+Next, we are ready to establish the global error bound. Using the result (49) and Lemma 5, we obtain
+$\mathbb{E}[\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}]$ $\leq\frac{4}{\mu}\left(\frac{1}{\mu+\lambda}+\frac{L}{(\mu+\lambda)^{2}}\right)\left\{4L^{2}\mathbb{E}[\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}]+4L^{2}R^{2}+2\frac{\rho^{2}}{n}\right\}$
+$\quad+\frac{8\rho^{2}}{\mu^{2}}\frac{1}{N}+\frac{8L^{2}}{\mu^{2}}R^{2}.$ (50)
+Let’s define $g(\lambda)=\frac{4}{\mu}\left(\frac{1}{\mu+\lambda}+\frac{L}{(\mu+\lambda)^{2}}\right)$ and reorganize (A.3.2), which yields
+$\mathbb{E}[\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}]$ $\leq\frac{g(\lambda)\left(4L^{2}R^{2}+2\frac{\rho^{2}}{n}\right)+\frac{8\rho^{2}}{\mu^{2}}\frac{1}{N}+\frac{8L^{2}}{\mu^{2}}R^{2}}{1-4L^{2}g(\lambda)}.$ (51)
+If we assume
+$g(\lambda)$ $\leq\frac{1}{8L^{2}}\land\left(\left(\frac{4\rho^{2}}{\mu^{2}N}+\frac{4L^{2}R^{2}}{\mu^{2}}\right)\Big{/}\left(2L^{2}R^{2}+\frac{\rho^{2}}{n}\right)\right),$ (52)
+we have
+$\mathbb{E}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}$ $\leq 32\left(\frac{\rho^{2}}{\mu^{2}}\frac{1}{N}+\frac{L^{2}}{\mu^{2}}R^{2}\right).$ (53)
+Furthermore, we can get a simplified condition for $\lambda$ as
+$\lambda\geq\frac{8\kappa}{\mu}\left\{8L^{2}\vee\left(\left(2L^{2}R^{2}+\frac{\rho^{2}}{n}\right)\Big{/}\left(\frac{4\rho^{2}}{\mu^{2}N}+\frac{4L^{2}R^{2}}{\mu^{2}}\right)\right)\right\}-\mu$ (54)
+where $\kappa=\frac{L}{\mu}\geq 1$ and the condition that $\lambda$ is non-negative holds trivially as $\frac{\kappa L^{2}}{\mu}=\kappa^{2}L\geq\mu$.
+Combine two-case arguments, we have
+$\Rightarrow\mathbb{E}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}\leq\begin{cases}\left(\frac{(1+\sqrt{3})\rho}{\mu}\right)^{2}\sum_{i\in[m]}\frac{p_{i}}{n_{i}}&R>\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}\\ 32\left(\frac{\rho^{2}}{\mu^{2}}\frac{1}{N}+\frac{L^{2}}{\mu^{2}}R^{2}\right)&R\leq\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}.\end{cases}$ (55)
 Thus it yields a one-line rate
-$\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|^{2}\leq C_{1}\sum_{i}\frac{p_{i}^{2}}{n_{i}}+C_{2}\left(\sum_{i\in[m]}\frac{p_{i}}{n_{i}}\wedge R\left(\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}+\sum_{i}\frac{p_{i}}{\sqrt{n_{i}}}\right)\right),$ (46)
+$\mathbb{E}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}\leq C_{1}\frac{1}{N}+C_{2}(\sum_{i\in[m]}\frac{p_{i}}{n_{i}}\wedge R^{2}),$ (56)
 where
-$\begin{split}C_{1}&=\frac{\rho^{2}}{\mu^{2}},\\ C_{2}&=\left(\frac{(1+\sqrt{1+2\mu})\rho}{2\mu}\right)^{2}\vee\frac{2\rho}{\mu}.\end{split}$ (47)
+$\begin{split}C_{1}&=32\frac{\rho^{2}}{\mu^{2}},\\ C_{2}&=\left(\frac{(1+\sqrt{3})\rho}{\mu}\right)^{2}\vee\frac{32L^{2}}{\mu^{2}}.\end{split}$ (57)
 Moreover, if we assume data balanced, i.e $n_{i}=n,\forall i\in[m]$ and $p_{i}=\frac{1}{m}$, we obtain
-$\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|^{2}\leq C_{1}\frac{1}{N}+C_{2}\left(\frac{1}{n}\wedge\frac{R}{\sqrt{n}}\right).$ (48)
-A.3 Proof of local statistical Error Bound
-Proof.
-Firstly, we will provide some basic results for the local optimality.
-To prove the local statistical error bound, we start with the optimality condition of ${\tilde{\bm{w}}^{(i)}}$ and ${\tilde{\bm{w}}^{(g)}}$ for a single client, which yields
+$\mathbb{E}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}\leq C_{1}\frac{1}{N}+C_{2}\left(\frac{1}{n}\wedge R^{2}\right),$ (58)
+as claimed.
+A.3.3 Proof of Local Statistical Error Bound in Theorem 1
+Case 1: first we consider the case when $R>\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$.
+To prove the local statistical error bound, we start with the optimality condition of ${\tilde{\bm{w}}^{(i)}}$ and ${\tilde{\bm{w}}^{(g)}}$ for a single client, which yields that for $i\in[m]$,
 00 0 $\geq L_{i}\left({\tilde{\bm{w}}^{(i)}},S_{i}\right)+\frac{\lambda}{2}\left\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\right\|^{2}-L_{i}\left({{\bm{w}}_{\star}^{(i)}},S_{i}\right)-\frac{\lambda}{2}\left\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\right\|^{2}.$ (59)
@@
 00 0 $\geq\sum_{j\in[n_{i}]}\frac{1}{n_{i}}\left(\ell\left({\tilde{\bm{w}}^{(i)}},z_{ij}\right)-\ell\left({{\bm{w}}_{\star}^{(i)}},z_{ij}\right)\right)+\frac{\lambda}{2}\left\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\right\|^{2}-\frac{\lambda}{2}\left\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\right\|^{2}.$ (60)
-For the first term on the RHS, apply the $\mu$-strongly convex of the loss function $\ell$ (Assumption 2)
-$\begin{split}\sum_{j\in[n_{i}]}\frac{1}{n_{i}}\left(\ell({\tilde{\bm{w}}^{(i)}},z_{ij})-\ell({{\bm{w}}_{\star}^{(i)}},z_{ij})\right)&\geq-\sum_{j\in[n_{i}]}\frac{1}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle+\mu\sum_{j\in[n_{i}]}\frac{1}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2},\end{split}$ (51)
+For the first term on the RHS, applying the $\mu$-strongly convex of the loss function $\ell$ (cf. Assumption 3) yields
+$\begin{split}\sum_{j\in[n_{i}]}\frac{1}{n_{i}}\left(\ell({\tilde{\bm{w}}^{(i)}},z_{ij})-\ell({{\bm{w}}_{\star}^{(i)}},z_{ij})\right)&\geq-\sum_{j\in[n_{i}]}\frac{1}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle+\frac{\mu}{2}\sum_{j\in[n_{i}]}\frac{1}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2},\end{split}$ (61)
 where we use the definition of $\tilde{\Delta}^{(i)}_{\text{stat}}$ in (16).
-Plugging (51) back into (50), it yields
-$\begin{split}0&\geq-\sum_{j}\frac{1}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle+\mu\sum_{j}\frac{1}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\|^{2}-\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}.\end{split}$ (52)
+Pluggging (61) back into (60), it yields
+$\begin{split}0&\geq-\sum_{j}\frac{1}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle+\frac{\mu}{2}\sum_{j}\frac{1}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\|^{2}-\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}.\end{split}$ (62)
 Apply Cauchy inequality on the first term of (62)
@@
 Combining (63) and (64), it yields
-$\begin{split}\sum_{j}\frac{1}{n_{i}}\mathbb{E}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle\leq\rho\frac{1}{\sqrt{n_{i}}}\|\tilde{\Delta}^{(i)}_{\text{stat}}\|_{2}.\\ \end{split}$ (55)
+$\begin{split}\sum_{j}\frac{1}{n_{i}}\mathbb{E}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle\leq\rho\frac{1}{\sqrt{n_{i}}}\mathbb{E}\|\tilde{\Delta}^{(i)}_{\text{stat}}\|_{2}.\\ \end{split}$ (65)
 Plugging (65) back to (62) yields
-$0\geq-\rho\frac{1}{\sqrt{n_{i}}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\mu\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\|^{2}-\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}.$ (56)
-To utilize the statistical global convergence rate shown in Appendix A.2, we need to argue $R$ in the same way.
-Recalling the upper bound of $\tilde{\Delta}^{(g)}_{\text{stat}}$ if data is balanced (48)
-$\Rightarrow\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}\leq\begin{cases}\frac{C_{2}}{n}&R>\frac{1}{\sqrt{n}}\\ \frac{C_{1}}{N}+\frac{C_{2}R}{{n}^{\frac{1}{2}}}&R\leq\frac{1}{\sqrt{n}}.\end{cases}$ (57)
-For simplicity, we denote the upper bound of $\|\tilde{\Delta}^{(g)}_{\text{stat}}\|_{2}$ as $U_{0}$
-Secondly, we need to discuss the upper bound of $\tilde{\Delta}^{(i)}_{\text{stat}}$ with two cases below.
-Case 1: For the first case, we consider the statistical heterogeneity is relative, i.e., $R>\frac{1}{\sqrt{n_{i}}}$.
+$0\geq-\rho\frac{1}{\sqrt{n_{i}}}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\frac{\mu}{2}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\mathbb{E}\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\|^{2}-\frac{\lambda}{2}\mathbb{E}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}.$ (66)
+To utilize the statistical global convergence rate shown in Appendix A.3.2, we need to discuss $R$ in the same way. Recall the upper bound of $\tilde{\Delta}^{(g)}_{\text{stat}}$ in (58):
+$\mathbb{E}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}\leq\begin{cases}\frac{C_{2}}{n}&R>\frac{1}{\sqrt{n}}\\ \frac{C_{1}}{N}+C_{2}R^{2}&R\leq\frac{1}{\sqrt{n}}.\end{cases}$ (67)
+For simplicity, in the following arguments, we denote the upper bound of $\mathbb{E}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|_{2}$ as $U_{0}$.
 Note that
-$\begin{split}-\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}&=-\left\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}+{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\right\|^{2}\\ &=-\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}-\|{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}-2\|{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|_{2}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2}\\ &\geq-\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}-R^{2}-2R\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2}.\end{split}$ (58)
+$\begin{split}-\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}&=-\left\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}+{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\right\|^{2}\\ &\geq-\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}-\|{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}-2\|{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|_{2}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2}\\ &\geq-\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}-R^{2}-2R\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2}.\end{split}$ (68)
 Plug (68) into (66)
-$\begin{split}0\geq&-\frac{\rho}{\sqrt{n_{i}}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\mu\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\left\|\tilde{\bm{\delta}}^{(i)}\right\|^{2}-\frac{\lambda}{2}\left\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}+{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\right\|^{2}\\ =&-\frac{\rho}{\sqrt{n_{i}}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\mu\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\left\|\tilde{\bm{\delta}}^{(i)}\right\|^{2}-\frac{\lambda}{2}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|^{2}-\frac{\lambda}{2}R^{2}-\lambda R\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2}.\\ \end{split}$ (59)
-Put (57) into (59) and reorganize
-$\begin{split}0\geq&-\frac{\rho}{\sqrt{n_{i}}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\mu\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\left\|\tilde{\bm{\delta}}^{(i)}\right\|^{2}-\frac{\lambda}{2}U_{0}^{2}-\frac{\lambda}{2}R^{2}-\lambda R\;U_{0}\\ \geq&\mu\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}-\frac{\rho}{\sqrt{n_{i}}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}-\frac{\lambda}{2}\left(U_{0}+R\right)^{2}.\end{split}$ (60)
-In addition, recall that in this case we set $\lambda^{\star}=\frac{\rho^{2}}{R^{2}}\sum_{i\in[m]}\frac{p_{i}}{n_{i}}$ in the statistical convergence analysis of the global model.
+$\begin{split}0\geq&-\frac{\rho}{\sqrt{n_{i}}}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\frac{\mu}{2}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\mathbb{E}\left\|\tilde{\bm{\delta}}^{(i)}\right\|^{2}-\frac{\lambda}{2}\mathbb{E}\left\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}+{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\right\|^{2}\\ \geq&-\frac{\rho}{\sqrt{n_{i}}}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\frac{\mu}{2}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\mathbb{E}\left\|\tilde{\bm{\delta}}^{(i)}\right\|^{2}-\frac{\lambda}{2}\mathbb{E}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|^{2}-\frac{\lambda}{2}R^{2}-\lambda R\mathbb{E}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2}.\\ \end{split}$ (69)
+Putting (67) into (69) and reorganizing, we obtain
+$\begin{split}0\geq&-\frac{\rho}{\sqrt{n_{i}}}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\frac{\mu}{2}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\mathbb{E}\left\|\tilde{\bm{\delta}}^{(i)}\right\|^{2}-\frac{\lambda}{2}U_{0}^{2}-\frac{\lambda}{2}R^{2}-\lambda R\;U_{0}\\ \geq&\frac{\mu}{2}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}-\frac{\rho}{\sqrt{n_{i}}}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}-\frac{\lambda}{2}\left(U_{0}+R\right)^{2}.\end{split}$ (70)
+In addition, recall that in this case we set $\lambda^{\star}\leq\frac{\rho^{2}}{\mu R^{2}}\sum_{i\in[m]}\frac{p_{i}}{n_{i}}$ in the statistical convergence analysis of the global model.
 If we assume data balanced, i.e.
-$p_{i}=\frac{1}{m}$ and $n_{i}=n,\forall i\in[m]$, it yields $\lambda^{\star}=\frac{1}{nR^{2}}$
-$\begin{split}\|\tilde{\Delta}^{(i)}_{\text{stat}}\|_{2}&\leq\frac{1}{2\mu}\left[\frac{\rho}{\sqrt{n}}+\left(\frac{\rho^{2}}{n}+\frac{2\mu\rho^{2}}{nR^{2}}\left(U_{0}+R\right)^{2}\right)^{\frac{1}{2}}\right].\\ &=\frac{1}{\sqrt{n}}\frac{1}{2\mu}\left(\rho+\left(\rho^{2}+\frac{2\mu\rho^{2}}{R^{2}}(U_{0}+R)^{2}\right)^{\frac{1}{2}}\right)\\ &\leq\frac{1}{\sqrt{n}}\frac{\rho+(\rho^{2}+4\mu\rho^{2}(C_{2}+1))^{\frac{1}{2}}}{2\mu}\\ &\leq\frac{1}{\sqrt{n}}\frac{\rho(1+(1+4\mu(C_{2}+1))^{\frac{1}{2}})}{2\mu}.\end{split}$ (61)
-Case 2: On the other hand, we will focus on the case that $R$ is relatively small, which means $R\leq\sqrt{\frac{m}{N}}$. From (56), note the definition of $\bm{\delta}_{\star}^{(i)}$ in (20)
-$\begin{split}0&\geq-\frac{\rho}{\sqrt{n_{i}}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\mu\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\left\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}+\delta_{\star}^{(i)}+{{\bm{w}}_{\star}^{(i)}}-{\tilde{\bm{w}}^{(i)}}\right\|^{2}\\ &-\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}.\end{split}$ (62)
-Note
-$2\left\|-\tilde{\Delta}^{(g)}_{\text{stat}}+\tilde{\Delta}^{(i)}_{\text{stat}}+\delta_{\star}^{(i)}\right\|^{2}+2\left\|\tilde{\Delta}^{(g)}_{\text{stat}}-\delta_{\star}^{(i)}\right\|^{2}\geq\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}.$ (63)
-Thus
-$\left\|-\tilde{\Delta}^{(g)}_{\text{stat}}+\tilde{\Delta}^{(i)}_{\text{stat}}+\delta_{\star}^{(i)}\right\|^{2}\geq\frac{1}{2}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}-\left\|\tilde{\Delta}^{(g)}_{\text{stat}}-\delta_{\star}^{(i)}\right\|^{2}.$ (64)
-Furthermore, note
-$\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}+\|{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}=\|\tilde{\Delta}^{(i)}_{\text{stat}}\|^{2}+\|\bm{\delta}_{\star}^{(i)}\|^{2}\geq\frac{1}{2}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}.$ (65)
-Plugging (64) and (66) into (62), it yields
-$0\geq-\frac{\rho}{\sqrt{n_{i}}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\mu\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{4}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}-\frac{\lambda}{2}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}-\delta_{\star}^{(i)}\right\|^{2}-\lambda\|\tilde{\Delta}^{(i)}_{\text{stat}}\|^{2}-\lambda\|\bm{\delta}_{\star}^{(i)}\|^{2},$ (66)
-which implies
-$0\geq-\frac{\rho}{\sqrt{n_{i}}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\mu\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{4}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}-2\lambda\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|^{2}-2\lambda\|\bm{\delta}_{\star}^{(i)}\|^{2}.$ (67)
-Apply Assumption 1 and (57) to (67)
-$0\geq\left(\mu+\frac{\lambda}{4}\right)\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}-\frac{\rho}{\sqrt{n_{i}}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}-2\lambda U_{0}^{2}-2\lambda R^{2}.$ (68)
-Solving (68) as a function of $\|\tilde{\Delta}^{(i)}_{\text{stat}}\|_{2}$, it yields
-$\begin{split}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}&\leq\frac{1}{2\mu+\lambda/2}\left\{\frac{\rho}{\sqrt{n_{i}}}+\left[\frac{\rho^{2}}{n_{i}}+4\left(\mu+\frac{\lambda}{4}\right)\lambda\left(U_{0}^{2}+R^{2}\right)\right]^{\frac{1}{2}}\right\}.\end{split}$ (69)
-If we assume data size is same among different clients, i.e.
-$n=n_{i},\forall i\in[m]$ and choose $\lambda=\frac{\rho}{R}\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}=\frac{\rho}{R}\frac{1}{\sqrt{n}}$, (69) can be reformulated as
-$\begin{split}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}&\leq\frac{2}{4\mu+\frac{\rho}{\sqrt{n}R}}\left\{\frac{\rho}{\sqrt{n}}+\left[\frac{\rho^{2}}{n}+\left(4\mu+\frac{\rho}{\sqrt{n}R}\right)\frac{\rho}{\sqrt{n}R}\left(U_{0}^{2}+R^{2}\right)\right]^{\frac{1}{2}}\right\}.\end{split}$ (70)
-Therefore, we have
-$\|\tilde{\Delta}^{(i)}_{\text{stat}}\|^{2}\leq\frac{16\rho^{2}}{\left((4\mu+\frac{\rho}{R\sqrt{n}})\sqrt{n}\right)^{2}}+\frac{8}{4\mu+\frac{\rho}{R\sqrt{n}}}\frac{\rho}{R\sqrt{n}}\left(\frac{C_{1}}{N}+\frac{C_{2}R}{\sqrt{n}}+R^{2}\right).$ (71)
-If $\mu<\frac{\rho}{R\sqrt{n}}$
-$\frac{\rho}{(4\mu+\frac{\rho}{R\sqrt{n}})\sqrt{n}}\leq\frac{\rho}{\frac{\rho}{R\sqrt{n}}\sqrt{n}}\leq R.$ (72)
-If $\mu\geq\frac{\rho}{R\sqrt{n}}$
-$\frac{\rho}{(4\mu+\frac{\rho}{R\sqrt{n}})\sqrt{n}}\leq\frac{\rho}{\mu\sqrt{n}}.$ (73)
-Combining the discussion in (72) and (73), it yields
-$\frac{16\rho^{2}}{\left((4\mu+\frac{\rho}{R\sqrt{n}})\sqrt{n}\right)^{2}}\leq 16\min{\{R^{2},\frac{\rho^{2}}{\mu^{2}n}\}}.$ (74)
-Then for the second term on the RHS of (71)
-$\begin{split}\frac{8}{4\mu+\frac{\rho}{R\sqrt{n}}}\frac{\rho}{R\sqrt{n}}\left(\frac{C_{1}}{N}+\frac{C_{2}R}{\sqrt{n}}+R^{2}\right)\leq\frac{8C_{1}}{N}+\frac{8C_{2}R}{\sqrt{n}}+\frac{8\rho R}{4\mu\sqrt{n}}.\end{split}$ (75)
-Plugging (74) and (75) into (71), we obtain
-$\begin{split}\|\tilde{\Delta}^{(i)}_{\text{stat}}\|^{2}&\leq 16\min{\{R^{2},\frac{\rho^{2}}{\mu^{2}n}\}}+\frac{8C_{1}}{N}+\frac{8C_{2}R}{\sqrt{n}}+\frac{8\rho R}{4\mu\sqrt{n}}\\ &\leq 16\min{\{R^{2},\frac{\rho^{2}}{\mu^{2}n}\}}+\frac{8C_{1}}{N}+(8C_{2}+\frac{2\rho}{\mu})\frac{R}{\sqrt{n}}\end{split}$ (76)
-Combining two-case arguments in (61) and (76), we have:
-$\Rightarrow\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\leq\begin{cases}\frac{1}{n}\left(\frac{\rho(1+(1+4\mu(C_{2}+1))^{\frac{1}{2}})}{2\mu}\right)^{2}&R>\sqrt{\frac{m}{N}}\\ 16(R^{2}\wedge\frac{\rho^{2}}{\mu^{2}n})+\frac{8C_{1}}{N}+(8C_{2}+\frac{2\rho}{\mu})\frac{R}{\sqrt{n}}&R\leq\sqrt{\frac{m}{N}}\end{cases}.$ (77)
+$p_{i}=\frac{1}{m}$ and $n_{i}=n,\forall i\in[m]$, it yields $\lambda^{\star}\leq\frac{\rho^{2}}{\mu nR^{2}}$ and
+$\begin{split}\mathbb{E}\|\tilde{\Delta}^{(i)}_{\text{stat}}\|_{2}&\leq\frac{1}{\mu}\left[\frac{\rho}{\sqrt{n}}+\left(\frac{\rho^{2}}{n}+\frac{2\rho^{2}}{nR^{2}}\left(U_{0}+R\right)^{2}\right)^{\frac{1}{2}}\right].\\ &=\frac{1}{\sqrt{n}}\frac{1}{\mu}\left(\rho+\left(\rho^{2}+\frac{2\rho^{2}}{R^{2}}(U_{0}+R)^{2}\right)^{\frac{1}{2}}\right)\\ &\leq\frac{1}{\sqrt{n}}\frac{\rho+(\rho^{2}+4\rho^{2}(C_{2}+1))^{\frac{1}{2}}}{\mu}\\ &=\frac{1}{\sqrt{n}}\frac{\rho(1+(1+4(C_{2}+1))^{\frac{1}{2}})}{\mu}.\end{split}$ (71)
+Case 2: we then consider the case when $R\leq\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$.
+$\mathbb{E}[\|{\tilde{\bm{w}}^{(i)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}]$
+$\leq 3\mathbb{E}[\|{\tilde{\bm{w}}^{(i)}}-{\tilde{\bm{w}}^{(g)}}\|^{2}]+3\mathbb{E}[\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}]+3\mathbb{E}[\|{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}]$
+$\overset{(a)}{\leq}\frac{12}{(\mu+\lambda)^{2}}\mathbb{E}[\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}})\|^{2}]+3\mathbb{E}[\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}]+3R^{2}$
+$\overset{(b)}{\leq}\frac{12}{(\mu+\lambda)^{2}}\left[4L^{2}\mathbb{E}[\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}]+4L^{2}R^{2}+2\frac{\rho^{2}}{n}\right]+3\mathbb{E}[\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}]+3R^{2}$
+$\overset{(c)}{\leq}\frac{12}{(\mu+\lambda)^{2}}\left[128L^{2}\left(\frac{\rho^{2}}{\mu^{2}}\frac{1}{N}+\frac{L^{2}}{\mu^{2}}R^{2}\right)+4L^{2}R^{2}+2\frac{\rho^{2}}{n}\right]+96\left(\frac{\rho^{2}}{\mu^{2}}\frac{1}{N}+\frac{L^{2}}{\mu^{2}}R^{2}\right)+3R^{2}$
+$=\frac{12}{(\mu+\lambda)^{2}}\left[\frac{128L^{2}\rho^{2}}{\mu^{2}}\frac{1}{N}+(128L^{4}/\mu^{2}+4L^{2})R^{2}+2\frac{\rho^{2}}{n}\right]+\frac{96\rho^{2}}{\mu^{2}}\frac{1}{N}+(96\frac{L^{2}}{\mu^{2}}+3)R^{2}$ (72)
+where step (a) comes from Assumption 1 and Lemma 4, step (b) comes from result (A.3.2) and step (c) is a direct result of applying the error bound of $\mathbb{E}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}$.
+If we assume $\lambda$ satisfies the following condition
+$\frac{12}{(\mu+\lambda)^{2}}\leq\frac{96\rho^{2}/(\mu^{2}N)+99\kappa^{2}R^{2}}{128\rho^{2}/N+132L^{2}R^{2}+2\rho^{2}/n}$ (73)
+we have
+$\mathbb{E}[\|{\tilde{\bm{w}}^{(i)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}]\leq\frac{192\rho^{2}}{\mu^{2}}\frac{1}{N}+198\kappa^{2}R^{2}$ (74)
+Furthermore, we can get a simplified condition for $\lambda$ as
+$\lambda\geq\sqrt{\frac{8(64\rho^{2}/N+66L^{2}R^{2}+\rho^{2}/n)}{32\rho^{2}/(\mu^{2}N)+33\kappa^{2}R^{2}}}-\mu$ (75)
+Combining two-case arguments, we have:
+$\Rightarrow\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\leq\begin{cases}\frac{1}{n}\left(\frac{\rho(1+(1+4(C_{2}+1))^{\frac{1}{2}})}{\mu}\right)^{2}&R>\sqrt{\frac{m}{N}}\\ \frac{192\rho^{2}}{\mu^{2}}\frac{1}{N}+198\kappa^{2}R^{2}&R\leq\sqrt{\frac{m}{N}}\end{cases}.$ (76)
 And it yields a one-line rate
-$\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\leq\frac{C_{3}}{N}+16(R^{2}\wedge\frac{\rho^{2}}{\mu^{2}n})+C_{4}(\frac{R}{\sqrt{n}}\wedge\frac{1}{n}),$ (78)
+$\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\leq C_{3}\frac{1}{N}+C_{4}(R^{2}\wedge\frac{1}{n}),$ (77)
 where
-$C_{3}$ $=8C_{1},$ (79)
-$C_{4}$ $=\left(\frac{\rho(1+(1+4\mu(C_{2}+1))^{\frac{1}{2}})}{2\mu}\right)^{2}\wedge(8C_{2}+\frac{2\rho}{\mu}),$ (80)
-and $C_{1},C_{2}$ are specified in (47). ∎
+$C_{3}$ $=\frac{192\rho^{2}}{\mu^{2}},$ (78)
+$C_{4}$ $=\left(\frac{\rho(1+(4C_{2}+5)^{\frac{1}{2}})}{\mu}\right)^{2}\vee 198\kappa^{2},$ (79)
+and $C_{2}$ are specified in (57).
+Combining the results in A.3.2 and A.3.3, we leave a remark below to further interpret the effect of personalization in obj (1) and discuss how to incorporate the conditions of $\lambda$ when establishing the global and local statistical accuracy.
+Remark 2.
+To better understand how such an interpolation is achieved through an adaptive personalization degree, we examine the role of $\lambda$ in navigating between the two extreme cases: GlobalTrain and LocalTrain.
+As the statistical heterogeneity $R\to\infty$, we have $\lambda\to 0$, leading to a high degree of personalization.
+The statistical error bound becomes $\mathcal{O}(n^{-1})$, matching the rate of LocalTrain; as the statistical heterogeneity $R\to 0$, $\lambda$ increases, leading to a low degree of personalization.
+The statistical error bound eventually converges to $\mathcal{O}(N^{-1})$, matching the rate of GlobalTrain under a homogeneous setting.
+As $R\in(0,\infty)$, $\lambda$ transits between the two extremes, making FedProx perform no worse than LocalTrain while benefiting from global training when clients share similarities.
+To make the global statistical error bound and local statistical error bound hold simultaneously, we should be aware about the condition of the adaptive strategy of $\lambda$ specified in our proof.
+It could be summarized as
+$\begin{cases}\lambda\geq max(a_{1},a_{2})&R<\sqrt{\frac{m}{N}}\\ \lambda\leq\frac{\rho^{2}}{n\mu R^{2}}&R\geq\sqrt{\frac{m}{N}}\end{cases},$ (81)
+where
+$a_{1}=\frac{8\kappa}{\mu}\left\{8L^{2}\vee\left(\frac{2L^{2}R^{2}+\rho^{2}/n}{4\rho^{2}/(\mu^{2}N)+4\kappa^{2}R^{2}}\right)\right\}-\mu,$
+and
+$a_{2}=\sqrt{\frac{8(64\rho^{2}/N+66L^{2}R^{2}+\rho^{2}/n)}{32\rho^{2}/(\mu^{2}N)+33\kappa^{2}R^{2}}}-\mu.$
+We can induce a sufficient condition of $\lambda$ for $R<\frac{1}{\sqrt{n}}$ as
+$\lambda\geq\text{max}\left\{64\kappa^{2}L,\left(2\kappa\vee 5\right)\mu\frac{2L^{2}R^{2}+\rho^{2}/n}{L^{2}R^{2}+\rho^{2}/N}\right\}-\mu.$ (82)
 Appendix B Optimization Convergence Rate
-B.1 Proof of lemmas
-B.1.1 Proof of Lemma 2
+B.1 Algorithm
+Input: Initial global model ${\bm{w}}_{1}^{(g)}$, initial local models $\{{\bm{w}}_{0,K}^{(i)}\}_{i\in[m]}$, global rounds $T$, global step sizes $\gamma$, local rounds $K$, local step sizes $\eta$
+Output: Local models $\{{\bm{w}}_{T,K}^{(i)}\}_{i\in[m]}$ and global model ${\bm{w}}_{T}^{(g)}$
+for $t=1,\dots,T$ do
+       The server sends ${\bm{w}}_{t}^{(g)}$ to client $i$, $\forall i\in[m]$
+       Set ${\bm{w}}_{t,0}^{(i)}={\bm{w}}_{t-1,K}^{(i)}$;
+       for $k=0,\dots,K-1$ do
+             ${\bm{w}}_{t,k+1}^{(i)}={\bm{w}}_{t,k}^{(i)}-\frac{\eta}{n_{i}}\sum_{j\in[n_{i}]}\left\{\nabla\ell({\bm{w}}_{t,k}^{(i)},z_{i,j})+\lambda\left({\bm{w}}_{t,k}^{(i)}-\bm{w}_{t}^{(g)}\right)\right\}$
+      Push ^$\widehat{\nabla}F_{i}(\bm{w}_{t}^{(g)})=\lambda(\bm{w}_{t}^{(g)}-{\bm{w}}_{t,K}^{(i)})$ to the server;
+      $\bm{w}_{t+1}^{(g)}\leftarrow\bm{w}_{t}^{(g)}-\frac{\gamma}{m}\sum_{i\in[m]}\widehat{\nabla}F_{i}(\bm{w}_{t}^{(g)})$;
+Algorithm 1 One-stage FedProx
+B.2 Proof of lemmas
+B.2.1 Proof of Lemma 2
 To facilitate our analysis, we state a claim first.
@@
 Now, we are ready to prove Lemma 2.
-Proof.
 First, from mishchenko2023convergence, we can directly obtain the analytic properties of $F_{i}$.
-Next, the analytic properties of $h_{i}$ are direct results from Assumption 2 and 4.
+Next, the analytic properties of $h_{i}$ are direct results from Assumption 3 and 2.
 Finally, to analyze the property of the mapping $w_{\star}^{(i)}$, we just need to use the result from the above claim as we know the definition of ${\bm{w}}_{\star}^{(i)}(\cdot)$ in (9).
-B.1.2 Inner Loop Convergence
+B.2.2 Inner Loop Convergence
 Lemma 6 ((Ji et al., 2022), Inner Loop Error Contraction).
-Suppose Assumption 2 and 4 hold, then if $\eta\leq 1/L_{\ell}$, for all $t\geq 0$ and $\epsilon>0$, the inner loop error bound of all clients can be formulated as:
+Suppose Assumption 3 and 2 hold, then if $\eta\leq 1/L_{\ell}$, for all $t\geq 0$ and $\epsilon>0$, the inner loop error bound of all clients can be formulated as:
 $e_{t+1}^{(i)}:=\|{\bm{w}}_{t+1,K}^{(i)}-{\bm{w}}_{t+1,\star}^{(i)}\|^{2}\leq(1+\epsilon)(1-\eta\mu_{\ell})^{K}e_{t}^{(i)}+\left(1+\frac{1}{\epsilon}\right)(1-\eta\mu_{\ell})^{K}\|\bm{w}^{(i)}_{t,\star}-\bm{w}^{(i)}_{t-1,\star}\|^{2}.$ (89)
 Proof.
-The proof can be found in Lemma 2 (Ji et al., 2022).
-B.1.3 Outer Loop convergence
+The proof can be found in Lemma 2 (Ji et al., 2022). ∎
+B.2.3 Outer Loop convergence
 Lemma 7 (Outer Loop Error Contraction).
-Suppose Assumption 2 and 4 hold, then if $\gamma\leq L_{g}$, we have for all $t>0$ and $\epsilon_{g}>0$:
+Suppose Assumption 3 and 2 hold, then if $\gamma\leq L_{g}$, we have for all $t>0$ and $\epsilon_{g}>0$:
 $\begin{split}\|{\bm{w}}_{t+1}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}&\leq(1-\gamma\mu_{g}+\gamma\epsilon_{g})\|{\bm{w}}_{t}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}-\gamma^{2}\|\nabla F({\bm{w}}_{t}^{(g)})\|^{2}\\ &+(\gamma^{3}L_{g})\|\widehat{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}+\left(\frac{\gamma}{\epsilon_{g}}+\gamma^{2}\right)\|\nabla F({\bm{w}}_{t}^{(g)})-\widehat{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}.\end{split}$ (90)
@@
 Substituting (94) and (95) into (93) leads to
-$\begin{split}&\|{\bm{w}}_{t+1}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}\\ \leq&(1+\epsilon_{g}\gamma)\|{\bm{w}}_{t}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}+\gamma^{2}\|\widetilde{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}+\left(\frac{\gamma}{\epsilon_{g}}\right)\|\widetilde{\nabla}F({\bm{w}}_{t}^{(g)})-\nabla F({\bm{w}}_{t}^{(g)})\|^{2}\\ &-2\gamma(F({\bm{w}}_{t+1}^{(g)})-\widehat{F})-\gamma\mu_{g}\|{\bm{w}}_{t}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}\\ &-\gamma^{2}\|\nabla F({\bm{w}}_{t}^{(g)})\|^{2}-\gamma\left(\gamma-L_{g}\gamma^{2}\right)\|\widetilde{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}+\gamma^{2}\|\nabla F({\bm{w}}_{t}^{(g)})-\widetilde{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}\\ \leq&(1-\gamma\mu_{g}+\gamma\epsilon_{g})\|{\bm{w}}_{t}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}-\gamma^{2}\|\nabla F({\bm{w}}_{t}^{(g)})\|^{2}+(\gamma^{3}L_{g})\|\widetilde{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}\\ &+\left(\frac{\gamma}{\epsilon_{g}}+\gamma^{2}\right)\|\nabla F({\bm{w}}_{t}^{(g)})-\widetilde{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}.\end{split}$ (95)
-B.2 Proof of Theorem 2
+$\begin{split}&\|{\bm{w}}_{t+1}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}\\ \leq&(1+\epsilon_{g}\gamma)\|{\bm{w}}_{t}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}+\gamma^{2}\|\widehat{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}+\left(\frac{\gamma}{\epsilon_{g}}\right)\|\widehat{\nabla}F({\bm{w}}_{t}^{(g)})-\nabla F({\bm{w}}_{t}^{(g)})\|^{2}\\ &-2\gamma(F({\bm{w}}_{t+1}^{(g)})-\widehat{F})-\gamma\mu_{g}\|{\bm{w}}_{t}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}\\ &-\gamma^{2}\|\nabla F({\bm{w}}_{t}^{(g)})\|^{2}-\gamma\left(\gamma-L_{g}\gamma^{2}\right)\|\widehat{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}+\gamma^{2}\|\nabla F({\bm{w}}_{t}^{(g)})-\widehat{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}\\ \leq&(1-\gamma\mu_{g}+\gamma\epsilon_{g})\|{\bm{w}}_{t}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}-\gamma^{2}\|\nabla F({\bm{w}}_{t}^{(g)})\|^{2}+(\gamma^{3}L_{g})\|\widehat{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}\\ &+\left(\frac{\gamma}{\epsilon_{g}}+\gamma^{2}\right)\|\nabla F({\bm{w}}_{t}^{(g)})-\widehat{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}.\end{split}$ (96)
+B.3 Proof of Theorem 2
 Recall the definition of $e_{t}^{(i)}$ in Lemma 6, we define
 $e_{t}=\frac{1}{m}\sum_{i\in[m]}e_{t}^{(i)}.$ (97)
-Recall the definition $\nabla F({\bm{w}}_{t}^{(g)}),\widehat{\nabla}F({\bm{w}}_{t}^{(g)})$ in Lemma and eq (91), the gradient approximation error is
+Recall the definition $\nabla F({\bm{w}}_{t}^{(g)}),\widehat{\nabla}F({\bm{w}}_{t}^{(g)})$ in Lemma 3 and eq (92), the gradient approximation error is
 $\begin{split}\|\nabla F({\bm{w}}_{t}^{(g)})-\widehat{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}=\lambda^{2}\Big{\|}\frac{1}{m}\sum_{i=1}^{m}\big{(}{\bm{w}}_{t,K}^{(i)}-{\bm{w}}_{t,\star}^{(i)}\big{)}\Big{\|}^{2}\leq\lambda^{2}e_{t}.\\ \end{split}$ (98)
@@
 $\begin{split}&\|{\bm{w}}_{t+1}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}\\ \overset{(a)}{\leq}&(1-\gamma\mu_{g}+\gamma\epsilon_{g})\|{\bm{w}}_{t}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}-\left(\gamma^{2}-\gamma^{3}L_{g}(1+\zeta)\right)\|\nabla F({\bm{w}}_{t}^{(g)})\|^{2}\\ &+\left(\frac{\gamma}{\epsilon_{g}}+\gamma^{2}+L_{g}\gamma^{3}(1+\zeta^{-1})\right)\|\nabla F({\bm{w}}_{t}^{(g)})-\widehat{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}\\ \overset{(b)}{\leq}&(1-\gamma\mu_{g}+\gamma\epsilon_{g})\|{\bm{w}}_{t}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}-\gamma^{2}\Big{(}1-\gamma L_{g}(1+\zeta)\Big{)}\|\nabla F({\bm{w}}_{t}^{(g)})\|^{2}\\ &+\left(\frac{\gamma}{\epsilon_{g}}+\gamma^{2}(2+\zeta^{-1})\right)\lambda^{2}e_{t},\end{split}$ (99)
-where step (a) comes from Young’s inequality with $\zeta>0$ to be chosen and step (b) follows from the step size condition $\gamma\leq 1/L_{g}$.
+where the step (a) comes from the Young’s inequality with $\zeta>0$ to be chosen and the step (b) follows from the step size condition $\gamma\leq 1/L_{g}$.
 Invoking Lemma 2 and Lemma 6, the inner loop optimization error is updated as
@@
 $\gamma<1/L_{g},\quad\Big{(}2+64L_{w}^{2}(1/\mu_{g})^{2}\lambda^{2}\Big{)}(1-\eta\mu_{\ell})^{K}\leq(1-\gamma L_{g})^{4}.$ (114)
-B.3 Proof of Corollary 1
+B.4 Proof of Corollary 1
 With $\eta=1/L_{\ell}=(\lambda+L)^{-1}$, $\gamma=1/(2L_{g})=\frac{\lambda+L}{2\lambda L}$, and using the fact that $L_{w}=\frac{\lambda}{\lambda+\mu}$ and $\mu_{g}=\frac{\lambda\mu}{\lambda+\mu}$ (cf.
@@
 $K\geq 2+\frac{\lambda+L}{\lambda+\mu}\cdot\log(1056\kappa^{2}).$ (118)
-B.4 Proof of The Local Model Convergence Rate in Theorem 2
+B.5 Proof of The Local Model Convergence Rate in Theorem 2
 Proof.
@@
 In this section, we provide more details on the implementation of the methods used in the empirical analysis.
+Real Dataset.
+The MNIST dataset consists of 70,000 grayscale images of handwritten digits, each 28x28 pixels, classified into 10 classes (digits 0-9).
+The EMNIST dataset extends MNIST, including handwritten letters and digits.
+The Balanced split contains 131,600 28x28 grayscale images across 47 balanced classes (digits and uppercase/lowercase letters).
+And the CIFAR-10 dataset contains 60,000 color images (32x32 pixels, 3 RGB channels), categorized into 10 classes representing real-world objects (e.g., airplane, car, and dog).
+Algorithms.
+To fully investigate the effect of personalization in FedProx with the Algorithm B.1, we compare the two extreme algorithms (GlobalTraining and LocalTraining).
+Furthermore, since there exist other methods solving the objective (1) to do personalized federated learning, we also try to investigate the effect of personalization in other methods, like pFedMe (T Dinh et al., 2020).
+The algorithm pFedMe has a similar design with a double loop to solve the subproblem in each communication round.
 Selection of $\lambda$.
@@
 Such a choice is to make sure that each line in the figure is clearly separable.
-Additionally, we implemented an adaptive choice of $\lambda=\frac{\rho}{\sqrt{n}R}\mathbb{I}\{R\leq n^{-1/2}\}+\frac{\rho^{2}}{nR^{2}}\mathbb{I}\{R>n^{-1/2}\}$ following the formula outlined in Theorem 1, with $\rho$ set to 2.
+Additionally, we implemented an adaptive choice of $\lambda$ following the formula outlined in Theorem 1, with $\rho$ set to 2.
 The value of $\rho$ was determined through grid search over the range $[1,10]$ to ensure optimal performance.
+For other real datasets, the regularization term $\lambda$ is set as small $\lambda=0.1$, medium $\lambda=$ 0.5, and large $\lambda=$ 1.
 Selection of Step Size.
@@
 The local step size for the synthetic dataset was further tuned through grid search in $\{0.001,0.01,0.1\}$.
-For the MNIST dataset, the global learning rate was set to $\frac{1}{\lambda}$, while the local learning rate was set to 0.01, chosen via grid search.
+For the real dataset, the global learning rate was set to $\frac{1}{\lambda}$, while the local learning rate was set to 0.01, chosen via grid search.
 For both GlobalTrain and LocalTrain training baselines, the learning rate was also chosen via grid search within the range $\{0.001,0.01,0.1\}$.
 LocalTrain is implemented using a local (S)GD, and GlobalTrain is implemented using FedAvg (McMahan et al., 2017).
-Selection of Hyperparameters.
-For the MNIST dataset, we implemented logistic regression using the SGD solver with 5 epochs, a batch size of 32, and 20 total runs.
-For the CNN model on the MNIST dataset, the same hyperparameter setup was used, except the batch size was reduced to 16 to ensure stable convergence.
+Selection of Hyperparameters and Convolutional Neural Network Model.
+For the MNIST dataset, we implemented logistic regression using the SGD solver with 5 epochs, a batch size of 32, and 20 total runs, and the same hyperparameter setup was used, except the batch size was reduced to 16 to ensure stable convergence in CNN with two convolution layers.
+To make our experiment more comprehensive, we also utilized a five-layer CNN to train the MNIST dataset with 10 epochs, a batch size of 16, and 50 total runs.
+For the EMNIST dataset, we used the same five-layer CNN with the same settings.
+And for CIFAR10, we implemented a three-layer CNN with 10 epochs, a batch size of 16, and 50 total runs.
+To study the data heterogeneity, we set the number of classes for each client in the MNIST, EMNIST, and CIFAR10 datasets as $\{2,6,10\},\{30,40,47\},$ and $\{2,6,10\}$, respectively.
 Optimizer.
 For the synthetic dataset, following the analysis in the paper, the loss function is optimized using gradient descent (GD).
-For the MNIST dataset, all models are implemented in PyTorch, and optimization is performed using the stochastic gradient descent (SGD) solver.
+For the real dataset, all models are implemented in PyTorch, and optimization is performed using the stochastic gradient descent (SGD) solver.
 Evaluation Criterion.
@@
 Figure 3: Performance of the FedProx algorithm with an adaptive choice of $\lambda$ (from Theorem 1) under varying levels of statistical heterogeneity.
-In Figure 3, Statistical error of the FedProx algorithm with an adaptive choice of $\lambda$ (from Theorem 1) under varying levels of statistical heterogeneity.
+In Figure 3, statistical error of the FedProx algorithm with an adaptive choice of $\lambda$ under varying levels of statistical heterogeneity.
 With the adaptive choice of $\lambda$ as defined in Theorem 1, FedProx consistently achieves the lowest statistical error across most scenarios, outperforming both GlobalTrain and LocalTrain baselines.
@@
 This highlights the trade-off between local computation and communication efficiency under different levels of personalization, validating our theoretical findings in Corollary 1.
+Dataset Client Class One-stage FedProx GlobalTrain LocalTrain pFedMe
+(small) (medium) (large) (small) (medium) (large)
+MNIST CNN2 2 0.3567 0.4721 0.6019 0.6123 0.3019 0.3973 0.5828 0.6172
+6 0.7001 0.7833 0.7843 0.8092 0.6231 0.7918 0.8093 0.8312
+10 0.8753 0.8893 0.9032 0.9098 0.8194 0.8756 0.8771 0.8749
+MNIST CNN5 2 0.8891 0.9000 0.9333 0.9377 0.6536 0.9000 0.9003 0.9380
+6 0.8451 0.8941 0.9392 0.9446 0.7333 0.9288 0.9186 0.9306
+10 0.9091 0.9102 0.9421 0.9433 0.7633 0.9340 0.9385 0.9440
+EMNIST CNN5 30 0.6084 0.6089 0.6011 0.6102 0.5583 0.6133 0.6129 0.6122
+40 0.6097 0.6123 0.6357 0.6423 0.5644 0.6012 0.6102 0.6134
+47 0.6278 0.6415 0.6672 0.6611 0.6111 0.6345 0.6532 0.6717
+CIFAR10 CNN3 2 0.6101 0.6292 0.6340 0.6444 0.5801 0.6056 0.6012 0.6033
+6 0.6211 0.6311 0.6712 0.6949 0.5623 0.6163 0.6163 0.6439
+10 0.6102 0.6712 0.7302 0.8041 0.6238 0.6744 0.7443 0.7732
+[…]
+To further validate our conclusions about the effect of personalization on the statistical accuracy of FedProx (Objective 1), we conduct additional experiments on real datasets, including EMNIST and CIFAR-10.
+For each dataset, we explore three levels of statistical heterogeneity, controlled by varying the number of classes each client has access to, which follows a widely used setup in federated learning literature.
+The models are trained using different convolutional neural network (CNN) architectures, denoted as CNN2 and CNN5, where the numbers indicate the number of convolutional layers in the network.
+As the true underlying model is not accessible for these real datasets, we evaluate model performance based on classification accuracy on the test set, as shown in Tables 3 and 4.
+We also conduct experiments another regularization-based personalization method, pFedMe (T Dinh et al., 2020), to demonstrate that our results could be generalized to other regularization-based personalized federated learning methods as well.
+For the columns corresponding to One-stage FedProx and pFedMe, we conduct experiments with varying regularization parameter $\lambda$, investigating the role of personalization in the algorithms.
+Specifically, we experiment with small, medium, and large $\lambda$ values (see Appendix C.1 for detailed parameter settings).
+In Table 3, we can see that consistently across datasets, as $\lambda$ decreases (indicating a higher degree of personalization), the accuracy approaches that of LocalTrain, which trains separate models for each client.
+Conversely, as $\lambda$ increases (indicating lower personalization and more collaboration), the accuracy approaches that of GlobalTrain, which represents purely global training.
+This interpolation behavior between LocalTrain and GlobalTrain, enabled by FedProx’s personalized objective, aligns well with our theoretical results in Theorem 1 and highlights the adaptive capability of the FedProx method.
+Moreover, as dataset homogeneity increases (i.e., clients share more similar data distributions), we observe a general improvement in accuracy across all methods, regardless of the level of personalization.
+This result demonstrates that while personalization offers significant benefits in highly heterogeneous settings, its necessity diminishes as heterogeneity decreases.
+Additionally, experiments with different CNN architectures confirm the robustness of these findings across model designs.
+Although we did not provide a theoretical analysis for pFedMe in this paper, our results suggest that the empirical behavior of pFedMe is consistent with the theoretical insights derived for FedProx.
+In particular, we observe that pFedMe also exhibits interpolation between local and global learning under varying levels of data heterogeneity.
+Dataset Client Class One-stage FedProx (small $\lambda$) One-stage FedProx (large $\lambda$) Dichotomous Strategy
+MNIST CNN2 2 0.4721 0.6019 0.2567
+MNIST CNN2 6 0.7833 0.7843 0.7001
+MNIST CNN2 10 0.8893 0.9032 0.8753
+MNIST CNN5 2 0.8891 0.9333 0.7032
+MNIST CNN5 6 0.8451 0.8941 0.7941
+MNIST CNN5 10 0.9091 0.9102 0.8922
+EMNIST CNN5 30 0.6084 0.6089 0.5291
+EMNIST CNN5 40 0.6097 0.6123 0.5732
+EMNIST CNN5 47 0.6278 0.6415 0.5856
+Table 4: Comparison of evaluation accuracy for One-stage FedProx with different values of $\lambda$ and Dichotomous Strategy across various datasets and client classes.
+C.3 Comparison of FedProx with A Dichotomous Strategy
+A simple baseline in personalized federated learning is the dichotomous strategy, where the dataset on each client is split into two subsets: one is used for global training (GlobalTrain) with pooled client data, while the other is used for purely local training (LocalTrain).
+After training, the two models are compared on the test set, and the best-performing model is selected for each client.
+Such a strategy is known to achieve minimax-optimal statistical accuracy under certain conditions; see Chen et al.
+(2023b) for a detailed discussion.
+Even though the convergence rate coincides with such a dichotomous strategy, we may reiterate that the key contribution of our work lies in the fine-grained analysis of the FedProx method—a method that has demonstrated strong empirical performance in federated learning [2].
+Through our analysis, we have, for the first time, established the minimax optimality of FedProx, showing that its statistical performance matches the lower bound of the estimation problem.
+If one seeks to understand why FedProx could be advantageous over the dichotomous strategy in practice, here are some of the reasons:
+  • Stability in Heterogeneous Environments: Empirical studies [1] demonstrate that FedProx achieves more stable convergence compared to methods such as FedAvg, especially in settings with significant in a heterogeneous network.
+  • Adaptivity and Flexibility: FedProx seamlessly interpolates between the extremes of global training and local training by tuning a single regularization parameter, $\lambda$.
+This adaptability allows practitioners to dynamically balance the trade-off between bias and variance based on the level of statistical heterogeneity, without the need for fixed, hard decisions inherent in the dichotomous strategy.
+To further compare FedProx with the dichotomous strategy, we conduct a simple experiment.
+In the dichotomous setup, each client’s dataset is partitioned into two subsets: one subset is used for local training (LocalTrain), where each client trains independently, and the other subset is pooled across all clients for global training (GlobalTrain).
+The accuracy of the two resulting models is evaluated on a test set, and the better-performing model is reported for each client.
+By contrast, FedProx is implemented using the experimental setup described earlier, leveraging the full dataset.
+As shown in Table 4, FedProx consistently outperforms the dichotomous strategy across various configurations.
+This performance gap can be attributed to the inefficiency of the dichotomous approach, which sacrifices valuable data during the sample splitting process, resulting in reduced sample efficiency.
+Even in a hypothetical scenario where the full sample is used for both LocalTrain and GlobalTrain in the dichotomous strategy, the results in Table 3 demonstrate that FedProx achieves a continuous interpolation between the two extremes of LocalTrain and GlobalTrain by adjusting its regularization parameter $\lambda$.
+This aligns with our theoretical findings.

### v4

@@
-The Effect of Personalization in FedProx: A Fine-grained Analysis on Statistical Accuracy and Communication Efficiency
+Understanding the Statistical Accuracy-Communication Trade-off in Personalized Federated Learning with Minimax Guarantees
 Abstract
-FedProx is a simple yet effective federated learning method that enables model personalization via regularization.
-Despite remarkable success in practice, a rigorous analysis of how such a regularization provably improves the statistical accuracy of each client’s local model hasn’t been fully established.
-Setting the regularization strength heuristically presents a risk, as an inappropriate choice may even degrade accuracy.
-This work fills in the gap by analyzing the effect of regularization on statistical accuracy, thereby providing a theoretical guideline for setting the regularization strength for achieving personalization.
-We prove that by adaptively choosing the regularization strength under different statistical heterogeneity, FedProx can consistently outperform pure local training and achieve a minimax-optimal statistical rate.
-In addition, to shed light on resource allocation, we design an algorithm, provably showing that stronger personalization reduces communication complexity without increasing the computation cost overhead.
-Finally, our theory is validated on both synthetic and real-world datasets and its generalizability is verified in a non-convex setting.
+Personalized federated learning (PFL) offers a flexible framework for aggregating information across distributed clients with heterogeneous data.
+This work considers a personalized federated learning setting that simultaneously learns global and local models.
+While purely local training has no communication cost, collaborative learning among the clients can leverage shared knowledge to improve statistical accuracy, presenting an accuracy-communication trade-off in personalized federated learning.
+However, the theoretical analysis of how personalization quantitatively influences sample and algorithmic efficiency and their inherent trade-off is largely unexplored.
+This paper makes a contribution towards filling this gap, by providing a quantitative characterization of the personalization degree on the tradeoff.
+The results further offers theoretical insights for choosing the personalization degree.
+As a side contribution, we establish the minimax optimality in terms of statistical accuracy for a widely studied PFL formulation.
+The theoretical result is validated on both synthetic and real-world datasets and its generalizability is verified in a non-convex setting.
+Personalized Federated Learning, Statistical Complexity, Communication Complexity
 1 Introduction
-Federated Learning (FL) has emerged as an attractive framework for aggregating distributed data, enabling clients to collaboratively train a shared global model while preserving data privacy.
-In the currently prevalent paradigm (McMahan et al., 2017), FL is formulated as a finite sum minimization problem focusing on a single shared model.
-Nevertheless, it has been well recognized that one of the key challenges in FL is the statistical heterogeneity of the client datasets.
-As the participants collect their own local data, it often reflects client-specific characteristics and is not identically distributed.
-With high statistical heterogeneity, training a single model for all clients by minimizing their average in-sample loss becomes questionable.
-To address this challenge, one solution is to relax the common model constraint and solve alternatively the following objective in FedProx (Li et al., 2020a):
-$\min_{{{\bm{w}}^{(g)}},\{{{\bm{w}}^{(i)}}\}_{i\in[m]}}\sum_{i\in[m]}p_{i}\left({L}_{i}({{\bm{w}}^{(i)}},S_{i})+\frac{\lambda}{2}\|{{\bm{w}}^{(g)}}-{{\bm{w}}^{(i)}}\|^{2}\right),$ (1)
-where $L_{i}({\bm{w}}^{(i)},S_{i})$ is the empirical risk of the $i$-th client, and $S_{i}$ is the local dataset of the $i$-th client.
-The set $\{p_{i}\}_{i\in[m]}$ is a collection of weights for the clients satisfying $p_{i}\geq 0$ and $\sum_{i}p_{i}=1$.
-In (1), each client $i$ learns a local personalized model ${\bm{w}}^{(i)}$ by fitting its local data $S_{i}$, while enabling collaborative learning by shrinking all local models towards a common global one ${\bm{w}}^{(g)}$ with intensity controlled by $\lambda$.
-The smaller $\lambda$ is, the weaker the coupling of the local models the formulation would enforce thus the higher personalization is.
-Adjusting the strength of regularization controls the extent to which the ${\bm{w}}^{(i)}$’s is adapted to client $i$’s local data, thus the degree of personalization.
-The promising empirical performance of FedProx in learning under heterogeneous data (Li et al., 2020a; Donevski et al., 2021; He et al., 2021; An et al., 2023) has led to a sequence of works studying the problem from a pure computation perspective, aiming at developing efficient algorithms under various problem assumptions to reach a (local) solution (Hanzely & Richtárik, 2020; T Dinh et al., 2020; Li et al., 2021; Acar et al., 2021; Yuan & Li, 2022).
-However, the fundamental question of how to characterize the accuracy of the models produced by (1) is largely unaddressed.
-As such, for a given set of clients, it is not even clear if introducing personalization will be beneficial, outperforming the baselines given by training a single global model or pure local training.
-In addition, communication cost in FL is a critical bottleneck in resource-constrained environments.
-Therefore, understanding when personalization will improve both training efficiency and statistical accuracy is of paramount importance.
-We address these questions by providing a fine-grained analysis of the effect of personalization in FedProx.
-Our contribution is summarized below:
-$\bullet$ Minimax-optimal Statistical Accuracy.
-We analyze the statistical accuracy of the objective in (1), taking statistical heterogeneity into account.
-For each of the models, we establish a state-of-the-art convergence rate $\mathcal{O}(1/(mn))+\mathcal{O}(1/n\wedge R^{2})$, where $R$ characterizes the statistical heterogeneity and $n$ is the sample size on each client.
-The first term $\mathcal{O}(1/(mn))$ represents the sample complexity of training a single global model under homogeneity, which is attained by (1) when $R$ is close to zero, highlighting the benefit of FL.
-For the second term, the first component $\mathcal{O}(1/n)$ is the sample complexity of purely local training, and the second component $\mathcal{O}(R^{2})$ demonstrates the flexibility of FedProx under different statistical heterogeneity, implying the statistical accuracy of FedProx never gets worse than purely local training’s.
-Such a rate matches the existing lower bound for the problem.
-Therefore, to the best of our knowledge, this is the first work that demonstrates that with a theoretically-informed adaptive choice of personalization degree characterized by $\lambda$, the objective in (1) achieves a minimax-optimal statistical accuracy across different levels of statistical heterogeneity.
-The theoretical analysis we use for objective (1) would extend to different objectives in regularization-based personalized federated learning (Smith et al., 2017; Li et al., 2020a; Hanzely & Richtárik, 2020; Hanzely et al., 2020; 2023).
-$\bullet$ Provably Enhanced Communication Efficiency.
-As the objective (1) could be reformulated as a bi-level problem, we propose an algorithm for (1) with an explicit computation cost assessment.
-We establish the communication complexity as $\mathcal{O}(\kappa\frac{\lambda+\mu}{\lambda+L}\log\frac{1}{\varepsilon})$ and the local computation complexity in each communication round as $\mathcal{\widetilde{O}}({\frac{\lambda+L}{\lambda+\mu}})$.
-With higher personalization degree, communication complexity would decrease and the computation complexity per round would increase, demonstrating that there is a trade-off between computation and communication costs.
-Our analysis concludes that model personalization can reduce communication overhead without incurring additional computation expenses.
-$\bullet$ Empirical Validation.
-We conduct numerical studies on logistic regression with synthetic data and real-world data under our problem assumptions.
+Federated Learning (FL) (McMahan et al., 2017) has emerged as a promising learning framework for aggregating information from distributed data, allowing clients to collaboratively train a shared global model in a communication-efficient manner.
+However, a critical challenge arising in the presence of various data across clients is data heterogeneity.
+In such cases, a single global model may fail to generalize well to all clients, motivating the need for model personalization.
+PFL enhances FL by learning personalized models tailored to individual clients, and has demonstrated strong empirical performance in various applications, such as driver monitoring (Yuan et al., 2023), and mobile computing (Zhang et al., 2024).
+An important question in PFL is determining the degree of personalization, which controls the transition between fully collaborative training and pure local training.
+A higher degree of collaboration (less personalization) typically requires more frequent information exchange, potentially improving learning accuracy when client data distributions are similar.
+Conversely, increasing personalization reduces communication costs by prioritizing localized training, but may lead to higher generalization errors due to the limited size of client datasets (Paragliola, 2022).
+Understanding this trade-off is essential for optimizing model performance under communication constraints.
+Most existing works focus purely on the algorithmic perspective of PFL (Lin et al., 2022; Li et al., 2024b; Wang et al., 2022, 2024b).
+However, the statistical accuracy of the solutions obtained in PFL remains largely unexplored.
+As a result, the connections between statistical accuracy, communication efficiency, and their trade-offs are not well understood, leaving a theoretical gap in understanding how to select the optimal personalization degree.
+In this paper, we fill in this gap by providing a fine-grained theoretical analysis of a widely adopted PFL problem formulation given by  (2) that trains simultaneously local and global models.
+We quantitatively analyze the influence of personalization on both statistical and optimization convergence rates for each of the local models.
+Specifically, our contribution can be summarized as follows:
+$\bullet$ Statistical Accuracy with Optimal Guarantee.
+We provide a non-asymptotic statistical convergence rate of the solution of Problem (2), revealing how personalization degree influences the statistical accuracy of each local model.
+In particular, as the personalization degree increases, the rate approaches to that of pure local training, $\mathcal{O}(1/n)$, where $n$ is the sample size per client.
+Conversely, decreasing the personalization degree allows the models to utilize information across all clients, achieving a rate closer to $\mathcal{O}(1/(mn)+R^{2})$, where $mn$ is the total sample size of all clients and $R$ quantifies the statistical heterogeneity of the local datasets.
+We further establish the minimax optimality of the derived statistical rates, demonstrating the tightness of our analysis.
+To the best of our knowledge, this is the first work to achieve such an optimality.
+$\bullet$ Communication Efficiency under Personalization.
+On the optimization aspect, we treat Problem (2) as a bi-level problem and propose (stochastic) algorithms with an explicit characterization of computation and communication complexity.
+We establish that the communication cost of finding an $\varepsilon$-solution is $\mathcal{O}(\kappa\frac{\lambda+\mu}{\lambda+L}\log\frac{1}{\varepsilon})$ and the computation cost in terms of gradient evaluations is $\mathcal{O}(\kappa\log\frac{1}{\varepsilon})$, where $\lambda$ is a parameter determining the personalization degree.
+This demonstrates that a smaller $\lambda$, corresponding to a higher degree of personalization, reduces communication complexity without incurring additional computational overhead.
+$\bullet$ Accuracy-Communication Tradeoff with Empirical Validation.
+Building on our theoretical results for statistical and optimization convergence, we quantitatively characterize the trade-off between statistical accuracy and communication efficiency in PFL and discuss practical insights for selecting the personalization degree.
+We then conduct numerical studies on logistic regression with synthetic data and real-world data under our assumptions.
 The results corroborate our theoretical findings.
-We further test the numerical performance of (1) for training CNNs, showing empirically the results, although proved under the convexity assumption, generalize to nonconvex losses.
+We further test the numerical performance with CNN models, showing that the theoretical results, though established under convexity assumptions, generalize to non-convex settings.
 2 Related Works
-In recent years, federated learning has become an attractive solution for training models locally on distributed clients, rather than transferring data to centralized data centers (Mammen, 2021; Wen et al., 2023; Beltrán et al., 2023).
-As each client generates its local data, statistical heterogeneity is common with data being non-identically distributed between clients (Li et al., 2020a; Ye et al., 2023).
-Given the variability of data in network, model personalization is a natural strategy used to improve learning accuracy for each client.
+In recent years, FL has become an attractive solution for training models locally on distributed clients, rather than transferring data to a single node for centralized processing (Mammen, 2021; Wen et al., 2023; Beltrán et al., 2023).
+As each client generates its local data, statistical heterogeneity naturally arises with data being non-identically distributed between clients (Li et al., 2020; Ye et al., 2023).
+Given the variability of data in a network, model personalization is an appealing strategy used to improve statistical accuracy for each client.
 Formulations enabling model personalization have been studied independently from multiple fields.
-For example, meta-learning (Chen et al., 2018; Jiang et al., 2019; Khodak et al., 2019; Fallah et al., 2020) assumes all models follow a common distribution.
-By minimizing the average loss, these methods aim to learn a meta-model that generalizes well to unseen new tasks.
-Multi-task learning methods (Agarwal et al., 2020; Liang et al., 2020) and transfer learning methods (Li et al., 2022; He et al., 2024a; b) study formulation similar to (1), but the statistical rate is established for either the average of all models or only the target model.
-Therefore, it is not clear if all clients can benefit from the procedure and learn a more accurate model than purely local training.
-Representation learning (Zhou et al., 2020; Wang et al., 2024), on the other hand, focuses on a different model structure where the local models can be represented as the composition of two parts, with one common to all clients and the other specific to each client.
+For example, meta-learning (Chen et al., 2018; Jiang et al., 2019; Khodak et al., 2019; Fallah et al., 2020) assumes all local models follow a common distribution.
+By minimizing the average validation loss, these methods aim to learn a meta-model that generalizes well to unseen new tasks.
+Representation learning (Zhou et al., 2020; Wang et al., 2024a) focuses on a setting where the local models can be represented as the composition of two parts, with one common to all clients and the other specific to each client.
+Our work, however, differs in that we focus on a personalized FL setting with a mixture of global and local models.
+Closely related are multi-task learning methods (Liang et al., 2020) and transfer learning methods (Li et al., 2022; He et al., 2024a, b), but the statistical rate is established for either the average of all models or only the target model.
+Therefore, it remains unclear how personalization influences the statistical accuracy of each individual local model in these settings.
 In the FL community, personalized FL methods can be broadly divided into two main strands.
-Different from the works mentioned previously, studies here focus primarily on the properties of the iterates computed by the algorithms.
-One line of works (Arivazhagan et al., 2019; Liang et al., 2020; Singhal et al., 2021; Collins et al., 2021) is based on the representation learning formulation, where both the algorithmic convergence and statistical accuracy of the iterates have been investigated.
-Another line of works (Smith et al., 2017; Li et al., 2020a; Hanzely & Richtárik, 2020; Hanzely et al., 2020; 2023), including FedProx, achieves personalization by relaxing the requirement of learning a common global model through regularization techniques.
-In particular, algorithms and complexity lower bounds specific to the FedProx objective were studied in works (Li et al., 2020a; Hanzely et al., 2020; T Dinh et al., 2020; Hanzely & Richtárik, 2020; Li et al., 2021), see Table 2 for a detailed comparison.
-These works study the problem from a pure optimization perspective and cannot show personalization improves learning accuracy.
-There are only a few recent works we are aware of studying the statistical accuracy of regularization-based PFL in the form of (1).
+Different from the works mentioned previously, studies here primarily focus on the properties of the iterates generated by the algorithms.
+One line of work (Arivazhagan et al., 2019; Liang et al., 2020; Singhal et al., 2021; Collins et al., 2021) is based on the representation learning formulation.
+Another line of work (Smith et al., 2017; Li et al., 2020; Hanzely & Richtárik, 2020; Hanzely et al., 2020; Li & Richtárik, 2024) achieves personalization by relaxing the requirement of learning a common global model through regularization techniques.
+In particular, algorithms and complexity lower bounds specific to Problem (2) were studied in (Li et al., 2020; Hanzely et al., 2020; T Dinh et al., 2020; Hanzely & Richtárik, 2020; Li et al., 2021), see Table 1 for a detailed comparison.
+These works study Problem (2) from a pure optimization perspective and have not provided how personalization influences statistical accuracy and, consequently, the trade-off between statistical accuracy and communication efficiency.
+There are only a few recent works we are aware of that study the statistical accuracy of regularization-based PFL.
 Specifically, Cheng et al.
-(2023) investigates the asymptotic behavior of the FedProx estimator under an over-parameterized linear regression model; neither a finite-sample rate nor algorithm is provided.
+(2023) investigates the asymptotic behavior of the (personalized) federated learning under an over-parameterized linear regression model; neither a finite-sample rate nor algorithms are provided.
 Chen et al.
-(2023b) studies the FedProx estimator under rather strong assumptions like bounded parameter space.
-By defining a new notion of algorithmic stability, they establish the non-asymptotic sample complexity for FedProx and induce minimax bound.
+(2023c) studies Problem (2) and establishes a non-asymptotic sample complexity.
 However, even with some overly strong assumptions, their analysis cannot match the statistical lower bound in most cases.
+Furthermore, neither of these works reveals the trade-off.
+While (Bietti et al., 2022) studies the influence of personalization in PFL, their focus is on the trade-off between privacy and optimization error, leaving the accuracy-communication trade-off unexplored.
 Notation.
-$[n]$ denotes the set of integers $\{1,2,\cdots,n\}$ for any positive integer $n$.
-Let $\|\cdot{}\|_{2}$ denote the $\ell_{2}$ vector norm and the Frobenius matrix norm.
-For two non-negative sequences $a_{n}$, $b_{n}$, we denote $a_{n}=\mathcal{O}(b_{n})$ as $a_{n}/b_{n}\xrightarrow{}{1}$.
+Denote $[n]:=\{1,2,\cdots,n\}$.
+$\|\cdot{}\|_{2}$ denotes the $\ell_{2}$ norm for vectors and the Frobenius norm for matrices.
+For two non-negative sequences $\{a_{n}\},\{b_{n}\}$, we denote $a_{n}\lesssim b_{n}$ if $a_{n}\leq Cb_{n}$ for some constant $C>0$ when $n$ is sufficiently large.
+We also use $a_{n}=\mathcal{O}\left(b_{n}\right)$, whose meaning is the same as $a_{n}\lesssim b_{n}$.
 $\tilde{\mathcal{O}}(\cdot)$ hides logarithmic factors.
+For two real numbers $a$, $b$, we let $a\wedge b=\min\{a,b\}$ and $a\vee b=\max\{a,b\}$.
 3 Preliminaries
-Considering there are $m$ clients, each client has a collection of $i.i.d$ data $S_{i}=\{z_{ij}\}_{j\in[n_{i}]}$ over possibly different data distributions $\mathcal{D}_{i}$.
-For each client, the ultimate goal is to find the solution that minimizes the local risk, i.e.
-${{\bm{w}}_{\star}^{(i)}}\in\arg\min_{{\bm{w}}}\>\mathbb{E}_{z\sim\mathcal{D}_{i}}\ell({\bm{w}},z)$.
-Here, $\ell({\bm{w}},z)$ denotes the loss function of interest and ${{\bm{w}}_{\star}^{(i)}}\in\mathbb{R}^{d}$ is called the ground truth local model for client $i$.
-In this paper, we assume the model dimension $d$ is finite.
-Under such a setup, two commonly adopted approaches are pure local training (LocalTrain) and pure global training (GlobalTrain):
+We consider an FL setting with $m$ clients.
+Each client $i$ has a collection of data $S_{i}=\{z_{ij}\}_{j\in[n_{i}]}$, where $n_{i}$ is the sample size of client $i$, with elements i.i.d.
+drawn from distribution $\mathcal{D}_{i}$.
+The data distributions among the clients are possibly heterogeneous.
+For each client $i$, the ultimate goal is to find a model that minimizes its local risk at the population level, defined as:
+${{\bm{w}}_{\star}^{(i)}}\in\arg\min_{{\bm{w}}}\>\mathbb{E}_{z\sim\mathcal{D}_{i}}\ell({\bm{w}},z),$ (1)
+where $\ell({\bm{w}},z)$ is a loss function measuring the fitness of ${\bm{w}}$ to data point $z$.
+We call ${{\bm{w}}_{\star}^{(i)}}\in\mathbb{R}^{d}$ the ground truth model for client $i$.
+To understand the influence of model personalization, we study the following widely adopted PFL problem (Hanzely & Richtárik, 2020; Mishchenko et al., 2023; Li et al., 2020):
+$\min_{\begin{subarray}{c}{{\bm{w}}^{(g)}}\\ \{{{\bm{w}}^{(i)}}\}_{i=1}^{m}\end{subarray}}\sum_{i\in[m]}p_{i}\left({L}_{i}({{\bm{w}}^{(i)}},S_{i})+\frac{\lambda}{2}\|{{\bm{w}}^{(g)}}-{{\bm{w}}^{(i)}}\|^{2}\right),$ (2)
+where ${{\bm{w}}^{(g)}}$ and ${{\bm{w}}^{(i)}}$ are respectively the global and $i$-th local model to be learned, $L_{i}({\bm{w}},S_{i}):=\sum_{j=1}^{n_{i}}\ell({\bm{w}},z_{ij})/n_{i}$ is the empirical risk of client $i$ on its local data $S_{i}$, and the last term is a regularization term with parameter $\lambda$ controlling the personalization degree.
+The set $\{p_{i}\}_{i\in[m]}$ is a collection of nonnegative weights with $\sum_{i=1}^{m}p_{i}=1$.
+In Problem (2), each client $i$ learns a personalized model ${{\bm{w}}^{(i)}}$ by fitting to its local data $S_{i}$, while collaboration is achieved by shrinking the local models ${{\bm{w}}^{(i)}}$ towards a common global model ${{\bm{w}}^{(g)}}$.
+As $\lambda\to 0$, the influence of the global model diminishes, and Problem (2) increasingly behaves like the LocalTrain problem:
 […]
-where $L_{i}({\bm{w}},S_{i}):=\sum_{j=1}^{n_{i}}\ell({\bm{w}},z_{ij})/n_{i}$ denotes the empirical risk of client $i$.
-In LocalTrain, each client individually minimizes its loss, whereas in GlobalTrain, all clients collaboratively minimize their average loss.
-Compared to LocalTrain, GlobalTrain integrates information from all clients to improve the learning performance of each client.
-When there are sufficient similarities between the data distribution $\mathcal{D}_{i}$ of different clients, GlobalTrain would be a better choice as more samples are included.
-However, if $\mathcal{D}_{i}$’s are highly unrelated, LocalTrain will be a more robust choice.
-In practice, the data distribution is neither totally the same nor completely unrelated across clients. To handle the wide middle ground of the two extremes, FedProx solves the following problem
-$\min_{{{\bm{w}}^{(g)}},\{{{\bm{w}}^{(i)}}\}_{i\in[m]}}\sum_{i=1}^{m}p_{i}\left({L}_{i}({{\bm{w}}^{(i)}},S_{i})+\frac{\lambda}{2}\|{{\bm{w}}^{(g)}}-{{\bm{w}}^{(i)}}\|^{2}\right).$
-The key feature of Fedprox is the incorporation of a proximal (regularization) term, which controls the personalization degree.
-As $\lambda\rightarrow 0$, the FedProx objective increasingly behaves like LocalTrain with higher personalization degree and reduces to $m$ separate LocalTrain problems in the end.
-On the other hand, as $\lambda\rightarrow\infty$, the local models ${\bm{w}}^{(i)}$ are shrunk towards the global model ${\bm{w}}^{(g)}$ with lower personalization degree and eventually degenerate to optimize a single shared model.
-Thus, by adjusting $\lambda$, FedProx interpolates between LocalTrain and GlobalTrain, allowing for a flexible choice of personalization degree to achieve higher modeling accuracy.
-It’s noticeable that this choice of $\lambda$ also affects the optimization process of FedProx.
-With larger $\lambda$, more frequent communication between clients is expected to ensure faster convergence, as local updates are more tightly coupled to the global model.
-In contrast, smaller $\lambda$ value allows for more independent updates, reducing the need for communication.
-In the next few sections, we provide a theoretical guarantee for the intuition provided above.
-In Section 4, we analyze when the personalization level, characterized by $\lambda$, would improve the statistical accuracy of both the global model ${\bm{w}}^{(g)}$ and the local models $\{{{\bm{w}}^{(i)}}\}_{i\in[m]}$.
-This is in sharp contrast to most of the existing works where the error is established for the sum of them.
-In Section 5, we explore how such a degree of personalization will affect communication complexity.
-4 When will Personalization Improve Statistical Accuracy
-In this section, we analyze how the choice of $\lambda$, or the degree of personalization, would improve the statistical accuracy of the solution obtained from the FedProx problem (1).
-Following the existing literature (Li et al., 2023; Chen et al., 2023a; b; Duan & Wang, 2023), we define the statistical heterogeneity as the Euclidean distance between the true local models ${{\bm{w}}_{\star}^{(i)}}$ and the true global model ${{\bm{w}}_{\star}^{(g)}}:=\sum_{i=1}^{m}p_{i}{{\bm{w}}_{\star}^{(i)}}$.
-Specifically, we adopt the following assumption:
-Assumption 1 (Statistical Heterogeneity).
-There exists a nonnegative constant $R$ such that we have $\|{{\bm{w}}_{\star}^{(i)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}\leq R^{2},\forall i\in[m]$.
-Intuitively, a larger $R$ indicates a greater divergence between the clients’ data distributions, thus stronger statistical heterogeneity, and vice versa.
-Therefore, in the FedProx objective, the choice of $\lambda$ should be adjusted according to $R$.
-More discussion about Assumption 1 can be found in Appendix A.1.
-Next, we introduce three standard assumptions.
+where each client independently trains its model using only local data.
+In this case, Problem (2) is fully decoupled into $m$ separate problems, achieving the maximum degree of personalization.
+On the other hand, as $\lambda\to\infty$, the regularization term shrinks all local models ${{\bm{w}}^{(i)}}$ towards the global model ${{\bm{w}}^{(g)}}$.
+This corresponds to reducing the personalization degree and eventually pushes Problem (2) to another extreme given by:
+[…]
+where the knowledge from all clients is pooled to train a single global model.
+Adjusting $\lambda$, therefore, controls the degree of personalization among clients and thus affects the statistical accuracy of the resulting solution.
+Noticeably, the choice of $\lambda$ also potentially affects solving Problem (2) algorithmically in the FL setting.
+Intuitively, for large $\lambda$, we anticipate more frequent communications among the clients to facilitate collaboration as the local models are more tightly coupled to the global model.
+In contrast, a smaller $\lambda$ encourages more independent updates and reduces the need for communication, with LocalTrain (cf.(3)) being an edge case where all clients train independently without any communication.
+We quantify both the statistical and optimization error of a PFL algorithm applied to Problem (2) by evaluating the Euclidean distance between the algorithm’s output after $t$ communication rounds, ${\bm{w}}_{t}^{(i)}$, and the ground truth local model ${{\bm{w}}_{\star}^{(i)}}$ for each client $i\in[m]$.
+Let $\widetilde{\bm{w}}^{(i)}$ denote the solution of $i$-th local model in Problem (2).
+We have the following decomposition:
+$\|{\bm{w}}_{t}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}\leq 2\underbrace{\|{\bm{w}}_{t}^{(i)}-\widetilde{\bm{w}}^{(i)}\|^{2}}_{\text{optimization error}}+2\underbrace{\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}}_{\text{statistical error}}.$
+In the rest of the paper, we first establish the statistical and optimization convergence rates independently.
+We then integrate these results to formally provide the trade-off between statistical accuracy and communication efficiency under different levels of personalization.
+4 Convergence Rate Analysis
+4.1 Effect of Personalization on Statistical Accuracy
+In this section, we analyze how the choice of $\lambda$ influences the statistical accuracy of the solution to Problem (2).
+Recall the ground truth local model ${{\bm{w}}_{\star}^{(i)}}\in\mathbb{R}^{d}$ defined in (1).
+Next, we introduce the measure of statistical heterogeneity and define the parameter space for the statistical estimation problem.
+Specifically, we consider estimating ${{\bm{w}}_{\star}^{(i)}}$ from the following parameter space:
+$\mathcal{P}(R)=\left\{\{{{\bm{w}}_{\star}^{(i)}}\}_{i=1}^{m}:\left\|{{\bm{w}}_{\star}^{(i)}}-\sum_{i=1}^{m}p_{i}{{\bm{w}}_{\star}^{(i)}}\right\|^{2}\leq R^{2},\forall i\in[m]\right\}.$ (5)
+In (5), the statistical heterogeneity is measured as the Euclidean distance between the ground truth local models and their weighted average.
+A larger $R$ indicates a larger difference among the clients’ local models, hence stronger statistical heterogeneity.
+Such a measure is commonly imposed in the existing literature (Li et al., 2023; Chen et al., 2023b, c; Duan & Wang, 2023).
+More discussion about the parameter space $\mathcal{P}(R)$ can be found in Appendix A.1.
+We assume the model dimension $d$ is finite.
+Under such a parameter space, we aim to investigate the statistical accuracy of the solution to Problem (2) measured by $\mathbb{E}\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}$, where the expectation is taken with respect to the joint distribution.
+Remark 1.
+Note that some existing literature adopts alternative metrics, such as individual excess risk (Chen et al., 2023c) to measure the statistical error.
+These two metrics are equivalent under the strong convexity and smoothness conditions on the loss functions.
+We study Problem (2) under the following regularity conditions.
 Assumption 1 (Smoothness).
-The loss function $\ell(\cdot,z)$ is L-smooth, i.e. for any $x,y\in\mathbb{R}^{d}$
-$\|\nabla\ell(x,z)-\nabla\ell(y,z)\|\leq L\|x-y\|,\quad\forall z.$ (3)
+The loss function $\ell(\cdot,z)$ is L-smooth, i.e. for any $x,y\in\mathbb{R}^{d}$:
+$\|\nabla\ell(x,z)-\nabla\ell(y,z)\|\leq L\|x-y\|,\ \forall z.$ (6)
 Assumption 2 (Strong Convexity).
-The loss function $\ell(\cdot,z)$ is almost surely $\mu$-strongly convex: $\ell(x,z)\geq\ell(y,z)+\langle\nabla\ell(y,z),y-x\rangle+\frac{\mu}{2}\|x-y\|^{2},\quad\forall x,y\in\mathbb{R}^{d}.$
+The empirical loss $L_{i}(\cdot,S_{i})$ is $\mu$-strongly convex, i.e., for any $x,y\in\mathbb{R}^{d}$ and $i\in[m]$:
+$L_{i}(x,S_{i})\geq L_{i}(y,S_{i})+\langle\nabla L_{i}(y,S_{i}),x-y\rangle+\frac{\mu}{2}\|x-y\|^{2}.$ (7)
 Assumption 3 (Bound Gradient Variance at Optimum).
-There exists a nonnegative constant $\rho$ such that $\forall i\in[m]$, we have $\mathbb{E}_{z\sim\mathcal{D}_{i}}\|\nabla\ell({{\bm{w}}_{\star}^{(i)}},z)\|^{2}\leq\rho^{2}$.
-Here, the strongly convex and smooth assumption is used to establish the estimation error and is widely adopted in the theoretical analysis of regularization-based personalized federated learning (T Dinh et al., 2020; Deng et al., 2020; Hanzely & Richtárik, 2020; Hanzely et al., 2020) and theoretical analysis in federated learning (Chen et al., 2023b; Cheng et al., 2023).
-Assumption 4 is used to quantify the observation noise, and is also standard in the literature (Duan & Wang, 2023; Chen et al., 2023b).
-Next, we establish a statistical rate of the optimal solution to the Fedprox problem (1) with a fine-grained analysis.
-To facilitate the analysis, we first define some notations.
-Define $\{\widetilde{\bm{w}}^{(i)}\}_{i=1}^{m}$ as the optimal solution for local models $\bm{w}^{(i)}$ in objective (1).
-A check on the optimality condition implies that the optimal solution for the global model is $\widetilde{\bm{w}}^{(g)}=\sum^{m}_{i=1}p_{i}\widetilde{\bm{w}}^{(i)}$.
-In this paper, we use global statistical error $\mathbb{E}\|\widetilde{\bm{w}}^{(g)}-{{\bm{w}}_{\star}^{(g)}}\|^{2}$ and local statistical error $\mathbb{E}\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}$ as the measure of statstical accuracy.
-Here the expectation is taken with respect to all data.
-Note that some existing literature adopts alternative metrics, such as average excess risk and individual excess risk (Chen et al., 2023b) to measure the statistical error.
-The convergence rate of these two definitions is closely related up to the smoothness constant and strong convexity constant.
-Define $\tilde{\bm{w}}_{\text{GT}}$ as the optimal solution for the GlobalTrain problem in (2).
-We first provide a lemma to cast light on the effect of $\lambda$ on the statistical error.
+There exists a nonnegative constant $\rho$ such that $\mathbb{E}_{z\sim\mathcal{D}_{i}}\|\nabla\ell({{\bm{w}}_{\star}^{(i)}},z)\|^{2}\leq\rho^{2}$ for all $i\in[m]$.
+The strong convexity and smoothness assumptions are used to establish the estimation error and are widely adopted in the theoretical analysis of regularization-based PFL (T Dinh et al., 2020; Deng et al., 2020; Hanzely & Richtárik, 2020; Hanzely et al., 2020) and FL (Chen et al., 2023c; Cheng et al., 2023).
+Assumption 3 is used to quantify the observation noise, and is also standard in the literature (Duan & Wang, 2023; Chen et al., 2023c).
+For simplicity, we assume $p_{i}=1/m$ and $n_{i}=n$ for all $i\in[m]$.
+The following theorem provides an upperbound of the estimation error $\mathbb{E}\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}$ as a function of $\lambda$.
+Theorem 1.
+Suppose Assumption 1, 2 and 3 hold and consider the parameter space $\mathcal{P}(R)$ given by (5). The local models $\{\widetilde{\bm{w}}^{(i)}\}_{i=1}^{m}$ obtained by solving Problem (2) satisfy
+$\begin{split}&\mathbb{E}\left\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\right\|^{2}\\ &\leq\min\Bigg{\{}\frac{48\rho^{2}}{\mu^{2}}\frac{1}{N}+\left(\frac{48L^{2}}{\mu^{2}}+3\right)R^{2}+\frac{1}{q_{1}(\lambda)},\\ &\frac{4\rho^{2}}{\mu^{2}}\frac{1}{n}+\left[\left(\frac{4\rho^{2}}{\mu^{3}}\frac{1}{n}+\frac{4}{\mu}R^{2}\right)\lambda+\frac{8}{\mu^{2}}R^{2}\lambda^{2}\right]\Bigg{\}},\end{split}$ (8)
+where $N=mn$ and $q_{1}(\lambda)$, defined in (70), is a monotonically increasing function of $\lambda$ with $\lim_{\lambda\to\infty}q_{1}(\lambda)=\infty$.
 See Appendix A.3 for the proof.
-Lemma 1.
-Suppose Assumptions 1, 2 and 3 hold. We have
-$\|\widetilde{\bm{w}}^{(g)}-\tilde{\bm{w}}_{\text{GT}}\|^{2}\leq\frac{4\left(1+L/\mu\right)}{\mu(\mu+\lambda)}\sum_{i\in[m]}p_{i}\|\nabla L_{i}(\widetilde{\bm{w}}^{(g)},S_{i})\|^{2},$ (4)
-$\sum_{i\in[m],j\in[n_{i}]}\frac{p_{i}}{n_{i}}$ $\left(\mu\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}+\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\rangle\right)\leq\frac{\lambda}{2}R^{2}.$ (5)
-Lemma 1 establishes a relationship between the model estimation error and the tuning parameter $\lambda$.
-From the bound (4), we can see that a larger $\lambda$ will push the global model solution $\widetilde{\bm{w}}^{(g)}$ towards the GlobalTrain solution.
-On the other hand, from the bound (5), we can set a smaller $\lambda\propto 1/nR^{2}$ to obtain a rate $O(1/n)$, independent of $R$.
-This suggests that $\lambda$ can be chosen adaptively w.r.t.
-$R$ to achieve the optimal statistical rate.
-A detailed discussion can be found in Appendix A.3.
-In the following theorem, we provide such a choice and establish a state-of-the-art statistical rate.
-Theorem 1.
-Suppose Assumptions 1, 2, 3 and 4 hold and $n_{i}=n,\forall i\in[m]$.
-Consider the FedProx objective in (1) with $p_{i}=\frac{1}{m},\forall i\in[m]$.
-If it takes $\lambda\geq\max\left\{64\kappa^{2}L,\left(2\kappa\vee 5\right)\mu\frac{2L^{2}R^{2}+\rho^{2}/n}{L^{2}R^{2}+\rho^{2}/N}\right\}-\mu$ when $R\leq\frac{1}{\sqrt{n}}$, and $\lambda\leq\frac{\rho^{2}}{n\mu R^{2}}$ when $R>\frac{1}{\sqrt{n}}$, the global model $\widetilde{\bm{w}}^{(g)}$ and local models $\{\widetilde{\bm{w}}^{(i)}\}_{i=1}^{m}$ obtained from solving the objective (1) satisfy
-$\mathbb{E}\left\|\widetilde{\bm{w}}^{(g)}-{{\bm{w}}_{\star}^{(g)}}\right\|^{2}\leq\frac{C_{1}}{N}+C_{2}\left(\frac{1}{n}\wedge R^{2}\right),$ (6)
-$\mathbb{E}\left\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\right\|^{2}\leq\frac{C_{3}}{N}+C_{4}(\frac{1}{n}\wedge R^{2})\quad\text{ forall }i\in[m],$ (7)
-where $N=mn$ and
-$C_{1}=32\frac{\rho^{2}}{\mu^{2}},C_{2}=\big{(}\frac{(1+\sqrt{3})\rho}{\mu}\big{)}^{2}\vee\frac{32L^{2}}{\mu^{2}},C_{3}=\frac{192\rho^{2}}{\mu^{2}},C_{4}=\left(\frac{\rho(1+(4C_{2}+5)^{\frac{1}{2}})}{\mu}\right)^{2}\vee\frac{198L^{2}}{\mu^{2}}.$
-Proof.
-See Appendix A.3. ∎
-Theorem 1 shows that with an appropriate choice of $\lambda$, both the global and local models achieve a fast statistical rate across a wide range of $R$.
-To interpret the result, we first analyze the error bound provided in (6) and (7).
-The two error bounds have the same order and can be decomposed into two key components.
-The first term, of order $\mathcal{O}(N^{-1})$, captures the benefit of collaborative learning, as $N=mn$ is the total sample size across $m$ clients.
-This term reflects the sample efficiency gained through collaborative learning.
-The second term, of order $\mathcal{O}(n^{-1}\wedge R^{2})$, measures the cost of statistical heterogeneity.
-As $R$ increases, the convergence rate gradually degrades, transitioning from $\mathcal{O}(N^{-1})$ to $\mathcal{O}(n^{-1})$.
-Notably, the guarantee is minimax-optimal, as the derived bound matches the problem’s lower bound under the given assumptions (Chen et al., 2023b).
-Remark 2 in Appendix A.3 discusses about how such an interpolation is achieved through different personalization degrees.
-Table 1: Results Comparison.
-$D:=\sup_{\bm{w},\bm{w}^{\prime}\in\mathcal{D}}\|\bm{w}-\bm{w}^{\prime}\|$ is diameter of the parameter space ($D\geq R$) and $\|\ell\|_{\infty}=\inf\{M\in\mathbb{R}:\ell(\cdot,z)\leq M,\text{for any }z\}$ is an uniform upper bound on the loss function.
-Statistical rate refers to $\mathbb{E}\|{{\bm{w}}_{\star}^{(i)}}-\widetilde{\bm{w}}^{(i)}\|^{2}$, with constants neglected.
-Source Assumption Paradigm Statistical Rate
-$R>\frac{1}{\sqrt{n}}$ $\frac{1}{m\sqrt{n}}<R\leq\frac{1}{\sqrt{n}}$ $R\leq\frac{1}{m\sqrt{n}}$
-Chen et al. (2023b) A1,2,3,4, $D\vee\|\ell\|_{\infty}\leq C$ Two-stage $\frac{D^{2}+\|\ell\|_{\infty}}{n}$ 1 $\frac{\|\ell\|_{\infty}}{\sqrt{n}}R$ $\frac{\|\ell\|_{\infty}}{\sqrt{m}n}$
-Ours A1,2,3,4 One-stage $\frac{1}{n}$ $R^{2}\wedge\frac{1}{mn}$ $\frac{1}{mn}$
-Lower Bound - - $\frac{1}{n}$ $R^{2}\wedge\frac{1}{mn}$ $\frac{1}{mn}$
-We now compare our theoretical results with the result up to date (Chen et al., 2023b), as summarized in Table 1.
-For a fair comparison, it is noticeable that we have transformed their result into our metrics $\mathbb{E}\|{{\bm{w}}_{\star}^{(i)}}-\widetilde{\bm{w}}^{(i)}\|^{2}$, by removing their optimization error and problem-dependent constant.
-Our contribution can be summarized into three aspects:
-$\bullet$ Our results are derived under more realistic assumptions.
-Chen et al.
-(2023b) imposes additional stringent conditions including a bounded parameter space $D$ and a uniform bound on the loss function $\ell(\cdot,z)$.
-Such conditions are difficult to satisfy in practice.
-For example, even a simple linear regression model with sub-Gaussian covariates would violate these assumptions.
-$\bullet$ Our result is established for both global and local models with the same choice of $\lambda$.
-In contrast, Chen et al.
-(2023b) proposes a two-stage process where the global model is first estimated, followed by an additional local training phase with a different choice of $\lambda$.
-This two-step approach is likely the artifact of their analysis, as existing work on FedProx also shows that a single-stage optimization with a chosen $\lambda$ suffices to achieve the desired performance (T Dinh et al., 2020).
-$\bullet$ We establish a state-of-the-art minimax statistical rate.
-As shown in Table 1, the established upper bound matches the lower bound established in Chen et al.
-(2023b).
-In a heterogeneous case when $R\geq n^{-1/2}$, Chen et al.
-(2023b) establishes a rate that is at least of the order $\mathcal{O}\left(n^{-1}R^{2}\right)$ as $D\geq R$.
-This bound becomes increasingly loose with larger $R$.
-In contrast, our bound is $\mathcal{O}\left(n^{-1}\right)$, which shows that properly-tuned FedProx is always no worse than LocalTrain, independent of $R$.
-In a homogeneous case when $R\leq m^{-1}n^{-1/2}$, they can only establish a rate slower than $\mathcal{O}\left(1/(\sqrt{m}n)\right)$, while our result achieves a rate of $\mathcal{O}\left(1/(mn)\right)$, leveraging all $mn$ samples and matching the rate of GlobalTrain on the IID data.
-As the client number could be extremely large ($10^{5}$ devices in chen2023fs), order of $m$ is non-trivial.
-During the transition period when $m^{-1}n^{-1/2}<R<n^{-1/2}$, we achieve a rate of $\mathcal{O}(R^{2})$, again matching the lower bound, and is strictly faster than the rate $\mathcal{O}(n^{-1/2}R)$ established in Chen et al.
-(2023b).
-5 When will Personalization Improve Communication Efficiency
-Notice that if we define the local objective of each client $i$ as $h_{i}({{\bm{w}}^{(i)}},{\bm{w}}^{(g)}):={L}_{i}({{\bm{w}}^{(i)}},S_{i})+\frac{\lambda}{2}\|{{\bm{w}}^{(g)}}-{{\bm{w}}^{(i)}}\|^{2}$, then the FedProx problem given by (1) can be rewritten in the following bilevel (iterated minimization) form:
-$\min_{{\bm{w}}^{(g)}}F({\bm{w}}^{(g)}):=\frac{1}{m}\sum_{i=1}^{m}F_{i}({\bm{w}}^{(g)}),\quad\text{where}\quad F_{i}({\bm{w}}^{(g)}):=\min_{{{\bm{w}}^{(i)}}}~{}h_{i}({{\bm{w}}^{(i)}},{\bm{w}}^{(g)}).$ (8)
-The reformulation (8) is a finite-sum minimization problem, with each component $F_{i}$ being the Moreau envelope (Moreau, 1965; Yosida, 1964) of $L_{i}$.
+The bound given by (8) consists of two terms.
+The first term decreases as $\lambda$ increases.
+When $\lambda\to\infty$, the term $[q_{1}(\lambda)]^{-1}\to 0$ and the first term approaches $\mathcal{O}(1/N+R^{2})$.
+Here, $\mathcal{O}(1/N)$ corresponds to the sample complexity one could obtain if the data distributions of the $m$ clients are homogeneous, reflecting the benefit of collaborative training.
+Term $\mathcal{O}(R^{2})$ reflects the negative impact due to the bias introduced by statistical heterogeneity.
+In contrast, the second term in (8) increases with $\lambda$, showing that a smaller $\lambda$ leads to a rate closer to $\mathcal{O}(1/n)$ corresponding to that of pure local training.
+This rate is independent of $R$, making it robust to high data heterogeneity but at the expense of reduced sample efficiency.
+Together, these results provide a continuous and quantitative characterization of how the personalization degree, governed by $\lambda$, determines the statistical accuracy of the solution $\widetilde{\bm{w}}^{(i)}$.
+Minimax-optimal Statistical Accuracy.
+To demonstrate the tightness of our analysis, we now show that as a direct implication of Theorem 1, the local models $\widetilde{\bm{w}}^{(i)}$ are rate-optimal.
+Denoting $\kappa=L/\mu$ as the conditional number, we have the following corollary.
+Corollary 1.
+Suppose Assumption 1, 2 and 3 hold and consider the parameter space $\mathcal{P}(R)$ given by (5).
+If setting $\lambda\geq\max\left\{64\kappa^{2}L,\left(2\kappa\vee 5\right)\mu\frac{2L^{2}R^{2}+\rho^{2}/n}{L^{2}R^{2}+\rho^{2}/N}\right\}-\mu$ when $R\leq\frac{1}{\sqrt{n}}$, and $\lambda\leq\frac{\rho^{2}}{n\mu R^{2}}$ when $R>\frac{1}{\sqrt{n}}$, then for all $i\in[m]$, the local model $\widetilde{\bm{w}}^{(i)}$ obtained by solving Problem (2) satisfy
+$\begin{split}\mathbb{E}\left\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\right\|^{2}\leq C_{3}\frac{1}{N}+C_{4}\left(R^{2}\wedge\frac{1}{n}\right),\quad\end{split}$ (9)
+where constants $C_{3}$ and $C_{4}$ are defined in (91).
+See Appendix A.3 for the proof.
+In addition, Theorem 8 in Chen et al.
+(2023c) states for all $i\in[m]$ and any estimator $\hat{\bm{w}}^{(i)}$ that is a measurable function of data $\{S_{i}\}_{i=1}^{m}$, we have
+$\begin{split}&\inf_{\hat{\bm{w}}^{(i)}}\sup_{\{{{\bm{w}}_{\star}^{(i)}}\}_{i=1}^{m}\in\mathcal{P}(R)}\mathbb{E}\left\|\hat{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\right\|^{2}\gtrsim\frac{1}{N}+R^{2}\wedge\frac{1}{n}.\end{split}$
+Therefore, Corollary 1 shows that with an appropriate choice of $\lambda$, the established rate achieves the minimax lower bound.
+A detailed comparison with prior results can be found in Appendix A.5, along with a discussion of several novel techniques developed to derive our result.
+While this result focuses on statistical accuracy and is not directly tied to the accuracy-communication trade-off, it highlights the tightness of our analysis.
+Furthermore, the techniques introduced to obtain the results are of independent interest.
+To the best of our knowledge, this is the first work to establish minimax-optimality for the solutions of Problem (2).
+Remark 2 (Technical Novelty of The Analysis).
+Unlike previous results (Chen et al., 2023c) which are algorithm-dependent, our analysis does not rely on any algorithm but directly tackles the objective function, establishing rates using purely the properties of the loss.
+Specifically, using the strong convexity of the loss function, we first proved the estimation error is bounded by the gradient norm and an extra statistical heterogeneity term controlled by $\lambda$, as detailed in Equation (33).
+This allows us to show for large $R$, a small $\lambda$ can be chosen to yield rate $O(1/n)$ and match one of the worst cases in the lower bound.
+For the complementary case $R\leq 1/\sqrt{n}$, we leveraged the GlobalTrain solution $\widetilde{\bm{w}}_{GT}$ as a bridge and proved the solutions of Problem (2) to $\widetilde{\bm{w}}_{GT}$ are bounded by a term inversely proportional to $\lambda$ (cf.
+Lemma 4), implying that they can be made arbitrarily close to $\widetilde{\bm{w}}_{GT}$ by setting $\lambda$ small.
+This, together with the rate of $\widetilde{\bm{w}}_{GT}$, yields a rate of $O(1/N+R^{2})$.
+Combining the two cases, we proved minimax optimality.
+Not only each piece of the above results are new, but more importantly, identifying they are the key ingredients to show the solution of (2) is minimax optimal, are the technical novelties of this proof.
+Notice that Theorem 1 and Corollary 1 focus on establishing the statistical rate for the solution in Problem (2), $\widetilde{\bm{w}}^{(i)}$.
+To fully quantify the error of the output of PFL algorithms, we need to establish the optimization error from the algorithm as well, as detailed in the next section.
+4.2 Effect of Personalization on Communication Efficiency
+In this section, we provide an FL algorithm for solving Problem (2) along with its convergence analysis, showing the impact of the personalization degree on communication and computation complexity.
+Notice that if we define the local objective of each client $i$ as
+$h_{i}({{\bm{w}}^{(i)}},{\bm{w}}^{(g)}):={L}_{i}({{\bm{w}}^{(i)}},S_{i})+\frac{\lambda}{2}\|{{\bm{w}}^{(g)}}-{{\bm{w}}^{(i)}}\|^{2},$
+then Problem (2) can be rewritten in the following bilevel (iterated minimization) form:
+$\begin{split}&\min_{{\bm{w}}^{(g)}}F({\bm{w}}^{(g)}):=\frac{1}{m}\sum_{i=1}^{m}F_{i}({\bm{w}}^{(g)}),\\ &\text{where}\quad F_{i}({\bm{w}}^{(g)}):=\min_{{{\bm{w}}^{(i)}}}~{}h_{i}({{\bm{w}}^{(i)}},{\bm{w}}^{(g)}).\end{split}$ (10)
+The reformulation (10) has a finite-sum minimization structure, with each component $F_{i}$ being the Moreau envelope (Moreau, 1965; Yosida, 1964) of $L_{i}$.
 Let ${\bm{w}}_{\star}^{(i)}({\bm{w}}^{(g)})$ be the minimizer of $h_{i}(\,\cdot\,,{\bm{w}}^{(g)})$, i.e.,
-${\bm{w}}_{\star}^{(i)}({\bm{w}}^{(g)}):=\operatorname{prox}_{L_{i}/\lambda}({\bm{w}}^{(g)})=\operatorname*{arg\,min}_{{{\bm{w}}^{(i)}}}~{}h_{i}({{\bm{w}}^{(i)}},{\bm{w}}^{(g)}).$ (9)
-Based on Assumption 2 and the new notations in (8) and (9), the following lemma describes the optimizing properties for $F_{i}$, $h_{i}$ and ${\bm{w}}_{\star}^{(i)}$.
-The proof can be found in Appendix B.2.
+$\begin{split}{\bm{w}}_{\star}^{(i)}({\bm{w}}^{(g)})&=\operatorname{prox}_{L_{i}/\lambda}({\bm{w}}^{(g)})\\ &:=\operatorname*{arg\,min}_{{{\bm{w}}^{(i)}}}~{}h_{i}({{\bm{w}}^{(i)}},{\bm{w}}^{(g)}).\end{split}$ (11)
+The following lemmas provide properties of $F_{i}$, $h_{i}$ and ${\bm{w}}_{\star}^{(i)}$, instrumental to the algorithm design and rate analysis. The proof can be found in Appendix B.2.
 Lemma 1.
 Under Assumption 1 and 2, $F_{i}$ is $\mu_{g}$-strongly convex and $L_{g}$-smooth, with $\mu_{g}=\frac{\lambda\mu}{\lambda+\mu}$ and $L_{g}=\frac{\lambda L}{\lambda+L}$, each $h_{i}$ is $\mu_{\ell}$-strongly convex and $L_{\ell}$-smooth, with $\mu_{\ell}=\mu+\lambda$ and $L_{\ell}=L+\lambda$, and the mapping $\bm{w}_{\star}^{(i)}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is $L_{w}$-Lipschitz with $L_{w}=\frac{\lambda}{\lambda+\mu}$.
-Next, we recall the properties of the Moreau envelope to guide the algorithm design.
 Lemma 2 (Lemaréchal & Sagastizábal (1997)).
 Under Assumption 2, each $F_{i}:\mathbb{R}^{d}\to\mathbb{R}$ is continuously differentiable, and the gradient is given by $\nabla F_{i}({\bm{w}}^{(g)})=\lambda({\bm{w}}^{(g)}-{\bm{w}}_{\star}^{(i)}({\bm{w}}^{(g)})).$
-From Lemma 2 and Lemma 3, it suggests applying a simple gradient algorithm to optimize ${\bm{w}}^{(g)}$:
-$\begin{split}{\bm{w}}_{t+1}^{(g)}&={\bm{w}}_{t}^{(g)}-\gamma\cdot\nabla F({\bm{w}}_{t}^{(g)})={\bm{w}}_{t}^{(g)}-\gamma\lambda\cdot\frac{1}{m}\sum_{i=1}^{m}\left({\bm{w}}_{t}^{(g)}-{\bm{w}}_{\star}^{(i)}({\bm{w}}_{t}^{(g)})\right),\end{split}$ (10)
+Lemma 1 and Lemma 2 suggest applying a simple gradient algorithm to optimize ${\bm{w}}^{(g)}$:
+$\begin{split}{\bm{w}}_{t+1}^{(g)}&={\bm{w}}_{t}^{(g)}-\gamma\cdot\nabla F({\bm{w}}_{t}^{(g)})\\ &={\bm{w}}_{t}^{(g)}-\gamma\lambda\cdot\frac{1}{m}\sum_{i=1}^{m}\left({\bm{w}}_{t}^{(g)}-{\bm{w}}_{\star}^{(i)}({\bm{w}}_{t}^{(g)})\right),\end{split}$ (12)
 where $\bm{w}_{t}^{(g)}$ is the update at the communication round $t$, $\gamma>0$ is a step size to be properly set (cf.
 Theorem 2).
-To implement (10), in each communication round $t$ the server broadcasts ${\bm{w}}_{t}^{(g)}$ to all clients, then each client $i$ updates its local model ${\bm{w}}_{\star}^{(i)}({\bm{w}}_{t}^{(g)})$ and upload to the server.
-Note that executing (10) requires each client $i$ computing the minimizer ${\bm{w}}_{\star}^{(i)}({\bm{w}}_{t}^{(g)})$.
+To implement (12), in each communication round $t$ the server broadcasts ${\bm{w}}_{t}^{(g)}$ to all clients, then each client $i$ updates its local model ${\bm{w}}_{\star}^{(i)}({\bm{w}}_{t}^{(g)})$ and uploads to the server.
+Note that executing the update (12) requires each client $i$ computing the minimizer ${\bm{w}}_{\star}^{(i)}({\bm{w}}_{t}^{(g)})$.
 In general, the subproblem (11) does not have a closed-form solution.
@@
 Per communication round $t$, ${\bm{w}}_{t,k}^{(i)}$ is initialized to be ${\bm{w}}_{t,0}^{(i)}={\bm{w}}_{t-1,K}^{(i)}$ as a warm start, and updated according to
-$\begin{split}{\bm{w}}_{t,k+1}^{(i)}={\bm{w}}_{t,k}^{(i)}-\eta\nabla h_{i}({\bm{w}}_{t,k}^{(i)},{\bm{w}}_{t}^{(g)}),\quad\forall\,k=0,\ldots,K-1,\end{split}$ (11)
-where $\nabla h_{i}({\bm{w}}_{t,k}^{(i)},{\bm{w}}_{t}^{(g)})$, for notation simplicity, denotes the partial gradient of $h_{i}$ with respect to ${\bm{w}}^{(i)}$.
+$\begin{split}{\bm{w}}_{t,k+1}^{(i)}={\bm{w}}_{t,k}^{(i)}-\eta\nabla h_{i}({\bm{w}}_{t,k}^{(i)},{\bm{w}}_{t}^{(g)}),\end{split}$ (13)
+for $k=0,\ldots,K-1$, where $\nabla h_{i}({\bm{w}}_{t,k}^{(i)},{\bm{w}}_{t}^{(g)})$, for notation simplicity, denotes the partial gradient of $h_{i}$ with respect to ${\bm{w}}^{(i)}$.
 The overall procedure is summarized in Alg.
 1, see Appendix B.1.
-Remark 1.
-Alg.
-1 is a customization of the general bi-level problem (Ji et al., 2022).
-While the algorithm design is not new, our contribution lies in analyzing how the personalization degree influences the trade-off between communication and local computation leveraging the FedProx problem structure.
-Extending to noisy settings to consider stochastic gradient or variance reduction would be interesting future work, which requires additional technical effort.
 Theorem 2.
@@
 then ${\bm{w}}_{t}^{(g)}$ converges to $\widetilde{\bm{w}}^{(g)}$ linearly at rate $21-(\gamma\mu_{g})/2-(\gamma\mu_{g})^{2}/$ 2, and ${\bm{w}}_{t,0}^{(i)}$ converges to $\widetilde{\bm{w}}^{(i)}$ linearly at the same rate for any $i\in[m]$.
-Proof.
-See Appendix B.3 and B.5. ∎
-Theorem 2 shows under the step size constraint $\gamma<1/L_{g}$, ${\bm{w}}_{t}^{(g)}$ converges to $\tilde{{\bm{w}}}^{(g)}$ at the same rate as the iteration (10) with exact gradient computation.
-Furthermore, the larger $\gamma$ is, the larger $K$ should be chosen so as to satisfy (12).
-This indicates a tradeoff between communication and computation: the fewer communication rounds are performed, then the more local computation is required for the algorithm to converge.
-To better interpret the result and obtain insights on the influence of $\lambda$ on the convergence, we provide the communication and computation complexity under specific choices the algorithm’s tuning parameters, as described in the following Corollary 1.
-See Appendix B.4 for the proof.
+The proof of Theorem 2 can be found in Appendix B.3 and B.3.2.
+Theorem 2 shows that with sufficiently small step sizes, both the local and global models converge linearly.
+The personalization degree, controlled by $\lambda$, plays a critical role in influencing the algorithm’s communication efficiency.
+A small $\lambda$ (indicating higher personalization) yields a small $L_{g}$, and thus permits a larger choice of the step size $\gamma$.
+By the expression of the convergence rate $21-(\gamma\mu_{g})/2-(\gamma\mu_{g})^{2}/$ 2, we can see less communication round would be required to reach an $\varepsilon$-optimal solution.
+However, increasing $\gamma$ will also decrease the right hand side of (14), implying that a larger number of local gradient steps $K$ should be taken to fulfill the condition.
+To provide a more concrete characterization of these dynamics, we present Corollary 2, which quantifies the communication and computation complexities under specific choices of tuning parameters.
+The proof of Corollary 2 can be found in Appendix B.4.
 Corollary 2.
-In the same setting as Theorem 2, if choosing the step size $\eta=1/L_{\ell}=(\lambda+L)^{-1}$, $\gamma=1/(2L_{g})=\frac{\lambda+L}{2\lambda L}$, and the inner loop iteration number $K=\tilde{\mathcal{O}}\left(\frac{\lambda+L}{\lambda+\mu}\right)$, then the communication complexity and computation complexity for Algorithm 1 to find an $\varepsilon$-solution (i.e., $\|{\bm{w}}_{T}^{(g)}-\tilde{{\bm{w}}}^{(g)}\|^{2}\leq\varepsilon$ and $\|{\bm{w}}_{T,0}^{(i)}-\tilde{{\bm{w}}}^{(i)}\|^{2}\leq\varepsilon$) are as follows ($\kappa:=L/\mu$):
-(i) # communication $\begin{split}&\text{(i) \# communication rounds}=\mathcal{O}\big{(}\kappa_{g}\log(1/\varepsilon)\big{)}=\mathcal{O}\left(\kappa\cdot\frac{\lambda+\mu}{\lambda+L}\cdot\log\frac{1}{\varepsilon}\right),\end{split}$ (13)
-(ii) # gradient $\begin{split}&\text{(ii) \# gradient evaluations}=\tilde{\mathcal{O}}\left(\kappa\cdot\log\frac{1}{\varepsilon}\right).\end{split}$ (14)
-$\bullet$ Communication cost.
-Corollary 1 shows as $\lambda$ increases from 00 to $\infty$, the communication complexity increases from $\tilde{\mathcal{O}}(1)$ to $\tilde{\mathcal{O}}(\kappa)$.
-The result reveals a higher personalization degree requires less communication cost to collaboratively learn the models.
-In the full personalization limit with $\lambda\rightarrow 0$, the order of the communication cost reduces to near constant $\tilde{\mathcal{O}}(1)$, which corresponds to pure local training.
-It is worth noting that even personalization can reduce communication complexity, as our ultimate goal is to accurately estimate each model so we can not simply pursue higher communication efficiency.
-$\bullet$ Computation cost.
-In contrast to the communication complexity, the gradient evaluations per round $K$ decrease with $\lambda$ increasing.
-Therefore, a higher degree of personalization requires a more intensive update of the local models once receiving a new global model ${\bm{w}}^{(g)}$, so as to adapt to each client’s local data set.
-Note that although a small $\lambda$ requires a larger $K$, the overall computation cost given by (LABEL:eq:comp-cost) is independent of $\lambda$.
-As such, model personalization (small $\lambda$) can provably reduce the communication cost without any extra computation overhead.
-[b] Methods Convexity Bounded Domain Computation Cost per Round Communication Cost FedProx (Li et al., 2020a) cvx (1) $\mathcal{O}(\Delta/(\rho\varepsilon))$ pFedMe (T Dinh et al., 2020) s-cvx (2) $\mathcal{O}(1/{(mR\varepsilon)})$ Ditto (Li et al., 2021) s-cvx $\mathcal{O}(1)$ $\mathcal{O}(1/\varepsilon)$ L2GD (Hanzely & Richtárik, 2020) s-cvx $\mathcal{O}(1+\frac{L}{\lambda})$ $\mathcal{O}({\frac{2\lambda\kappa}{\lambda+L}\log(\frac{1}{\varepsilon})})$ Algorithm 3 (Chen et al., 2023b) s-cvx $\mathcal{O}(\lambda/\varepsilon)$ $\mathcal{O}((\lambda\vee 1)/\varepsilon)$ Our Result s-cvx $\mathcal{O}(\frac{\lambda+L}{\lambda+\mu})$ $\mathcal{O}(\kappa\frac{\lambda+\mu}{\lambda+L}\log{\frac{1}{\varepsilon}})$
-Table 2: Comparison of the communication and computation complexity in existing works of analyzing the objective in (1) (cvx denotes convexity and s-cvx denotes strong convexity).
+In the setting of Theorem 2, if we further choose the step size $\eta=(\lambda+L)^{-1}$, $\gamma=(\lambda+L)/(2\lambda L)$, and the inner loop iteration number $K=\tilde{\mathcal{O}}((\lambda+L)/(\lambda+\mu))$, then ${\bm{w}}_{t}^{(g)}$ converges to $\widetilde{\bm{w}}^{(g)}$ linearly at rate $1-(\lambda+L)/(4\kappa(\lambda+\mu))$, and ${\bm{w}}_{t,0}^{(i)}$ converges to $\widetilde{\bm{w}}^{(i)}$ linearly at the same rate for any $i\in[m]$.
+Thus the communication and computation complexity for Algorithm 1 to find an $\varepsilon$-solution (i.e., $\|{\bm{w}}_{T}^{(g)}-\tilde{{\bm{w}}}^{(g)}\|^{2}\leq\varepsilon$ and $\|{\bm{w}}_{T,0}^{(i)}-\tilde{{\bm{w}}}^{(i)}\|^{2}\leq\varepsilon$) are as follows ($\kappa:=L/\mu$):
+(i) # communication $\begin{split}&\text{(i) \# communication rounds}=\mathcal{O}\Big{(}\kappa\cdot\frac{\lambda+\mu}{\lambda+L}\cdot\log\frac{1}{\varepsilon}\Big{)},\end{split}$
+(ii) # gradient $\begin{split}&\text{(ii) \# gradient evaluations}=\tilde{\mathcal{O}}\Big{(}\kappa\cdot\log\frac{1}{\varepsilon}\Big{)}.\end{split}$
+Corollary 2 clearly shows the influence of the personalization degree $\lambda$ on the communication cost.
+Specifically, as $\lambda$ increases from 00 to $\infty$, the communication complexity increases from $\mathcal{O}(\log(1/\epsilon))$ to $\mathcal{O}(\kappa\log(1/\epsilon))$.
+This fact corroborates our intuition that a higher degree of collaboration requires more communication resources.
+Moreover, the result given by Corollary 2 also indicates that the total computation cost is independent of $\lambda$.
+As such, we can see that model personalization can provably reduce the communication cost without any extra computation overhead.
+Notice that although personalization indeed enhances communication efficiency, doing so may deteriorate the statistical accuracy of the local models.
+In the next section, we will discuss in detail the trade-off between communication and statistical accuracy.
+We also provide an extension of Alg.
+1 to the stochastic setting using mini-batch stochastic gradients for the local updates.
+The proposed algorithm, termed FedCLUP, is given in Appendix B.5.
+Following similar argument to Theorem 2, Theorem 3 in Appendix B.5 establishes the convergence rate of FedCLUP, showing that it takes $\mathcal{O}\left(\kappa\frac{\lambda+\mu}{\lambda+L}\cdot\log\frac{1}{\varepsilon}\right)$ rounds of communication to obtain an output ${\bm{w}}_{t}^{(g)}$ within an error ball of size $\left(\mathcal{O}(\varepsilon)+\frac{\sigma^{2}}{\mu^{2}Bm}\right)$ around $\widetilde{\bm{w}}^{(g)}$.
+Here $\sigma^{2}$ is the variance of the stochastic gradient defined in Assumption 4 in Appendix B.5 and $B$ is the mini-batch size.
+Therefore, the influence of $\lambda$ on communication complexity is consistent with the noiseless setting presented in Corollary 2.
+Additionally, one may note that in the stochastic setting, the error due to the stochastic noise scales inversely with the number of participating clients, showing the advantage of client collaboration in reducing the variance of ${\bm{w}}_{t}^{(g)}$.
+5 Communication-Accuracy Trade-off
+Building on the results established in Section 4.1 and 4.2, we cast insights on the trade-off between communication efficiency and statistical accuracy, and discuss its implications on real-world practice.
+Corollary 3.
+Under the conditions of Theorem 1 and Corollary 2, ${\bm{w}}_{T,K}^{(i)}$ generated by Algorithm 1 satisfies
+$\begin{split}&\mathbb{E}\left\|{\bm{w}}_{T,K}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\right\|^{2}\\ &\leq 2\left(1-\frac{1}{4\kappa}-\frac{L-\mu}{4\kappa(\lambda+\mu)}\right)^{T}\mathbb{E}\left\|{\bm{w}}_{0,0}^{(i)}-\widetilde{\bm{w}}^{(i)}\right\|^{2}\\ &+\mathcal{O}\left[\left(\frac{1}{N}+R^{2}+\frac{1}{q_{1}(\lambda)}\right)\wedge\left(\frac{1}{n}+q_{2}(\lambda)\right)\right]\end{split}$ (15)
+for any $i\in[m]$, where $q_{1}(\lambda)$ and $q_{2}(\lambda)$, defined in (70) and (67), respectively, are monotonically increasing functions of $\lambda$ with $\lim_{\lambda\to\infty}q_{1}(\lambda)=\infty$ and $\lim_{\lambda\to 0}q_{2}(\lambda)=0$.
+Methods Convexity Bounded Domain Computation Cost per Round Communication Cost Statistical Error
+FedProx (Li et al., 2020) cvx (1) $\mathcal{O}(\Delta/(\rho\varepsilon))$
+pFedMe (T Dinh et al., 2020) s-cvx (2) $\mathcal{O}(1/{(mR\varepsilon)})$
+L2GD (Hanzely & Richtárik, 2020) s-cvx $\mathcal{O}(1+\frac{L}{\lambda})$ $\mathcal{O}({\frac{2\lambda\kappa}{\lambda+L}\log(\frac{1}{\varepsilon})})$
+AL2SGD (Hanzely et al., 2020) s-cvx $\mathcal{O}\left(\frac{(m+\sqrt{m(L+\lambda)/\mu})}{\sqrt{\min\{L,\lambda\}/\mu}}\right)$ $\mathcal{O}(\sqrt{\min\{L,\lambda\}/\mu}\log\frac{1}{\varepsilon})$
+Algorithm 3 (Chen et al., 2023c) s-cvx $\mathcal{O}(\lambda/\varepsilon)$ $\mathcal{O}((\lambda\vee 1)/\varepsilon)$
+FedCLUP s-cvx $\mathcal{O}(\frac{\lambda+L}{\lambda+\mu})$ $\mathcal{O}(\kappa\frac{\lambda+\mu}{\lambda+L}\log{\frac{1}{\varepsilon}})$
   • (1)
@@
     Related to the subiteration number $R$ and the precision $v$.
-We end this section with a comparison between our results and the existing algorithms for solving FedProx, as reported in Table 2.
-Our result shows how the personalization degree trades off the communication cost and local computation cost, whereas most of the existing works either don’t explicitly analyze the local computational cost (Li et al., 2020a; T Dinh et al., 2020), or fail to demonstrate such a trade-off (Chen et al., 2023b).
-Hanzely & Richtárik (2020) also reveals similar tradeoffs, but their work is a randomized method that requires more coordination between server and clients, and more importantly, when $\lambda$ is small, the computation cost per round is unbounded.
-6 Simulation
-In this section, we provide empirical validation for our theoretical results, evaluating FedProx on both synthetic and real datasets with convex and non-convex loss functions.
-We first examine the effect of personalization on statistical accuracy under different levels of heterogeneity (Section 6.2), followed by an analysis of its impact on communication efficiency (Section 6.3).
-Detailed descriptions of the experimental setup are available in Section 6.1 and Appendix C, and the complete anonymized codebase is accessible at https://anonymous.4open.science/r/fedprox-bilevel.
+Table 1: Comparison of the existing works analyzing Problem (2) (cvx denotes convexity and s-cvx denotes strong convexity).
+See Appendix C for the proof.
+When collaborative learning is beneficial, i.e., when $R^{2}\lesssim 1/n$, in this case, Corollary 1 establishes that the minimax-optimal statistical rate is $\mathcal{O}(1/N+R^{2})$.
+As $\lambda$ increases, $[q_{1}(\lambda)]^{-1}$ monotonically decreases to zero, leading to improved statistical accuracy, approaching the optimal accuracy.
+However, since the optimization error in (15) is monotonically increasing with $\lambda$, a higher degree of collaboration will also results in slower convergence over communication rounds.
+Consequently, Corollary 3 implies that when the data heterogeneity is relatively low, increasing personalization will improve communication efficiency but at the expense of lower statistical accuracy, and vice versa.
+Such an opposite effect of the personalization degree on the statistical accuracy and communication efficiency leads to an accuracy-communication trade-off.
+As the optimization error vanishes when $T\to\infty$, the statistical error dominates the total error in Corollary 3.
+However, the associated communication cost becomes prohibitively large to achieve such an error.
+Therefore, balancing the optimization and statistical errors is crucial for achieving a target total error with efficient resource usage.
+Under the influence of the personalization degree, a practical guideline emerges: when collaborative learning is beneficial (i.e., $R^{2}\leq 1/\sqrt{n}$), to achieve a target magnitude of total error, we control the optimization errors at the same magnitude as the statistical error.
+Specifically, we increase $\lambda$ to reduce the statistical error to the same magnitude of the total error, and then gradually increase the communication rounds $T$ until the optimization error reaches the same magnitude as well.
+Since the communication efficiency will decrease will an increasing $\lambda$, the above strategy can achieve a given target total error at the highest possible communication efficiency.
+Later, we will also show that, with the optimal choice of personalization degree, FedCLUP improves communication efficiency significantly compared with GlobalTrain and achieve a total error smaller than LocalTrain in Section 6.
+Comparison with Existing Literature.
+As reported in Table 1, our work is the first to quantitatively characterize how changing the personalization degree leads to the trade-off between communication cost and statistical accuracy.
+In contrast, most prior studies either do not explicitly analyze statistical convergence or fail to provide a tight convergence guarantee for optimization and statistical error.
+For example, Chen et al.
+(2023c) attempted to establish both optimization and statistical rates simultaneously; however, when $\lambda\to\infty$, Problem (2) reduces to GlobalTrain, yet their results incorrectly imply that the communication and computation costs diverge to infinity.
+6 Empirical Study
+In this section, we provide empirical validation for our theoretical results, evaluating Problem (2) on both synthetic and real datasets with convex and non-convex loss functions.
+We first present the experimental setup, and then analyze the impact of personalization on statistical accuracy and communication efficiency, and conclude by analyzing the trade-off between the two.
+Details about the experimental setup are available in Section 6.1 and Appendix D, and the complete anonymized codebase is accessible at https://github.com/ZLHe0/fedclup.
 6.1 Experimental Details
 Synthetic Dataset.
-To generate the synthetic data, we follow a similar procedure to prior works (Li et al., 2020a) but with some modifications to align with the setup in this paper.
+We generate two cases of synthetic dataset: (1)As our theoretical analysis is established under strong convexity, first, we consider an overdetermined linear regression task, where the choice of hyparamter is strictly followed Corollary 3.
+(2) We follow a similar procedure to prior works (Li et al., 2020) but with some modifications to align with the setup in this paper.
 Specifically, for each client we generate samples $(\bm{X}_{k},\bm{y}_{k})$, where the labels $\bm{y}_{k}$ are produced by a logistic regression model $\bm{y}_{k}=\text{argmax}(\sigma({\bm{w}}_{k}^{\top}\bm{X}_{k}))$, with $\sigma$ being the sigmoid function.
 The feature vectors $\bm{X}_{k}$ are drawn from a multivariate normal distribution $\mathcal{N}(\bm{v}_{k},\bm{\Sigma})$, where $v_{k}\sim\mathcal{N}(0,1)$ and the covariance matrix $\bm{\Sigma}$ follows a diagonal structure $\bm{\Sigma}_{j,j}=j^{-1.2}$.
-The heterogeneity is introduced by sampling the model weights ${\bm{w}}_{k}$ for each client from a normal distribution $\mathcal{N}(0,R)$, where $R\in[0,3]$ controls the degree of variation (statistical heterogeneity) across clients’ data distributions.
+The heterogeneity is introduced by sampling the model weights ${\bm{w}}_{k}$ for each client from a normal distribution $\mathcal{N}(0,R)$, where $R\in[0,3]$ controls the statistical heterogeneity across clients’ data.
+Dataset LocalTrain FedCLUP pFedMe GlobalTrain
+High Medium Low High Medium Low
+MNIST Logit 0.7828 0.8123 0.8213 0.8312 0.8001 0.8239 0.8291 0.8391
+MNIST CNN2 0.8194 0.8753 0.8893 0.9032 0.8756 0.8771 0.8749 0.9098
+MNIST CNN5 0.7633 0.9091 0.9102 0.9421 0.9340 0.9385 0.9440 0.9433
+EMNIST CNN5 0.6111 0.6278 0.6415 0.6672 0.6345 0.6532 0.6717 0.6611
+CIFAR10 CNN3 0.6238 0.6102 0.6712 0.7302 0.6744 0.7443 0.7732 0.8041
+Table 2: Test accuracy of algorithms across datasets under varying personalization degrees.
+“Logit” refers to logistic regression, while “CNN-k” denotes a convolutional neural network with k convolutional layers.
+“Low,” “Medium,” and “High” indicate low, moderate, and high personalization levels, respectively.
+As the personalization degree decreases, the statistical accuracy of FedCLUP and pFedMe transitions from resembling LocalTrain to GlobalTrain.
 Real Dataset.
-We use the MNIST, EMNIST, and CIFAR10 datasets for the real dataset analysis.
-We first formulate the problem as a convex classification task using multinomial logistic regression, then test the generalization of our theory to a more general, non-convex scenario by conducting experiments using different CNN classifiers.
-Following Li et al.
-(2020a), to impose statistical heterogeneity, we distribute the data across clients in the way that each client only has access to a fixed number of classes.
+We use the MNIST, EMNIST, CIFAR10, Sent140 an CelebA datasets for real data analysis.
+For MNIST, EMNIST and CIFAR10, they aren’t naturally partitioned datasets.
+Therefore, following Li et al.
+(2020), to impose statistical heterogeneity, we distribute the data across clients in a way that each client only has access to a fixed number of classes.
 The fewer classes each client has access to, the higher the statistical heterogeneity.
-The detail of data processing to set data heterogeneity for each dataset can be found in Appendix C.
+Details on data preprocessing and heterogeneity settings for each dataset are provided in Appendix D.
+For Sent140 and CelebA datasets, as they are naturally partitioned, we don’t need to create the data heterogeneity and directly apply our algorithm to these datasets with advanced models, like CNN and LSTM model (Sak et al., 2014).
+We strictly follow the previous work (Li et al., 2021) and (Duan et al., 2021) to pre-process the Sent140 and CelebA dataset, respectively.
 Implementation and Evaluation.
-For each setup, we evaluate the FedProx algorithm with three different choices of $\lambda$: small, medium, and large, with specific values varying based on the dataset (details provided in Appendix C).
-For the synthetic dataset, in alignment with theorem 2, we implement Algorithm 1 using a global step size $\gamma=(\lambda+L)/(\lambda L)$ and a local step size $\eta=(L+\lambda)^{-1}$.
-For the real dataset, since $L$ is unknown, we implement the same algorithm with a global step size $\gamma$ set to $1/(\lambda)$, while the local step size is determined via grid search.
-In terms of evaluation, for the synthetic dataset, we evaluate by tracking the ground-truth models and measuring error as the distance from these true models.
-For the real dataset, we report training loss and testing accuracy.
-Additional details, including hyperparameter setting and evaluation metric definition, can be found in Appendix C.
-Figure 1: Statistical and total error of different methods.
-In the top row, subfigures (a), (b), and (c) represent statistical error, showing the difference between the minimizer and the underlying true model for Synthetic-Logit (logistic regression on synthetic data), MNIST-Logit (logistic regression on MNIST), and MNIST-CNN (CNN on MNIST), respectively.
-In the bottom row, subfigures (d), (e), and (f) represent total error, showing the error between the model at each round and the true model for the same datasets.
-(1) FedProx adapts between GlobalTrain and LocalTrain, achieving low error across varying levels of heterogeneity with a properly chosen $\lambda$; (2) Larger $\lambda$ (less personalization) requires more communication to converge, and vice versa.
-6.2 Statistical Aspect: Balancing Local and Global Learning
-We now explore how FedProx balances between GlobalTrain and LocalTrain by achieving different levels of personalization.
-For comparison, in addition to implementing FedProx with different choices of $\lambda$, we also implemented LocalTrain and GlobalTrain, with details specified in Appendx C.
-We run all methods to a stable convergence and measure the statistical error.
-The top row of Figure 1 shows the performance of these methods under different impact of statistical heterogeneity.
-From Figure 1(a), in a controlled convex setting, we observe that FedProx consistently interpolates between GlobalTrain and LocalTrain.
-As $\lambda$ increases, FedProx transitions from behaving similarly to LocalTrain towards GlobalTrain.
-When statistical heterogeneity is low, both GlobalTrain and FedProx with larger $\lambda$ values achieve better statistical performance.
-However, as heterogeneity increases, LocalTrain and FedProx with smaller $\lambda$ values show better results.
-The figure also demonstrates that with an adaptive choice of $\lambda$ in theoretical analysis outlined in Theorem 1, FedProx consistently achieves a low error across varying levels of heterogeneity.
-See more results and discussion in Appendix C.
-Note that as statistical heterogeneity increases, LocalTrain experiences a slight rise in error due to the increased difficulty of the statistical problem.
-However, it remains the least affected by heterogeneity since it relies solely on local data.
-In Figure 1(b) and 1(c), for the MNIST dataset, which includes both convex logistic regression and non-convex CNN settings, we observe similar patterns.
-FedProx again strikes a balance between GlobalTrain and LocalTrain, adapting its performance as $\lambda$ changes.
-While GlobalTrain performs best in this scenario due to the relatively weak statistical heterogeneity, FedProx, with increasing $\lambda$, moves closer to GlobalTrain performance.
-In Figure 4 in Appendix C, we also evaluate the local model’s statistical error and observe that as statistical heterogeneity increases, FedProx with smaller $\lambda$ values can surpass GlobalTrain.
-Figure 2: Effect of local updates on FedProx performance.
-Smaller $\lambda$ reduces communication with more local updates, while larger $\lambda$ benefits marginally from increasing local updates and requires more frequent communication.
-6.3 Optimization Aspect: Communication-Computation Trade-off
-In this section, we explore the trade-off between communication and computation in FedProx, controlled by varying the personalization degree.
-Communication Cost.
-In the bottom row of Figure 1, we show the change in the total error of the algorithm (i.e., the difference of the model in $t$th round between the true underlying model) over communication rounds.
-The first observation is that as $\lambda$ increases, the convergence rate w.r.t.
-communication rounds slows across all settings.
-FedProx with a large $\lambda$ (closer to GlobalTrain) needs the most communication rounds to converge, while smaller $\lambda$ values (closer to LocalTrain) result in faster convergence.
-This result aligns with our theory (c.f.
-Corollary 1), showing that Fedprox with less personalization relies more heavily on communication.
-On the statistical side, Figure 1(d) shows that when statistical heterogeneity is low, larger $\lambda$ leads to the smallest total error.
-As $R$ increases, smaller $\lambda$ (more personalization) results in better performance.
-Computation Cost.
-In Figure 2, we investigate how the number of local updates per communication round impacts the performance of FedProx at different personalization degree.
-Additional results in the non-convex setting can be found in Appendix C.
-Notice that for models with smaller $\lambda$ (higher personalization), increasing the number of local updates significantly improves convergence, making the algorithm less dependent on frequent communication.
-This aligns with our analysis: more personalized models rely less on communication and benefit more from local computation.
+For each setup, we evaluate FedCLUP with three different degree of personalization: low, medium, and high, with specific personalization degree varying based on the dataset (details provided in Appendix D).
+For the synthetic dataset, in alignment with theorem 2, we implement FedCLUP using a global step size $\gamma=(\lambda+L)/(\lambda L)$ and a local step size $\eta=(L+\lambda)^{-1}$.
+For the real dataset, since $L$ is unknown, we implement the same algorithm with a global step size $\gamma$ set to $1/\lambda$, while the local step size is determined via grid search.
+In terms of evaluation, for the synthetic dataset, we evaluate by tracking the ground truth models and measuring error as the distance from these ground true models.
+For the real datasets, we report training loss and testing accuracy.
+Additional details, including hyperparameter settings and evaluation metric definition, can be found in Appendix D.
+Figure 1: Total error and training loss of FedCLUP with varying personalization degrees.
+Total error (y-axis in (a)) quantifies the distance between the model at each communication round and the ground truth model, i.e.
+$\|{\bm{w}}_{T,K}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}$.
+(a) compares total error under low ($R=0.5$) and high ($R=2.0$) heterogeneity.
+(b) presents training loss across different datasets and models under low heterogeneity.
+Synthetic-Logit and MNIST-Logit represent logistic regression on synthetic and MNIST data, respectively, while MNIST-CNN, CIFAR10-CNN, and EMNIST-CNN represent CNN models trained on MNIST, CIFAR-10, and EMNIST datasets.
+In a low-heterogeneity setting, higher personalization (high per) accelerates convergence, while lower personalization (low per) improves final error.
+In a high-heterogeneity setting, higher personalization achieves both smaller error and faster convergence.
+6.2 Results
+Effect of Personalization on Statistical Accuracy.
+We compare FedCLUP, under different personalization degrees, with GlobalTrain, LocalTrain and pFedMe (T Dinh et al., 2020), an alternative algorithm for solving Problem (2).
+Implementation details are provided in Appendix D.
+All methods are run until stable convergence, and their test accuracy is reported in Table 2 in the low heterogeneity setting.
+Table 2 shows that as the personalization degree decreases and collaborative learning increases, the solution of Problem (2) becomes closer to GlobalTrain, leading to improved statistical accuracy due to increased information sharing across clients.
+A similar trend is observed for both FedCLUP and pFedMe across different datasets and models, showing that under low heterogeneity, increased collaboration improves statistical accuracy.
+Figure 2: Effect of personalization on communication efficiency in FedCLUP for Synthetic-Logit (left) and MNIST-Logit (right).
+Each subfigure compares the impact of varying the number of local updates per communication round under different personalization levels.
+Effect of Personalization on Communication Efficiency.
+Figure 2 investigates how personalization impacts communication efficiency in FedCLUP by analyzing the benefit of increasing local updates under different personalization levels.
+In the low-personalization setting (left column), more local updates significantly accelerate convergence, reducing the reliance on frequent communication.
+However, in the high-personalization setting (right column), increasing local updates has a limited effect on convergence, indicating that frequent communication is essential for effective learning.
+This demonstrates that higher personalization requires more communication rounds to achieve comparable performance, aligning with our theoretical findings, showing that increased personalization improves communication efficiency.
+The Trade-off under Personalization.
+The above two observations imply that there shall be a accuracy-communication trade-off in choosing the personalization degree to achieve the smallest total error under given communication budget (See Figure 1).
+To further demonstrate this point, in the left part of Figure 3, we compare the total error over iterations among LocalTrain, GlobalTrain, and FedCLUP under varying levels of personalization.
+Once the algorithms converge, for any given total error, there exists a specific degree of personalization in FedCLUP that incurs the least communication cost among all considered methods.
+This observation supports the conclusion in Corollary 3 that a unique personalization degree exists in FedCLUP to achieve communication efficiency for a fixed total error.
+Furthermore, we observe that no fixed personalization degree consistently outperforms others across all error levels.
+This implies that adaptively adjusting the personalization degree is essential for FedCLUP to maintain communication efficiency across different total error regimes.
+To further validate this finding, we conduct the same experiment on real-world datasets.
+Results on the CelebA and Sent140 datasets are presented in Figures 9 and 10, respectively.
+Additionally, we observe a similar phenomenon in pFedMe, another method solving Problem 2, when we vary its personalization degree.
+Witnessing the insightful results, it motivates us to adaptively change the personalization degree in FedCLUP for achieving communication efficiency over different total errors.
+Potential Solution in Practice.
+After understanding the trade-off under personalization, we propose a dynamic tuning strategy for the optimal $\lambda$ with communication efficiency(See Figure 3 Right ).
+Beginning with a small $\lambda$ for efficiency, as soon as the validation performance plateaus or the statistical error stops improving significantly, we gradually increase $\lambda$.
+This allows the model to benefit from enhanced generalization through increased collaboration at the expense of slightly higher communication costs.
+By progressively adjusting in this way, one can finally identify the optimal personalization degree that balances the trade-off and minimizes the total communication cost required to meet a desired performance threshold.
+Figure 3: Total error versus communication rounds.
+Left: Comparison among LocalTrain, GlobalTrain, and FedCLUP with varying levels of personalization.
+For a given target total error, an optimal $\lambda$ exists that has the minimum communication rounds.
+Right: A dynamic $\lambda$ strategy: we begin with local training (highest personalization degree), and once the decrease of the validation error slows down, we gradually decrease the personalization degree.
+Comparing the right figure with the left one shows that the dynamic strategy approximates the optimal personalization level needed to meet a desired error threshold with minimal communication rounds, demonstrating the practical effectiveness of tuning personalization dynamically according to the accuracy-communication trade-off.
 7 Conclusion
-In this paper, we analyzed of the effect of personalization in FedProx.
-With a theoretically-informed adaptive choice of personalization degree, FedProx achieves a minimax-optimal statistical accuracy under different statistical heterogeneity degrees and consistently outperforms pure local training.
-As a minor contribution, we provably show that personalization reduces communication costs without increasing the computation overhead.
-The results are validated through extensive experiments.
+In this paper, we provide a precise theoretical characterization of the statistical and optimization convergence of a widely used personalized federated learning problem.
+Our analysis reveals that when collaborative learning is beneficial, increasing personalization reduces communication complexity but comes at the cost of statistical accuracy due to limited information sharing across clients.
+We then validate our theoretical findings across convex and non-convex settings, multiple datasets, and different model architectures.
+Acknowledgment
+This work was supported in part by the U.S. National Science Foundation (NSF) grant CCF-2007823 and National Institutes of Health (NIH) grants R01AI170249 and R01GM152812.
+Impact Statement
+This paper aims to contribute to the advancement of the field of Machine Learning.
+While our work has the potential for various societal implications, we believe none require specific emphasis in this context.
 Appendix A Proof of statistical converngence
@@
 $\tilde{\Delta}^{(i)}_{\text{stat}}={{\bm{w}}_{\star}^{(i)}}-\widetilde{\bm{w}}^{(i)}\quad\forall i\in[m],$ (17)
-where $\tilde{\Delta}^{(g)}_{\text{stat}}$ is the difference between the true global model and the optimal solution of the global model in (1) and $\tilde{\Delta}^{(i)}_{\text{stat}}$ is the difference between the true local model and the optimal solution of the local model in (1) for any $i\in[m]$.
+where ${{\bm{w}}_{\star}^{(g)}}:=\sum_{i\in[m]}p_{i}{{\bm{w}}_{\star}^{(i)}}$, $\tilde{\Delta}^{(g)}_{\text{stat}}$ is the difference between the true global model and the optimal solution of the global model in Problem (2) and $\tilde{\Delta}^{(i)}_{\text{stat}}$ is the difference between the true local model and the optimal solution of the local model in (2) for any $i\in[m]$.
 The statistical error bound is established as $\mathbb{E}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}$ and $\mathbb{E}\|\tilde{\Delta}^{(i)}_{\text{stat}}\|^{2}$, where the expectation is taken over all the data.
-And we also denote
+Notably, we should stress that the definition of ${{\bm{w}}_{\star}^{(g)}}$ is just to serve as a bridge to establish the convergence rate of local models, and it is not our focus in the theoretical analysis.
+We also denote
 $\bm{\delta}_{\star}^{(i)}:={{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}},\quad\tilde{\bm{\delta}}^{(i)}:={\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}},$ (18)
 where $\bm{\delta}_{\star}^{(i)}$ measures the difference between the true global model ${{\bm{w}}_{\star}^{(g)}}$ and the true local model ${{\bm{w}}_{\star}^{(i)}}$, and $\tilde{\bm{\delta}}^{(i)}$ is the estimator of such a difference.
-A.1 Discussion on Assumption 1
+A.1 Discussion on the Parameter Space
 When assuming the statistical heterogeneity of each client is different ($R$ is related to the client index $i$), it would provide a more delicate description of statistical heterogeneity, and it also requires different personalization degree per client.
-Therefore, it is better to consider the objective (18), where the $i$-th client will be shrunk to the global with strength $\lambda_{i}$.
+Therefore, let’s consider the following problem with different heterogeneity in personalized federated learning
+$\min_{{{\bm{w}}^{(g)}},\{{{\bm{w}}^{(i)}}\}_{i\in[m]}}\sum_{i=1}^{m}p_{i}\left({L}_{i}({{\bm{w}}^{(i)}},S_{i})+\frac{\lambda_{i}}{2}\|{{\bm{w}}^{(g)}}-{{\bm{w}}^{(i)}}\|^{2}\right),$ (19)
+where the $i$-th client will be shrunk to the global with strength $\lambda_{i}$.
 Therefore, solving (19) will consider different personalization degree for different clients.
-Solving the objective (18) would be open for future work.
-$\min_{{{\bm{w}}^{(g)}},\{{{\bm{w}}^{(i)}}\}_{i\in[m]}}\sum_{i=1}^{m}p_{i}\left({L}_{i}({{\bm{w}}^{(i)}},S_{i})+\frac{\lambda_{i}}{2}\|{{\bm{w}}^{(g)}}-{{\bm{w}}^{(i)}}\|^{2}\right),$ (18)
-Under Assumptions 2 and 3 as we assume, Assumption 1 about statistical heterogeneity would be equivalent to many other similar assumptions for statistical heterogeneity, like $B$-dissimilarity (Li et al., 2020a), parameter difference (Chen et al., 2023b) and gradient diversity (T Dinh et al., 2020; Deng et al., 2020).
-A.2 Lemmas and the Proof
-To facilitate the proof of the theorem, we first present two additional lemmas and provide their proof.
+Establish the minimax statistical error bound for Problem (19) would be open for future work.
+Under Assumptions 1 and 2 as we assume, statistical heterogeneity in parameter space (5) would be equivalent to many other similar assumptions for statistical heterogeneity, like $B$-dissimilarity (Li et al., 2020), parameter difference (Chen et al., 2023c) and gradient diversity (T Dinh et al., 2020; Deng et al., 2020).
+A.2 Useful Lemmas
+To facilitate our analysis, we first present two important lemmas and provide their proof.
 Lemma 3.
@@
 $\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\|_{2}\leq\frac{2\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})\|_{2}}{\mu+\lambda}.$
+Proof.
 By the $\mu$-strongly convexity of $\nabla L_{i}$, we have that for any ${\bm{w}}\in\mathbb{R}^{d}$,
@@
 $\geq L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})+\|{\bm{w}}-{\tilde{\bm{w}}^{(g)}}\|\left(\frac{\mu+\lambda}{2}\|{\bm{w}}-{\tilde{\bm{w}}^{(g)}}\|-\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})\|\right),$ (20)
-where in the last step we used the Cauchy inequality.
-In addition, notice that ${\tilde{\bm{w}}^{(i)}}$ is the optimal solution for the local model ${{\bm{w}}^{(i)}}$ in the FedProx problem (1), thus we have
+where in the last step, we used the Cauchy inequality.
+In addition, notice that ${\tilde{\bm{w}}^{(i)}}$ is the optimal solution for the local model ${{\bm{w}}^{(i)}}$ in Problem (2), thus we have
 ${\tilde{\bm{w}}^{(i)}}=\underset{{\bm{w}}}{\operatorname{argmin}}\left\{L_{i}({\bm{w}},S_{i})+\frac{\lambda}{2}\|{\bm{w}}-{\tilde{\bm{w}}^{(g)}}\|^{2}\right\}.$
@@
 $\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\|\leq\frac{2\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})\|}{\mu+\lambda},$ (21)
-as claimed.
+as claimed. ∎
+Remark: Lemma 3 is crucial to bridge the solution of the global model with the local model in Problem 2.
+Note that the difference is bounded by the gradient norm of local loss functions evaluated at the optimal solution of the global model, which would be decomposed further and controlled by the global model’s statistical accuracy, as we will show later.
 Lemma 4.
-Under Assumption 1, 2, 3, 4, for $\tilde{\bm{w}}_{\text{GT}}$, the solution of the GlobalTrain problem in (2), we have
+Under Assumption 1, 2, 3, for $\tilde{\bm{w}}_{\text{GT}}$, the solution of the GlobalTrain problem in (4), we have
 $\mathbb{E}\|\tilde{\bm{w}}_{\text{GT}}-{{\bm{w}}_{\star}^{(g)}}\|\leq\frac{\rho\sqrt{\sum_{i\in[m]}\frac{p_{i}^{2}}{n_{i}}}+LR}{\mu/2}.$
+Proof.
 By the optimality condition of the GlobalTrain problem in (4) and the strong convexity of $L_{i}$’s, we have
@@
 $\|\tilde{\bm{w}}_{\text{GT}}-{{\bm{w}}_{\star}^{(g)}}\|$ $\leq\frac{\Big{\|}\sum_{i\in[m]}p_{i}\nabla L_{i}({{\bm{w}}_{\star}^{(g)}},S_{i})\Big{\|}}{\mu/2}$
-$\leq\frac{\Big{\|}\sum_{i\in[m]}p_{i}\nabla L_{i}({{\bm{w}}_{\star}^{(i)}},S_{i})\Big{\|}+\sum_{i\in[m]}p_{i}\Big{\|}\nabla L_{i}({{\bm{w}}_{\star}^{(g)}},S_{i})-\nabla L_{i}({{\bm{w}}_{\star}^{(i)}},S_{i})\Big{\|}}{\mu/2}$ (21)
+$\leq\frac{\Big{\|}\sum_{i\in[m]}p_{i}\nabla L_{i}({{\bm{w}}_{\star}^{(i)}},S_{i})\Big{\|}+\sum_{i\in[m]}p_{i}\Big{\|}\nabla L_{i}({{\bm{w}}_{\star}^{(g)}},S_{i})-\nabla L_{i}({{\bm{w}}_{\star}^{(i)}},S_{i})\Big{\|}}{\mu/2}.$ (22)
 By the $L$-smoothness property of $L_{i}$’s in Assumption 1 and the bounded gradient property listed in Assumption 3, we have
@@
 $\mathbb{E}\|\tilde{\bm{w}}_{\text{GT}}-{{\bm{w}}_{\star}^{(g)}}\|\leq\frac{\rho\sqrt{\sum_{i\in[m]}\frac{p_{i}^{2}}{n_{i}}}+LR}{\mu/2}$
-as claimed.
-A.3 Proof of Lemma 1 and Theorem 1
-A.3.1 Proof of Lemma 1
-Lemma 1 is a direct consequence of the proof of Theorem 1. To see this, from (A.3.2) and the fact that $(\mu+\lambda)^{2}\geq\mu(\mu+\lambda)$, we have
-$\|{\tilde{\bm{w}}^{(g)}}-\tilde{\bm{w}}_{\text{GT}}\|^{2}$ $\leq\frac{4}{\mu}\left(\frac{1}{\mu+\lambda}+\frac{L}{(\mu+\lambda)^{2}}\right)\sum_{i}p_{i}\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})\|^{2}$
-$\leq\frac{4(1+L/\mu)}{\mu(\mu+\lambda)}\sum_{i}p_{i}\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})\|^{2},$
-thus (4) holds. In addition, with the fact that $\|\tilde{\bm{\delta}}^{(i)}\|_{2}\geq 0$, from (28) we find
-00 0 $\geq-\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle+\frac{\mu}{2}\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\sum_{i}p_{i}\left\|\tilde{\bm{\delta}}^{(i)}\right\|^{2}-\frac{\lambda}{2}R^{2}$
-$\geq-\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle+\frac{\mu}{2}\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}-\frac{\lambda}{2}R^{2}.$
-which implies the result in (5).
-A.3.2 Proof of Global Statistical Error Bound in Theorem 1
+as claimed. ∎
+Remark: Lemma 4 implies that if we solve the GlobalTrain problem exactly, the solution will be close to the ground truth of the global model with a certain error, which is independent of $\lambda$.
+After presenting the two useful lemmas in Appendix A.2, we will start to prove the theoretical results in Section 4.1.
+In Section A.3, we will prove Theorem 1 first, then we will prove the statistical error bound in Problem 2 as demonstrated in Corollary 1.
+A.3 Proof of Statistical Convergence in Theorem 1
+In the discuss below, we always assume ${{\bm{w}}_{\star}^{(i)}}$ discussed below comes from the parameter space (5).
+As the local model is shrunk towards the global model under the influence of personalization degree, how fast the global model ${\tilde{\bm{w}}^{(g)}}$ convergence to ${{\bm{w}}_{\star}^{(g)}}$ can also influence the convergence rate of the local model.
+In Appendix A.3.1, we will establish the statistical error bound of the global model first.
+Then in Appendix A.3.2, we will provide the local model error bound, where we explicitly show that the statistical accuracy of the local model depends on the statistical accuracy of the global model.
+Finally, in Appendix A.3.3, we combine the results of local models and the global model to establish the one-line rate for the local model explicitly stated in Theorem 1.
+A.3.1 Global Statistical Error Bound
+To analyze the statistical error bound, we consider two separate cases: $R>\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$ and $R\leq\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$.
+The motivation for this distinction lies in whether collaboration among different clients is beneficial.
 Case 1: We first consider the case when $R>\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$.
-Recall that ${\tilde{\bm{w}}^{(g)}}$ and $\{{\tilde{\bm{w}}^{(i)}}\}_{i\in[m]}$ are the minimizers of the objective (1). According to the first-order condition, we have
+Recall that ${\tilde{\bm{w}}^{(g)}}$ and $\{{\tilde{\bm{w}}^{(i)}}\}_{i\in[m]}$ are the minimizers of Problem (2). According to the first-order condition, we have
 $\nabla_{{{\bm{w}}^{(i)}}}L_{i}\left({{\bm{w}}^{(i)}},S_{i}\right)\big{|}_{{{\bm{w}}^{(i)}}={\tilde{\bm{w}}^{(i)}}}=\lambda\left({\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\right),$ (23)
@@
 00 0 $\geq\sum_{i\in[m]}\sum_{j\in[n_{i}]}\frac{p_{i}}{n_{i}}\left(\ell({\tilde{\bm{w}}^{(i)}},z_{ij})-\ell({{\bm{w}}_{\star}^{(i)}},z_{ij})\right)+\sum_{i\in[m]}p_{i}\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\|^{2}-\sum_{i\in[m]}p_{i}\frac{\lambda}{2}\|{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}.$ (26)
-For the first term on the RHS in (25), apply the $\mu$-strongly convexity of the loss function $\ell$ (cf. Assumption 3), we then have
+For the first term on the R.H.S. in (26), apply the $\mu$-strongly convexity of the loss function $\ell$ (cf. Assumption 2), we then have
 $\begin{split}\sum_{i\in[m]}\sum_{j\in[n_{i}]}\frac{p_{i}}{n_{i}}\left(\ell({\tilde{\bm{w}}^{(i)}},z_{ij})-\ell({{\bm{w}}_{\star}^{(i)}},z_{ij})\right)&\geq-\sum_{i\in[m]}\sum_{j\in[n_{i}]}\frac{p_{i}}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle\\ &+\frac{\mu}{2}\sum_{i\in[m]}\sum_{j\in[n_{i}]}\frac{p_{i}}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2},\\ \end{split}$ (27)
@@
 $\begin{split}0&\geq-\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle+\frac{\mu}{2}\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\\ &+\sum_{i\in[m]}p_{i}\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\|^{2}-\sum_{i\in[m]}p_{i}\frac{\lambda}{2}\|{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}.\end{split}$ (28)
-Recall the definition in (17) and the Assumption 1 about statistical heterogeneity
+Recall the definition in (18) and the assumption about statistical heterogeneity in (5)
 00 0 $\geq-\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle+\frac{\mu}{2}\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\sum_{i}p_{i}\left\|\tilde{\bm{\delta}}^{(i)}\right\|^{2}-\frac{\lambda}{2}R^{2}.$ (29)
@@
 $0\geq-\rho\left(\Big{(}\sum_{i\in[m]}\frac{p_{i}}{n_{i}}\Big{)}\Big{(}\sum_{i}p_{i}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\Big{)}\right)^{\frac{1}{2}}+\frac{\mu}{2}\sum_{i}p_{i}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}-\frac{\lambda}{2}R^{2}.$ (34)
-The equation (33) would be regarded as a quadratic function of $\sum_{i}p_{i}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}$. If we set the $\lambda$ as
-$\lambda\leq\frac{\rho^{2}}{\mu R^{2}}\sum_{i\in[m]}\frac{p_{i}}{n_{i}},$ (34)
-then solving the inequality in (33) w.r.t $\sum_{i}p_{i}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}$ yields
-$\begin{split}\mathbb{E}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2}&\leq\left(\sum_{i}p_{i}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\right)^{\frac{1}{2}}\leq\frac{1}{\mu}\left(\rho\left(\sum_{i\in[m]}\frac{p_{i}}{n_{i}}\right)^{\frac{1}{2}}+\left(\rho^{2}\sum_{i\in[m]}\frac{3p_{i}}{n_{i}}\right)^{\frac{1}{2}}\right)\\ &=\frac{(1+\sqrt{3})\rho}{\mu}\left(\sum_{i\in[m]}\frac{p_{i}}{n_{i}}\right)^{\frac{1}{2}}.\\ \end{split}$ (35)
+If we assume $p_{i}=\frac{1}{m}$ and $n_{i}=n_{j}\forall i\neq j$ (assumptions used in Theorem 1), we have
+$\mathbb{E}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|^{2}\leq\frac{1}{u^{2}}\left(\frac{4\rho^{2}}{n}+2\mu\lambda R^{2}\right).$ (35)
 Case 2: Then we consider $R\leq\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$.
@@
 On the other hand, note that the optimality condition of $\tilde{\bm{w}}_{\text{GT}}$ is
-$\sum_{i}p_{i}\nabla L_{i}(\tilde{\bm{w}}_{\text{GT}},S_{i})=0,$ (42)
+$\sum_{i}p_{i}\nabla L_{i}(\tilde{\bm{w}}_{\text{GT}},S_{i})=0.$ (42)
 Using the strong-convexity assumption in Assumption 2, we can show that
@@
 $\leq 2L^{2}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}+2\|\nabla L_{i}({{\bm{w}}_{\star}^{(i)}},S_{i})\|^{2}.$ (47)
-Take expectation w.r.t $S_{i}$ and use Assumption 1 and 4, then we have
+Take expectation w.r.t all data and use Assumption 3 in the parameter space (5), then we have
 $\mathbb{E}[\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})\|^{2}]$ $\leq 2L^{2}\mathbb{E}[\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}]+2\mathbb{E}\left[\left\|\frac{1}{n_{i}}\sum_{j=1}^{n_{i}}\nabla L({{\bm{w}}_{\star}^{(i)}},z_{ij})\right\|^{2}\right]$
@@
 $\mathbb{E}[\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}]$ $\leq\frac{g(\lambda)\left(4L^{2}R^{2}+2\frac{\rho^{2}}{n}\right)+\frac{8\rho^{2}}{\mu^{2}}\frac{1}{N}+\frac{8L^{2}}{\mu^{2}}R^{2}}{1-4L^{2}g(\lambda)}.$ (51)
-If we assume
-$g(\lambda)$ $\leq\frac{1}{8L^{2}}\land\left(\left(\frac{4\rho^{2}}{\mu^{2}N}+\frac{4L^{2}R^{2}}{\mu^{2}}\right)\Big{/}\left(2L^{2}R^{2}+\frac{\rho^{2}}{n}\right)\right),$ (52)
-we have
-$\mathbb{E}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}$ $\leq 32\left(\frac{\rho^{2}}{\mu^{2}}\frac{1}{N}+\frac{L^{2}}{\mu^{2}}R^{2}\right).$ (53)
-Furthermore, we can get a simplified condition for $\lambda$ as
-$\lambda\geq\frac{8\kappa}{\mu}\left\{8L^{2}\vee\left(\left(2L^{2}R^{2}+\frac{\rho^{2}}{n}\right)\Big{/}\left(\frac{4\rho^{2}}{\mu^{2}N}+\frac{4L^{2}R^{2}}{\mu^{2}}\right)\right)\right\}-\mu$ (54)
-where $\kappa=\frac{L}{\mu}\geq 1$ and the condition that $\lambda$ is non-negative holds trivially as $\frac{\kappa L^{2}}{\mu}=\kappa^{2}L\geq\mu$.
-Combine two-case arguments, we have
-$\Rightarrow\mathbb{E}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}\leq\begin{cases}\left(\frac{(1+\sqrt{3})\rho}{\mu}\right)^{2}\sum_{i\in[m]}\frac{p_{i}}{n_{i}}&R>\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}\\ 32\left(\frac{\rho^{2}}{\mu^{2}}\frac{1}{N}+\frac{L^{2}}{\mu^{2}}R^{2}\right)&R\leq\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}.\end{cases}$ (55)
-Thus it yields a one-line rate
-$\mathbb{E}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}\leq C_{1}\frac{1}{N}+C_{2}(\sum_{i\in[m]}\frac{p_{i}}{n_{i}}\wedge R^{2}),$ (56)
-where
-$\begin{split}C_{1}&=32\frac{\rho^{2}}{\mu^{2}},\\ C_{2}&=\left(\frac{(1+\sqrt{3})\rho}{\mu}\right)^{2}\vee\frac{32L^{2}}{\mu^{2}}.\end{split}$ (57)
-Moreover, if we assume data balanced, i.e $n_{i}=n,\forall i\in[m]$ and $p_{i}=\frac{1}{m}$, we obtain
-$\mathbb{E}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}\leq C_{1}\frac{1}{N}+C_{2}\left(\frac{1}{n}\wedge R^{2}\right),$ (58)
-as claimed.
-A.3.3 Proof of Local Statistical Error Bound in Theorem 1
+If we assume $p_{i}=\frac{1}{m}$, $n_{i}=n_{j}\;\forall i\neq j$ (assumptions used in Theorem 1) and $g(\lambda)\leq 1/(8L)^{2}$, we have
+$\mathbb{E}[\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}]$ $\leq\frac{g(\lambda)}{1-4L^{2}g(\lambda)}\left(4L^{2}R^{2}+2\frac{\rho^{2}}{n}\right)+\frac{16\rho^{2}}{\mu^{2}}\frac{1}{N}+\frac{16L^{2}}{\mu^{2}}R^{2}.$ (52)
+A.3.2 Local Statistical Error Bound
+Analogous to the analysis of the global model, we examine the statistical accuracy of the local model in a similar manner.
+For simplicity, in the following arguments, we denote the upper bound of $\mathbb{E}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|_{2}$ as $U_{0}$.
+The expliciate expression would be found in results (35) for $R>\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$ and (52) for $R\leq\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$.
 Case 1: first we consider the case when $R>\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$.
@@
 00 0 $\geq\sum_{j\in[n_{i}]}\frac{1}{n_{i}}\left(\ell\left({\tilde{\bm{w}}^{(i)}},z_{ij}\right)-\ell\left({{\bm{w}}_{\star}^{(i)}},z_{ij}\right)\right)+\frac{\lambda}{2}\left\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\right\|^{2}-\frac{\lambda}{2}\left\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\right\|^{2}.$ (54)
-For the first term on the RHS, applying the $\mu$-strongly convex of the loss function $\ell$ (cf. Assumption 3) yields
+For the first term on the R.H.S., applying the $\mu$-strongly convex of the loss function $\ell$ (cf. Assumption 2) yields
 $\begin{split}\sum_{j\in[n_{i}]}\frac{1}{n_{i}}\left(\ell({\tilde{\bm{w}}^{(i)}},z_{ij})-\ell({{\bm{w}}_{\star}^{(i)}},z_{ij})\right)&\geq-\sum_{j\in[n_{i}]}\frac{1}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle+\frac{\mu}{2}\sum_{j\in[n_{i}]}\frac{1}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2},\end{split}$ (55)
 where we use the definition of $\tilde{\Delta}^{(i)}_{\text{stat}}$ in (17).
-Pluggging (61) back into (60), it yields
+Plugging (55) back into (54), it yields
 $\begin{split}0&\geq-\sum_{j}\frac{1}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle+\frac{\mu}{2}\sum_{j}\frac{1}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\|^{2}-\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}.\end{split}$ (56)
@@
 $0\geq-\rho\frac{1}{\sqrt{n_{i}}}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\frac{\mu}{2}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\mathbb{E}\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\|^{2}-\frac{\lambda}{2}\mathbb{E}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}.$ (60)
-To utilize the statistical global convergence rate shown in Appendix A.3.2, we need to discuss $R$ in the same way. Recall the upper bound of $\tilde{\Delta}^{(g)}_{\text{stat}}$ in (58):
-$\mathbb{E}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}\leq\begin{cases}\frac{C_{2}}{n}&R>\frac{1}{\sqrt{n}}\\ \frac{C_{1}}{N}+C_{2}R^{2}&R\leq\frac{1}{\sqrt{n}}.\end{cases}$ (67)
-For simplicity, in the following arguments, we denote the upper bound of $\mathbb{E}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|_{2}$ as $U_{0}$.
 Note that
@@
 $\begin{split}0\geq&-\frac{\rho}{\sqrt{n_{i}}}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\frac{\mu}{2}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\mathbb{E}\left\|\tilde{\bm{\delta}}^{(i)}\right\|^{2}-\frac{\lambda}{2}\mathbb{E}\left\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}+{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\right\|^{2}\\ \geq&-\frac{\rho}{\sqrt{n_{i}}}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\frac{\mu}{2}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\mathbb{E}\left\|\tilde{\bm{\delta}}^{(i)}\right\|^{2}-\frac{\lambda}{2}\mathbb{E}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|^{2}-\frac{\lambda}{2}R^{2}-\lambda R\mathbb{E}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2}.\\ \end{split}$ (62)
-Putting (67) into (69) and reorganizing, we obtain
+Once we denote the upper bound of $\mathbb{E}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|^{2}$ as $U_{0}$, it yields
 $\begin{split}0\geq&-\frac{\rho}{\sqrt{n_{i}}}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\frac{\mu}{2}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\mathbb{E}\left\|\tilde{\bm{\delta}}^{(i)}\right\|^{2}-\frac{\lambda}{2}U_{0}^{2}-\frac{\lambda}{2}R^{2}-\lambda R\;U_{0}\\ \geq&\frac{\mu}{2}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}-\frac{\rho}{\sqrt{n_{i}}}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}-\frac{\lambda}{2}\left(U_{0}+R\right)^{2}.\end{split}$ (63)
-In addition, recall that in this case we set $\lambda^{\star}\leq\frac{\rho^{2}}{\mu R^{2}}\sum_{i\in[m]}\frac{p_{i}}{n_{i}}$ in the statistical convergence analysis of the global model.
-If we assume data balanced, i.e.
-$p_{i}=\frac{1}{m}$ and $n_{i}=n,\forall i\in[m]$, it yields $\lambda^{\star}\leq\frac{\rho^{2}}{\mu nR^{2}}$ and
-$\begin{split}\mathbb{E}\|\tilde{\Delta}^{(i)}_{\text{stat}}\|_{2}&\leq\frac{1}{\mu}\left[\frac{\rho}{\sqrt{n}}+\left(\frac{\rho^{2}}{n}+\frac{2\rho^{2}}{nR^{2}}\left(U_{0}+R\right)^{2}\right)^{\frac{1}{2}}\right].\\ &=\frac{1}{\sqrt{n}}\frac{1}{\mu}\left(\rho+\left(\rho^{2}+\frac{2\rho^{2}}{R^{2}}(U_{0}+R)^{2}\right)^{\frac{1}{2}}\right)\\ &\leq\frac{1}{\sqrt{n}}\frac{\rho+(\rho^{2}+4\rho^{2}(C_{2}+1))^{\frac{1}{2}}}{\mu}\\ &=\frac{1}{\sqrt{n}}\frac{\rho(1+(1+4(C_{2}+1))^{\frac{1}{2}})}{\mu}.\end{split}$ (71)
+If we assume $p_{i}=\frac{1}{m}$, $n_{i}=n_{j}\;\forall i\neq j$ (assumptions used in Theorem 1), we can obtain
+$\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\leq\frac{1}{u^{2}}\left(\frac{4\rho^{2}}{n}+4u\lambda\left(U_{0}^{2}+R^{2}\right)\right).$ (64)
 Case 2: we then consider the case when $R\leq\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$.
+If we assume $p_{i}=\frac{1}{m}$, $n_{i}=n_{j}\;\forall i\neq j$ (assumptions used in Theorem 1), we can obtain
 $\mathbb{E}[\|{\tilde{\bm{w}}^{(i)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}]$
@@
 $\overset{(a)}{\leq}\frac{12}{(\mu+\lambda)^{2}}\mathbb{E}[\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}})\|^{2}]+3\mathbb{E}[\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}]+3R^{2}$
-$\overset{(b)}{\leq}\frac{12}{(\mu+\lambda)^{2}}\left[4L^{2}\mathbb{E}[\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}]+4L^{2}R^{2}+2\frac{\rho^{2}}{n}\right]+3\mathbb{E}[\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}]+3R^{2}$
-$\overset{(c)}{\leq}\frac{12}{(\mu+\lambda)^{2}}\left[128L^{2}\left(\frac{\rho^{2}}{\mu^{2}}\frac{1}{N}+\frac{L^{2}}{\mu^{2}}R^{2}\right)+4L^{2}R^{2}+2\frac{\rho^{2}}{n}\right]+96\left(\frac{\rho^{2}}{\mu^{2}}\frac{1}{N}+\frac{L^{2}}{\mu^{2}}R^{2}\right)+3R^{2}$
-$=\frac{12}{(\mu+\lambda)^{2}}\left[\frac{128L^{2}\rho^{2}}{\mu^{2}}\frac{1}{N}+(128L^{4}/\mu^{2}+4L^{2})R^{2}+2\frac{\rho^{2}}{n}\right]+\frac{96\rho^{2}}{\mu^{2}}\frac{1}{N}+(96\frac{L^{2}}{\mu^{2}}+3)R^{2}$ (72)
-where step (a) comes from Assumption 1 and Lemma 4, step (b) comes from result (A.3.2) and step (c) is a direct result of applying the error bound of $\mathbb{E}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}$.
+$\overset{(b)}{\leq}\frac{12}{(\mu+\lambda)^{2}}\left[4L^{2}\mathbb{E}[\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}]+4L^{2}R^{2}+2\frac{\rho^{2}}{n}\right]+3\mathbb{E}[\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}]+3R^{2},$ (65)
+where step (a) comes from the assumption of parameter space in (5) and Lemma 3 and step (b) comes from result (A.3.1).
+A.3.3 Proof the Statistical Error Bound in Theorem 1
+For $R>\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$, putting the statistical error of global model (35) into the statistical error of local models (64), we have
+$\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\leq\left(\frac{4\rho^{2}}{u^{2}}+\frac{4\rho^{2}}{u^{3}}\lambda\right)\frac{1}{n}+\frac{8}{u^{2}}\lambda^{2}R^{2}+\frac{4}{u}\lambda R^{2},$ (66)
+therefore the second part of (8) in Theorem 1 is established. For clarity, we can define
+$q_{2}(\lambda)=\frac{8}{u^{2}}\lambda^{2}R^{2}+\frac{4}{u}\lambda R^{2}.$ (67)
+For $R>\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$, plugging the result (52) of the global model into the local model’s error bound (A.3.2) and reorganizing the terms yields
+$\mathbb{E}\left[\|\tilde{\Delta}^{(i)}_{\text{stat}}\|^{2}\right]\leq$ $\frac{12}{(\mu+\lambda)^{2}}\left[4L^{2}R^{2}+2\frac{\rho^{2}}{n}\right]+{\left[\frac{48L^{2}}{(\mu+\lambda)^{2}}+3\right]U_{0}^{2}+3R^{2}}$ (68)
+$\leq$ $\frac{48\rho^{2}}{\mu^{2}}\frac{1}{N}+\left[\frac{48L^{2}}{\mu^{2}}+3\right]R^{2}+\frac{1}{q_{1}(\lambda)},$ (69)
+where
+$\begin{split}[q_{1}(\lambda)]^{-1}=&\frac{12}{(\mu+\lambda)^{2}}\left[4L^{2}R^{2}+2\frac{\rho^{2}}{n}\right]+\frac{48L^{2}}{(\mu+\lambda)^{2}}\left[\frac{g(\lambda)}{1-4L^{2}g(\lambda)}\left(4L^{2}R^{2}+2\frac{\rho^{2}}{n^{2}}\right)\right.\\ &\left.+2\left(\frac{8\rho^{2}}{\mu^{2}}\frac{1}{N}+\frac{8L^{2}}{\mu^{2}}R^{2}\right)\right]+\frac{3g(\lambda)}{1-4L^{2}g(\lambda)}\left(4L^{2}R^{2}+2\frac{\rho^{2}}{n^{2}}\right).\end{split}$ (70)
+This combining with the definition of $g(\lambda)$ implies that $q_{1}(\lambda)$ is a monotonically increasing function of $\lambda$ and $\lim_{\lambda\to\infty}q_{1}(\lambda)=\infty$, as claimed.
+A.4 The Proof of Corollary 1
+Analogous to the proof of Theorem 1, we will discuss the two cases of $R^{2}$ separately as well. For clarity of our analysis, we will assume $p_{i}=\frac{1}{m}$ and $n_{i}=n_{j}\>\forall i\neq j$.
+A.4.1 The Statistical Convergence of the Global Model with a Choice of $\lambda$
+Case 1: For $R>\frac{1}{\sqrt{n}}$, we have derived the global model error bound in (34) as
+$0\geq-\rho\left(\Big{(}\sum_{i\in[m]}\frac{p_{i}}{n_{i}}\Big{)}\Big{(}\sum_{i}p_{i}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\Big{)}\right)^{\frac{1}{2}}+\frac{\mu}{2}\sum_{i}p_{i}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}-\frac{\lambda}{2}R^{2}.$ (71)
+If we set the $\lambda$ as
+$\lambda\leq\frac{\rho^{2}}{\mu nR^{2}},$ (72)
+then solving the inequality in (35) w.r.t $\sum_{i}p_{i}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}$ cound yield
+$\begin{split}\mathbb{E}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|\leq\left(\sum_{i}p_{i}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\right)^{\frac{1}{2}}\leq\frac{(1+\sqrt{3})\rho}{\mu\sqrt{n}}.\end{split}$ (73)
+Case 2: For $R\leq\frac{1}{\sqrt{n}}$, we have obtain the results in (52) as
+$\mathbb{E}[\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}]$ $\leq\frac{g(\lambda)}{1-4L^{2}g(\lambda)}\left(4L^{2}R^{2}+2\frac{\rho^{2}}{n}\right)+\frac{16\rho^{2}}{\mu^{2}}\frac{1}{N}+\frac{16L^{2}}{\mu^{2}}R^{2}.$ (74)
+If we assume
+$g(\lambda)$ $\leq\frac{1}{8L^{2}}\land\left(\left(\frac{4\rho^{2}}{\mu^{2}N}+\frac{4L^{2}R^{2}}{\mu^{2}}\right)\Big{/}\left(2L^{2}R^{2}+\frac{\rho^{2}}{n}\right)\right),$ (75)
+we have
+$\mathbb{E}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}$ $\leq 32\left(\frac{\rho^{2}}{\mu^{2}}\frac{1}{N}+\frac{L^{2}}{\mu^{2}}R^{2}\right).$ (76)
+Furthermore, we can get a simplified condition for $\lambda$ as
+$\lambda\geq\frac{8\kappa}{\mu}\left\{8L^{2}\vee\left(\left(2L^{2}R^{2}+\frac{\rho^{2}}{n}\right)\Big{/}\left(\frac{4\rho^{2}}{\mu^{2}N}+\frac{4L^{2}R^{2}}{\mu^{2}}\right)\right)\right\}-\mu,$ (77)
+where $\kappa=\frac{L}{\mu}\geq 1$ and the condition that $\lambda$ is non-negative holds trivially as $\frac{\kappa L^{2}}{\mu}=\kappa^{2}L\geq\mu$.
+Combine two-case arguments, we have
+$\mathbb{E}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}\leq\begin{cases}\left(\frac{(1+\sqrt{3})\rho}{\mu\sqrt{n}}\right)^{2}&R>\frac{1}{\sqrt{n}}\\ 32\left(\frac{\rho^{2}}{\mu^{2}}\frac{1}{N}+\frac{L^{2}}{\mu^{2}}R^{2}\right)&R\leq\frac{1}{\sqrt{n}}.\end{cases}$ (78)
+Thus it yields a one-line rate
+$\mathbb{E}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}\leq C_{1}\frac{1}{N}+C_{2}(\frac{1}{n}\wedge R^{2}),$ (79)
+where
+$\begin{split}C_{1}&=32\frac{\rho^{2}}{\mu^{2}},\\ C_{2}&=\left(\frac{(1+\sqrt{3})\rho}{\mu}\right)^{2}\vee\frac{32L^{2}}{\mu^{2}}.\end{split}$ (80)
+A.4.2 The Statistical Convergence of the Local Model with a Choice of $\lambda$
+Case 1: For $R>\frac{1}{\sqrt{n}}$, we have derived the result in (64) as
+$\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\leq\frac{1}{u^{2}}\left(\frac{4\rho^{2}}{n}+4u\lambda\left(U_{0}^{2}+R^{2}\right)\right).$ (81)
+In addition, recall that, when $R>\frac{1}{\sqrt{n}}$, we set $\lambda\leq\frac{\rho^{2}}{\mu nR^{2}}$ in the statistical convergence analysis of the global model.
+Therefore, putting the global model’s error bound (See result in (78)) into (64), which yields
+$\begin{split}\mathbb{E}\|\tilde{\Delta}^{(i)}_{\text{stat}}\|_{2}&\leq\frac{1}{\mu}\left[\frac{\rho}{\sqrt{n}}+\left(\frac{\rho^{2}}{n}+\frac{2\rho^{2}}{nR^{2}}\left(U_{0}+R\right)^{2}\right)^{\frac{1}{2}}\right].\\ &=\frac{1}{\sqrt{n}}\frac{1}{\mu}\left(\rho+\left(\rho^{2}+\frac{2\rho^{2}}{R^{2}}(U_{0}+R)^{2}\right)^{\frac{1}{2}}\right)\\ &\leq\frac{1}{\sqrt{n}}\frac{\rho+(\rho^{2}+4\rho^{2}(C_{2}+1))^{\frac{1}{2}}}{\mu}\\ &=\frac{1}{\sqrt{n}}\frac{\rho(1+(1+4(C_{2}+1))^{\frac{1}{2}})}{\mu},\end{split}$ (82)
+where $C_{2}$ has been defined in (80).
+Case 2: For $R\leq\frac{1}{\sqrt{n}}$, from the result (A.3.2), we know
+$\quad\mathbb{E}[\|{\tilde{\bm{w}}^{(i)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}]\leq\frac{12}{(\mu+\lambda)^{2}}\left[4L^{2}U_{0}^{2}+4L^{2}R^{2}+2\frac{\rho^{2}}{n}\right]+3U_{0}^{2}+3R^{2}.$ (83)
+Then putting the global model’s results (78) into (A.3.2), it yields
+$\quad\mathbb{E}[\|{\tilde{\bm{w}}^{(i)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}]$ (84)
+$\leq\frac{12}{(\mu+\lambda)^{2}}\left[128L^{2}\left(\frac{\rho^{2}}{\mu^{2}}\frac{1}{N}+\frac{L^{2}}{\mu^{2}}R^{2}\right)+4L^{2}R^{2}+2\frac{\rho^{2}}{n}\right]+96\left(\frac{\rho^{2}}{\mu^{2}}\frac{1}{N}+\frac{L^{2}}{\mu^{2}}R^{2}\right)+3R^{2}$
+$=\frac{12}{(\mu+\lambda)^{2}}\left[\frac{128L^{2}\rho^{2}}{\mu^{2}}\frac{1}{N}+(128L^{4}/\mu^{2}+4L^{2})R^{2}+2\frac{\rho^{2}}{n}\right]+\frac{96\rho^{2}}{\mu^{2}}\frac{1}{N}+(96\frac{L^{2}}{\mu^{2}}+3)R^{2}.$ (85)
 If we assume $\lambda$ satisfies the following condition
-$\frac{12}{(\mu+\lambda)^{2}}\leq\frac{96\rho^{2}/(\mu^{2}N)+99\kappa^{2}R^{2}}{128\rho^{2}/N+132L^{2}R^{2}+2\rho^{2}/n}$ (73)
-we have
-$\mathbb{E}[\|{\tilde{\bm{w}}^{(i)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}]\leq\frac{192\rho^{2}}{\mu^{2}}\frac{1}{N}+198\kappa^{2}R^{2}$ (74)
+$\frac{12}{(\mu+\lambda)^{2}}\leq\frac{96\rho^{2}/(\mu^{2}N)+99\kappa^{2}R^{2}}{128\rho^{2}/N+132L^{2}R^{2}+2\rho^{2}/n},$ (86)
+then we will have
+$\mathbb{E}[\|{\tilde{\bm{w}}^{(i)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}]\leq\frac{192\rho^{2}}{\mu^{2}}\frac{1}{N}+198\kappa^{2}R^{2}.$ (87)
 Furthermore, we can get a simplified condition for $\lambda$ as
-$\lambda\geq\sqrt{\frac{8(64\rho^{2}/N+66L^{2}R^{2}+\rho^{2}/n)}{32\rho^{2}/(\mu^{2}N)+33\kappa^{2}R^{2}}}-\mu$ (75)
-Combining two-case arguments, we have:
+$\lambda\geq\sqrt{\frac{8(64\rho^{2}/N+66L^{2}R^{2}+\rho^{2}/n)}{32\rho^{2}/(\mu^{2}N)+33\kappa^{2}R^{2}}}-\mu.$ (88)
+Combining two-case arguments, we have
 $\Rightarrow\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\leq\begin{cases}\frac{1}{n}\left(\frac{\rho(1+(1+4(C_{2}+1))^{\frac{1}{2}})}{\mu}\right)^{2}&R>\sqrt{\frac{m}{N}}\\ \frac{192\rho^{2}}{\mu^{2}}\frac{1}{N}+198\kappa^{2}R^{2}&R\leq\sqrt{\frac{m}{N}}\end{cases}.$ (89)
@@
 where
-$C_{3}$ $=\frac{192\rho^{2}}{\mu^{2}},$ (78)
-$C_{4}$ $=\left(\frac{\rho(1+(4C_{2}+5)^{\frac{1}{2}})}{\mu}\right)^{2}\vee 198\kappa^{2},$ (79)
+$\begin{split}C_{3}&=\frac{192\rho^{2}}{\mu^{2}},\\ C_{4}&=\left(\frac{\rho(1+(4C_{2}+5)^{\frac{1}{2}})}{\mu}\right)^{2}\vee 198\kappa^{2},\\ \end{split}$ (91)
 and $C_{2}$ are specified in (80).
-Combining the results in A.3.2 and A.3.3, we leave a remark below to further interpret the effect of personalization in obj (1) and discuss how to incorporate the conditions of $\lambda$ when establishing the global and local statistical accuracy.
-Remark 2.
-To better understand how such an interpolation is achieved through an adaptive personalization degree, we examine the role of $\lambda$ in navigating between the two extreme cases: GlobalTrain and LocalTrain.
-As the statistical heterogeneity $R\to\infty$, we have $\lambda\to 0$, leading to a high degree of personalization.
-The statistical error bound becomes $\mathcal{O}(n^{-1})$, matching the rate of LocalTrain; as the statistical heterogeneity $R\to 0$, $\lambda$ increases, leading to a low degree of personalization.
-The statistical error bound eventually converges to $\mathcal{O}(N^{-1})$, matching the rate of GlobalTrain under a homogeneous setting.
-As $R\in(0,\infty)$, $\lambda$ transits between the two extremes, making FedProx perform no worse than LocalTrain while benefiting from global training when clients share similarities.
+Therefore, combining results from Section A.3.1 and A.3.2 we can obtain the upper bound in (9).
 To make the global statistical error bound and local statistical error bound hold simultaneously, we should be aware about the condition of the adaptive strategy of $\lambda$ specified in our proof.
@@
 $\lambda\geq\text{max}\left\{64\kappa^{2}L,\left(2\kappa\vee 5\right)\mu\frac{2L^{2}R^{2}+\rho^{2}/n}{L^{2}R^{2}+\rho^{2}/N}\right\}-\mu.$ (93)
-Appendix B Optimization Convergence Rate
+To sum up, we give the one-line rate in (90) for Corollary 1 and provide the potential solution of $\lambda$ with a closed-form solution in (92).
+A.5 Discussion on the Statistical Convergence Rate and Comparison with Existing Results
+Combining the results in A.3.1 and A.3.2, we leave a remark below to further interpret the effect of personalization in Problem (2) and discuss how to incorporate the conditions of $\lambda$ when establishing the global and local statistical accuracy.
+To better understand how such an interpolation is achieved through an adaptive personalization degree, we examine the role of $\lambda$ in navigating between the two extreme cases: GlobalTrain and LocalTrain.
+As the statistical heterogeneity $R\to\infty$, we have $\lambda\to 0$, leading to a high degree of personalization.
+The statistical error bound becomes $\mathcal{O}(n^{-1})$, matching the rate of LocalTrain; as the statistical heterogeneity $R\to 0$, $\lambda$ increases, leading to a low degree of personalization.
+The statistical error bound eventually converges to $\mathcal{O}(N^{-1})$, matching the rate of GlobalTrain under a homogeneous setting.
+As $R\in(0,\infty)$, $\lambda$ transits between the two extremes, making PFL Problem (2) perform no worse than LocalTrain while benefiting from global training when clients share similarities.
+Our contribution can be summarized into three aspects :
+$\bullet$ Our results are derived under more realistic assumptions.
+Chen et al.
+(2023c) imposes additional stringent conditions including a bounded parameter space $D$ and a uniform bound on the loss function $\ell(\cdot,z)$.
+Such conditions are difficult to satisfy in practice.
+For example, even a simple linear regression model with sub-Gaussian covariates would violate these assumptions.
+$\bullet$ Our result is established for both global and local models with the same choice of $\lambda$.
+In contrast, Chen et al.
+(2023c) proposes a two-stage process where the global model is first estimated, followed by an additional local training phase with a different choice of $\lambda$.
+This two-step approach is likely the artifact of their analysis, as existing work on Problem (2) also shows that a single-stage optimization with a chosen $\lambda$ suffices to achieve the desired performance (T Dinh et al., 2020).
+$\bullet$ We establish a state-of-the-art minimax statistical rate.
+As shown in Table 3, the established upper bound matches the lower bound established in (Chen et al., 2023c).
+In a heterogeneous case when $R\geq n^{-1/2}$, Chen et al.
+(2023c) establishes a rate that is at least of the order $\mathcal{O}\left(n^{-1}R^{2}\right)$ as $D\geq R$.
+This bound becomes increasingly loose with larger $R$.
+In contrast, our bound is $\mathcal{O}\left(n^{-1}\right)$, which shows that properly-tuned Problem 2 is always no worse than LocalTrain, independent of $R$.
+In a homogeneous case when $R\leq m^{-1}n^{-1/2}$, they can only establish a rate slower than $\mathcal{O}\left(1/(\sqrt{m}n)\right)$, while our result achieves a rate of $\mathcal{O}\left(1/(mn)\right)$, leveraging all $mn$ samples and matching the rate of GlobalTrain on the IID data.
+As the client number could be extremely large ($10^{5}$ devices in (Chen et al., 2023a)), order of $m$ is non-trivial.
+During the transition period when $m^{-1}n^{-1/2}<R<n^{-1/2}$, we achieve a rate of $\mathcal{O}(R^{2})$, again matching the lower bound, and is strictly faster than the rate $\mathcal{O}(n^{-1/2}R)$ established in Chen et al.
+(2023c).
+Table 3: Results Comparison.
+$D:=\sup_{\bm{w},\bm{w}^{\prime}\in\mathcal{D}}\|\bm{w}-\bm{w}^{\prime}\|$ is the diameter of the parameter space ($D\geq R$) and $\|\ell\|_{\infty}=\inf\{M\in\mathbb{R}:\ell(\cdot,z)\leq M,\text{for any }z\}$ is a uniform upper bound on the loss function.
+Statistical rate refers to $\mathbb{E}\|{{\bm{w}}_{\star}^{(i)}}-\widetilde{\bm{w}}^{(i)}\|^{2}$, with constants neglected.
+Source Assumption Paradigm Statistical Rate
+$R>\frac{1}{\sqrt{n}}$ $\frac{1}{m\sqrt{n}}<R\leq\frac{1}{\sqrt{n}}$ $R\leq\frac{1}{m\sqrt{n}}$
+(Chen et al., 2023c) A1,2,3, $D\vee\|\ell\|_{\infty}\leq C$ Two-stage $\frac{D^{2}+\|\ell\|_{\infty}}{n}$ 1 $\frac{\|\ell\|_{\infty}}{\sqrt{n}}R$ $\frac{\|\ell\|_{\infty}}{\sqrt{m}n}$
+Ours A1,2,3 One-stage $\frac{1}{n}$ $R^{2}\wedge\frac{1}{mn}$ $\frac{1}{mn}$
+Lower Bound - - $\frac{1}{n}$ $R^{2}\wedge\frac{1}{mn}$ $\frac{1}{mn}$
+Next, we discuss our theoretical contributions.
+Prior work (Chen et al., 2023c) established the rate through algorithmic stability, and to obtain bounded stability, the boundedness of the loss function is needed.
+This assumption (or bounded gradients) is, in fact, commonly imposed for analysis based on the tool of algorithmic stability.
+To relax this assumption and obtain a bound independent of the diameter of the domain/gradient norm/loss function norm, we need to jump out of the stability analysis framework and develop new techniques.
+Our analysis does not rely on any algorithm but directly tackles the objective function, establishing rates using purely the properties of the loss.
+Specifically, using the strong convexity of the loss function, we first proved the estimation error is bounded by the gradient variance and an extra statistical heterogeneity term controlled by $\lambda$, as detailed in Lemma 1.
+This allows us to show for large $R$, a small $\lambda$ can be chosen to yield rate $O(1/n)$ and match one of the worst cases in the lower bound.
+For the complementary case $R\leq 1/\sqrt{n}$, we leveraged the GlobalTrain solution $\widetilde{\bm{w}}_{GT}$ as a bridge and proved the solutions of Problem (2) to $\widetilde{\bm{w}}_{GT}$ are bounded by a term inversely proportional to $\lambda$ (cf.
+Lemma 4), implying that they can be made arbitrarily close to $\widetilde{\bm{w}}_{GT}$ by setting $\lambda$ small.
+This, together with the rate of $\widetilde{\bm{w}}_{GT}$ as proved in Lemma 5, yields a rate of $O(1/N+R^{2})$.
+Combining the two cases, we proved minimax optimality.
+Appendix B Algorithmic Convergence Rate
+In this section, we will start to prove the algorithmic convergence rate rigorously.
+First, we present our Algorithm 1 in Appendix 1.
+In Appendix B.2, we provide some useful lemmas to facilitate our analysis, including the proof of Lemma 1 presented in the main paper and the inner loop and outer loop error contractions in Algorithm 1.
+Next, in Appendix B.3, we are ready to prove Theorem 2 with the local and global model convergence.
+In Appendix B.4, to show the communication cost and computation cost of our method, the proof of Corollary 2 is established.
+Finally, in Appendix B.5, as stochastic gradient descent would reduce the computation cost over the full sample, we present Algorithm 2 for the noise setting and provably show the benefit of collaboration to reduce the noise.
 B.1 Algorithm
+Algorithm 1 Federated Gradient Descent with K-Step Local Optimization
 Input: Initial global model ${\bm{w}}_{1}^{(g)}$, initial local models $\{{\bm{w}}_{0,K}^{(i)}\}_{i\in[m]}$, global rounds $T$, global step sizes $\gamma$, local rounds $K$, local step sizes $\eta$
@@
        The server sends ${\bm{w}}_{t}^{(g)}$ to client $i$, $\forall i\in[m]$
-       Set ${\bm{w}}_{t,0}^{(i)}={\bm{w}}_{t-1,K}^{(i)}$;
-       for $k=0,\dots,K-1$ do
+Set ${\bm{w}}_{t,0}^{(i)}={\bm{w}}_{t-1,K}^{(i)}$for $k=0,\dots,K-1$ do
              ${\bm{w}}_{t,k+1}^{(i)}={\bm{w}}_{t,k}^{(i)}-\frac{\eta}{n_{i}}\sum_{j\in[n_{i}]}\left\{\nabla\ell({\bm{w}}_{t,k}^{(i)},z_{i,j})+\lambda\left({\bm{w}}_{t,k}^{(i)}-\bm{w}_{t}^{(g)}\right)\right\}$
-      Push ^$\widehat{\nabla}F_{i}(\bm{w}_{t}^{(g)})=\lambda(\bm{w}_{t}^{(g)}-{\bm{w}}_{t,K}^{(i)})$ to the server;
-      $\bm{w}_{t+1}^{(g)}\leftarrow\bm{w}_{t}^{(g)}-\frac{\gamma}{m}\sum_{i\in[m]}\widehat{\nabla}F_{i}(\bm{w}_{t}^{(g)})$;
-Algorithm 1 One-stage FedProx
+      Push ^$\widehat{\nabla}F_{i}(\bm{w}_{t}^{(g)})=\lambda(\bm{w}_{t}^{(g)}-{\bm{w}}_{t,K}^{(i)})$ to the server  $\bm{w}_{t+1}^{(g)}\leftarrow\bm{w}_{t}^{(g)}-\frac{\gamma}{m}\sum_{i\in[m]}\widehat{\nabla}F_{i}(\bm{w}_{t}^{(g)})$
 B.2 Proof of lemmas
@@
 Now, we are ready to prove Lemma 1.
-First, from mishchenko2023convergence, we can directly obtain the analytic properties of $F_{i}$.
-Next, the analytic properties of $h_{i}$ are direct results from Assumption 3 and 2.
+First, from (Mishchenko et al., 2023), we can directly obtain the analytic properties of $F_{i}$.
+Next, the analytic properties of $h_{i}$ are direct results from Assumption 2 and 1.
 Finally, to analyze the property of the mapping $w_{\star}^{(i)}$, we just need to use the result from the above claim as we know the definition of ${\bm{w}}_{\star}^{(i)}(\cdot)$ in (11).
@@
 Lemma 5 ((Ji et al., 2022), Inner Loop Error Contraction).
-Suppose Assumption 3 and 2 hold, then if $\eta\leq 1/L_{\ell}$, for all $t\geq 0$ and $\epsilon>0$, the inner loop error bound of all clients can be formulated as:
+Suppose Assumption 2 and 1 hold, then if $\eta\leq 1/L_{\ell}$, for all $t\geq 0$ and $\epsilon>0$, the inner loop error bound of all clients can be formulated as:
 $e_{t+1}^{(i)}:=\|{\bm{w}}_{t+1,K}^{(i)}-{\bm{w}}_{t+1,\star}^{(i)}\|^{2}\leq(1+\epsilon)(1-\eta\mu_{\ell})^{K}e_{t}^{(i)}+\left(1+\frac{1}{\epsilon}\right)(1-\eta\mu_{\ell})^{K}\|\bm{w}^{(i)}_{t,\star}-\bm{w}^{(i)}_{t-1,\star}\|^{2}.$ (100)
@@
 Lemma 6 (Outer Loop Error Contraction).
-Suppose Assumption 3 and 2 hold, then if $\gamma\leq L_{g}$, we have for all $t>0$ and $\epsilon_{g}>0$:
+Suppose Assumption 2 and 1 hold, then if $\gamma\leq L_{g}$, we have for all $t>0$ and $\epsilon_{g}>0$:
 $\begin{split}\|{\bm{w}}_{t+1}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}&\leq(1-\gamma\mu_{g}+\gamma\epsilon_{g})\|{\bm{w}}_{t}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}-\gamma^{2}\|\nabla F({\bm{w}}_{t}^{(g)})\|^{2}\\ &+(\gamma^{3}L_{g})\|\widehat{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}+\left(\frac{\gamma}{\epsilon_{g}}+\gamma^{2}\right)\|\nabla F({\bm{w}}_{t}^{(g)})-\widehat{\nabla}F({\bm{w}}_{t}^{(g)})\|^{2}.\end{split}$ (101)
@@
 B.3 Proof of Theorem 2
+B.3.1 Proof of The Global Model Convergence Rate in Theorem 2
 Recall the definition of $e_{t}^{(i)}$ in Lemma 5, we define
@@
 It remains to specify the free parameters $(\epsilon_{g},\epsilon,\zeta$) in (112), (115) and others listed in Lemma 6 and Lemma 5.
-$\bullet$ Convergence Rate For the connection term in (103), if we set
+$\bullet$ Convergence Rate. For the connection term in (114), if we set
 $\epsilon_{g}=\frac{\mu_{g}}{2}\cdot(1-\gamma\mu_{g})>0,$ (116)
@@
 $\gamma<1/L_{g},\quad\Big{(}2+64L_{w}^{2}(1/\mu_{g})^{2}\lambda^{2}\Big{)}(1-\eta\mu_{\ell})^{K}\leq(1-\gamma L_{g})^{4}.$ (125)
+B.3.2 Proof of The Local Model Convergence Rate in Theorem 2
+Proof.
+Note that
+$\|{\bm{w}}_{t,K}^{(i)}-\widetilde{\bm{w}}^{(i)}\|^{2}\leq 2\|{\bm{w}}_{t,K}^{(i)}-\bm{w}^{(i)}_{t,\star}\|^{2}+2\|\bm{w}^{(i)}_{t,\star}-\widetilde{\bm{w}}^{(i)}\|^{2}.$ (126)
+For the first part on the right-hand side, recalling the definition of $e_{t}^{(i)}$ in Lemma 5
+$2\|{\bm{w}}_{t,K}^{(i)}-\bm{w}^{(i)}_{t,\star}\|^{2}=2\|e_{t}^{(i)}\|^{2}.$ (127)
+For the second term on the right-hand side, first, note the property in (11)
+$\bm{w}^{(i)}_{t,\star}=\operatorname{prox}_{L_{i}/\lambda}({\bm{w}}_{t}^{(g)}).$ (128)
+Based on the first order condition of (2)
+$\widetilde{\bm{w}}^{(i)}=\operatorname{prox}_{L_{i}/\lambda}(\widetilde{\bm{w}}^{(g)}).$ (129)
+Next, plugging (128) and (129) into (126), it yields
+$\begin{split}\|{\bm{w}}_{t,K}^{(i)}-\widetilde{\bm{w}}^{(i)}\|^{2}&\leq 2\|e_{t}^{(i)}\|^{2}+2\|\operatorname{prox}_{L_{i}/\lambda}({\bm{w}}_{t}^{(g)})-\operatorname{prox}_{L_{i}/\lambda}(\widetilde{{\bm{w}}}^{(g)})\|^{2}\\ &\leq 2\|e_{t}^{(i)}\|^{2}+2L_{w}^{2}\|\bm{w}^{(i)}_{t,\star}-\widetilde{\bm{w}}^{(i)}\|^{2},\end{split}$ (130)
+where the last we use Lemma 1.
+In the proof of Theorem 2, we have established that both $\|\bm{w}_{t}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}$ and $e_{t}$ converges to zero linearly at rate $21-\gamma\mu_{g}/2-(\gamma\mu_{g})^{2}/$ 2, combining with (130), the proof is finished.
+∎
 B.4 Proof of Corollary 2
@@
 $K\geq 2+\frac{\lambda+L}{\lambda+\mu}\cdot\log(1056\kappa^{2}).$ (134)
-B.5 Proof of The Local Model Convergence Rate in Theorem 2
+B.5 Global Convergence Rate with Stochastic Noise
+Similar to the noiseless case, we propose FedCLUP below which uses stochastic gradient descent to solve Problem (2).
+Algorithm 2 FedCLUP: Federated Learning with Constant Local Update Personalization
+Input: Initial global model ${\bm{w}}_{1}^{(g)}$, initial local models $\{{\bm{w}}_{0,K}^{(i)}\}_{i\in[m]}$, global rounds $T$, global step sizes $\gamma$, local rounds $K$, local step sizes $\eta$, stochastic batch size $B$
+Output: Local models $\{{\bm{w}}_{T,K}^{(i)}\}_{i\in[m]}$ and global model ${\bm{w}}_{T}^{(g)}$
+for $t=1,\dots,T$ do
+       The server sends ${\bm{w}}_{t}^{(g)}$ to client $i$, $\forall i\in[m]$;
+Each client randomly draw $B$ data without replacement;
+Set ${\bm{w}}_{t,0}^{(i)}={\bm{w}}_{t-1,K}^{(i)}$for $k=0,\dots,K-1$ do
+             ${\bm{w}}_{t,k+1}^{(i)}={\bm{w}}_{t,k}^{(i)}-\frac{\eta}{B}\sum_{j\in[B]}\left\{\nabla\ell({\bm{w}}_{t,k}^{(i)},z_{i,j})+\lambda\left({\bm{w}}_{t,k}^{(i)}-\bm{w}_{t}^{(g)}\right)\right\}$
+      Push ^$\widehat{\nabla}F_{i}(\bm{w}_{t}^{(g)})=\lambda(\bm{w}_{t}^{(g)}-{\bm{w}}_{t,K}^{(i)})$ to the server  $\bm{w}_{t+1}^{(g)}\leftarrow\bm{w}_{t}^{(g)}-\frac{\gamma}{m}\sum_{i\in[m]}\widehat{\nabla}F_{i}(\bm{w}_{t}^{(g)})$
+We use the following assumption to characterize the stochastic noise when sampling new data in each iteration over all clients without replacement.
+Assumption 4 (Stochastic noise).
+For any $i\in[m],j\in[n_{i}],k\in[K],t>0$, we assume
+$\mathbb{E}\left[\left\|\nabla_{\bm{w}_{i}}\hat{h}_{i}(\bm{w}_{t,k}^{(i)},\bm{w}_{t}^{(i)},s_{ij})-\nabla_{\bm{w}_{i}}h_{i}(\bm{w}_{t,\star}^{(i)},\bm{w}_{t}^{(i)})\right\|^{2}\right]\leq\sigma^{2}$ (135)
+Let we define $\mathcal{F}_{t,k}$ to be the sigma algebra generated by the randomness by Algorithm 2 up to ${\bm{w}}_{t,k}^{(i)}$.
+Lemma 7 (Inner Loop Error Contraction with stochastic).
+Suppose Assumption 1 and 2 hold, then if $\eta\leq 1/L_{\ell}$, for all $t\geq 0$ and $\epsilon>0$, the inner loop error bound of all clients can be formulated as:
+$\begin{split}g_{t+1}=&\;\mathbb{E}\left(\left\|\frac{1}{m}\sum_{i\in[m]}\bm{w}_{t+1,k}^{(i)}-\bm{w}_{t+1,*}^{(i)}\right\|^{2}\Big{|}\mathcal{F}_{t+1,k}\right)\\ &\leq\left[1+\epsilon+2\left(1+\frac{1}{\epsilon}\right)L_{w}^{2}\gamma^{2}\lambda^{2}\right]\left(1-\eta_{t}\mu_{\ell}\right)^{k}g_{t}\\ &\quad+2\left(1+\frac{1}{\epsilon}\right)\left(1-\mu_{\ell}\eta_{t}\right)^{k}L_{w}^{2}\gamma^{2}\|\nabla F(\bm{w}_{t}^{(g)})\|^{2}\\ &\quad+\frac{2\eta_{t}}{\mu_{\ell}}\frac{\sigma^{2}}{Bm}.\end{split}$ (136)
 Proof.
-Note that
-$\|{\bm{w}}_{t,K}^{(i)}-\widetilde{\bm{w}}^{(i)}\|^{2}\leq 2\|{\bm{w}}_{t,K}^{(i)}-\bm{w}^{(i)}_{t,\star}\|^{2}+2\|\bm{w}^{(i)}_{t,\star}-\widetilde{\bm{w}}^{(i)}\|^{2}.$ (119)
-For the first part on the right-hand side, recalling the definition of $e_{t}^{(i)}$ in Lemma 6
-$2\|{\bm{w}}_{t,K}^{(i)}-\bm{w}^{(i)}_{t,\star}\|^{2}=2\|e_{t}^{(i)}\|^{2}.$ (120)
-For the second term on the right-hand side, first, note the property in (9)
-$\bm{w}^{(i)}_{t,\star}=\operatorname{prox}_{L_{i}/\lambda}({\bm{w}}_{t}^{(g)}).$ (121)
-Based on the first order condition of (1)
-$\widetilde{\bm{w}}^{(i)}=\operatorname{prox}_{L_{i}/\lambda}(\widetilde{\bm{w}}^{(g)}).$ (122)
-Next, plugging (121) and (122) into (119), it yields
-$\begin{split}\|{\bm{w}}_{t,K}^{(i)}-\widetilde{\bm{w}}^{(i)}\|^{2}&\leq 2\|e_{t}^{(i)}\|^{2}+2\|\operatorname{prox}_{L_{i}/\lambda}({\bm{w}}^{(g)})-\operatorname{prox}_{L_{i}/\lambda}(\widetilde{{\bm{w}}}^{(g)})\|^{2}\\ &\leq 2\|e_{t}^{(i)}\|^{2}+2L_{w}^{2}\|\bm{w}^{(i)}_{t,\star}-\widetilde{\bm{w}}^{(i)}\|^{2},\end{split}$ (123)
-where the last we use Lemma 2.
-In the proof of Theorem 2, we have established that both $\|\bm{w}_{t}^{(g)}-\widetilde{\bm{w}}^{(g)}\|^{2}$ and $e_{t}$ converges to zero linearly at rate $21-\gamma\mu_{g}/2-(\gamma\mu_{g})^{2}/$ 2, combining with (123), the proof is finished.
-∎
-Appendix C Additional Experiment Details and Results
-C.1 Additional Experiment Details
+$\begin{split}g_{t+1}=&\;\mathbb{E}\left(\left\|\frac{1}{m}\sum_{i\in[m]}\bm{w}_{t+1,k}^{(i)}-\bm{w}_{t+1,*}^{(i)}\right\|^{2}\Big{|}\mathcal{F}_{t+1,k}\right)\\ =&\;\mathbb{E}\left(\left\|\frac{1}{m}\sum_{i\in[m]}\left[\bm{w}_{t+1,k-1}^{(i)}-\frac{\eta_{t}}{B}\sum_{j}\nabla_{w_{i}}\hat{h}_{i}(\bm{w}_{t+1,k-1}^{(i)},\bm{w}_{t+1}^{(i)},s_{ij})-\bm{w}_{t+1,*}^{(i)}\right]\right\|^{2}\Big{|}\mathcal{F}_{t+1,k}\right)\\ =&\;\mathbb{E}\left(\left\|\frac{1}{m}\sum_{i}\bm{w}_{t+1,k-1}^{(i)}-\bm{w}_{t+1,*}^{(i)}\right\|^{2}\Big{|}\mathcal{F}_{t+1,k}\right)\\ &+\eta_{t}^{2}\mathbb{E}\left(\left\|\frac{1}{mB}\sum_{i}\sum_{j}\nabla_{w_{i}}\hat{h}_{i}(\bm{w}_{t+1,k-1}^{(i)},\bm{w}_{t+1}^{(i)},s_{ij})\right\|^{2}\Big{|}\mathcal{F}_{t+1,k}\right)\\ &-2\eta_{t}\mathbb{E}\left[\left\langle\frac{1}{m}\sum_{i}\bm{w}_{t+1,k-1}^{(i)}-\bm{w}_{t+1,*}^{(i)},\frac{1}{Bm}\sum_{j}\nabla_{w_{i}}\hat{h}_{i}(\bm{w}_{t+1,k-1}^{(i)},\bm{w}_{t+1}^{(i)},s_{ij})\right\rangle\Big{|}\mathcal{F}_{t+1,k}\right].\end{split}$ (137)
+To simplify the result, we shall note
+$\mathbb{E}\left(\left\|\frac{1}{mB}\sum_{i}\sum_{j}\nabla_{\bm{w}_{i}}\hat{h}_{i}(\bm{w}_{t+1,k-1}^{(i)},\bm{w}_{t+1}^{(i)},s_{ij})\right\|^{2}\Big{|}\mathcal{F}_{t+1,k}\right)$
+$=$ $\;\mathbb{E}\left(\left\|\frac{1}{mB}\sum_{i}\sum_{j}\nabla_{\bm{w}_{i}}\hat{h}_{i}(\bm{w}_{t+1,k-1}^{(i)},\bm{w}_{t+1}^{(i)},s_{ij})-\nabla_{\bm{w}_{i}}h_{i}(\bm{w}_{t+1,k-1}^{(i)},\bm{w}_{t+1}^{(i)})\right\|^{2}\Big{|}\mathcal{F}_{t+1,k}\right)$
+$+\left\|\frac{1}{m}\sum_{i}\nabla_{\bm{w}_{i}}h_{i}(\bm{w}_{t+1,k-1}^{(i)},\bm{w}_{t+1}^{(i)})\right\|^{2}$
+$\overset{(a)}{\leq}$ $\;\frac{\sigma^{2}}{mB}+\left\|\frac{1}{m}\sum_{i}\left(\nabla_{\bm{w}_{i}}h_{i}(\bm{w}_{t+1,k-1}^{(i)},\bm{w}_{t+1}^{(i)})-\nabla_{\bm{w}_{i}}h_{i}(\bm{w}_{t+1,*}^{(i)},\bm{w}_{t+1}^{(i)})\right)\right\|^{2}$
+$\overset{(b)}{\leq}$ $\;\frac{\sigma^{2}}{mB}+2L_{\ell}\frac{1}{m}\sum_{i}\left(h_{i}(\bm{w}_{t+1,k-1}^{(i)},\bm{w}_{t+1}^{(i)})-h_{i}(\bm{w}_{t+1,*}^{(i)},\bm{w}_{t+1}^{(i)})\right).$ (138)
+where step (a) stands due to Assumption 4 and the optimality of $\bm{w}_{t+1,\star}^{(i)}$, step (b) uses the Lemma 2.29 in (Garrigos & Gower, 2023). Moreover, for the third term in (137), it yields
+$\begin{split}&\mathbb{E}\left[\left\langle\frac{1}{m}\sum_{i}\bm{w}_{t+1,k-1}^{(i)}-\bm{w}_{t+1,*}^{(i)},\frac{1}{Bm}\sum_{i,j}\nabla_{\bm{w}_{i}}\hat{h}_{i}(\bm{w}_{t+1,k-1}^{(i)},\bm{w}_{t+1}^{(i)},s_{j})\right\rangle\Big{|}\mathcal{F}_{t+1,k}\right]\\ =&\left\langle\frac{1}{m}\sum_{i}\bm{w}_{t+1,k-1}^{(i)}-\bm{w}_{t+1,*}^{(i)},\frac{1}{m}\sum_{i}\nabla_{\bm{w}_{i}}h_{i}(\bm{w}_{t+1,k-1}^{(i)},\bm{w}_{t+1}^{(i)})\right\rangle\\ \geq&\frac{1}{m}\sum_{i}\left(h_{i}(\bm{w}_{t+1,k-1}^{(i)},\bm{w}_{t+1}^{(i)})-h_{i}(\bm{w}_{t+1,*}^{(i)},\bm{w}_{t+1}^{(i)})\right)+\frac{\mu_{\ell}}{2}\left\|\frac{1}{m}\sum_{i}\bm{w}_{t+1,k-1}^{(i)}-\bm{w}_{t+1,*}^{(i)}\right\|^{2},\\ \end{split}$ (139)
+where the last line comes from the strong convexity of $h_{i}$ in Assumption 2.
+Plugging the results (B.5) and (139) into (137), we obtain
+$\begin{split}&\mathbb{E}\left[\left\|\frac{1}{m}\sum_{i}\bm{w}_{t+1,k}^{(i)}-\bm{w}_{t+1,*}^{(i)}\right\|^{2}\Big{|}\mathcal{F}_{t+1,k}\right]\\ &\leq\left(1-\eta_{t}\mu_{\ell}\right)\left\|\frac{1}{m}\sum_{i}\bm{w}_{t+1,k-1}^{(i)}-\bm{w}_{t+1,*}^{(i)}\right\|^{2}\\ &+2\eta_{t}(L_{\ell}\eta_{t}-1)\left(\frac{1}{m}\sum_{i}\left(h_{i}(\bm{w}_{t+1,k-1}^{(i)},\bm{w}_{t+1}^{(i)})-h_{i}(\bm{w}_{t+1,*}^{(i)},\bm{w}_{t+1}^{(i)})\right)\right)\\ &+\eta_{t}^{2}\frac{\sigma^{2}}{Bm}.\end{split}$ (140)
+If we assume $\eta_{t}\leq\frac{1}{L_{\ell}}$ we can simplify (140) as
+$\begin{split}&\mathbb{E}\left[\left\|\frac{1}{m}\sum_{i}\bm{w}_{t+1,k}^{(i)}-\bm{w}_{t+1,*}^{(i)}\right\|^{2}\Big{|}\mathcal{F}_{t+1,k}\right]\\ &\leq\left(1-\mu_{\ell}\eta_{t}\right)\left\|\frac{1}{m}\sum_{i}\bm{w}_{t+1,k-1}^{(i)}-\bm{w}_{t+1,*}^{(i)}\right\|^{2}+\eta_{t}^{2}\frac{\sigma^{2}}{Bm}\\ &\leq\left(1-\mu_{\ell}\eta_{t}\right)^{k}\left\|\frac{1}{m}\sum_{i}\bm{w}_{t,k}^{(i)}-\bm{w}_{t,*}^{(i)}+\bm{w}_{t,*}^{(i)}-\bm{w}_{t+1,*}^{(i)}\right\|^{2}+\sum_{l=0}^{k-1}\left(1-\mu_{\ell}\eta_{t}\right)^{k-l}\eta_{t}^{2}\frac{\sigma^{2}}{Bm}\\ &\leq(1+\epsilon)\left(1-\mu_{\ell}\eta_{t}\right)^{k}g_{t}+\sum_{l=0}^{k-1}\left(1-\mu_{\ell}\eta_{t}\right)^{k-l}\eta_{t}^{2}\frac{\sigma^{2}}{Bm}+\left(1+\frac{1}{\epsilon}\right)\left(1-\mu_{\ell}\eta_{t}\right)^{k}L_{w}^{2}\gamma^{2}\left(2\|\nabla F(\bm{w}_{t}^{(g)})\|^{2}+2\lambda^{2}g_{t}\right).\end{split}$ (141)
+Reorganizing all the terms in (141), we get
+$\begin{split}&\mathbb{E}\left[\left\|\frac{1}{m}\sum_{i}\bm{w}_{t+1,k}^{(i)}-\bm{w}_{t+1,*}^{(i)}\right\|^{2}\Big{|}\mathcal{F}_{t+1,k}\right]\\ &\leq\left[1+\epsilon+2\left(1+\frac{1}{\epsilon}\right)L_{w}^{2}\gamma^{2}\lambda^{2}\right]\left(1-\eta_{t}\mu_{\ell}\right)^{k}g_{t}+2\left(1+\frac{1}{\epsilon}\right)\left(1-\mu_{\ell}\eta_{t}\right)^{k}L_{w}^{2}\gamma^{2}\|\nabla F(\bm{w}_{t}^{(g)})\|^{2}+\frac{2\eta_{t}}{\mu_{\ell}}\frac{\sigma^{2}}{Bm}.\end{split}$ (142)
+Theorem 3.
+Suppose Assumptions 1, 2 and 4 hold.
+Let $\{{\bm{w}}_{t}^{(g)}\}_{t\geq 0}$ and $\{{\bm{w}}_{t,K}^{(i)}\}_{t\geq 0}$ be the sequence generated by Algorithm 2 with $\gamma<1/L_{g}$, $\eta\leq 1/L_{\ell}$, and the inner loop iteration number satisfying
+$\Big{(}2+64L_{w}^{2}(1/\mu_{g})^{2}\lambda^{2}\Big{)}(1-\eta\mu_{\ell})^{K}\leq(1-\gamma L_{g})^{4},$ (143)
+then ${\bm{w}}_{t}^{(g)}$ converges to $\widetilde{\bm{w}}^{(g)}$ linearly at rate $21-(\gamma\mu_{g})/2-(\gamma\mu_{g})^{2}/$ 2 within a neighbor of
+$\frac{8}{\textit{min}\{\frac{1}{2}\gamma\mu_{g},\mu_{\ell}\eta\}}\frac{1}{\mu_{g}^{2}}\lambda^{2}\eta^{2}\frac{\sigma^{2}}{Bm}.$
+Proof.
+Analogy to the proof of Theorem 2, we define
+$q:=\left(1+\epsilon+2\left(1+\frac{1}{\epsilon}\right)L_{w}^{2}\gamma^{2}\lambda^{2}\right)\left(1-\eta_{t}\mu_{\ell}\right)^{k}.$ (144)
+If we assume
+$q$ $\leq 1-\gamma\mu_{g}\leq 1-\gamma\mu_{g}+\gamma\epsilon_{g},\quad$ (145)
+and define
+$c_{1}=\frac{1}{1-\gamma\mu_{g}+\gamma\epsilon_{g}-q}\left(\frac{\gamma}{\epsilon_{g}}+\gamma^{2}(2+\zeta^{-1})\right)\lambda^{2},$ (146)
+combining Lemma 7 and Lemma 6, it yields
+$\begin{split}&\mathbb{E}\left[\left\|\bm{w}_{t+1}^{(g)}-\tilde{\bm{w}}^{(g)}\right\|^{2}\Big{|}\mathcal{F}_{t+1,k}\right]+c_{1}g_{t+1}\\ &\leq(1-\gamma\mu_{g}+\gamma\epsilon_{g})\left(\left\|\bm{w}^{(g)}_{t}-\tilde{\bm{w}}^{(g)}\right\|^{2}+c_{1}g_{t}\right)-\gamma^{2}(1-\gamma L_{g}(1+\zeta))\left\|\nabla F(\bm{w}_{t}^{(g)})\right\|^{2}\\ &\quad+\left(\frac{2}{1-\gamma\mu_{g}+\gamma\epsilon_{g}-q}\left(\frac{\gamma}{\epsilon_{g}}+\gamma^{2}(2+\zeta^{-1})\right)\lambda^{2}\right)2\left(1+\frac{1}{\epsilon}\right)\left(1-\mu_{\ell}\eta_{t}\right)^{k}L_{w}^{2}\gamma^{2}\left\|\nabla F(\bm{w}_{t}^{(g)})\right\|^{2}\\ &\quad+c_{1}\frac{2\eta_{t}}{\mu_{\ell}}\frac{\sigma^{2}}{Bm}.\end{split}$
+For $\left\|\nabla F(\bm{w}_{t}^{(g)})\right\|^{2}$, it suffices to show
+$\frac{4\gamma}{1-\gamma\mu_{g}+\gamma\epsilon_{g}-q}\left(\frac{1}{\epsilon_{g}}+\gamma(2+\zeta^{-1})\right)\lambda^{2}(1+\frac{1}{\epsilon})\left(1-\mu_{\ell}\eta_{t}\right)^{k}L_{w}^{2}\leq 1-\gamma L_{g}(1+\zeta).$
+If we set $\epsilon_{g}=\frac{\mu_{g}}{2}$, it yields the convergence rate
+$1-\frac{\gamma\mu_{g}}{2}.$ (148)
+And the condition (B.5) can be reformulated as
+$\frac{4}{\epsilon_{g}}\left(\frac{1}{\epsilon_{g}}+\gamma(2+\zeta^{-1})\right)\lambda^{2}(1+\epsilon^{-1})\left(1-\mu_{\ell}\eta_{t}\right)^{k}L_{w}^{2}\leq 1-\gamma L_{g}(1+\zeta).$
+Restrict $\gamma$
+$\gamma:$ $\quad\gamma(2+\zeta^{-1})\leq\frac{1}{\epsilon_{g}}.$ (150)
+Then condition (B.5) yields
+$\frac{8}{\epsilon_{g}^{2}}\lambda^{2}(1+\frac{1}{\epsilon})\left(1-\mu_{\ell}\eta_{t}\right)^{k}L_{w}^{2}\leq 1-\gamma L_{g}(1+\zeta).$ (151)
+Then if we assume $\epsilon=1$ and $\zeta=1-\gamma L_{g}$, the conditions (145), (150) and (151) can be satisfied with the following sufficient conditions
+$\begin{split}&\gamma<\frac{1}{L_{g}},\\ &\eta_{t}\leq\frac{1}{L_{\ell}},\\ &(2+4L_{w}^{2}\gamma^{2}\lambda^{2})\left(1-\eta_{t}\mu_{\ell}\right)^{k}\leq 1-\gamma\mu_{g},\\ &\frac{64}{\mu_{g}^{2}}\left(1-\eta_{t}\mu_{\ell}\right)^{k}L_{w}^{2}\lambda^{2}\leq(1-\gamma L_{g})^{2}.\end{split}$ (152)
+And we can get a more straightforward sufficient condition
+$\gamma<\frac{1}{L_{g}},$ (153)
+$\eta_{t}\leq\frac{1}{L_{\ell}},$ (154)
+$\left(2+64\frac{L_{w}^{2}}{\mu_{g}^{2}}\lambda^{2}\right)\left(1-\eta_{t}\mu_{\ell}\right)^{k}\leq(1-\gamma L_{g})^{2}.$ (155)
+Under such a sufficient condition, the contraction of the combination of inner loop and the outer loop would be rewritten as
+$\begin{split}&\mathbb{E}\left[\left\|\bm{w}_{t+1}^{(g)}-\tilde{\bm{w}}^{(g)}\right\|^{2}\Big{|}\mathcal{F}_{t+1,k}\right]+c_{1}g_{t+1}\\ &\leq(1-\gamma\mu_{g}+\gamma\epsilon_{g})\left(\left\|\bm{w}^{(g)}_{t}-\tilde{\bm{w}}^{(g)}\right\|^{2}+c_{1}g_{t}\right)\quad+\sum_{l=0}^{t}\left(1-\frac{1}{2}\gamma\mu_{g}\right)^{t-l}\frac{2}{\epsilon_{g}^{2}}\lambda^{2}\sum_{l=0}^{k-1}\left(1-\mu_{\ell}\eta\right)^{k-l}\eta^{2}\frac{\sigma^{2}}{Bm}\\ &\leq(1-\frac{1}{2}\gamma\mu_{g})^{t+1}\left(\left\|\bm{w}^{(g)}_{0}-\tilde{\bm{w}}^{(g)}\right\|^{2}+c_{1}g_{0}\right)+\sum_{l=0}^{t}\left(1-\frac{1}{2}\gamma\mu_{g}\right)^{t-l}\frac{2}{\epsilon_{g}^{2}}\lambda^{2}\sum_{l=0}^{k-1}\left(1-\mu_{\ell}\eta\right)^{k-l}\eta^{2}\frac{\sigma^{2}}{Bm}\\ &\leq(1-\frac{1}{2}\gamma\mu_{g})^{t+1}\left(\left\|\bm{w}^{(g)}_{0}-\tilde{\bm{w}}^{(g)}\right\|^{2}+c_{1}g_{0}\right)+\sum_{l=0}^{tk-1}2\left(\text{max}(1-\frac{1}{2}\gamma\mu_{g},1-\mu_{\ell}\eta)\right)^{l}\frac{1}{\epsilon_{g}^{2}}\lambda^{2}\eta^{2}\frac{\sigma^{2}}{Bm}\\ &\leq(1-\frac{1}{2}\gamma\mu_{g})^{t+1}\left(\left\|\bm{w}^{(g)}_{0}-\tilde{\bm{w}}^{(g)}\right\|^{2}+c_{1}g_{0}\right)+\frac{8}{\text{min}\{\frac{1}{2}\gamma\mu_{g},\mu_{\ell}\eta\}}\frac{1}{\mu_{g}^{2}}\lambda^{2}\eta^{2}\frac{\sigma^{2}}{Bm}.\end{split}$ (156)
+With $\eta=\text{min}\{\frac{\mu(\lambda+L)}{4L(\lambda+\mu)^{2}},\frac{1}{\lambda+L}\}$, $\gamma=1/(2L_{g})=\frac{\lambda+L}{2\lambda L}$, we can get the solution within an error ball $\left(\mathcal{O}(\varepsilon)+\text{min}\{\frac{1}{4\mu L},\frac{\lambda+\mu}{\mu^{2}(\lambda+L)}\}\frac{\sigma^{2}}{Bm}\right)$ with the communication cost $\mathcal{O}\big{(}\kappa_{g}\log(1/\varepsilon)\big{)}=\mathcal{O}\left(\kappa\cdot\frac{\lambda+\mu}{\lambda+L}\cdot\log\frac{1}{\varepsilon}\right)$ and the computation cost $\tilde{\mathcal{O}}\left(\kappa\cdot\log\frac{1}{\varepsilon}\right)$.
+As stated in Theorem 4.2 of (Hanzely & Richtárik, 2020), when $p^{\star}=\frac{\lambda}{\lambda+L}$, the stochastic noise term is independent of the number of clients.
+Consequently, it does not theoretically demonstrate the expected benefit of collaboration in reducing noise.
+In contrast, our analysis reveals that even when adjusting $\lambda$, the noise remains of the same order, specifically $\mathcal{O}\left(\frac{1}{Bm}\right)$.
+Appendix C Proof in Section 5
+Corollary 4.
+Under the assumptions of Theorem 1 and Corollary 2, ${\bm{w}}_{T,K}^{(i)}$ generated by Algorithm 1 satisfies
+$\begin{split}&\mathbb{E}\left\|{\bm{w}}_{T,K}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\right\|^{2}\\ &\leq\left(1-\frac{1}{\kappa}-\frac{L-\mu}{\kappa(\lambda+\mu)}\right)^{T}\mathbb{E}\left\|{\bm{w}}_{0,0}^{(i)}-\widetilde{\bm{w}}^{(i)}\right\|^{2}\\ &+\mathcal{O}\left[\left(\frac{1}{N}+R^{2}+\frac{1}{q_{1}(\lambda)}\right)\wedge\left(\frac{1}{n}+q_{2}(\lambda)\right)\right]\end{split}$ (157)
+for any $i\in[m]$, where $q_{1}(\lambda)$ and $q_{2}(\lambda)$ are monotone increasing functions of $\lambda$ with $\lim_{\lambda\to+\infty}q_{1}(\lambda)=+\infty$ and $\lim_{\lambda\to 0}q_{2}(\lambda)=0$.
+Proof.
+The corollary is a direct result of Theorem 1 and Theorem 2, where $q_{1}(\lambda)$ and $q_{2}(\lambda)$ are defined in (70) and (67).
+Appendix D Additional Experiment Details and Results
+D.1 More Experiment Details
+Federated Learning (FL) has emerged as a widely adopted approach with numerous real-world applications, like Large Language Model (LLM) (Bai et al., 2024; Wu et al., 2024; Zhao et al., 2023), Reinforcement Learning (Zheng et al., 2024, 2025; Zhang et al., 2025), Diffusion Model (Li et al., 2024a).
 In this section, we provide more details on the implementation of the methods used in the empirical analysis.
@@
 The Balanced split contains 131,600 28x28 grayscale images across 47 balanced classes (digits and uppercase/lowercase letters).
-And the CIFAR-10 dataset contains 60,000 color images (32x32 pixels, 3 RGB channels), categorized into 10 classes representing real-world objects (e.g., airplane, car, and dog).
+The CIFAR-10 dataset contains 60,000 color images (32x32 pixels, 3 RGB channels), categorized into 10 classes representing real-world objects (e.g., airplane, car, and dog).
 Algorithms.
-To fully investigate the effect of personalization in FedProx with the Algorithm B.1, we compare the two extreme algorithms (GlobalTraining and LocalTraining).
-Furthermore, since there exist other methods solving the objective (1) to do personalized federated learning, we also try to investigate the effect of personalization in other methods, like pFedMe (T Dinh et al., 2020).
+To fully investigate the effect of personalization in Problem (2) with the Algorithm B.1, we compare the two extreme algorithms (GlobalTraining and LocalTraining).
+Furthermore, since there exist other methods solving Problem (2) to do personalized federated learning, we also try to investigate the effect of personalization in other methods, like pFedMe (T Dinh et al., 2020).
 The algorithm pFedMe has a similar design with a double loop to solve the subproblem in each communication round.
@@
 Such a choice is to make sure that each line in the figure is clearly separable.
-Additionally, we implemented an adaptive choice of $\lambda$ following the formula outlined in Theorem 1, with $\rho$ set to 2.
+Additionally, we implemented an adaptive choice of $\lambda$ following the formula outlined in Corrollary 1, with $\rho$ set to 2.
 The value of $\rho$ was determined through grid search over the range $[1,10]$ to ensure optimal performance.
@@
 Selection of Step Size.
-For the synthetic dataset, aligned with Theorem 2, we set the global step size $\gamma=\frac{\lambda+L}{\lambda L}$ and the local step size $\eta=\frac{1}{L+\lambda}$.
+For the synthetic dataset, aligned with Theorem 2, we set the global step size $\gamma=(\lambda+L)/(\lambda L)$ and the local step size $\eta=1/(L+\lambda)$.
 The smoothness constant $L$ is computed as the upper bound of the $L$-smoothness constant for logistic regression, specifically $4^{-1}\Lambda_{\text{max}}$, where $\Lambda_{\text{max}}$ denotes the largest eigenvalue of the Gram matrix $\bm{X}^{\top}\bm{X}$.
 The local step size for the synthetic dataset was further tuned through grid search in $\{0.001,0.01,0.1\}$.
-For the real dataset, the global learning rate was set to $\frac{1}{\lambda}$, while the local learning rate was set to 0.01, chosen via grid search.
+For the real dataset, the global learning rate was set to $1/\lambda$, while the local learning rate was set to 0.01, chosen via grid search.
 For both GlobalTrain and LocalTrain training baselines, the learning rate was also chosen via grid search within the range $\{0.001,0.01,0.1\}$.
@@
 For the synthetic dataset, following the analysis in the paper, the loss function is optimized using gradient descent (GD).
-For the real dataset, all models are implemented in PyTorch, and optimization is performed using the stochastic gradient descent (SGD) solver.
+For the MNIST dataset, all models are implemented in PyTorch, and optimization is performed using the stochastic gradient descent (SGD) solver.
 Evaluation Criterion.
-In terms of evaluation, for the synthetic dataset, we track the ground-truth global model ${{\bm{w}}_{\star}^{(g)}}$ and local model ${{\bm{w}}_{\star}^{(i)}}$.
-We compute the global statistical error as $\|\widetilde{\bm{w}}^{(g)}-{{\bm{w}}_{\star}^{(g)}}\|^{2}$ and the local statistical error as $\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}$.
-To capture the total error at the $t$-th iteration, we report both the global total error $\|\bm{w}_{t}^{(g)}-{{\bm{w}}_{\star}^{(g)}}\|^{2}$ and the local total error $\|{\bm{w}}_{t,0}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}$, which combine both optimization and statistical errors.
+In terms of evaluation, for the synthetic dataset, we track the ground truth local model ${{\bm{w}}_{\star}^{(i)}}$.
+We compute the local statistical error as $\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}$.
+To capture the total error at the $t$-th iteration, we report the local total error $\|{\bm{w}}_{t,0}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}$, which combine both optimization and statistical errors.
 For the MNIST dataset, we report test accuracy on the MNIST testing set as a measure of the statistical error, and training loss for each client’s training set as total error at the $t$-th iteration.
-C.2 Additional Experiment Results
-Figure 3: Performance of the FedProx algorithm with an adaptive choice of $\lambda$ (from Theorem 1) under varying levels of statistical heterogeneity.
-In Figure 3, statistical error of the FedProx algorithm with an adaptive choice of $\lambda$ under varying levels of statistical heterogeneity.
-With the adaptive choice of $\lambda$ as defined in Theorem 1, FedProx consistently achieves the lowest statistical error across most scenarios, outperforming both GlobalTrain and LocalTrain baselines.
-This is particularly evident under moderate and high levels of statistical heterogeneity.
-However, as $R$ approaches very small values, the global baseline performs slightly better, which aligns with the slower rate of FedProx with small $R$ described in our theoretical results.
-Overall, these results validate our choice of adaptive $\lambda$ and confirm the statistical rate provided by our analysis.
-Figure 4: (a) Testing accuracy of different methods (FedProx, LocalTrain, GlobalTrain) versus statistical heterogeneity levels on the MNIST dataset (logistic regression).
-(b) Training loss versus communication rounds for the same methods.
-In Figure 4(a), we compare FedProx, LocalTrain, and GlobalTrain on the MNIST dataset by plotting testing accuracy against varying levels of statistical heterogeneity.
-As heterogeneity increases, each client has access to fewer classes, resulting in fewer classification tasks.
-This change simplifies the problem, which explains why all methods show improved evaluation accuracy as the number of classes decreases.
-However, this does not diminish the insights from the relative comparison between methods.
-GlobalTrain and FedProx with larger $\lambda$ perform better at lower heterogeneity levels, while LocalTrain and FedProx with smaller $\lambda$ perform better as heterogeneity increases.
-In Figure 4(b), we track testing loss as a function of communication rounds.
-Across all sub-figures, FedProx with smaller $\lambda$ (higher personalization) demonstrates faster convergence, especially in the early stages of training.
-The final training loss achieved by different methods also aligns with the results shown in Figure 4(a) and Figure 1.
-Figure 5: Performance of FedProx with 100 clients (top row) and 1,000 clients (bottom row) with 10% client sampling under different levels of statistical heterogeneity ($R$).
-In Figure 5, the top row shows results for 100 clients, and the bottom row shows results for 1,000 clients, and in both results we adopt a client sampling of 10% in each communication round.
-This tests the scalability and generalizability of the our results under typical federated learning setups.
-Consistent with the findings from Figure 1, FedProx with smaller $\lambda$ (greater personalization) consistently achieves faster convergence at the early stages of training, while FedProx with larger $\lambda$ (less personalization) takes more communication rounds to converge.
-As statistical heterogeneity increases (measured by $R$), FedProx with smaller $\lambda$ gradually outperforms models with larger $\lambda$, showcasing the adaptability of FedProx to different levels of heterogeneity.
-These results further validate the theoretical insights from Theorem 1 and Theorem 2, highlighting the influence of personalization on statistical and optimization performance in federated setups with larger numbers of clients.
-Figure 6: Effect of local updates on FedProx performance for local models in MNIST logistic regression and global models ni MNIST CNN, complementing the results shown in Figure 2.
-Figure 6 demonstrates the impact of varying the number of local updates on FedProx performance, complementing the results shown in Figure 2.
-For FedProx with smaller $\lambda$ (higher personalization), increasing the number of local updates notably reduces the need for communication, leading to faster convergence.
-In contrast, for FedProx with larger $\lambda$ (lower personalization), increasing local updates yields minimal benefit and only adds to computation costs without reducing communication demands.
-The results indicate that for larger $\lambda$ , the optimal strategy is to prioritize frequent communication rather than increasing local computation.
-This highlights the trade-off between local computation and communication efficiency under different levels of personalization, validating our theoretical findings in Corollary 1.
-Dataset Client Class One-stage FedProx GlobalTrain LocalTrain pFedMe
-(small) (medium) (large) (small) (medium) (large)
+D.2 Additional Experiment Results
+The Effect of Personalization on Accuracy over different data heterogeneity To further validate our conclusions about the effect of personalization on solving Problem 2, we conduct additional experiments on real datasets, including EMNIST and CIFAR-10.
+For each dataset, we explore three levels of statistical heterogeneity, controlled by varying the number of classes each client has access to, which follows a widely used setup in federated learning literature.
+The models are trained using different convolutional neural network (CNN) architectures, denoted as CNN2 and CNN5, where the numbers indicate the number of convolutional layers in the network.
+As the ground truth underlying model is not accessible for these real datasets, we evaluate model performance based on classification accuracy on the test set, as shown in Tables 4.
+We also conduct experiments another regularization-based personalization method focusing on Problem (2), pFedMe (T Dinh et al., 2020), to demonstrate that our results could be generalized to other regularization-based personalized federated learning methods as well.
+For the columns corresponding to FedCLUP and pFedMe, we conduct experiments with varying regularization parameter $\lambda$, investigating the role of personalization in the algorithms.
+Specifically, we experiment with small, medium, and large $\lambda$ values (see Appendix D.1 for detailed parameter settings).
+Dataset Client Class FedCLUP GlobalTrain LocalTrain pFedMe
+high median low high median low
+MNIST Logit 2 0.5001 0.6011 0.6109 0.7001 0.3214 0.5534 0.6312 0.6712
+6 0.6813 0.7001 0.7241 0.7718 0.6000 0.6312 0.7123 0.7681
+10 0.8123 0.8213 0.8312 0.8391 0.7828 0.8001 0.8239 0.8291
 MNIST CNN2 2 0.3567 0.4721 0.6019 0.6123 0.3019 0.3973 0.5828 0.6172
@@
 […]
-To further validate our conclusions about the effect of personalization on the statistical accuracy of FedProx (Objective 1), we conduct additional experiments on real datasets, including EMNIST and CIFAR-10.
-For each dataset, we explore three levels of statistical heterogeneity, controlled by varying the number of classes each client has access to, which follows a widely used setup in federated learning literature.
-The models are trained using different convolutional neural network (CNN) architectures, denoted as CNN2 and CNN5, where the numbers indicate the number of convolutional layers in the network.
-As the true underlying model is not accessible for these real datasets, we evaluate model performance based on classification accuracy on the test set, as shown in Tables 3 and 4.
-We also conduct experiments another regularization-based personalization method, pFedMe (T Dinh et al., 2020), to demonstrate that our results could be generalized to other regularization-based personalized federated learning methods as well.
-For the columns corresponding to One-stage FedProx and pFedMe, we conduct experiments with varying regularization parameter $\lambda$, investigating the role of personalization in the algorithms.
-Specifically, we experiment with small, medium, and large $\lambda$ values (see Appendix C.1 for detailed parameter settings).
-In Table 3, we can see that consistently across datasets, as $\lambda$ decreases (indicating a higher degree of personalization), the accuracy approaches that of LocalTrain, which trains separate models for each client.
-Conversely, as $\lambda$ increases (indicating lower personalization and more collaboration), the accuracy approaches that of GlobalTrain, which represents purely global training.
-This interpolation behavior between LocalTrain and GlobalTrain, enabled by FedProx’s personalized objective, aligns well with our theoretical results in Theorem 1 and highlights the adaptive capability of the FedProx method.
+In Table 4, we can see that consistently across datasets, as personalization degree increase, the accuracy approaches that of LocalTrain, which trains separate models for each client.
+Conversely, as personalization degree decreases, the accuracy approaches that of GlobalTrain, which represents purely global training.
+This interpolation behavior between LocalTrain and GlobalTrain, enabled by Problem (2), aligns well with our theoretical results in Theorem 1.
 Moreover, as dataset homogeneity increases (i.e., clients share more similar data distributions), we observe a general improvement in accuracy across all methods, regardless of the level of personalization.
@@
 Additionally, experiments with different CNN architectures confirm the robustness of these findings across model designs.
-Although we did not provide a theoretical analysis for pFedMe in this paper, our results suggest that the empirical behavior of pFedMe is consistent with the theoretical insights derived for FedProx.
+Although we did not provide a theoretical analysis for pFedMe in this paper, our results suggest that the empirical behavior of pFedMe is consistent with the theoretical insights derived for FedCLUP.
 In particular, we observe that pFedMe also exhibits interpolation between local and global learning under varying levels of data heterogeneity.
-Dataset Client Class One-stage FedProx (small $\lambda$) One-stage FedProx (large $\lambda$) Dichotomous Strategy
-MNIST CNN2 2 0.4721 0.6019 0.2567
-MNIST CNN2 6 0.7833 0.7843 0.7001
-MNIST CNN2 10 0.8893 0.9032 0.8753
-MNIST CNN5 2 0.8891 0.9333 0.7032
-MNIST CNN5 6 0.8451 0.8941 0.7941
-MNIST CNN5 10 0.9091 0.9102 0.8922
-EMNIST CNN5 30 0.6084 0.6089 0.5291
-EMNIST CNN5 40 0.6097 0.6123 0.5732
-EMNIST CNN5 47 0.6278 0.6415 0.5856
-Table 4: Comparison of evaluation accuracy for One-stage FedProx with different values of $\lambda$ and Dichotomous Strategy across various datasets and client classes.
-C.3 Comparison of FedProx with A Dichotomous Strategy
-A simple baseline in personalized federated learning is the dichotomous strategy, where the dataset on each client is split into two subsets: one is used for global training (GlobalTrain) with pooled client data, while the other is used for purely local training (LocalTrain).
-After training, the two models are compared on the test set, and the best-performing model is selected for each client.
-Such a strategy is known to achieve minimax-optimal statistical accuracy under certain conditions; see Chen et al.
-(2023b) for a detailed discussion.
-Even though the convergence rate coincides with such a dichotomous strategy, we may reiterate that the key contribution of our work lies in the fine-grained analysis of the FedProx method—a method that has demonstrated strong empirical performance in federated learning [2].
-Through our analysis, we have, for the first time, established the minimax optimality of FedProx, showing that its statistical performance matches the lower bound of the estimation problem.
-If one seeks to understand why FedProx could be advantageous over the dichotomous strategy in practice, here are some of the reasons:
-  • Stability in Heterogeneous Environments: Empirical studies [1] demonstrate that FedProx achieves more stable convergence compared to methods such as FedAvg, especially in settings with significant in a heterogeneous network.
-  • Adaptivity and Flexibility: FedProx seamlessly interpolates between the extremes of global training and local training by tuning a single regularization parameter, $\lambda$.
-This adaptability allows practitioners to dynamically balance the trade-off between bias and variance based on the level of statistical heterogeneity, without the need for fixed, hard decisions inherent in the dichotomous strategy.
-To further compare FedProx with the dichotomous strategy, we conduct a simple experiment.
-In the dichotomous setup, each client’s dataset is partitioned into two subsets: one subset is used for local training (LocalTrain), where each client trains independently, and the other subset is pooled across all clients for global training (GlobalTrain).
-The accuracy of the two resulting models is evaluated on a test set, and the better-performing model is reported for each client.
-By contrast, FedProx is implemented using the experimental setup described earlier, leveraging the full dataset.
-As shown in Table 4, FedProx consistently outperforms the dichotomous strategy across various configurations.
-This performance gap can be attributed to the inefficiency of the dichotomous approach, which sacrifices valuable data during the sample splitting process, resulting in reduced sample efficiency.
-Even in a hypothetical scenario where the full sample is used for both LocalTrain and GlobalTrain in the dichotomous strategy, the results in Table 3 demonstrate that FedProx achieves a continuous interpolation between the two extremes of LocalTrain and GlobalTrain by adjusting its regularization parameter $\lambda$.
-This aligns with our theoretical findings.
+We observe in Table 4 that LocalTrain consistently underperforms across all levels of statistical heterogeneity.
+This is because, while statistical heterogeneity is introduced by restricting each client’s exposure to only a subset of classes in the training process, evaluation is conducted on the entire dataset, including classes not seen during training.
+As a result, LocalTrain’s performance remains lower than that of other methods.
+The Effect of Personalization on Statistical Error over different data heterogeneity Figure 4 demonstrates the effect of personalization on statistical error across varying levels of statistical heterogeneity $R$.
+As demonstrated in the theory, increasing personalization (moving from FedCLUP (low per) to FedCLUP (high per)) shifts the solution closer to LocalTrain, while decreasing personalization makes the model behave more like GlobalTrain.
+For low heterogeneity, methods with lower personalization (i.e., greater collaboration) achieve lower statistical error, as information can be effectively shared across clients to improve learning.
+GlobalTrain achieves the lowest error in this regime, while FedCLUP with less personalization also performs well.
+However, as statistical heterogeneity increases, collaborative learning becomes less effective.
+In this setting, LocalTrain outperforms other methods.
+Highly personalized FedCLUP also performs better than less personalized variants, aligning with the theory we established in Theorem 1.
+Figure 4: Impact of personalization on statistical error of FedCLUP solution across different levels of heterogeneity (high per means training FedCLUP with high personalization degree).
+In Figure 5, we track testing loss as a function of communication rounds.
+Across all sub-figures, FedCLUP with smaller $\lambda$ (higher personalization) demonstrates faster convergence, especially in the early stages of training.
+Thus, it shows that the trade-off between statistical accuracy and computation cost under the influence of personalization widely exists over statistical heterogeneity.
+Figure 5: Logistic Regression training loss versus communication rounds for the FedCLUP methods on MNIST dataset (high per refers to training FedCLUP with high personalization degree).
+Figure 6: Performance of FedCLUP on the Synthetic dataset with 100 clients (top row) and 1,000 clients (bottom row) with 10% client sampling under different levels of statistical heterogeneity ($R$).
+Different Number of Clients In Figure 6, the top row shows results for 100 clients, and the bottom row shows results for 1,000 clients, and in both results we adopt a client sampling of 10% in each communication round.
+This tests the scalability and generalizability of the our results under typical federated learning setups.
+Consistent with the findings from Figure 1, FedCLUP with smaller $\lambda$ (higher personalization) consistently achieves faster convergence at the early stages of training, while FedCLUP with larger $\lambda$ (lower personalization) takes more communication rounds to converge.
+As statistical heterogeneity increases (measured by $R$), FedCLUP with smaller $\lambda$ gradually outperforms models with larger $\lambda$, showcasing the adaptability of solving Problem 2 to different levels of heterogeneity.
+These results further validate the theoretical insights from Theorem 1 and Theorem 2, highlighting the influence of personalization on statistical and optimization performance in federated setups with larger numbers of clients.
+The Effect of Personalization on Communication Cost Figure 2 investigates how personalization impacts communication efficiency in FedCLUP by analyzing the benefit of increasing local updates under different personalization levels.
+In the low-personalization setting (left column), more local updates significantly accelerate convergence, reducing the reliance on frequent communication.
+However, in the high-personalization setting (right column), increasing local updates has a limited effect on convergence, indicating that frequent communication is essential for effective learning.
+This demonstrates that higher personalization requires more communication rounds to achieve comparable performance, aligning with our theoretical findings.
+D.3 Additional Experiment for Rebuttal
+D.3.1 Strongly Convex Setting with Synthetic Data
+As the theoretical results is established under strongly-convex settings, hence some expiremnts under strongly convex setting is expected with discussion related to the conditional number $\kappa$.
+We conduct an experiment on strongly convex problem: an overdeterminded linear regression task.
+We strictly follow the choice of local step size, local computation rounds, and global step size as specified in Corollary 2.
+As shown in Table 7 and 8, for a fixed personalization parameter $\lambda$, we observe that longer value of $\kappa$ result in slower convergence rates with respect tot he number of communication rounds.
+This expirical trend is consistent with our theoretical prediction in Corollary 2, where the number of communication rounds required to achieve a given target error $\epsilon$ scales with $\mathcal{O}(\kappa\frac{\lambda+L}{\lambda+\mu}\log(1/\varepsilon))$.
+Moreover, we observe that the impact of increasing $\kappa$ becomes stronger as $\lambda$ increases.
+This phenomenon aligns with our theoretical analysis, which shows that the sensitivity of communication complexity to $\kappa$ is amplified for larger values of $\lambda$.
+Similarly, we observe that increasing the condition number $\kappa$ also leads to a higher total number of gradient evaluations required to reach a target error.
+This observation is in agreement with our theoretical results where the total number of local gradient evaluations scales as $\mathcal{O}(\kappa\log(1/\varepsilon))$.
+Figure 7: Optimization error versus communication rounds under varying condition numbers $\kappa$.
+The left, middle, and right panels correspond to FedCLUP with small, medium, and large values of $\lambda$, respectively.
+Across all settings, we observe that larger $\kappa$ leads to slower convergence, with the effect becoming more pronounced as $\lambda$ increases.
+Figure 8: Optimization error versus total number of local gradient evaluations per client under varying condition numbers $\kappa$.
+As $\kappa$ increases, the number of gradient evaluations required to achieve a given optimization error also increases.
+D.3.2 Datasets with Natural Partitions
+(a) The loss of FedCLUP with CelebA
+(b) The loss of pFedMe with CelebA
+Figure 9: Loss of FedCLUP (a) and pFedMe (b) with different personalization degrees on the CelebA dataset.
+Low, median, and high personalization correspond to regularization terms 0.5, 0.1, and 0.001, respectively.
+For FedCLUP, the test accuracy is highest with low personalization (0.919), followed by median (0.915) and high (0.888).
+For pFedMe, the test accuracy is highest for low personalization (0.910), followed by median (0.901) and high (0.890).
+The test accuracy for GlobalTrain and LocalTrain is 0.911 and 0.561, respectively.
+(a) The loss of FedCLUP with Sent140
+(b) The loss of pFedMe with Sent140
+Figure 10: Loss of FedCLUP (a) and pFedMe (b) with different personalization degrees on the Sent140 dataset.
+Low, median, and high personalization correspond to regularization terms 0.1, 0.005, and 0.001, respectively.
+For FedCLUP, the test accuracy is highest for low personalization (0.734), followed by median (0.702) and high (0.696).
+For pFedMe, the test accuracy is highest for low personalization (0.727), followed by median (0.703) and high (0.701).
+The test accuracy for GlobalTrain and LocalTrain is 0.730 and 0.353, respectively.